A Complete Bibliography of Publications in Computational Statistics & Data Analysis (1990–1999)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

02 October 2019 Version 1.04

Title word cross-reference

$107.50 [406]. $125.00 [405]. $129 [285]. $138.00 [717]. $150.00 [766]. $165.00 [745]. 2 [1406, 1020]. $24.95 [34]. $29.95 [35]. 2^{m} × 3^{n} [47]. 2^{nd} [1002, 1003]. 2 × 2 [1229, 1316, 204]. 2 × J [1029]. 2 × K [771]. 3 [50, 312]. $332.50/£ [534]. $348 [284]. $49.95 [1017, 116]. $57.20 [631]. $59.95 [683]. $63.50. [559]. $65.00 [645]. 6′ × 9′ [35]. $70.00 [1096]. $75 [283]. $89.75 [406]. $90.00 [923]. > [349]. ^{nd} [1001]. ^{R} [1085]. (R) [741]. ^{th} [1098]. A [553]. C_{p} [902]. χ^{2} [397, 665, 656]. D [795, 753, 314, 553, 804]. E [553]. F [707, 687, 74, 975]. G [127, 553, 72]. G^{⋆} [1286]. h [448]. I [1010, 487, 1276]. k [518]. κ^{2} [1085]. L [154].

L_{1} [199, 957, 298, 917, 95, 142, 582, 277]. L_{p} [108]. M [90, 315]. N [1313]. n ≡ 2 (mod 4) [1119]. O(n^{−2}) [428]. O(n log n) [912]. P [1229, 975, 1376, 687, 683]. Φ_{p} [11]. Q [1338]. R [1338]. R/S [714]. S [354]. T [1011, 1132, 1341, 1241, 1261, 758]. T (h, a) [758]. T_{3} [1208]. × [34]. U + V [294]. U × V [294]. W_{u} [51]. X^{2} [397].

-D [50]. -designs [312]. -dimensional [804]. -distribution [1261, 1011]. -DM45.- [617]. -estimation [154, 315]. -inverse [72]. -mode [1338]. -norm [199, 298, 108, 95]. -optimal [795, 753, 314, 553]. -plot [1208]. -probabilities [758]. -statistic [687]. -tests [975]. -type [90]. -Value [683, 1229, 11, 1376]. -values [687, 975]. -way [1313].

.pp [925].

0 [131, 590, 1297, 505, 785, 193, 589, 745, 746, 924, 1296]. 0-12-304752-8 [116]. 0-13-454810-8 [823]. 0-13-850652-3 [631]. 0-19-852204-5 [248]. 0-19-874180-4 [1021]. 0-387-91443-9 [504]. 0-387-94570-9 [1050]. 0-387-94660-8 [967]. 0-387-94679-9 [1020]. 0-412-04061-1 [1297]. 0-412-39920-2 [589]. 0-412-56310-X [785]. 0-412-71050-1 [986]. 0-412-71630-5 [855]. 0-412-71640-2 [882]. 0-412-98281-1 [1018]. 0-444-80934-1 [115]. 0-444-80988-0 [115]. 0-444-87410-0 [32, 131]. 0-444-88029-1 [33]. 0-444-88096-8 [503]. 0-471-00710-2 [728]. 0-471-04222-6 [924]. 0-471-10647-X [821]. 0-471-11030-2 [884]. 0-471-11338-7 [923]. 0-471-14191-7 [1096]. 0-471-15407-5 [1072]. 0-471-18245-1 [1296]. 0-471-31073-5 [1073]. 0-471-50280-4 [358]. 0-471-54641-0 [559]. 0-471-54897-9 [467]. 0-471-55239-9 [883]. 0-471-55761-7 [683]. 0-471-55779-X [925]. 0-471-55951-2 [633]. 0-471-56881-3 [535]. 0-471-57150-4 [848]. 0-471-57152-0 [822]. 0-471-57427-9 [846]. 0-471-58494-0 [847]. 0-471-58495-9 [746]. 0-471-59282-X [1097]. 0-471-82221-3 [987]. 0-471-82986-2 [190]. 0-471-96688-6 [1019]. 0-534-92026-8 [193]. 0-631-15943-6 [370]. 0-7923-0233-8 [213]. 0-7923-0928-6 [285]. 0-7923-1634-7 [717]. 0-7923-2155-3 [534]. 0-7923-2156-1 [534]. 0-7923-2157-X [534]. 0-7923-2356-4 [616]. 0-7923-2491-9 [505]. 0-8039-4984-7 [786]. 0-8039-4985-5 [786]. 0-8247-7600-3 [283]. 0-8247-7980-0 [557]. 0-8247-8354-9 [406]. 0-8247-8437-5 [766]. 0-8247-8682-3 [405]. 0-8247-8888-5 [645]. 0-8247-9077-4 [590]. 0-8247-9169-X [745]. 0-8247-9338-2 [965]. 0-8493-0132-7 [1017].

1 [133, 466, 1297, 1296]. 1-56081-348-2 [466]. 1-85058-047-2 [133]. 1-85166-484-X [284]. 10 [1098]. 110 [743]. 11th [1417, 284]. 128.00 [632]. 135.00 [1074]. 168 [821]. 18.99 [882]. 19.99 [986]. 198 [192]. 1987 [1414]. 1989 [1418]. 1990 [1421, 113]. 1991 [218]. 1992 [338, 1422]. 1994 [646]. 1995 [744, 1430]. 1996 [796, 788, 789]. 1997 [1001, 1003, 1070]. 1998 [1098]. 1999 [1191, 1298]. 1st [1420].

2 [133, 466, 589, 266]. 20 [743, 1125]. 20/1 [743]. 21.- [115]. 227 [1049]. 24.80 [407]. 24.95 [633]. 250.- [117]. 277 [855]. 28 [54]. 29.80 [373]. 2nd [303, 883, 467, 633, 746, 847, 1072].

3 [658, 466]. 3-11-012107-7 [192].

3-437-40268-4 [558]. 3-437-40278-1 [373]. 3-437-40280-3 [466]. 3-437-40281-1 [408]. 3-437-50320-0 [266]. 3-450-52493-3 [536]. 3-540-54169-1 [371]. 3-540-55821-7 [468]. 3-540-60676-9 [966]. 3-540-61081-2 [1074]. 3-540-97211-0 [194]. 3-540-97884-4 [658]. 3-7643-2832-0 [454]. 3-7643-2874-6 [451]. 3-7643-2923-8 [657]. 3-7785-2157-8 [372]. 3-7908-0486-X [617]. 3-7908-0593-9 [632]. 3-7908-0677-3 [504]. 3-7908-0700-1 [618]. 3-8244-2027-9 [452]. 3-8252-1293-9 [453]. 3-8252-1602-0 [407]. 3-8252-1632-2 [558]. 3-89028-132-X [249]. 31 [1339]. 314 [884]. 32.50 [1072]. 325 [846]. 33.50 [967]. 34.80 [453]. 35.00 [1018]. 364 [823]. 38 [452]. 39 [657]. 3rd [966, 1212, 728, 1050].

4 [590, 658]. 4-914903-03-2 [191]. 42.00 [468]. 449 [848]. 48th [188]. 49. [408]. 49.- [408]. 49.95 [728]. 4th [425, 193, 617, 1414].

50 [1021]. 50.00 [987, 1019, 1073]. 51st [972]. 54 [454]. 55.00 [883, 1097]. 58 [451]. 58.00 [1050]. 5th [882, 1212].

6 [1356, 924, 408]. 60th [1431]. 62-02 [925]. 62F03 [925]. 62H15 [925]. 68 [372]. 6th [1423].

719 [847]. 76. [266]. 76.- [266]. 78.00 [1020].

8 [54, 193]. 88 [194]. ’89 [266]. 89.- [658].

9 [505, 746]. ’90 [113, 114, 1394, 1050]. 90-90-09022-3 [1049]. ’91 [1423]. ’92 [1427, 534]. ’93 [425, 389, 464]. ’93/94 [389, 464]. ’94 [560, 646, 604, 389, 464]. ’94/95 [604]. ’95 [1430, 604, 756]. ’95/96 [756]. ’96 [951, 922, 938, 756]. ’96/97 [938]. 97 [1095, 1303, 938]. ’98 [858, 1027, 371, 1232]. 98.00 [536].

= [966, 1405].

Abb [373]. abrupt [841]. absolute [675, 143, 982, 1333]. Academic [285, 213, 717, 505, 534, 116, 616]. acceleration [37]. acceptance [1012]. accepting [691]. Accuracy [802, 1288, 1284, 726, 1095, 1303, 1206]. Accurate [182, 1308]. accurately [206]. achtzehn [1363, 373]. Acknowledgement [104, 177, 732]. acoustic [316]. ACT [516]. active [1104, 272]. Activities [1052, 212, 188]. Adams [1050]. Adapting [23]. Adaptive [828, 1421, 852, 1094, 1046, 813, 1082, 143, 1267, 565, 81]. adaptive-mixtures [1082]. adaptive-mixtures/plug-in [1082]. add [55]. add-ons [55]. additional [298]. additive [329, 958, 1320, 1183]. additivity [363, 382]. adequacy [261]. Adèr [1049]. adjusted [887, 62]. adjusting [655]. adjustment [155, 1376, 683]. adjustments [1132]. Advanced [1382, 20, 1421, 745]. advancement [118]. Advances [1417, 466, 1423, 284, 1414]. advisory [542, 541]. Æ’ [187]. affine [144]. AFMULT [517]. after [181]. against [1046, 295, 803, 1280, 47, 700]. age [1273]. aggregation [1247]. agreeable [209, 845]. Agresti [923]. Agriculture [251]. Aims [1069, 1070]. airways [715]. Aitkin [248]. alarm [79]. Alessandra [54]. Alfred [618]. Algebra [1348, 192, 1348]. algorithm [1313, 291, 650, 293, 245, 1244, 299, 708, 1302, 384, 462, 463, 639, 484, 912, 310, 385, 682, 601, 57, 312, 110, 1278, 1312, 808, 872, 772, 1160, 1029, 95, 77, 111, 1120, 1328, 40, 333, 582, 400, 1144]. Algorithms [750, 966, 854, 245, 1028, 1333, 1258, 771, 1032, 1117, 12, 1405, 314, 665, 1250, 262, 279, 758, 419]. alias [341]. aligned [1239]. all-pairwise [666]. Allen [358]. allocation [1321, 1033]. Almost [366]. Almost-exact [366]. alpha [821, 1389]. alpha-stable [821, 1389]. alternating [385]. Alternative [1364, 1195, 1243, 329, 1236, 803, 1280, 623, 679, 1317, 1106, 535]. alternatives [1180, 1046, 688, 698]. Alvin [822]. American [1191]. among [420, 608, 308, 552]. amplitude [1106]. AMS [925]. Amsterdam [32, 33, 115, 131, 1049, 503]. analyses [1185, 1379, 1395, 986, 1338, 785]. Analysis [54, 405, 1127, 396, 766, 1001, 887, 1125, 1096, 430, 465, 443, 48, 819, 532, 1408, 316, 1120, 1020, 1097, 398, 112, 1392, 220, 779, 33, 41, 199, 421, 743, 1026, 1287, 1286, 1284, 339, 450, 1385, 221, 1266, 667, 1377, 362, 1355, 865, 1123, 222, 1411, 1387, 891, 50, 1365, 1301, 28, 1345, 512, 1223, 1336, 519, 1430, 1416, 429, 609, 866, 223, 867, 517, 654, 224, 1339, 638, 713, 1397, 868, 804, 810, 322, 514, 771, 1032, 1400, 846, 1431, 340, 472, 226, 1388, 414, 984, 540, 1200, 1424, 1103, 882, 750, 515, 626]. analysis [820, 227, 1135, 657, 485, 812, 1239, 1083, 967, 905, 1426, 738, 65, 851, 692, 1406, 1270, 677, 473, 724, 671, 1361, 872, 1390, 1340, 105, 10, 70, 186, 345, 697, 71, 531, 717, 395, 689, 1407, 764, 194, 567, 365, 3, 1146, 890, 263, 204, 342, 1081, 784, 923, 616, 32, 131, 823, 632, 1074, 822, 965, 970, 1297]. analysis. [884]. analyze [725]. analyzes [1363]. anchor [900]. Anderson [590, 248]. Andrews [1422]. Angewandte [1348, 192]. animal [1312]. annealing [184, 317]. Announcement [303, 943, 798]. Announcements [31]. announces [15, 16, 17, 18]. Announcing [246, 389]. ANOVA [198, 46, 228, 49]. Anscombe [696]. Anwender [1375, 618]. Anwendung [1414]. Anwendungsbeispielen [1359]. any [295]. Anzstat [703]. appearance [900]. Application [1277, 1414, 1313, 115, 46, 610, 1273, 1359, 1103, 1343, 240, 613, 1033, 1146, 191, 758, 627, 1186]. application-oriented [115, 1343]. Applications [1333, 1275, 1097, 425, 1026, 1141, 1420, 527, 1422, 854, 51, 50, 773, 759, 933, 983, 1400, 817, 1135, 1371, 1350, 1389, 805, 1168, 1254, 66, 69, 39, 1382, 1283, 1422, 821, 658, 193, 194, 505]. Applied [284, 797, 1023, 1231, 192, 1348]. approach [596, 1329, 675, 184, 638, 1028, 754, 1158, 140, 623, 1343, 1370, 1134, 692, 978, 48, 473, 814, 914, 1338, 79, 9, 1245, 1314, 640, 115, 590]. approaches [224, 10, 784]. Approximate [947, 428, 866, 1147]. Approximating [355, 206]. Approximation [752, 397, 1308, 1341, 109, 551, 1241, 366, 316, 386, 555, 1048, 582]. Approximations [278, 802, 722, 977, 669]. April [1423]. AR-models [691]. arbitrary [111]. ARC [422]. ARC/INFO [422]. area [212]. areal [1357]. areas [1304]. ARES [476]. arguments [932]. arising [1251]. Arlinghaus [1017]. ARMA [445]. Armin [604, 938, 618]. Armitage [1281]. arrangement [1229]. array [1184]. arrays [1313, 1277, 156, 725]. art [1424, 717]. Artificial [371, 286]. Asano [191]. Ascona [425]. ASK [1015]. aspects [595, 1069, 362, 143, 913, 229, 1070, 3, 125]. Assessing [1117, 383]. assessment [1284, 1267, 626, 761]. associated [1280]. Associates [421]. Association [798, 1191, 1213, 221, 387, 36, 1209, 601, 961, 1290, 788]. assumed [125]. assuming [1335]. assumptions [692]. assurance [1349, 190].

Asymmetric [1132, 1332]. asymmetry [962]. Asymptotic [792, 387, 71]. asymptotics [1402, 987, 69]. Athens [1069, 1070]. attendance [268]. atypical [259]. Auflage [453, 468]. augmentation [1302, 1285]. August [199, 858, 743, 744, 1027, 1213, 646, 1070, 359, 1069]. Auswertung [468]. Author [43, 84, 178, 438, 376, 440, 574, 509, 577, 731, 662, 733, 860, 800, 861, 1055, 929, 991, 1060, 1216, 1115, 1165, 1221, 1300, 1340]. autoCAD [35]. autocorrelated [701]. autocorrelations [428, 273]. autoLISP [35]. Automatic [1247]. autoregressive [1199, 655, 1182, 1317, 489, 79, 488]. autoregressive-integrated [1199]. auxiliary [90, 762]. available [1225, 1011]. Average [1337, 1199, 1257, 1107]. Avoiding [567]. axes [1148].

B [986, 248, 192, 633, 737]. B-splines [737]. B.V [115]. back [115]. Badi [632]. Balakrishnan [746, 847]. balance [60, 1242]. Balanced [80, 156]. Baldev [632]. Baltagi [632]. Band [536]. bandwidth [783]. Barcelona [797, 951, 922]. Bartle [924]. Bartlett [1311, 328]. Bartoszynski [1073]. base [19]. based [947, 199, 371, 1266, 447, 412, 720, 815, 1327, 122, 782, 185, 544, 540, 1034, 1104, 791, 332, 1239, 1285, 624, 1206, 241, 1376, 1246, 582, 683]. Basel [451, 454, 657]. Basics [1353, 249]. Bayes [220, 166, 381, 960, 1269, 73]. Bayesian [285, 1418, 360, 1335, 1170, 678, 308, 224, 413, 352, 681, 983, 1171, 1305, 1090, 1135, 1306, 977, 79, 9, 1260, 640]. BDS [832]. be [421, 191]. bearing [1250, 792]. Bechhofer [846]. behavioral [1352]. behaviour [394]. Behavioural [358]. Behrens [1010, 487]. Beispiele [249, 1353]. Belsley [616]. benchmark [186]. benefit [892]. Bensberg [618]. Benzécri [766]. Bergum [505]. Berlin [371, 536, 468, 231, 192, 194].

Bernoulli [1321, 651]. Best [1225, 74, 668]. Best- [1225]. Beta [1307, 554]. better [354]. Betty [557]. between [1121, 1231, 1209, 1265, 240, 319, 1338, 1262]. Bhattacharya [96]. Bias [157, 1264, 1090, 567]. BIAS. [112]. biased [155, 60]. BIB [295]. BIFROST [599]. bilinear [958]. Billingsley [728]. binary [1304, 868, 1305, 791, 888, 1249, 1093, 979, 725]. binned [802]. Binning [1034]. binomial [361, 29, 551]. bioassays [128]. Bioavailability [405, 1365]. Bioequivalence [405, 1365]. biomedical [542, 541]. biometrical [112]. biotechnology [370]. Biplot [851]. Biplots [1398, 855]. Birkhäuser [451, 454, 657]. Birnbaum [360]. birthday [912, 1431]. births [1329]. Bivariate [770, 737, 1170, 1065, 66, 739, 874, 758, 1196]. black [186]. Bleicher [372]. Block [1315, 599, 688, 295]. Blosoft [117]. BLUEs [947]. BMDP [55, 114, 228, 263]. BMDP5V [784]. board [2, 267, 327, 441, 45, 86, 135, 179, 254, 377, 510, 579, 663, 734, 801, 862, 930, 992, 1061, 1116, 1166, 1222]. BoD [267, 857]. Bolasco [32, 131]. Book [32, 33, 34, 35, 115, 266, 285, 370, 371, 372, 373, 405, 406, 407, 408, 439, 451, 452, 453, 454, 575, 604, 938, 1056, 1218, 925, 131, 133, 536, 966, 213, 766, 468, 855, 884, 590, 846, 683, 249, 728, 821, 823, 882, 883, 1096, 557, 231, 631, 632, 645, 657, 658, 986, 1049, 1050, 1074, 967, 987, 923, 248, 466, 558, 190, 1018, 717, 1297, 504, 505, 559, 283, 284, 534, 785, 786, 1017, 1020, 1021, 193, 116, 192, 194, 467, 503, 535, 589, 617, 633, 745, 746, 822, 847, 848]. Book [924, 965, 1019, 1097, 358, 616, 1072, 1073, 191, 1296, 618, 1294, 1293]. bookstore [591, 634]. Bootstrap [1089, 1315, 1240, 675, 37, 810, 1080, 379, 1310, 366, 182]. Bootstrapping [538, 91]. Boston [285, 213, 657, 505, 193, 116]. both [609, 933]. Bouleau [559]. bound [35, 1259]. bounded [475, 207]. bounds [891, 1225, 380]. box [186, 676]. boxplots [1196]. BOYS [400]. Brace [116]. Brainerd [1050]. break [242]. breakdown [1258, 1271]. breaks [26, 25]. breakthroughs [433]. breast [1273]. Brian [785]. bridge [1231]. Bristol [858, 1027, 356]. British [589]. Brosius [249]. brushing [16]. Buch [372]. Bugaj [1073]. Build [402]. Bulgarian [232]. Burman [341]. Burr [1328]. business [225].

C [855, 821, 882, 1050, 1018, 503, 822, 965]. CA [798, 1002, 249, 248]. Cairo [188]. calculated [603]. calculating [665, 40, 1186]. calculation [708, 366, 1013, 171]. calculations [803, 1155]. Calendar [102, 119, 134, 148, 164, 175, 196, 217, 233, 252, 270, 287, 305, 325, 336, 348, 375, 392, 410, 426, 436, 456, 482, 494, 507, 524, 548, 561, 572, 621, 648, 660, 685, 730, 593, 636, 674, 704, 606, 768, 799, 885, 896, 748, 777, 826, 838, 859, 974, 989, 1041, 1076, 1088, 910, 927, 941, 954, 1005, 1053, 1114, 1130, 1153, 1164, 1193, 1203, 1100, 1142, 1177, 1214, 1299]. calibration [121, 640]. Call [42, 268, 267, 858, 1038, 1127, 1141, 1211, 1098, 798, 1051]. Cambridge [117]. Canada [406]. Canadian [832]. cancer [1273, 892]. candidates [267, 1211]. canonical [540, 105, 518]. Capri [1098]. carcinogenicity [1312]. Carlo [1306, 1254, 1010, 444, 123, 52, 1045, 475, 1271]. case [157, 976, 743, 274, 457, 307, 1225, 656, 671, 700, 239, 415]. Casebook [884, 1387]. cases [1064, 26, 25]. casual [18]. Categorical [923, 1392, 1185, 1181, 1078, 1270, 77, 124]. categories [1247, 819]. categorization [519]. CATSCALE [519]. cause [830]. cautionary [1138, 611]. CBR [948]. cells [609]. censored [947, 166, 917, 1256, 1168, 1120]. censoring [200, 914]. census [690]. center [556]. Centro [425]. certain [26, 25]. Ch [191]. chain [1309]. chains [1251]. Challenge [1212]. Challenges [1001, 646]. change [829, 447, 1135, 805, 841]. change-point [447, 1135, 805]. changed [900]. changepoint [1273]. chaotic [1265]. Chapman [855, 882, 986, 1018, 1297, 785, 589]. characteristic [59]. characteristics [708, 1276, 603]. Characterization [782, 76, 67]. Characterizations [75]. Characterizing [70]. Charles [1050]. charts [721, 1337]. Chatterjee [884]. Chattfield [882]. check [1105]. Checking [692]. Chernoff [331]. Chernoff-type [331]. Chi [925, 790, 1244, 1399, 380]. chi-square [380]. Chi-Squared [925, 790, 1244, 1399]. Choice [501, 220, 1102, 238, 1285, 229]. Choosing [500, 552]. Chow [405, 745]. Christensen [27]. Chung [405, 745]. Chytil [133]. circular [585]. CL8 [894]. Clarendon [248, 1021]. class [296, 380, 550, 94]. classes [998, 433, 66]. classical [23, 259, 149, 78]. Classics [924]. Classification [925, 842, 864, 1284, 1181, 293, 569, 736, 868, 344, 681, 998, 871, 89, 240, 875]. classifications [363, 382, 1146]. classifying [791]. Clifford [557]. clinical [708, 332, 1276, 7]. close [832]. closely [667]. closest [583]. cloth [786]. clouds [477, 874]. CLUSPLOT [1279]. Cluster [1304, 865, 1301, 866, 1339, 868, 1340]. clustered [1093]. Clustering [344, 948, 949, 293, 709, 1279, 873, 331, 1160, 1063, 330]. clusters [934, 569, 869]. Cochran [1281]. Coding [741]. coefficient [827, 918, 106]. coefficient-estimation [918]. coefficients [828, 1118, 442, 834, 914, 333]. Cognition [597]. Cohen [557]. cointegrating [913]. cointegration [1311]. Coleman [1339]. column [13, 760, 1029]. combination [143, 82]. Combinatorial [875]. combined [1082]. combining [1243]. Comer [284]. Comment [744]. Comments [1339, 951, 1340, 27]. Committee [304]. common [934, 420, 1159, 488]. communications [1000]. compact [1356]. compactly [1044]. Company [249, 193]. Comparative [1229, 460, 1199, 1347, 551, 586, 918, 52, 116]. compare [892, 978]. Comparing [258, 222, 553, 784, 500, 339, 1158, 235, 755, 568, 203]. Comparison [868, 261, 832, 331, 783, 1121, 740, 595, 420, 1185, 1122, 363, 763, 1123, 273, 274, 297, 774, 903, 654, 771, 228, 818, 1265, 74, 1312, 10, 262, 279, 125, 1107, 1271, 1009]. Comparisons [1018, 1147, 1385, 1156, 815, 245, 1149, 202, 1401, 666, 1085, 78, 846]. compatible [17, 18]. competing [843]. Competition [1191]. competitors [444]. complete [24, 642, 1363]. complex [237, 150, 566, 9]. complex-valued [566]. complexity [396, 1169]. Component [779, 290, 222, 429, 654, 830, 1103, 236]. components [740, 54, 654, 496, 1, 1167, 65, 624, 5]. compositional [1278]. Comprehensive [1192]. COMPSTAT [1213, 1230, 216, 837, 1027, 560, 951, 646, 922, 858, 1232]. COMPSTAT’92 [320]. COMPSTAT’96 [877]. Compstat’98 [1231, 1111]. Computation [1168, 187, 554, 827, 597, 993, 607, 656, 207]. Computational [54, 1295, 1377, 143, 1001, 1125, 876, 1372, 1352, 796, 971, 595, 879, 1069, 1420, 1301, 1339, 185, 1388, 1231, 823, 139, 1340, 1070, 503, 358, 191, 333, 1127, 229, 970, 616]. compute [1106]. Computer [1425, 892, 1023, 150, 399, 1431, 312, 967, 187, 1245, 504]. Computers [41, 922, 17, 18, 598, 133]. Computing [251, 798, 1191, 1213, 788, 402, 1077, 874, 317, 232, 212, 968, 1038, 907, 1141, 277, 1244, 1302, 356, 113, 876, 323, 985, 1079, 504, 77, 333]. concise [1410]. Conditional [687, 459, 1327, 937, 946, 640]. conditionally [1131, 94]. conditions [790, 71]. Condra [645]. conducting [484]. Conference [798, 968, 1213, 788, 1422, 1430, 1414, 1423, 1098, 199, 1141, 1001, 1002, 356, 231, 1070, 796]. Confidence [256, 807, 762, 995, 1089, 891, 29, 1080, 1280, 167, 666, 486, 834, 182, 1013]. configurations [957]. Confirmation [997]. confirmatory [472, 689]. Confusion [741]. congress [1003]. Conjugate [472, 1131]. connection [5]. consecutive [1077]. consensus [840, 1146]. conservative [643]. consistency [307, 1137]. consistent [309]. constant [37, 227, 1009]. constants [1013]. constrained [736, 340, 1143]. constraints [430, 719, 550, 95, 890]. constructing [314]. Construction [753, 759, 127]. Constructions [60]. consultation [985]. ContentsDirect [1109]. Contextual [1284]. contingency [258, 1180, 220, 397, 24, 771, 380, 937, 1247, 1029, 961, 1227]. continuation [944]. continuity [790]. Continuous [1380, 260, 699, 773, 1336, 1357, 791, 746, 847]. continuum [236]. contour [354]. contours [874]. contrasts [995, 1104, 272, 486]. contributed [798]. Control [406, 1421, 815, 394, 693, 721, 449, 1419, 1337]. controlled [682, 74]. Convergence [639, 381]. convergent [650]. conversion [418]. Converting [151]. convex [307, 111]. Convolution [551, 1186]. Cook [611]. Cook-type [611]. Coppi [32, 131]. copy [970, 319]. COR.ANALYSIS [465]. core [1313, 1200]. corrected [1241]. correcting [88]. Correction [205, 790, 523, 328]. corrections [1311]. correlated [996, 93, 1078, 710, 11, 1144]. correlation [350, 827, 608, 1179, 914, 889, 1159, 125, 518]. correlations [540]. Correspondence [1355, 465, 362, 766]. Corrigendum [198]. cost [417]. costs [433]. Council [972, 304]. count [887, 1275]. Counting [563]. Countries [251, 406]. Couper [1296]. coupled [280]. course [1387, 884]. Courses [103, 120, 165, 176, 197, 234, 253, 271, 288, 306, 326, 337, 349, 393, 411, 437, 483, 495, 508, 525, 549, 562, 573, 622, 637, 649, 661, 686, 705, 749, 778, 886, 897, 911, 928, 942, 955, 990, 1006, 1042, 1054, 1101, 1178, 1194, 1204, 1215, 797]. Covariance [155, 89, 342, 60, 463, 340, 806, 343, 1334, 317, 243, 1081, 784]. covariances [444]. covariate [812]. covariates [260, 1138, 791, 1172]. Cox [676, 1168, 1263]. Cramer [407]. CRC [1017]. criminal [1330]. criteria [157, 1169, 1085]. criterion [1327]. critical [544]. cross [201, 814]. cross-validation [201]. cross-variograms [814]. crossing [568]. CSDA [218, 338]. CSDA-1991 [218]. CSDA-1992 [338]. Current [1113]. curvature [152, 1272]. curve [1264, 57, 1172, 851, 806, 582]. curve-fitting [582]. curves [240]. cuscore [1337]. cutoff [761]. CWA [245].

D [35, 370, 1339, 855, 231, 248, 559, 193, 589, 358, 616, 50, 1406, 1020]. D. [1339]. Daly [589]. Dartmouth [1418]. Data [54, 970, 1127, 813, 1001, 1339, 1125, 923, 160, 805, 319, 532, 1020, 23, 1392, 780, 1313, 32, 33, 17, 199, 421, 1026, 1291, 1288, 1293, 259, 396, 809, 956, 1185, 131, 966, 329, 789, 1387, 50, 1345, 512, 519, 166, 1430, 1416, 866, 1302, 638, 1304, 868, 1080, 887, 1428, 884, 463, 1032, 1431, 1102, 225, 226, 414, 653, 515, 832, 601, 227, 1331, 632, 657, 1074, 917, 1268, 690, 332, 1405, 967, 1093, 1278, 1426, 47, 301, 1145, 1285, 1168, 641, 1406, 1195, 1270, 724, 6, 1361, 602, 10, 739, 531, 1408, 83, 1184, 39, 77]. data [1120, 1143, 321, 589, 365, 919, 1146, 107, 725, 1148, 342, 784, 1419, 1297, 34, 406]. data-based [1285]. Database [1098, 332, 92]. databases [1354, 631]. Datenanalyse [407]. David [846]. DBMS [319]. DBMS/COPY [319]. Deadline [1191]. deadlines [606]. Dec [231]. decade [544]. decomposition [850, 279, 612]. decompositions [221]. deconvolution [1102]. DEDICOM [385]. definite [94]. degree [382, 1285]. Degrees [24]. Dekker [370, 405, 406, 766, 590, 557, 645, 283, 745, 965]. deleting [107]. deletion [415]. demonstrated [613]. demonstration [290]. Dempster [82]. dendrograms [1146]. densities [1082, 898]. density [23, 157, 1121, 276, 460, 720, 1307, 1325, 852, 1267, 201, 804, 309, 88, 737, 170, 539, 1238, 311, 1253, 275, 1044, 138, 489, 416, 1260]. density-based [720]. Deny [145]. Department [971]. dependence [414, 783]. Dependent [491, 679, 805]. depiction [36]. DePriest [283]. depth [874]. Derived [603]. described [4]. description [364]. Descriptive [1369]. Design [1385, 1365, 1370, 1083, 1242, 290, 370, 1207, 1198, 795, 1366, 308, 444, 13, 645, 332, 1276, 697, 316, 567, 3, 9, 1246, 405, 590, 846]. designs [54, 1094, 688, 60, 815, 294, 417, 295, 296, 759, 1, 58, 585, 753, 760, 64, 831, 1158, 1104, 341, 312, 128, 47, 314, 700, 81, 263, 553, 1363, 784, 1363, 373]. Deskriptive [1369, 558]. detect [1208, 1261]. Detecting [713, 1172, 242, 1091]. Detection [958, 330, 259, 1237, 1092, 1196]. determinant [463, 317]. determinants [1346, 213]. determination [706, 1186]. determining [1121, 721, 400]. Deutscher [452]. Developing [251]. Development [542, 795, 982, 433]. developments [137]. deviances [1092]. deviates [38]. deviations [143, 1333]. Diagnostic [1105, 998, 851, 98]. Diagnostics [361, 329, 1255, 238, 108, 1064]. diagonal [60]. diagram [9]. Diday [1074]. difference [29, 613, 1262]. differences [1091, 365]. differencing [714]. different [784]. differing [274]. Digital [690]. digits [461]. dimension [477]. dimensional [157, 14, 866, 804, 552, 316]. Dimensionality [395, 1314]. Direct [1183]. directional [867, 227]. directions [657, 1426]. directors [267]. Dirichlet [609]. disambiguating [597]. discordant [1245]. Discrete [467, 396, 789, 1336, 1080, 1357, 1032, 238, 314, 1195, 1314]. discretization [1325]. Discriminant [804, 1336, 713, 810, 626, 905, 677, 872, 395]. discrimination [584, 308, 106, 475]. Discussion [26, 27, 646, 1291, 137]. Discussions [528]. disease [709]. Disfree [117]. Disjunctive [1021, 1381]. disk [658]. diskette [589]. diskettes [117]. diskreter [1357]. dispersion [782, 652, 751]. display [851]. Displaying [1279, 625]. distance [720, 1228, 240, 111]. distortion [641]. distributed [1141, 228, 61, 446]. Distribution [1343, 947, 59, 200, 707, 397, 554, 827, 1244, 299, 457, 1157, 933, 1225, 935, 379, 126, 695, 944, 1125, 1077, 1011, 1096, 656, 817, 139, 311, 1241, 381, 1261, 888, 1332, 478, 899, 1168, 473, 1067, 698, 76, 642, 52, 171, 1407, 183, 73, 386, 919, 1328, 1147, 1124, 762, 115]. Distribution-free [1343, 115]. Distributional [125]. Distributions [1367, 258, 688, 699, 610, 915, 782, 496, 1380, 726, 1095, 821, 1090, 235, 167, 1337, 898, 1389, 169, 487, 38, 66, 1079, 680, 752, 489, 355, 203, 75, 628, 1144, 467, 633, 746, 847]. disturbances [123]. disuniformity [397]. divergence [361, 1206]. DM [266, 371, 372, 373, 407, 408, 451, 452, 453, 454, 536, 468, 249, 632, 658, 1050, 1074, 248, 466, 1020, 192, 194, 966, 558, 504]. DM40. [618]. DM40.- [618]. DM45 [617]. DNA [1330]. Dodge [33]. Dong [1339]. Doordrecht [213]. DoPE [795]. Doppler [792]. Doppler-bearing [792]. Dordrecht [717, 505, 616]. DOS-Anwender [1375, 618]. dose [291, 1281, 1282]. dose-response [291]. dosimetric [983]. double [1240, 182]. double-bootstrap [182]. Doubletree [1002]. doubly [947, 707]. Douglas [283]. down [26, 25, 755]. dual [1198, 310]. duality [74]. Dunnett [274]. duplicates [92]. duration [1123]. Durbin [448]. during [188]. Dynamic [625, 1277, 959, 323, 241, 1272].

Eastern [894]. easy [1007, 167, 416, 22]. econometric [1123]. Econometrics [616, 1377, 1351]. Economic [616, 118, 1377]. economics [425]. Ed [1112, 285, 728, 882, 631, 1050, 466, 284, 503, 617, 616, 1072]. ed. [966]. edge [907]. edited [32, 33, 604, 938]. Edition [467, 633, 746, 847, 924, 883, 193, 358]. Editor [744, 967, 523, 743, 547, 546, 545]. Editorial [2, 327, 441, 1319, 56, 87, 526, 578, 289, 45, 86, 135, 179, 254, 377, 510, 579, 663, 734, 801, 862, 930, 992, 1061, 1116, 1166, 1222]. Editorials [1220]. eds [406, 131, 133, 231, 632, 657, 1074, 504, 505, 283, 191]. Education [922, 796]. Edward [283]. effect [294, 1273, 397, 761, 83, 5, 919]. effective [983, 600]. effectiveness [497]. Effects [530, 812, 107, 891, 655, 638, 1254, 83, 844, 919, 1263]. efficiency [1197, 1007, 993]. Efficient [185, 709, 171, 1244, 668, 1323, 1029, 182, 333, 1271]. efficiently [588]. Eigenprojections [889]. eigenstructure [107]. eigenvalues [343, 5, 415]. eigenvectors [1159]. eighteen [1363]. Einfuhrung [1359, 1353, 249, 1360, 372]. Einstieg [1375, 618]. Election [304]. Elections [857, 1211]. Electronic [644]. elemental [412]. Elements [924, 1384]. ellipsoid [1007]. ellipsoidal [144]. elliptical [530]. Ellis [631]. Elsevier [32, 33, 115, 284, 1109]. emission [1322]. Empirical [73, 166, 381, 1269, 275, 5]. employing [690]. endpoints [1156]. Engelbrecht [133]. engineering [1349, 1323, 190]. entries [621, 648, 660, 685, 730, 768, 799, 885, 896, 886, 897, 974, 989, 1041, 1076, 1088, 928, 942, 1042, 1114, 1130, 1153, 1164, 1193, 1203, 1194, 1204, 1299]. entropic [1314]. Entropy [461, 285, 1418]. environment [323]. environmental [9].

Epi [307]. Epi-consistency [307]. equal [1029]. equality [1236, 818, 889, 95, 890]. equation [430, 473, 1120, 1143, 890]. equations [850, 623, 4, 1184, 1243]. equiprobable [656]. Equivalence [668, 613]. equivalent [811]. ERC [1052]. ergänzte [453]. ergodic [309]. Erratum [1084, 30, 367, 1125, 1068, 856]. Error [791, 1010, 1238, 580, 487, 491, 183, 78, 807]. Errors [676, 121, 60, 996, 93, 225, 226, 711, 11, 77]. errors-in-latent-variable [225]. Errors-in-variables [676]. ERS [267, 857, 1212, 878, 1051]. ESBPDF [1287]. ESRI [422]. essence [1410]. estimability [24]. estimate [92, 1179, 126, 607, 139, 138, 914, 416]. estimated [713]. estimates [277, 1357, 74, 840, 612, 207]. Estimating [96, 1118, 539, 566, 888, 1249, 1145, 1044, 1159, 809, 699, 1329, 150, 387, 892, 812, 488]. Estimation [68, 200, 350, 1224, 714, 1274, 225, 653, 1331, 946, 723, 1253, 181, 689, 157, 499, 934, 291, 205, 54, 155, 530, 1122, 596, 447, 237, 1240, 460, 720, 915, 26, 1342, 90, 154, 144, 1154, 1226, 852, 959, 122, 982, 1267, 201, 37, 1, 185, 152, 804, 309, 1264, 1258, 25, 983, 1133, 126, 695, 944, 1125, 310, 340, 88, 936, 754, 1228, 137, 1167, 737, 170, 557, 180, 1238, 381, 1252, 918, 1256, 1257, 1134, 1312, 315, 275, 1195, 4, 343, 739, 772, 108, 498, 83, 5, 1143, 994, 73, 318, 330, 501, 1328, 1260]. estimator [23, 1121, 781, 1007, 1082, 1327, 166, 463, 993, 735, 62, 751, 106, 317, 489]. Estimators [6, 1197, 1121, 273, 1307, 143, 497, 496, 1347, 1028, 379, 311, 711, 1317, 1269, 665, 431, 52, 353, 78, 116, 40, 1271, 1009]. Estonia [796]. etc [231]. Europe [894]. European [267]. EUROSTAT [1023]. Evaluating [140, 761, 1158]. evaluation [934, 638, 127, 52]. Evans [633]. Everitt [986, 785]. evolution [966, 1405, 701]. Exact [803, 1309, 607, 680, 489, 1155, 643, 688, 484, 1065, 771, 1032, 1077, 753, 937, 449, 656, 445, 1281, 700, 366, 1029, 281]. example [26, 25]. Examples [683, 1353, 1028, 1359, 603, 1376]. Excel [1095, 1303]. exchange [880, 1265, 314, 319, 627, 34]. execution [709, 1323]. Executive [304, 972]. executives [1113]. exercises [186]. expectations [70]. experiences [263]. experiment [1083]. Experimental [54, 296, 1, 1198, 795, 1123, 370]. experimentation [795]. experiments [1385, 245, 1366, 590, 846, 226, 13, 1119, 272, 1370, 3, 645]. Expert [1049, 1383, 133, 1413]. expertise [556]. Explorapedia [1289]. exploration [813]. Exploratory [221, 140, 724]. exponential [527, 915, 457, 566, 1135, 169, 75, 1107]. exponentials [1251]. exponents [715]. Extended [145, 295, 126]. Extension [51, 274, 1085]. extra [183]. extraction [779]. extrapolation [58]. extreme [53, 1008, 231]. extremes [1077, 415]. ExVar [449]. EzPATH [18].

F [266, 285, 505, 589]. face [331]. facility [985]. Factor [220, 517, 514, 444, 472, 984, 1200, 750, 961, 689, 333]. factorial [417, 1104, 272, 1083, 47, 697]. factorials [1210]. factors [761, 328, 1242]. failure [1170, 830]. fairly [947]. familial [83]. families [1135]. family [350, 146, 1305, 710, 1263]. far [320]. FAST [1106, 664, 445, 1032, 1117, 708]. fatigue [360, 1226]. Faulbaum [266, 466]. feasible [384, 462, 463, 639, 1258]. features [20]. February [798, 1001, 1002, 1003]. feedback [994]. Feelings [1230]. few [740]. Fibonacci [1422, 505]. Fifth [1422]. Fig [504, 505]. files [589]. Filleben [123]. filter [352, 811]. filtering [681, 241]. filters [1045, 1243]. Financial [894]. Finding [841, 53]. Finetti [145]. fingerprints [1330]. Finite [1317, 1007, 1267, 804]. Finite-sample [1317, 1007]. Finland [973]. First [858, 39, 1387, 884, 1257, 642, 489]. First-order [39, 1257, 489]. Fischer [266, 373, 407, 408, 453, 466, 558]. fish [316]. Fisher [643, 1229, 1010, 127, 487, 1137]. fit [1266, 916, 583, 935, 1206]. fitness [1179]. fitted [612]. Fitting [260, 329, 235, 642, 307, 600, 94, 582]. Fixed [666, 1111]. Fixed-width [666, 1111]. flächenhafter [1357]. Flexible [154]. Florida [286]. fluorometric [1313]. following [972, 915, 1331]. font [1111]. Forecast [1148, 1252]. forecasting [1277, 627]. form [1111]. formal [638, 597]. former [894]. forming [279]. forms [72]. Fortran [1050, 1394]. Fortschritte [266, 1414]. forum [1037]. forward [687]. Fougere [285]. founded [232]. Fourier [708, 1198, 1032, 1117, 690, 1106]. Fourth [286]. Fractal [715]. fraction [1302]. fractional [417, 1210, 47, 697]. fractionally [714]. fractions [294]. framework [1210, 1090, 488]. France [1421]. Francis [248]. Frank [372, 466, 618]. Franscini [425]. free [1109, 417, 299, 609, 1343, 473, 739, 698, 627, 115]. freedom [24, 382]. frequencies [1262]. frequency [931]. Friedman [743, 420, 444, 775, 671, 1067]. Ft [286]. Fuller [883]. function [932, 307, 1225, 1302, 109, 721, 936, 656, 138, 111, 1338, 386, 758, 762]. functional [901, 802, 539, 890]. functions [520, 1307, 1325, 1335, 151, 1149, 1238, 311, 1253, 1168, 94, 489]. Further [156, 297]. future [879, 172, 999]. Fuzzy [641, 873, 330].

G [453, 249, 986, 505, 924]. gamma [1131, 1157, 695, 1125, 1147]. Garfinkel [658]. GAUSS [726]. Gaussian [712, 959, 224, 1182, 935, 1062]. GB [1297]. GCVPACK [354]. GEE [1137, 1195]. general [1313, 1286, 538, 307, 122, 387, 1270, 1242, 698, 94, 1283]. generalised [185, 1320, 775, 242]. Generalization [901, 980]. Generalized [610, 141, 429, 540, 398, 1266, 1255, 476, 611, 810, 912, 668, 997, 917, 236, 1183, 91, 1282, 241, 72]. Generally [294]. generate [1144]. Generating [341, 38, 933, 1078, 312, 1146]. generation [399]. generator [1161, 1014]. generators [146]. Genetic [854, 1405, 966]. genuine [569]. Geodesic [292]. geographical [1304]. geometric [901, 75, 1124]. Geometrical [871]. geostatistical [1357]. geostatistics [1381, 1427, 1021, 534]. geostatistischen [452, 1357]. German [1353, 1414, 1368, 1356, 1357, 1348, 1359, 1369, 1360, 1373, 1351, 1375, 1362, 1363]. Germany [743, 789, 113, 212, 744]. Geßler [451]. get [920]. Getting [354]. Ghosh [1208]. Gibbs [763, 1134]. Gini [781]. Giovagnoli [54]. Giri [965]. Girko [213]. GIS [422]. given [380, 416]. GLIM [1344, 260, 227, 248]. global [81]. GM [993]. Goel [717]. Goldberg [1050]. Goldsman [846]. Goodness [1266, 916, 935, 1206]. Goodness-of-fit [1266, 916, 935, 1206]. Göttsche [408]. Gower [855]. gradient [1123, 472]. graphic [249]. graphical [763, 1158, 601, 581, 331]. Graphically [447]. Graphics [214, 1353, 1368, 424, 1023, 323, 1412, 283]. Graphik [1368, 451]. Greece [1070, 1069]. Greenwood [925]. Grenoble [1421]. grid [344, 804]. Group [41, 708, 122, 1331, 475]. groups [274, 48, 568]. Groves [1296]. growth [57, 1172, 851, 806, 239, 784]. Gruber [116]. Gruyter [192]. GUHA [598]. Gui [745]. Guide [389, 604, 938, 1035, 925, 464, 756, 1050, 246, 303, 1394, 1399, 1410, 738]. Gunar [406]. Gustav [266, 373, 407, 408, 466, 558].

H [266, 743, 683, 632, 1049, 1050, 967, 116, 848, 1019]. Haag [604, 938]. Haaland [370]. half [947, 200, 983, 1104]. half-life [983]. half-normal [1104]. Hall [823, 882, 785, 855, 986, 1018, 1297, 589].

Hamburg [249]. Hand [855, 589].

Handbook

[1404, 1192, 1429, 1355, 1379, 1395, 1428, 986, 785, 1017, 589, 766, 503, 1019]. Handcock [884]. Hans [407, 617]. Harcourt [116]. hardback [115]. Hardcover [1020, 1021]. Hardin [1339]. Härdle [504]. Hastings [633]. Haux [266]. Havránek [594, 597]. hazards [914, 568]. HBP [993]. HBP-GM [993]. Health [1038]. healthcare [1038]. Heidelberg [371, 536, 1414, 632, 504, 617, 618, 1423]. Heinrich [452]. Heinsohn [371]. held [231]. Hellinger [1228]. Helmuth [744]. Helsinki [973]. heterogeneity [394, 1083]. heterogeneous [916, 224]. heteroscedastic [946]. heteroscedasticity [74]. Hidelberg [372]. hierarchial [351]. hierarchical [24, 784]. high [1141, 866, 1258, 1271]. high-dimensional [866]. highways [1023]. Hill [34, 35, 249]. Hinde [248]. histogram [900]. historical [982]. HLM [784]. Hocking [1097]. Holland [131, 503]. homogeneity [937, 679, 978]. honor [1431]. Hood [35]. Horadam [505]. Horwood [631]. Hotel [1002]. household [1409, 1296]. HPM [678]. Hsu [1018]. Huppert [192]. Hüsler [231]. Hüthig [372]. hybrid [763]. Hyperparameter [1252]. hyperplanes [583]. hypotheses [297, 1008]. hypothesis [474, 538, 675, 774, 903].

IAS [837]. IAS-News [837]. IASC

[1001, 1002, 1212, 101, 132, 163, 174, 304, 195, 324, 269, 359, 335, 347, 374, 409, 435, 455, 469, 481, 537, 592, 605, 619, 635, 647, 659, 673, 684, 694, 702, 729, 620, 857, 747, 757, 776, 824, 825, 849, 895, 972, 971, 909, 926, 939, 953, 988, 1004, 1025, 1040, 1057, 1075, 1087, 973, 1099, 1110, 1152, 1234, 391, 1202, 188, 547, 492, 571, 767, 1211, 1003, 1113, 787, 1176, 878, 1051]. IASC-News [592, 605, 619, 635, 647, 659, 673, 684, 694, 702, 729, 857, 747, 776, 824, 825, 849, 767, 787, 757]. IBM [17, 18].

IBM-compatible [17, 18]. ICM [1284]. ICM98 [894]. ICM98-CL8 [894]. ideas [880]. identical [527, 203]. Identification [957, 867, 272, 691, 563, 816]. identified [430]. Identifying [569, 1245, 1104, 1160]. IDIOSCAL [540]. if [125]. II [200, 166, 542, 794]. III [47]. illustrated [34, 35]. Image [1223, 1284, 352, 344, 681, 1102, 1322, 283, 1412, 670]. impact [152, 140]. Implementation [1049, 1383, 667, 1324, 600, 1322]. Implicit [127]. importance [1254]. impression [229]. Impressions [951]. Improved [499, 1258, 203]. improvement [1366, 693, 645, 313]. Improving [1206, 1081, 1288]. Imputation [667, 833]. imputed [1145]. Inc. [16, 17, 18]. including [1085, 589, 9]. Incoming [878]. Incomplete [13, 830, 6, 1408, 1297, 1143, 263]. inconsistency [379]. increase [1007]. increasing [652]. independence [843, 1327, 937, 6]. independent [378, 500, 894, 1030, 551, 581, 711]. Index [1055, 1216, 1326, 43, 84, 44, 85, 178, 438, 376, 440, 574, 509, 577, 731, 662, 733, 860, 800, 861, 929, 991, 1060, 1115, 1165, 1221, 1300, 380, 488, 781]. Indicator [257]. indices [1089, 1034]. indirect [1249]. individual [256]. induced [599]. Industrial [1212, 20, 1069, 1070]. Inference [142, 1180, 1393, 1335, 1309, 722, 1330, 1249, 82, 281, 1124, 1416, 33, 1073]. Inferences [360, 121]. infinite [157]. inflation [88]. Influence [105, 255, 984, 383, 1318, 1338, 475, 9, 207]. Influential [806, 713, 1172]. INFO [422]. Informatics [1035, 1038]. Information [157, 100, 130, 162, 173, 250, 212, 403, 434, 480, 506, 522, 718, 836, 1059, 952, 1039, 1086, 1220, 1151, 1175, 149, 1098, 493, 546, 533, 1023, 1327, 1302, 830, 127, 1247, 624, 978, 106, 762]. Informational [1169]. input [4]. Inputing [1291]. inside/outside [872]. inspection [830, 449]. Installed [19]. instrumental [429]. integral [4, 814, 316]. integrated [422, 1199, 75]. Integrating [1063]. Integration [241, 924, 1384]. Integration-based [241]. intellegence [286]. Intelligence [371, 1295]. intensity [1335]. intensive [150, 1425, 1431, 967, 504, 1245]. interaction [1064, 586, 49, 198]. interactions [1246]. intercept [68, 446]. interests [597]. interface [581, 764]. interlaboratory [153, 840]. International [251, 798, 1213, 1298, 788, 1422, 1421, 1430, 1098, 286, 199, 425, 423, 971, 1141, 186]. Internet [1024, 1035, 1291]. Interpolation [452, 1357]. interpretation [670]. interrelations [1402, 987]. Interscience [190]. Interuniversity [556]. Interval [915, 447, 29, 1168]. intervals [995, 1089, 1199, 678, 458, 1080, 1280, 256, 666, 486, 834, 1013, 807, 762]. interview [1409, 1296]. intraclass [350]. introduces [55, 20, 114]. Introduction [1353, 1396, 864, 923, 1381, 1375, 1392, 1411, 854, 512, 1388, 1359, 823, 882, 1360, 883, 1021]. invariant [363]. inventory [1277]. inverse [609, 109, 935, 3, 72]. invertible [386]. investigation [275, 345]. investigations [156]. involved [920]. irregular [1207]. ISAC [304]. ISBN [32, 33, 115, 266, 285, 370, 371, 372, 373, 405, 406, 407, 408, 451, 452, 453, 454, 925, 131, 133, 536, 966, 213, 766, 468, 855, 884, 590, 846, 683, 249, 728, 821, 823, 882, 883, 1096, 557, 631, 632, 645, 657, 658, 986, 1049, 1050, 1074, 967, 987, 923, 248, 466, 558, 190, 1018, 717, 1297, 504, 505, 559, 283, 284, 534, 785, 786, 1017, 1020, 1021, 193, 116, 192, 194, 467, 503, 535, 589, 617, 633, 745, 822, 847, 848, 924, 965, 1019, 1097, 358, 616, 1072, 1073, 191, 1296]. ISBN [618, 746]. ISI [972, 269, 188, 1070]. ISI’99 [973]. isotonic [110, 550, 111, 1031]. isotropic [398]. Issue [1127]. Issues [1421, 1250, 1301, 1339, 1340, 1001, 646].

Istanbul [972]. Italy [32, 1098, 971, 1026]. items [678]. iterated [697]. Iterative [850, 612, 959, 1308, 185, 310, 600, 555]. iteratively [946]. IV [971, 1293].

J [115, 371, 766, 1339, 855, 884, 590, 846, 231, 1049, 248, 117, 1297, 283, 116, 589, 848, 358]. J.-P [766]. Jackknife [37, 343, 1310, 138]. Jacques [1021]. Jan [618]. Jan-Armin [618]. Jana [987]. January [286, 1191]. Japan [191]. Jason [1018]. Jeanne [1050]. Jen [405]. Jen-Pei [405]. Jena [373, 453]. Jennrich [823]. Joeckel [266]. John [35, 925, 846, 190, 717, 559, 467, 535, 633, 746, 847, 848, 1019, 358, 1072, 1296]. Johnson [467, 746, 847]. joint [888, 603]. Jones [557]. journal [1038]. Journals [1109]. Jovanovich [116]. July [1141, 1422, 1339, 1098, 971]. June [796, 1430, 113, 971]. Jurecková [987]. Jürgen [451].

K.-H [266]. Kalman [959, 811, 241, 1243]. Kalman-filtering [241]. Katsuo [1096]. Kaufmann [454]. Kemp [467]. KENT [193]. Kernel [309, 23, 1307, 802, 88, 1034, 138]. kernel-based [1034]. keys [1242]. Klein [717]. Kluwer [285, 213, 717, 505, 534, 616]. knot [180]. Knowledge [544, 1049, 1383, 371, 599, 624]. Knowledge-based [544]. known [106]. Knut [743, 744]. Koch [604, 938]. Kolmogorov [980, 680]. kompakt [408, 1356]. komplette [373, 1363]. Konferenz [1414]. kontinuierlicher [1357]. Kotz [467, 746, 847]. Krauth [115, 117]. Kriging [1248, 1021, 1381]. Kronecker [80]. Kruse [371]. Kruskal [1085]. Kumar [987]. kurtosis [83, 752].

L [213, 1339, 590, 821, 658, 1297, 504, 193, 467, 746, 847]. Lach [1017]. lack [529, 75].

LAD [1118, 550]. lag [61]. Lagrange [1157]. Lamperti [1072]. Landsat [670]. landscapes [1179]. large [947, 258, 932, 866, 1251]. largest [1302]. last [1230, 489]. late [1091]. latent [1154, 225, 889]. later [117]. Latest [1235]. lattice [14]. Lattices [1198, 1062]. Lauderdale [286]. LAV [26, 297, 774, 903, 25]. layer [723]. lead [892]. leading [907]. Least [180, 1064, 706, 675, 143, 185, 152, 384, 462, 639, 1333, 607, 12, 385, 750, 1092, 946, 1257, 1278, 719, 446, 280, 419, 72]. least-squares [152, 750, 1257, 719]. Lebesgue [924, 1384]. Lechevallier [1074]. Length [1043, 1315, 1337]. lengths [1337]. Lépingle [559]. Letter [523, 743, 744]. level [417, 341, 1083, 851, 473, 689, 725, 342]. Leverage [1064, 98]. Library [924, 591, 634]. Liepins [406]. Life [830, 360, 1342, 983, 557]. lifetime [782]. like [172]. likelihood [530, 46, 237, 773, 959, 224, 151, 868, 810, 1028, 126, 695, 944, 1125, 1102, 310, 913, 1322, 984, 936, 445, 709, 1183, 1256, 1093, 1312, 624, 1270, 48, 999, 1328, 418]. likelihoods [497]. limit [182]. limited [624]. limiting [299]. limits [1121, 653, 478]. line [706, 515, 807]. Linear [596, 144, 1348, 1391, 1021, 745, 1097, 568, 68, 563, 520, 121, 1255, 476, 417, 538, 154, 1224, 982, 224, 1302, 152, 36, 255, 1400, 816, 1171, 607, 58, 668, 839, 180, 997, 235, 711, 236, 918, 1256, 74, 91, 62, 719, 979, 1282, 1270, 4, 240, 1381, 792, 71, 431, 241, 81, 1160, 1272, 242, 78, 318, 192, 919, 1382, 1009, 784, 848]. lineare [1348, 192]. linearities [427]. link [936, 1137]. Linkage [223, 866]. links [1305]. Lippe [558]. LISP [323]. LISP-Stat [323]. List [906, 1291, 1292]. listserv [1293]. Literature [1059]. Liu [405]. Lloyd [645]. LMS [596]. LMS-estimation [596]. Local [984, 81, 1318, 477, 1248, 383, 316, 1338, 1009]. locally [700]. Locating [1031]. location [412, 292, 1107]. locations [180]. Log [769, 979, 610, 28, 944, 235]. Log-linear [979]. Log-normal [769, 610, 944]. Logistic [442, 947, 969, 200, 474, 902, 328, 793, 794, 741]. logit [260]. loglinear [221, 1138, 223]. lognormal [52, 1124]. logrank [1091]. logspline [737, 170]. London [788, 213, 855, 631, 657, 986, 1018, 1297, 505, 503, 589]. long [1122, 167]. long-memory [1122]. long-tailed [167]. longitudinal [956, 1195, 919, 784]. longitudinally [443]. look [172, 832]. Looking [667]. Loosely [280]. Lorenzen [590]. loss [1132]. lot [678]. Lothar [468]. lower [1259, 555]. lungs [715]. Lunn [589]. Lütkepohl [1019]. lynx [832].

M [743, 744, 925, 133, 1339, 884, 846, 248, 190, 194, 633, 1296]. MA [735]. Mac [16]. Macintosh [15, 1286]. Macro [465, 1085]. macros [450, 820]. Magdalena [1073]. magnitude [433]. Mahoney [658]. mailing [1292]. main [294]. main-effect [294]. majorization [555]. mammalian [715]. Management [1098, 1410, 9]. manifold [933]. Manual [117]. many [1155]. many-one [1155]. Marcel [370, 405, 406, 766, 590, 557, 645, 283, 745, 965]. MAREG [950]. Marginal [224, 258, 397, 950, 937]. marginals [380, 1144]. Markov [1223, 1309, 1047, 772, 1251]. Markov-modulated [772]. MARS [1277]. Marvin [116]. master [1410]. matching [400]. Mathematica [1359, 454]. Mathematical [1289, 1350, 1403, 822, 1072, 1033, 193]. mathematicians [894]. matrices [1404, 160, 873, 1159, 1334, 80, 555, 72, 962, 784, 1019]. matrix [60, 540, 127, 806, 343, 751, 889, 1251, 1062, 78, 107]. matrixes [127]. matroids [36]. maximally [898]. maximally-skewed [898]. maximizing [497]. Maximum [1028, 695, 944, 1125, 1102, 936, 1256, 530, 868, 126, 310, 1322, 984, 380, 624, 1270, 48, 999, 1031, 1328, 1418, 285]. maximum-likelihood [1270]. McConway [589]. McCullagh [1185]. McCulloch [127]. MCD [317]. McGraw [34, 35, 249]. McGraw-Hill [34, 35, 249]. MCMC [1250]. Mean [78, 54, 901, 1, 1337, 1332, 613, 106, 1283]. mean-shift [1283]. means [378, 957, 121, 222, 122, 202, 679, 755, 486, 125]. measure [1384, 1386, 782, 255, 899, 728, 924]. measured [443]. Measurement [121, 580]. measurements [1357, 1331, 316, 365]. measures [815, 611, 387, 444, 414, 1268, 39, 784]. Measuring [477, 497, 900]. median [781, 384, 639, 12, 110, 446, 419]. median-unbiased [781]. medians [1149]. Medical [1035, 1413, 133]. Medicine [1024]. meeting [1069, 188, 951]. Meetings [1217, 788, 102, 119, 134, 148, 164, 175, 196, 217, 233, 252, 270, 287, 305, 325, 336, 348, 375, 392, 410, 426, 436, 359, 388, 456, 482, 494, 507, 524, 548, 561, 572, 470, 621, 648, 660, 685, 730, 593, 636, 674, 704, 768, 799, 885, 896, 748, 777, 826, 838, 859, 974, 989, 1041, 1076, 1088, 910, 927, 941, 954, 1005, 1053, 1114, 1130, 1153, 1164, 1193, 1203, 1100, 1142, 1177, 1214, 1299, 356, 646, 1232]. meetings-new [1153]. members [620]. Memoriam [471]. memory [529, 1122, 75]. Mendenhall [193]. mergers [169]. Messungen [1357]. Meta [820]. Meta-analysis [820]. Meteorological [8, 3]. method [205, 809, 1259, 1240, 699, 996, 474, 584, 387, 714, 944, 816, 340, 551, 1322, 516, 1104, 312, 74, 1312, 313, 624, 275, 697, 491, 182, 1143, 365, 1146, 206, 203, 333, 1186, 1260]. Methoden [468, 1373]. Methodik [452, 1357]. Methodological [1383, 1049]. methodologies [1026]. Methodology [1030, 519, 1357, 1410]. Methods [1400, 683, 1390, 1373, 498, 535, 360, 934, 740, 199, 371, 743, 1298, 396, 1364, 1378, 513, 460, 1123, 150, 28, 1273, 933, 1267, 654, 1418, 1425, 1357, 1401, 1431, 472, 1349, 668, 892, 904, 1078, 1330, 1306, 918, 1136, 190, 264, 671, 1079, 331, 504, 559, 612, 1045, 318, 194, 1376, 1048, 204, 1018, 285, 967, 822, 1097]. metric [1301, 1339, 1340]. Michael [372, 658]. Michalewicz [966, 631]. Mick [1296]. MicrOsiris [21]. Microsoft [1095, 1303]. Mid [1229]. Mid- [1229]. Mihalsin [421]. milestone [19]. Millennium [1212]. Min [821]. minimax [144, 982, 431]. minimax-estimation [144]. minimization [917, 95]. Minimum [720, 1228, 106, 417, 1007, 463, 665, 317, 72]. mining [1026]. Minitab [20, 450, 228]. Missing [1293, 1291, 1302, 601, 1256]. misspecification [530, 812, 1137]. Misspecifying [1093, 919]. mixed [763, 224, 626, 723, 692, 1047, 919, 725, 1048, 784]. mixed-level [725]. mixed-model [692]. mixed-variable [626]. Mixture [1250, 496, 868, 352, 1305, 1278, 1312]. mixtures [1228]. mixtures/plug [1082]. ML3 [784]. mobility [257]. mode [46, 1200, 1338]. Model [220, 839, 71, 627, 360, 68, 499, 260, 291, 843, 222, 1329, 1170, 153, 154, 1226, 638, 387, 352, 1171, 58, 721, 902, 754, 261, 625, 812, 1172, 1268, 332, 1257, 977, 1278, 229, 851, 692, 351, 806, 1168, 1270, 473, 186, 345, 697, 241, 83, 689, 612, 5, 489, 1283]. Model-free [627]. Modeling [1090, 769, 18, 1011, 168, 1183, 580, 814]. Modelling [1415, 1250, 701, 1344, 1324, 789, 394, 413, 248, 239, 1338, 183, 1148]. Models [1133, 1209, 1097, 54, 1198, 221, 361, 563, 813, 865, 1169, 329, 763, 292, 1255, 916, 150, 1199, 996, 1335, 1342, 476, 773, 1223, 1138, 296, 759, 144, 1154, 307, 223, 1224, 959, 308, 224, 1, 185, 24, 202, 1320, 1275, 1309, 1400, 599, 228, 225, 668, 950, 256, 1158, 623, 682, 832, 1305, 601, 227, 430, 485, 235, 1265, 57, 1252, 710, 711, 1249, 1317, 1256, 1134, 735, 91, 328, 238, 960, 723, 1047, 351, 979, 1137, 383, 1008, 1282, 61, 1254, 240, 81, 961, 1062, 1391, 242, 1120, 1143, 124, 844, 488, 919, 1382, 890, 243, 691]. models [1048, 158, 257, 784, 745, 557, 848]. Moderate [780, 427, 932]. moderately [167]. modification [82]. Modified [935, 999, 1043, 201, 340, 261]. modulated [772]. moment [497, 496]. moments [428, 278, 642, 489]. money [1052]. monotone [1180, 1105]. Monte [425, 1010, 444, 1306, 123, 1254, 52, 1045, 475, 1271]. Moore [72]. Morgenthaler [657]. moving [1199, 1257]. moving-average [1199]. MS [589, 1286]. MS-DOS [589, 1286]. MSAE [291]. multi [830, 160, 1144]. multi-component [830]. multi-variate [1144]. multi-way [160]. multicriteria [754]. multidimensional [519, 1148]. multidimentional [1026]. Multifactor [1239]. multilevel [1324, 185, 1249]. multimedia [1023]. multimodality [564]. Multinomial [158, 260, 1316, 399, 609, 387]. multinormality [1208, 1261]. Multiple [517, 202, 228, 1401, 683, 1092, 1013, 1107, 995, 1385, 809, 981, 420, 412, 274, 774, 903, 827, 713, 846, 997, 1085, 74, 443, 833, 1160, 841, 1245, 1376, 1196, 1018]. Multirelation [608]. Multiresponse [853]. multisample [430]. multistage [291]. Multivariate [1066, 1397, 414, 68, 499, 779, 780, 421, 1294, 450, 1266, 53, 1236, 1082, 1208, 413, 322, 463, 721, 722, 1237, 817, 485, 1085, 899, 1247, 38, 48, 6, 1390, 240, 70, 331, 1408, 1297, 39, 764, 822, 488, 206, 691, 640, 965]. Multiway [1345, 513, 512, 32, 131]. Multrus [372]. Munich [789, 744]. mutual [1327]. MYSTAT [16].

N [505, 559, 467, 633, 746, 847, 191]. Nair [1185]. name [590, 1370]. Naples [1026]. Narayan [965]. nasopharix [10]. National [212]. Natural [459]. Nearest [1181]. Nearest-neighbor [1181]. Nearly [937, 781]. nearness [137]. needed [420, 880]. negative [5]. neighbor [1181]. nested [226]. Net [1059, 964, 1022, 1220, 1108, 1129, 1140, 1150, 1163, 1189, 1190, 1174, 933]. Netherlands [1213]. NetMul [764]. nets [969]. network [994]. networks [1265, 793, 794, 8]. Neuchâtel [199, 359]. Neuchatel [33]. Neural [969, 793, 794, 1265, 994]. Nevada [1141]. News [99, 129, 147, 161, 189, 211, 230, 247, 265, 282, 334, 346, 357, 369, 390, 469, 537, 479, 502, 521, 1059, 926, 1025, 1040, 1057, 1075, 1087, 1022, 1220, 1099, 1140, 570, 545, 1176, 101, 132, 163, 174, 195, 324, 335, 347, 374, 409, 435, 455, 481, 576, 727, 757, 895, 909, 939, 953, 988, 1004, 964, 1110, 1152, 1234, 1235, 1108, 1129, 1150, 1163, 1189, 1190, 391, 1202, 547, 492, 571, 1174, 592, 605, 619, 635, 647, 659, 673, 684, 694, 702, 729, 837, 857, 747, 776, 824]. News [825, 849, 767, 787]. Newsletter [401, 97, 159, 208, 244, 300, 368, 432, 490, 543]. Newton [340]. NeXTSTEP [1371, 658]. Niewiadomska [1073]. Niewiadomska-Bugaj [1073]. Niki [191]. Nikias [821]. Nikulin [925]. nls2 [520]. No [558, 363, 1301, 1339, 590, 1370, 1340, 49, 198]. no-name [590, 1370]. noise [428, 352, 1102, 39]. nominal [1185]. nominal-ordinal [1185]. nominations [1051]. Non [520, 1021, 157, 427, 14, 527, 149, 956, 712, 959, 1208, 1308, 152, 1182, 139, 430, 1261, 711, 65, 487, 1381, 331, 5, 353, 203]. non-classical [149]. non-Gaussian [712, 959, 1182]. non-graphical [331]. non-identical [527, 203]. non-identified [430]. non-independent [711]. non-iterative [1308]. Non-Linear [1021, 520, 152, 1381]. non-linearities [427]. non-multinormality [1208, 1261]. non-negative [5]. non-normal [487]. non-parametric [157, 956, 139, 711, 353]. non-robust [65]. non-stationarity [14]. noncentral [707, 554, 1244, 187, 355, 758]. nonconforming [678]. nonconstant [807]. nonhomogeneous [1335]. nonignorable [979]. noninvertible [1096, 1407]. Nonlinear [1421, 773, 322, 1045, 518, 916, 759, 1158, 997, 946, 700, 108, 95, 280, 1048, 627, 582]. nonlinearity [1255]. nonmonotonic [1335]. nonnegative [94]. nonnormal [1045]. Nonparametric [743, 1240, 584, 394, 1149, 671, 1030, 805, 264, 281, 501, 1260, 783]. Nonresponse [1409, 979, 1296]. Nonstationary [1096, 1407]. norm [199, 298, 108, 95, 72]. normal [68, 769, 59, 378, 610, 496, 109, 202, 944, 817, 1104, 679, 1241, 755, 1337, 486, 899, 487, 6, 999, 240, 386, 206, 758]. normalising [227]. normality [1266, 123]. normally [228]. North [131, 503]. North-Holland [131]. Notation [1049, 1383]. note [1014, 277, 53, 1138, 774, 903, 201, 611, 1341, 40, 207, 333]. Notes [656]. Nov [1421]. November [425, 1055, 1056, 1026]. NPMLE [1168]. NT [933]. NT-net [933]. number [1161, 1014, 92, 295, 146, 678, 496, 869]. numbers [1422, 1186, 505]. Numerical [1378, 913, 532, 418, 371, 623, 127, 531, 1251, 206, 1161, 1014, 559].

O [1074]. Oberwolfach [231]. Obituary [404]. object [658, 1371, 323]. object-oriented [658, 1371, 323]. oblique [1200]. observation [806, 107]. observational [9]. observations [295, 298, 713, 599, 625, 1119, 679, 791, 331, 1245]. observed [151, 665]. obtaining [1313]. Occasion [646]. occurring [1091]. Oct [1298]. October [1216, 1218, 1220, 1217, 1219, 789, 1220, 1217, 1219]. odds [934, 28]. off [884, 312]. offered [421]. officers [304, 1113]. Okazaki [191]. Ökonometrie [1351, 617]. One [1282, 669, 803, 658, 1371, 1232, 83, 5, 589, 1155, 962]. One-sided [1282, 803]. One-step [669]. one-way [83, 5]. only [740, 621, 648, 660, 685, 730, 606, 768, 799, 885, 896, 886, 897, 974, 989, 1041, 1076, 1088, 928, 942, 1042, 1114, 1130, 1153, 1164, 1193, 1203, 1194, 1204, 1299, 1329, 191]. ons [55]. Operating [1276, 708]. Operational [1184]. operations [721]. operator [719]. operators [981]. Opitz [1074]. Opportunities [1001]. Optimal [790, 1316, 308, 184, 58, 318, 79, 500, 1198, 795, 1325, 294, 759, 585, 753, 314, 78, 553]. optimality [81]. optimization [42, 299, 445, 8]. optimized [761]. optimizing [540, 682]. Optimum [4, 290, 60, 700, 400]. Order [1180, 527, 14, 1257, 433, 550, 999, 39, 489]. order-of-magnitude [433]. Order-restricted [1180]. ordered [1046, 688, 1236, 1223, 202, 803, 679, 819]. Ordinal [1430, 1103, 1185, 1305, 1209, 844, 1430, 1074]. organized [971]. oriented [115, 513, 1343, 658, 1371, 323]. orthogonal [1277, 852, 725]. oscillation [1275]. OSDA [1430]. Ostrowski [589]. other [406, 245]. outcome [443]. outcomes [1256]. outlier [412, 457, 816, 181, 1283, 1196]. outliers [957, 958, 413, 1237, 997, 1092, 1317, 98, 1160, 330, 1227]. output [186, 345, 697]. overall [255]. OVERALS [518]. Overdispersion [1133, 710]. Overview [541, 513, 905]. Owen [758]. Oxford [131, 248].

P [370, 373, 925, 766, 717, 1296, 504, 505, 559, 785, 786, 1020, 1021, 503, 302, 465]. P-STAT [302, 465]. P.v.d [558]. Package [44, 85, 112, 55, 15, 17, 1287, 1324, 517, 264]. packages [743, 420, 433, 301, 671, 319]. Padua [971]. pages [32, 33, 34, 35, 590, 1096, 557, 631, 632, 645, 657, 923, 717, 535, 589, 633, 965]. paired [638, 125]. pairwise [666]. Panel [646, 1361, 137, 632]. Pannatier [1020]. Paola [54]. Paper [1191, 115]. paper-back [115]. Paperback [248, 34, 35, 786]. Papers [858, 1141, 1098, 42, 798, 1038, 1127, 1420]. PARAFAC [514]. PARAFAC2 [385]. Parallel [1324, 514, 419, 425, 1127, 1141, 1320, 1322, 1323]. Parameter [1342, 299, 1154, 1226, 24, 714, 126, 695, 944, 1125, 1102, 381, 1134, 4, 700, 52, 612, 183, 994, 73, 1328, 1147, 1081, 557]. parameters [947, 1131, 809, 292, 699, 222, 150, 915, 185, 936, 566, 6, 498, 83, 1107]. parametric [157, 1121, 956, 139, 1305, 711, 1269, 366, 833, 353]. Pareto [381, 73]. Paris [1430]. Part [793, 794]. partial [428, 642]. Partially [833, 1223, 1224, 710, 665]. Participant [1217, 1232, 388, 470, 951, 356]. particular [1236]. partition [48]. partitioned [80]. partitioning [769]. partitioning.com [1292]. partly [58]. Pasadena [798, 1001, 1002, 1003]. Pascal [146]. PASE [425]. passes [19]. Patrick [728]. pattern [1141]. Paul [285, 786]. PC [21, 114, 1353, 319, 491, 249]. PC-90 [114]. PCA [1313]. PDPTA’98 [1141]. Peacock [633]. peak [1262]. Pearson [397, 38]. Pei [405]. Penalized [959, 1183]. Penrose [72]. Percentile [126]. percentiles [46]. Performance [186, 697, 1141, 1267, 902, 626, 1317, 431]. performances [457]. period [857]. periodic [655, 1275]. periodogram [205]. permutation [1294, 722, 1067, 975, 171]. person [1232]. personal [133, 598]. Perturbational [362]. Peter [453, 683, 190, 284]. phase [1276]. PhD [1410]. Philippou [505]. Physica [632, 504, 617, 618]. Physica-Verlag [632, 504, 617, 618]. pitfall [1012]. Pitman [137]. Pittenger [358]. Plackett [341]. planar [152]. Planning [7, 795]. plans [290, 449]. Planung [468]. please [1111]. plot [276, 96, 1208, 13, 1104, 1261, 98, 1272, 423]. plots [354, 476, 831, 625]. PLS [1298]. PLS’99 [1298]. plug-in [1082]. plus [39, 520, 1379, 1063, 785]. PM [665]. point [447, 477, 1002, 126, 1135, 738, 805, 874, 8]. points [829, 53, 98, 242, 841]. Poisson [1335, 1102, 902, 1228, 710, 772]. policy [1415]. Poly [423]. polychotomous [474]. polygon [931]. Polynomial [580, 296, 945]. polynomials [1248]. polytomous [48]. pooling [89]. population [1197, 1329, 1274, 1145, 487]. populations [1236, 1321]. portmanteau [261]. position [900]. positron [1322]. possible [1012]. Posterior [220, 276, 1253, 70]. potential [255]. Pounds [589]. Power [1263, 1286, 361, 1065, 721, 487, 275, 448, 205, 1286]. pp [266, 285, 370, 371, 372, 405, 406, 407, 408, 451, 452, 453, 454, 131, 133, 966, 213, 766, 468, 855, 884, 846, 683, 249, 728, 821, 823, 882, 883, 658, 986, 1049, 1050, 1074, 967, 987, 248, 466, 558, 190, 1018, 1297, 283, 284, 1017, 193, 116, 192, 194, 467, 617, 745, 746, 847, 848, 924, 1019, 1097, 358, 616, 1072, 1073, 191, 1296, 618]. pp. [822]. Practical [1429, 1353, 1167, 738, 1360, 1017]. practice [552, 1362]. pragmatics [1419, 406]. praktische [1353, 249]. Pranab [987]. Praxis [1362, 453]. praxisbezogene [372, 1360]. pre [915]. pre-test [915]. Precise [898]. Prediction [998, 57, 478, 740, 1199, 458, 811, 831, 140, 1265, 1134]. predictive [1132, 1171]. predictors [1256, 999, 78]. Preface [511, 136, 219, 863]. Preference [613]. preliminary [735]. Prem [717]. Prentice [823]. prepared [41]. preprocessing [602]. presence [1083, 1317, 74, 914, 762]. presentations [113]. President [1112, 210, 878]. Press [133, 248, 283, 284, 1017, 1021, 116, 1268]. Price [32, 33, 746]. primal [310]. primal-dual [310]. Principal [1146, 222, 429, 654, 1103, 236, 65, 624, 1148]. print [1111]. prior [139]. priori [1301, 1339, 1281, 1340]. priors [1131, 61, 70]. prize [1015]. Probabilistic [865, 12]. probabilities [151, 713, 278, 904, 187, 206, 355, 758, 1147]. Probability [1393, 1386, 1403, 1097, 1197, 276, 1287, 63, 1325, 1082, 109, 864, 235, 1337, 138, 752, 822, 1186, 728, 1072, 1073]. probit [1259]. problem [1010, 152, 912, 310, 565, 1312, 487, 872, 698, 39, 398]. problems [550, 3]. PROC [741]. procedure [339, 420, 274, 122, 982, 687, 600, 1237, 167, 351]. procedures [1192, 185, 228, 1402, 1167, 140, 755, 1303, 987, 1195, 142, 1107]. Proceedings [231, 968, 1422, 1421, 1430]. process [428, 712, 655]. processes [14, 1122, 1378, 1335, 1357, 445, 738, 364, 772, 79, 1124, 1314, 559]. processing [1023, 1127, 1141, 821, 1389, 283, 1412]. Procrustes [398]. products [80]. Prof. [1112]. Profile [1112, 1154]. ProGAMMA [556]. prognosis [10]. prognostic [761]. program [112, 973, 1286, 1094, 420, 1321, 985, 319, 491, 187]. Programmer [1394, 1050]. programming [596, 658, 1371, 313, 1374, 786, 1033]. programs [1277, 515, 1405, 966]. Projection [427, 677, 724, 780, 1034]. proper [1014]. Properties [183, 59, 1285, 40, 553]. property [585, 75]. proportional [600, 914, 1334]. proportions [500, 29, 773, 1145]. proposals [277]. proposed [1341]. prototype [542, 541]. Prozesse [1357]. Pseudo [530, 237, 1242, 792]. pseudo-factors [1242]. Pseudo-likelihood [237]. pseudo-linear [792]. pseudorandom [1161]. pseudosample [140]. Pseudotable [204]. PSI [423]. PSI-Plot [423]. Publication [190, 1090]. Publications [786]. Publishers [32, 33, 115, 285, 213, 466, 717, 505, 116, 616]. Publishing [193]. pursuit [427, 780, 1034, 677, 724]. PWS [193]. PWS-KENT [193].

Q [128]. Q-designs [128]. QR [279, 612]. quadratic [1132, 395]. Quadrature [1079]. Quality [406, 693, 1349, 449, 1419, 190].

Quantile [831, 917, 650, 706, 1226, 782, 1238, 808, 1009]. quantiles [1266, 706, 699, 1244, 1308, 669]. quantitative [1104, 1306]. quantities [149]. Quasi [413, 237, 773, 1275, 550]. Quasi-Bayesian [413]. quasi-likelihood [237, 773]. quasi-order [550]. quasi-periodic [1275]. question [876]. Questionnaire [620]. quick [262]. quota [609]. quota-free [609].

R [32, 266, 371, 406, 451, 131, 133, 231, 193, 503, 1097]. R.-D [231]. Rahlfs [743, 744]. Raj [632]. ran1 [1161, 1014]. Random [844, 378, 1207, 1014, 527, 699, 399, 146, 299, 213, 461, 1346, 1179, 539, 682, 442, 38, 6, 1254, 83, 5, 919, 1263]. random-effects [919]. Randomization [1085, 93, 680]. randomized [892, 977]. range [421, 891, 1308, 803]. Rank [829, 815, 1246, 459, 688, 1010, 1236, 584, 913, 586, 1239, 568, 555, 40, 962]. Rank-based [815, 1246]. rank-one [962]. ranks [459]. Rao [503, 1206, 240]. Rapid [1000]. rate [1265]. rates [1010, 791, 381, 487, 627]. ratio [934, 46, 28, 810, 913, 709]. rational [109]. ratios [1131]. räumlichen [452, 1357]. raumzeitlicher [1357]. Re [1323]. Re-engineering [1323]. Recht [536]. Recipes [1161, 1014]. reciprocals [339]. Recognition [870, 1141]. reconsidered [931]. reconstruction [1322]. record [75]. rectangular [1184]. recurrent [394]. Recursive [1292, 595, 769, 599, 904, 168, 498]. Recursive-partitioning.com [1292]. Reducing [1264]. Reduction [962, 651, 395, 1314]. Reepmeyer [618]. referee [921, 920]. Referees [218, 338, 1126, 1205, 906]. Reference [1288, 653, 264]. Refined [722]. refinement [850, 612]. Reflections [646, 388, 470, 356]. reflective [813]. Reflextions [1217]. regional [267]. registration [1111]. Regression [1347, 823, 739, 1097, 365, 68, 499, 769, 427, 780, 1043, 969, 595, 1207, 1064, 667, 1364, 1259, 650, 520, 1169, 696, 957, 329, 916, 706, 770, 1066, 996, 141, 245, 474, 28, 1170, 296, 759, 297, 675, 774, 903, 298, 307, 959, 458, 1118, 1030, 255, 384, 462, 639, 1333, 850, 1400, 816, 607, 1388, 12, 902, 950, 625, 140, 623, 682, 839, 1305, 945, 917, 442, 711, 918, 110, 1256, 328, 960, 805, 851, 383, 1008, 123, 550, 700, 98, 262, 446, 491, 431, 793, 1160, 1272, 1318, 279, 612, 853, 142, 111, 353, 475, 78]. regression [318, 1031, 124, 844, 794, 415, 1283, 207, 807, 419, 1263, 1009, 783, 784, 535, 116, 194]. Regression-free [739]. regressions [956, 997, 236]. regressors [501]. Regularization [905]. Reinhard [407]. Reisensburg [113]. Reiss [231]. rejection [181]. Related [656, 99, 129, 147, 161, 189, 211, 230, 247, 265, 282, 199, 334, 346, 357, 369, 390, 479, 502, 521, 570, 545, 998]. relation [540]. relations [71]. relationship [1197, 901, 1238, 1338]. relationships [341]. relative [1197, 809, 497]. relevant [740, 599]. Reliability [1417, 532, 976, 290, 610, 1335, 1342, 557, 531, 284, 645]. Reliabilty [1366]. remote [352, 681, 3]. Rencher [822]. renewal [932, 1047]. Reparameterisation [1250]. reparameterization [976]. Repeated [760, 815, 444, 1331, 888, 1268, 365, 1155, 784]. replicated [1274]. replication [49, 198]. replications [363]. Reply [744]. Report [1052, 560, 532, 44, 85]. report/review [44, 85]. Representation [602, 628]. representatives [269]. requirements [412]. Rerandomization [396]. Resampled [1238, 1007]. Resampling [1376, 1293, 996, 683]. Resampling-based [1376, 683]. Research [1414, 1289, 795, 544, 1410]. researchers [1015, 542, 541, 1374, 786]. residual [1064, 394, 982, 224, 1257, 692]. residuals [612]. Resistant [555]. resolution [47]. respect [429]. Response [1340, 780, 291, 361, 948, 831, 485, 977, 1281, 651, 979, 1282, 240, 158]. responses [888, 603]. restoration [1102]. Restricted [945, 315, 1180, 307, 585, 383, 1008]. restrictions [144, 318]. Restrictive [565]. Results [857, 993, 888, 11, 1243]. Review [32, 33, 34, 35, 115, 266, 285, 370, 371, 372, 373, 405, 406, 407, 408, 451, 452, 453, 454, 604, 938, 925, 131, 133, 536, 966, 213, 766, 468, 855, 884, 590, 846, 693, 683, 249, 728, 821, 823, 882, 883, 1096, 557, 231, 631, 632, 645, 657, 658, 986, 1049, 1050, 1074, 967, 987, 923, 248, 466, 558, 190, 1018, 717, 1297, 504, 505, 559, 283, 284, 534, 785, 786, 1017, 1020, 1021, 193, 116, 192, 194, 467, 503, 535, 589, 617, 633, 745, 746, 822, 847, 848, 924, 965, 1019]. Review [1097, 358, 616, 1072, 1073, 191, 1296, 618, 44, 85, 544, 1032, 320, 314, 69, 7, 784]. Reviews [1056, 1218, 439, 575]. revisited [448]. Revisiting [1161]. reweighted [1092]. reweighting [1254]. ridge [1043, 1318]. Rieder [967]. right [1256]. risks [843]. rival [308]. Rivoirard [1021]. Robert [846, 823, 924, 1073, 1296]. robot [1421]. Robotics [1421]. Robust [259, 995, 458, 1402, 446, 475, 1196, 595, 696, 167, 918, 65, 69, 262, 739, 751, 498, 171, 1063, 567, 330, 207, 553, 1431, 967, 987]. Robustness [295, 47, 700, 71, 1426, 657]. Rocke [1339]. roles [625]. Rome [32]. Ronald [1097]. Ronchetti [657]. root [1285, 1044]. roots [889]. Rosenblatt [1121]. Ross [34]. rotation [1200]. rotations [1148]. roughening [139]. Rouhi [27]. row [13, 760]. row-column [13, 760]. rule [82]. rules [563, 400]. run [417, 1337].

S [32, 34, 373, 925, 131, 884, 683, 657, 986, 1050, 785, 467, 746, 847, 520, 1379, 1063, 785]. S-PLUS [520, 1379, 1063, 785]. S. [213].

Sachs [468]. saddlepoint [366, 669]. Safe [957]. Sage [786]. Salerno [32]. Sample [970, 412, 1185, 1010, 274, 28, 1007, 774, 903, 827, 1267, 638, 309, 721, 565, 812, 888, 1317, 1253, 1285, 487, 1206, 448, 69, 698, 171, 1033]. sampled [428]. samples [112, 947, 852, 999, 680, 1144]. Sampling [92, 1197, 763, 237, 1138, 1134, 1012]. Sandra [1017]. Santner [846]. SAS [1395, 1356, 228, 820, 1085, 1360, 741, 1374, 263, 986, 372, 408, 786]. SAS-System [1360]. Satellite [1069, 1070]. Saunders [360]. scale [1131, 915, 90, 565, 381, 739, 73]. scaled [200]. scaling [519, 398]. scan [278, 904]. scatter [873]. scatterplot [664]. scatterplots [50]. Schaeffer [193]. Schätzungen [1357]. Scheffé [74]. scheme [332]. schemes [237]. Schneeweiß [617]. School [1212, 971, 229]. Schools [349, 411, 437, 120, 165, 176, 197, 234, 253, 271, 288, 306, 326, 337, 393, 483, 495, 508, 525, 549, 562, 573, 622, 637, 649, 661, 686, 705, 749, 778, 886, 897, 911, 928, 942, 955, 990, 1006, 1042, 1054, 1101, 1178, 1194, 1204, 1215]. Schwecke [371]. Science [32, 33, 115, 971, 248, 284, 1410, 1109]. sciences [1352, 358]. Scienctific [1098]. Scientific [1423, 631, 1354]. Scientist [191]. scientists [1374, 786]. scores [689]. Scotland [1422]. screenable [1156]. screening [1385, 846, 892]. search [753, 682, 351]. seasonal [1275]. seasonally [655]. Sebastiani [54]. secant [40]. Second [788, 798, 14, 199, 798]. second-order [14]. Section [1191, 1293, 232, 267, 1000, 822]. seemingly [623]. segmentation [344]. Selection [1336, 654, 626, 1385, 1315, 184, 687, 816, 1171, 846, 668, 902, 839, 1268, 624]. selectors [783]. semantic [332]. semi [1269, 1260]. semi-Bayesian [1260]. semi-parametric [1269]. Sen [987, 194]. senses [597]. sensing [352, 681]. Sensitivity [891, 186, 345, 697, 890, 485, 1106, 39].

Separation [781]. September [788, 356, 1055, 1056]. sequences [417, 701]. Sequential [124, 708, 1321, 1331, 7]. serially [996]. Series [1096, 248, 504, 1097, 779, 290, 1315, 273, 1411, 958, 852, 1396, 1182, 1274, 1275, 946, 1407, 822, 1048, 627, 882, 883]. service [1109]. services [644]. session [1141, 188]. set [384, 602, 534]. sets [1288, 809, 412, 866, 1428, 1148, 518, 589]. setting [458]. settings [149]. several [339, 460, 222, 274, 668, 539, 978, 48, 1271]. SGCSA [356]. Shafer [1297]. shape [183, 1147]. shapes [1281, 203]. Sharp [380]. Shein [405, 745]. Shein-Chung [405, 745]. shift [1283]. shifted [275]. Shoa [821]. Short [797]. should [172]. showing [26, 25]. sided [803, 1282]. sight [320]. Sigma [133]. Signal [1389, 1023, 39, 821]. signal-plus-noise [39]. signals [566, 4]. Significance [834, 1210, 418]. Simar [504]. similar [603]. similarity [333]. Simon [884]. Simon-off [884]. Simple [150, 402, 68, 934, 1313, 664, 803, 607, 1200, 1078, 364, 751]. SIMPLIMAX [1200]. simply [679]. Simson [658]. simulated [184, 317]. Simulation [245, 433, 740, 1094, 28, 712, 993, 902, 892, 478, 651, 52, 752, 353, 475]. Simultaneous [956, 1199, 486, 339, 1280, 816, 1171, 666, 1013]. Singapore [631]. single [866, 444, 755]. single-stage [755]. singular [221, 46, 850, 817]. singular-value [850]. SIR [22]. SIR/Easy [22]. size [1197, 934, 412, 1329, 1065, 1206]. sizes [350, 274, 721, 812, 487]. skewed [898]. slope [807]. Small [1185, 1285, 69, 28, 774, 903, 852, 1428, 1275, 812, 888, 1206, 448, 589]. Small-sample [1185, 1285, 448]. SMART [1289]. Smirnov [484, 980, 680]. Smooth [311, 582]. Smoothed [1080]. smoothers [552]. Smoothing [1182, 139, 650, 460, 90, 959, 664, 1102, 690, 960, 808, 491]. smoothness [61]. snow [811]. social [41, 1374, 786, 257]. Society [118, 1024]. Softstat [1423, 466, 266]. Software [41, 266, 389, 401, 576, 604, 727, 938, 1071, 1015, 1128, 1233, 1219, 1414, 464, 756, 466, 1020, 1352, 17, 18, 99, 129, 147, 161, 114, 97, 159, 246, 303, 189, 211, 230, 247, 265, 282, 208, 244, 300, 422, 334, 346, 357, 369, 390, 423, 368, 432, 479, 502, 521, 490, 543, 588, 1026, 1288, 513, 1335, 172, 1423, 387, 570, 545, 1011, 693, 581, 1323, 1406, 209, 845, 764, 358, 113, 536]. sold [191]. Soliman [27]. Solo [55]. solution [462, 463, 639, 1258, 850, 623]. solutions [487]. solver [1320]. Some [59, 67, 50, 1301, 759, 933, 1339, 597, 27, 1167, 1078, 888, 11, 1340, 3, 72, 527, 273, 830, 444, 993, 755, 586, 918, 238, 314, 1285, 66, 262, 353, 603]. Song [745]. Song-Gui [745]. Sons [925, 884, 846, 923, 190, 559, 535, 1019, 358]. sorting [519]. sounding [3]. sources [1030]. Soviet [894]. space [53]. spaces [866]. spacings [1077]. span [1342, 557]. Sparse [1062, 258, 397, 24, 937, 1251]. Spatial [1020, 1429, 712, 519, 1357, 227, 738, 960, 1406, 364, 321, 1017, 243]. spatial-temporal [364]. Spatio [811, 1357]. Spatio-temporal [811, 1357]. Special [1127, 1141, 264]. specification [754, 1272]. Spector [786]. spectral [738, 1262]. spectrum [1274, 205]. spiral [35]. spiral-bound [35]. spline [429, 853, 316]. splines [650, 842, 770, 1066, 90, 737, 180, 808]. split [13, 760]. split-plot [13]. spread [1105]. spreadsheet [424]. spring [60]. Springer [371, 536, 966, 468, 231, 658, 1050, 1074, 967, 504, 1020, 194]. SPSS [15, 1353, 228, 407, 249]. SPSS/PC [1353, 249]. square [827, 380, 1044, 819, 78]. Squared [925, 790, 1244, 1399]. squares [1064, 706, 143, 185, 152, 384, 462, 639, 607, 12, 385, 750, 180, 946, 1257, 1278, 719, 446, 280, 489, 419, 72]. Srivastava [194]. SSDBM [1098, 215]. SSN [921].

SSNinCSDA [743, 438, 439, 574, 575, 576, 731, 860, 1055, 1056, 1216, 1218, 1217]. St [1422]. Stability [298, 65, 900]. stabilization [89]. stable [245, 821, 898, 1389]. stage [1094, 1309, 755, 487]. Stahel [657]. standard [743, 387, 671, 1276, 77]. standardized [613]. Stanley [683]. Stapleton [848]. starting [497]. STAT [302, 465, 323]. State [717, 1424]. Statement [878, 210]. states [894]. stationarity [14]. STATIS [516]. statistic [696, 51, 278, 687, 913, 904, 832]. Statistical [1344, 114, 97, 159, 251, 208, 244, 300, 368, 401, 432, 490, 543, 604, 798, 938, 971, 1191, 1213, 1290, 1289, 788, 1367, 1098, 402, 226, 1349, 1330, 1354, 248, 1136, 301, 466, 263, 1124, 41, 99, 129, 147, 161, 246, 303, 212, 189, 211, 230, 247, 265, 282, 199, 422, 334, 346, 357, 369, 390, 479, 502, 521, 588, 796, 907, 1192, 1288, 1393, 1385, 420, 292, 789, 854, 172, 1379, 1395, 1414, 1423, 184, 1396, 387, 1397, 542, 356, 570, 599, 545, 846, 1011, 693, 1402, 1231, 137, 726]. statistical [1095, 113, 449, 883, 876, 657, 986, 581, 1303, 987, 1426, 1323, 323, 978, 985, 264, 319, 541, 1391, 1338, 785, 848, 965, 1073, 40, 8, 1416, 1368, 1373, 1412, 33, 389, 464, 756, 631, 248, 190, 283, 633]. Statistics [1352, 54, 970, 1127, 1001, 1339, 1125, 1369, 922, 321, 1097, 55, 42, 20, 115, 232, 286, 425, 424, 796, 743, 797, 1023, 971, 879, 1069, 1429, 1420, 527, 361, 63, 1387, 867, 544, 1425, 864, 884, 1431, 1388, 1231, 823, 644, 1343, 656, 876, 1350, 967, 671, 1372, 999, 69, 1070, 504, 171, 193, 503, 822, 207, 1212, 1362, 1127, 504, 1017, 503, 191, 358]. Statistik [266, 1414, 1369, 1362, 558, 453]. Statistik-Software [266, 1414]. Statistische [1368, 1373, 451, 468]. StatSci [422, 421]. Stayers [1047]. Stefano [425]. step [658, 755, 1371, 792, 669]. step-down [755]. step-up [755]. Stephan [454]. Steven [34]. Stiefel [933]. Stieltjes [759]. stimulus [519]. stochastic [1378, 293, 1277, 153, 519, 430, 313, 559]. Stone [709]. strategies [68, 763, 830]. strategy [542]. stratification [168]. stratified [1309]. string [585]. Structural [82, 563, 430, 473, 1120, 1143, 890]. structure [1313, 1182, 340, 1200, 550, 66, 342, 1081]. structured [784]. structures [966, 789, 1405, 875]. Student [1191, 1002, 358, 1341, 1241]. Studentized [1308, 803]. students [1015, 1410]. Studies [1357, 1365, 332, 1312, 405]. Study [41, 1197, 740, 1131, 1010, 460, 1199, 28, 519, 1347, 830, 444, 551, 170, 832, 139, 586, 918, 123, 700, 239, 1251, 5, 353, 116]. studying [1273]. Stuttgard [453]. Stuttgart [266, 373, 407, 408, 466, 558]. sub [665]. sub-tables [665]. subject [1102, 719]. submanifolds [292]. subsets [1064, 668]. subvectors [6]. successive [365]. suite [515]. sum [459, 1333]. summarize [276]. Summer [349, 411, 1212, 120, 165, 176, 197, 234, 253, 271, 288, 306, 326, 337, 393, 483, 495, 508, 525, 549, 562, 573, 622, 637, 649, 661, 686, 705, 749, 778, 886, 897, 797, 911, 928, 942, 955, 990, 1006, 1042, 1054, 971, 1101, 1178, 1194, 1204, 1215, 229, 437]. sums [981, 461, 1077, 1029]. superiority [81]. supervised [681]. supplement [18]. support [894]. supported [1044]. supremum [1091]. surface [831]. Survey [788, 41, 736, 356, 225, 1403, 1072]. surveying [1136]. surveys [1409, 316, 1296]. Survival [168, 10, 1273, 1149, 1091, 812, 568, 1424, 717]. SWEEP [719]. Switzerland [33, 425, 199]. Sydney [631]. Symbolic [981, 1430, 1074]. symmetric [167]. symmetry [67, 1065]. symposium [251, 1298, 1417, 284]. synthesis [1032, 602]. SYSTAT [16, 17, 18]. System [402, 290, 372, 542, 921, 693, 581, 1360, 920, 541, 1360]. system-independent [581]. Systems [532, 1383, 99, 129, 147, 161, 189, 211, 230, 247, 265, 282, 334, 346, 357, 369, 390, 479, 502, 521, 588, 133, 1413, 597, 830, 544, 570, 545, 4, 531, 79, 1049, 371].

T [408, 590, 191]. Tab [504, 505]. table [220]. Tables [417, 258, 1180, 397, 24, 771, 380, 937, 160, 1247, 819, 665, 1029, 961, 49, 198, 204, 1227, 257]. tabu [753]. tabulation [898]. Tafeln [373]. tail [1089, 1240, 151, 278, 355, 1147]. tailed [167]. tails [169]. Tanaka [1096]. Tartu [796]. Taschenbücher [373, 558]. Taudes [618]. Taylor [1048]. Taylor-series [1048]. teaching [968]. technical [1141]. technique [26, 712, 25, 1239]. Techniques [1289, 42, 421, 1023, 1141, 667, 1377, 1122, 412, 599, 186, 345, 833, 1063, 616]. technologies [1023]. Technology [1417, 1289, 284]. Telegram [587, 614, 629, 742, 765, 881, 893, 908, 940, 963, 1016, 1036, 1058, 1219, 1139, 1162, 1173, 1187, 1188, 716, 1201, 835]. Telemedicine [1035, 1037]. temporal [1357, 811, 364]. Temporality [63]. tensor [737]. term [807]. ternary [1186]. test [1121, 290, 500, 643, 790, 1316, 459, 339, 1046, 420, 46, 1236, 915, 538, 457, 810, 444, 803, 1280, 1065, 775, 913, 656, 1091, 382, 1085, 1281, 123, 6, 448, 980, 1067, 1106, 698, 1334, 1029, 1262, 1155, 207]. Testing [378, 916, 843, 564, 496, 652, 683, 679, 711, 123, 531, 49, 1294, 96, 675, 774, 903, 1399, 830, 803, 1311, 1210, 1013, 1376, 198, 925]. tests [829, 688, 363, 474, 153, 297, 1304, 935, 93, 484, 722, 261, 937, 565, 1331, 818, 709, 586, 74, 1285, 1206, 1282, 834, 975, 680, 613, 281, 568, 1246, 1081]. their [667, 1422, 854, 512, 933, 387, 626, 240, 1250, 80]. theorem [145]. theorems [67]. theoretic [157]. Theory [406, 1346, 1018, 1143, 118, 530, 1401, 1231, 1103, 1403, 1419, 1247, 978, 792, 1407, 194, 1382, 1420, 1096, 231, 1072, 213, 191]. there [420]. Thomas [846]. Three [1200, 961, 46, 363, 417, 1321, 202, 126, 695, 944, 1125, 385, 515, 755, 382, 602, 10, 186, 680, 52, 1146, 1328, 342]. three-degree [382]. Three-factor [961]. three-level [417, 342]. three-mode [46, 1200]. three-parameter [695, 944, 1125, 52]. Three-way [1200, 46, 385, 515, 382, 602, 961, 1146]. threshold [226, 1310, 691]. thresholding [828, 805]. ties [420]. Time [1407, 779, 1315, 273, 1411, 958, 1170, 1396, 1182, 1274, 1275, 892, 883, 946, 1134, 627, 882, 1096]. time-series [882]. time-varying-parameter [1134]. Tips [588, 615, 630, 672]. today [588, 433]. Tokyo [131, 631, 503, 191]. tolerance [653]. Tomás [597, 594]. tomography [1322]. tool [795, 950, 1359]. tools [16, 871, 1062]. topical [606]. Toronto [631]. total [1197, 899]. tracker [792]. trade [1415, 312]. trades [64]. traits [1306]. transfer [17]. transform [1032, 1117, 1239, 814, 708]. transformation [1010, 676, 1171, 723, 275, 1263]. transformations [429]. transforms [759, 690]. Treating [1012]. Treatment [1156, 891, 638, 565]. treatments [339, 420, 815, 13]. trees [141, 949]. trend [1180, 417, 1281, 241]. trend-free [417]. trial [1276]. trials [1229, 1316, 708, 892, 7]. triangle [922]. triangular [723]. Tricks [588, 615, 630, 672]. trilateration [1136]. Trimmed [138, 1028, 462, 1333, 607]. trisection [152]. Troia [534, 1427]. truncated [1332, 1168, 1120]. Tübingen [743]. TUCKALS [515]. Turbo [146]. tutor [1352, 358]. tutorial [1352, 358]. twin [1186]. Two [1269, 316, 1048, 829, 500, 14, 1094, 363, 1010, 293, 1236, 915, 245, 29, 417, 608, 1309, 552, 256, 261, 565, 485, 341, 1083, 851, 487, 473, 6, 792, 698, 680, 642, 689, 171, 475, 568, 49, 198, 1246, 1227, 784]. two- [417]. Two-dimensional [316, 14, 552]. two-group [475]. two-level [341, 1083, 851, 473, 689]. two-sample [1010, 565, 171]. two-stage [1309, 487]. two-step [792]. two-way [363, 256, 49, 198, 1246, 1227]. Type [200, 1010, 487, 696, 90, 166, 611, 331, 75].

U [452]. U.S [406]. überarbeitete [453, 468]. UK [788, 858, 1027, 356]. umbrella [1280, 486]. unavailability [295, 47]. unbalanced [956, 202, 301, 263]. unbiased [781, 74]. Uncertainty [371]. unconditional [790, 1316]. unconditioned [500]. unequal [1197, 350, 444]. unequally [93]. Uni [373, 558]. Unified [628, 1329, 1028, 1210]. uniform [933]. unimodal [276]. Union [894]. UNISTAT [424]. unit [1285]. units [222, 294, 184, 760]. Univariate [1358, 429, 1380, 960, 746, 847, 1363, 373, 467]. universally [309]. Universitäts [452]. Universiteit [1049]. University [32, 33, 1213, 1422]. UNIX [1375, 618]. unknown [180, 1281, 1012]. Unmasking [1227]. unnecessary [567]. unordered [771, 1029]. unrelated [623]. unreplicated [1104, 272]. unsafe [957]. unstructured [806]. Untersuchungen [1357]. unweighted [750]. upper [46, 891, 715]. Uppuluri [406]. Urgent [620]. US$131.50/Dfl [33]. US$135.00 [965]. US$136.75/Dfl [32]. US$190.00 [503]. US$21.95 [358]. US$51.00 [1296]. US$57.50 [535]. US$59.75 [590]. US$68.95 [467]. US$92 [616]. US$99.75 [370, 557]. USA [1418, 1001, 1003, 286, 798, 1141]. Use [708, 1423, 1289, 1052, 1131, 1014, 611, 1179, 1264, 960, 279]. useful [976]. user [513, 581, 433, 624, 764]. user-knowledge-based [624]. user-oriented [513]. users [1375]. Using [276, 420, 1138, 93, 360, 220, 588, 221, 361, 133, 842, 1240, 223, 1379, 1395, 496, 387, 944, 811, 1011, 1090, 1330, 986, 917, 57, 711, 1247, 624, 319, 446, 873, 741, 612, 1045, 785, 365, 1031, 330, 1186, 784, 35]. UTB [373]. utility [985]. utilizing [478]. Utrecht [1213]. Uwe [604, 938].

V [406, 743, 213, 590]. v1.3 [1287]. vagueness [371]. validation [201, 281]. Validität [1357]. validity [790, 1357]. Value [683, 1229, 221, 675, 850, 231, 11, 1008, 752, 1376]. valued [566]. values [46, 497, 687, 761, 840, 975, 75]. variability [899, 602]. variable [1197, 934, 1132, 816, 1171, 225, 449, 626, 624, 752]. variables [378, 527, 1181, 699, 1336, 608, 429, 676, 551, 625, 539, 626, 1209, 791, 443, 651, 1012, 603, 518]. Variance [651, 157, 54, 1118, 1, 1400, 88, 1167, 831, 1239, 1083, 1257, 692, 5, 807, 1363, 1097]. variances [809, 916, 1225, 224, 387, 256, 818, 946, 978, 1012, 203]. Varianzanalysen [1363, 373]. variate [1144]. variates [399, 1078, 48, 105]. variation [700, 834, 181, 106]. Variogram [94]. variograms [814]. VARIOWIN [1406, 1020]. varying [1134]. VCH [466]. Vector [61, 68, 6]. Vegas [1141]. Verfahren [1357, 452]. Verita [425]. Verlag [266, 372, 373, 407, 408, 451, 452, 454, 536, 632, 657, 658, 466, 558, 504, 1020, 617, 618]. Version [408, 423, 1356, 1208, 1091, 319]. versions [293, 145]. versus [969, 259]. very [643, 751]. VI [215]. via [429, 184, 723, 1044, 48, 366, 145, 1120]. Vienna [646, 560]. view [1002, 1232, 8]. views [329, 951, 1232]. viii [504]. visualization [214, 1023]. Vol [534, 746, 847]. Volker [744]. Volume [731, 860, 1055, 1056, 43, 84, 44, 85, 178, 438, 376, 440, 439, 574, 509, 577, 575, 576, 662, 733, 800, 861, 929, 991, 1060, 1115, 1165, 1221, 1300, 1007, 505]. Volumes [1216, 1218, 1220, 1217, 1219]. vs [1236].

W [743, 744, 645, 657, 190, 504, 193, 467, 1072]. Wackerly [193]. walks [1179]. Wallis [1085]. Walter [1050]. Wang [745]. water [811]. Wavelet [1207, 828, 805]. wavelets [1044]. way [1313, 46, 363, 1007, 256, 385, 1200, 515, 382, 160, 602, 1106, 1012, 83, 961, 5, 49, 198, 1146, 1246, 1227]. Wayne [883]. weak [1197]. weakened [318]. Web [764]. website [1295, 1290]. Wegman [283, 1112]. Weibull [976, 932, 126, 478, 76]. Weighing [1119, 60]. Weighted [840, 296, 1090, 946, 1283]. weighting [809]. weights [1264]. Welcome [877]. Werkzeug [1359, 454]. Westfall [683]. Where [879, 529]. which [781]. white [428]. Whither [646]. Whitten [557]. whole [13]. whole-plot [13]. Wide [764]. width [1111, 666]. Wiesbaden [452]. Wilcoxon [459]. Wiley [925, 884, 846, 683, 728, 821, 883, 1096, 987, 923, 190, 559, 467, 535, 633, 746, 822, 847, 848, 924, 1019, 1097, 358, 1072, 1073, 1296]. Wiley-Interscience [190]. Wilmslow [133]. Win32 [1287]. windows [424]. WinMAREG [950]. Wissenschaft [373, 558]. Wissenschaftliche [1414]. without [1010]. Wittenberg [407]. Wittkowski [743, 744]. Woodruff [1339]. word [597]. Workshop [1026, 1420, 1421, 286, 393, 425, 789]. Workshops [349, 411, 437, 103, 120, 165, 176, 197, 234, 253, 271, 288, 306, 326, 337, 483, 495, 508, 525, 549, 562, 573, 622, 637, 649, 661, 686, 705, 749, 778, 886, 897, 911, 928, 942, 955, 990, 1006, 1042, 1054, 1101, 1178, 1194, 1204, 1215]. World [798, 1001, 1002, 1003, 764]. World-Wide [764]. worst [1225, 1302]. worst-case [1225]. wrote [743].

X [925, 1339, 249, 821, 284, 534, 617, 1097, 373, 1096, 717, 745, 785, 745]. XGobi [1034]. xi [923]. xii [33, 925, 965, 1328]. xiii [285, 848]. xiv [32, 213, 1213]. xix [746, 847]. xvi [589, 822, 1019]. xvii [284]. xx [467]. xxxiv [505].

Year [1038]. Yen [191]. York [370, 405, 406, 925, 131, 536, 966, 766, 590, 683, 728, 1096, 557, 631, 645, 658, 967, 987, 923, 466, 504, 559, 283, 284, 467, 503, 535, 745, 746, 847, 848, 924, 965, 1019, 358, 1072, 1073, 1296]. Young [683, 1015]. Yvan [1020].

Zbigniew [966, 631]. Zealand [1290]. zero [887, 125]. zero-adjusted [887]. zeros [1138]. Zimmermann [743, 744]. Zipf [628]. Zöfel [373, 453]. zur [1357, 1357, 452]. zygosity [1186].

References

Giovagnoli:1989:EDM

[1] Alessandra Giovagnoli and Paola Sebastiani. Experimental designs for mean and variance estimation in variance components models. Computational Statistics & Data Analysis, 8 (1):21–28, May 1989. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                              //www.sciencedirect.com/science/
                                              article/pii/0167947389900625. See

comments [54].

Anonymous:1990:EBa

[2] Anonymous. Editorial board. Computational Statistics & Data Analysis, 9(1):ii, January 1990. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                              //www.sciencedirect.com/science/
                                              article/pii/016794739090065P.

Uspensky:1990:SNA

[3] Alexandre B. Uspensky. Some new aspects of analysis and design of experiments in inverse problems of remote meteorological sounding. Computational Statistics & Data Analysis, 9(1):1–10, January 1990. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739090066Q.

Rafajlwicz:1990:OIS

[4] Ewaryst Rafajłwicz. Optimum input signals for parameter estimation in systems described by linear integral equations. Computational Statistics & Data Analysis, 9 (1):11–19, January 1990. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739090067R.

Sutradhar:1990:ESE

[5] B. C. Sutradhar and R. F. Bartlett. An empirical study of the estimation of eigenvalues in connection with non-negative estimation of variance components of a one-way random effect model. Computational Statistics & Data Analysis, 9(1):21– 35, January 1990. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739090068S.

Provost:1990:EPM

[6] Serge B. Provost. Estimators for the parameters of a multivariate normal random vector with incomplete data on two subvectors and test of independence. Computational Statistics & Data Analysis, 9(1):37– 46, January 1990. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739090069T.

Salvan:1990:PSC

[7] Alessandra Salvan. Planning sequential clinical trials: a review. Computational Statistics & Data Analysis, 9(1):47–56, January 1990. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739090070X.

derMegreditchian:1990:MNO

[8] Guy der Megreditchian. Meteorological networks optimization from a statistical point of view. Computational Statistics & Data Analysis, 9 (1):57–75, January 1990. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739090071O.

Varis:1990:BID

[9] Olli Varis, Juhani Kettunen, and Hannu Sirviö. Bayesian influence diagram approach to complex environmental management including observational design. Computational Statistics & Data Analysis, 9(1):77– 91, January 1990. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739090072P.

Rossi:1990:SAP

[10] Carla Rossi. Survival analysis and prognosis: a comparison of three approaches (for the analysis of nasopharix data). Computational Statistics & Data Analysis, 9(1):93– 111, January 1990. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739090073Q.

Martin:1990:SRV

[11] R. J. Martin. Some results on the Φ_{p}- value when errors are correlated. Computational Statistics & Data Analysis, 9(1):113–121, January 1990. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739090074R.

Joss:1990:PAL

[12] Johann Joss and Alfio Marazzi. Probabilistic algorithms for least median of squares regression. Computational Statistics & Data Analysis, 9(1):123– 133, January 1990. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739090075S.

Kachlicka:1990:ISP

[13] Danuta Kachlicka and Stanisław Mejza. Incomplete split-plot experiments — whole-plot treatments in a row-column design. Computational Statistics & Data Analysis, 9(1):135– 146, January 1990. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739090076T.

Arbia:1990:SON

[14] Giuseppe Arbia. On second-order non-stationarity in two-dimensional lattice processes. Computational Statistics & Data Analysis, 9(1):147– 160, January 1990. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739090077U.

Anonymous:1990:SAM

[15] Anonymous. SPSS announces Macintosh package. Computational Statistics & Data Analysis, 9(1): 161, January 1990. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739090078V.

Anonymous:1990:SIAa

[16] Anonymous. SYSTAT, Inc. announces new Mac MYSTAT with brushing tools. Computational Statistics & Data Analysis, 9(1):161– 162, January 1990. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739090079W.

Anonymous:1990:SIAb

[17] Anonymous. SYSTAT, Inc. announces new data transfer package for SYSTAT software on IBMcompatible computers. Computational Statistics & Data Analysis, 9(1):162– 163, January 1990. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947390900802.

Anonymous:1990:SIAc

[18] Anonymous. SYSTAT, Inc. announces EzPATH, a new casual modeling supplement to SYSTAT software on IBMcompatible computers. Computational Statistics & Data Analysis, 9 (1):163, January 1990. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739090081R.

Anonymous:1990:IBP

[19] Anonymous. Installed base passes 50,000 “milestone”. Computational Statistics & Data Analysis, 9(1):163– 164, January 1990. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739090082S.

Anonymous:1990:MIA

[20] Anonymous. Minitab introduces advanced features for industrial statistics. Computational Statistics & Data Analysis, 9(1):165– 166, January 1990. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739090083T.

Anonymous:1990:MP

[21] Anonymous. MicrOsiris for the PC. Computational Statistics & Data Analysis, 9(1):166–167, January 1990. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/016794739090084U.

Anonymous:1990:SE

[22] Anonymous. SIR/Easy. Computational Statistics & Data Analysis, 9 (1):167, January 1990. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739090085V.

Abdous:1990:ACK

[23] B. Abdous. Adapting the classical kernel density estimator to data. Computational Statistics & Data Analysis, 9(2):169–178, March 1990. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739090117Z.

Haslett:1990:DFP

[24] Stephen Haslett. Degrees of freedom and parameter estimability in hierarchical models for sparse complete contingency tables. Computational Statistics & Data Analysis, 9 (2):179–195, March 1990. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947390901182.

Herce:1990:ESN

[25] Miguel A. Herce. An example showing that a new technique for LAV estimation breaks down in certain cases. Computational Statistics & Data Analysis, 9(2):197–202, March 1990. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947390901193. See

discussion [26].

Christensen:1990:DES

[26] G. S. Christensen, S. A. Soliman, and A. Rouhi. Discussion of “An example showing that a new technique for LAV estimation breaks down in certain cases”. Computational Statistics & Data Analysis, 9 (2):203–213, March 1990. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947390901207. See

[25, 27].

Herce:1990:SCC

[27] Miguel A. Herce. Some comments on Christensen, Soliman and Rouhi’s “Discussion … ”. Computational Statistics & Data Analysis, 9(2):215– 216, March 1990. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739090121W. See

[26].

Cologne:1990:SSS

[28] John B. Cologne and Norman E. Breslow. A small sample simulation study of methods for log odds ratio regression analysis. Computational Statistics & Data Analysis, 9 (2):217–235, March 1990. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739090122X.

Conlon:1990:NCI

[29] Michael Conlon and Ronald G. Thomas. A new confidence interval for the difference of two binomial proportions. Computational Statistics & Data Analysis, 9(2):237– 241, March 1990. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739090123Y.

Anonymous:1990:E

[30] Anonymous. Erratum. Computational Statistics & Data Analysis, 9 (2):243, March 1990. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739090124Z.

Anonymous:1990:A

[31] Anonymous. Announcements. Computational Statistics & Data Analysis, 9(2):245–247, March 1990. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947390901252.

Anonymous:1990:BRBa

[32] Anonymous. Book review: Multiway data analysis: edited by R. Coppi, University of Rome, Italy and S. Bolasco, University of Salerno, Italy, 1989 xiv + 552 pages, Price: US$136.75/Dfl. 260.00, ISBN 0-444-87410-0 (Elsevier Science Publishers, Amsterdam). Computational Statistics & Data Analysis, 9(2):248, March 1990. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947390901274.

Anonymous:1990:BRBb

[33] Anonymous. Book review: Statistical data analysis and inference: edited by Y. Dodge, University of Neuchatel, Switzerland, 1989 xii + 618 pages, Price: US$131.50/Dfl. 250.00, ISBN 0-444-88029-1 (Elsevier Science Publishers, Amsterdam). Computational Statistics & Data Analysis, 9 (2):248, March 1990. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947390901263.

Anonymous:1990:BRBc

[34] Anonymous. Book review: Data exchange: by Steven S. Ross; 411 pages; illustrated; 6” × 9”; paperback; $24.95 (McGraw-Hill). Computational Statistics & Data Analysis, 9(2):248–249, March 1990. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947390901285.

Anonymous:1990:BRBd

[35] Anonymous. Book review: Using autoCAD with autoLISP: by John D. Hood; 239 pages; illustrated; 6′ × 9′; spiral-bound paperback: $29.95 (McGraw-Hill). Computational Statistics & Data Analysis, 9(2):249–250, March 1990. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947390901296.

Greene:1990:DLA

[36] Tom Greene. The depiction of linear association by matroids. Computational Statistics & Data Analysis, 9(3):251–269, May 1990. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739090108T.

Frangos:1990:JEB

[37] Chris C. Frangos and William R. Schucany. Jackknife estimation of the bootstrap acceleration constant. Computational Statistics & Data Analysis, 9(3):271–281, May 1990. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739090109U.

Parrish:1990:GRD

[38] Rudolph S. Parrish. Generating random deviates from multivariate Pearson distributions. Computational Statistics & Data Analysis, 9(3): 283–295, May 1990. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947390901104.

Spall:1990:FOD

[39] James C. Spall and Daniel C. Chin. First-order data sensitivity measures with applications to a multivariate signal-plus-noise problem. Computational Statistics & Data Analysis, 9(3):297–307, May 1990. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739090111T.

Welsh:1990:NSP

[40] A. H. Welsh. A note on the statistical properties of the secant algorithm for calculating rank estimators. Computational Statistics & Data Analysis, 9(3):309–315, May 1990. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739090112U.

Anonymous:1990:SSS

[41] Anonymous. Software for statistical and social survey analysis 1989–1990 prepared for the study group on computers in survey analysis. Computational Statistics & Data Analysis, 9(3):317–340, May 1990. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739090113V.

Anonymous:1990:CPO

[42] Anonymous. Call for papers on optimization techniques in statistics. Computational Statistics & Data Analysis, 9(3):341, May 1990. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739090114W.

Anonymous:1990:AIVa

[43] Anonymous. Author index volume 9 (1990). Computational Statistics & Data Analysis, 9(3):343– 344, May 1990. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739090115X.

Anonymous:1990:PRRa

[44] Anonymous. Package report/review index volume 9 (1990). Computational Statistics & Data Analysis, 9 (3):345, May 1990. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739090116Y.

Anonymous:1990:EBb

[45] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 10(1):ii, August 1990. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947390900972.

Boik:1990:LRT

[46] Robert J. Boik. A likelihood ratio test for three-mode singular values: upper percentiles and an application to three-way ANOVA. Computational Statistics & Data Analysis, 10(1):1–9, August 1990. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947390900983.

Namini:1990:RFF

[47] Hamid Namini and Subir Ghosh. On robustness of 2^{m} × 3^{n} fractional factorial resolution III designs against the unavailability of data. Computational Statistics & Data Analysis, 10(1):11–16, August 1990. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947390900994.

Poon:1990:AMP

[48] Wai-Yin Poon, Sik-Yum Lee, Abdelmonem A. Afifi, and Peter M.

Bentler. Analysis of multivariate polytomous variates in several groups via the partition maximum likelihood approach. Computational Statistics & Data Analysis, 10(1): 17–27, August 1990. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739090100V.

Tusell:1990:TIT

[49] Fernando Tusell. Testing for interaction in two-way ANOVA tables with no replication. Computational Statistics & Data Analysis, 10(1): 29–45, August 1990. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739090101M. See

corrigendum [198].

Cheng:1990:SAD

[50] S. W. Cheng and F. A. Spiring. Some applications of 3-D scatterplots in data analysis. Computational Statistics & Data Analysis, 10 (1):47–61, August 1990. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739090102N.

Cheng:1990:ESA

[51] S. W. Cheng and F. A. Spiring. Extension of the W_{u} statistic with applications. Computational Statistics & Data Analysis, 10(1): 63–69, August 1990. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739090103O.

Singh:1990:CEE

[52] V. P. Singh, J. F. Cruise, and Ming Ma. A comparative evaluation of the estimators of the threeparameter lognormal distribution by Monte Carlo simulation. Computational Statistics & Data Analysis, 10 (1):71–85, August 1990. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739090104P.

Chatterjee:1990:NFE

[53] Sangit Chatterjee and Samprit Chatterjee. A note on finding extreme points in multivariate space. Computational Statistics & Data Analysis, 10(1):87–92, August 1990. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739090105Q.

Anonymous:1990:EDM

[54] Anonymous. “Experimental designs for mean and variance estimation in variance components models”: by Alessandra Giovagnoli and Paola Sebastiani, Computational Statistics & Data Analysis 8, 21–28. Computational Statistics & Data Analysis, 10 (1):93, August 1990. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739090106R. See

[1].

Anonymous:1990:BIN

[55] Anonymous. BMDP introduces new add-ons for Solo statistics package. Computational Statistics & Data Analysis, 10(1):95– 98, August 1990. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739090107S.

Azen:1990:E

[56] Stanley P. Azen. Editorial. Computational Statistics & Data Analysis, 10(2):i–ii, October 1990. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739090053K.

Liski:1990:PGC

[57] Erkki P. Liski and Tapio Nummi. Prediction in growth curve models using the EM algorithm. Computational Statistics & Data Analysis, 10 (2):99–108, October 1990. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739090054L.

Huang:1990:OED

[58] Mong-Na Lo Huang. Optimal extrapolation designs for a partly linear model. Computational Statistics & Data Analysis, 10(2):109– 115, October 1990. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739090055M.

Ahsanullah:1990:SCP

[59] M. Ahsanullah. Some characteristic properties of normal distribution. Computational Statistics & Data Analysis, 10(2):117– 120, October 1990. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739090056N.

Ceranka:1990:COB

[60] Bronisław Ceranka and Krystyna Katulska. Constructions of optimum biased spring balance weighing designs with the diagonal covariance matrix of errors. Computational Statistics & Data Analysis, 10 (2):121–131, October 1990. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739090057O.

Polasek:1990:VDL

[61] Wolfgang Polasek. Vector distributed lag models with smoothness priors. Computational Statistics & Data Analysis, 10(2):133– 141, October 1990. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739090058P.

Naidu:1990:ALE

[62] L. Kannaiah Naidu. An adjusted linear estimator. Computational Statistics & Data Analysis, 10(2):143– 151, October 1990. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739090059Q.

Blyth:1990:TPS

[63] Colin R. Blyth. Temporality in probability and statistics. Computational Statistics & Data Analysis, 10(2):153– 162, October 1990. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739090060U.

Khosrovshahi:1990:TD

[64] G. B. Khosrovshahi. On trades and designs. Computational Statistics & Data Analysis, 10(2):163– 167, October 1990. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739090061L.

Naga:1990:SRN

[65] Ramses Abul Naga and Gérard Antille. Stability of robust and non-robust principal components analysis. Computational Statistics & Data Analysis, 10 (2):169–174, October 1990. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739090062M.

Rao:1990:SSC

[66] M. Bhaskara Rao and K. Subramanyam. The structure of some classes of bivariate distributions and some applications. Computational Statistics & Data Analysis, 10(2):175– 187, October 1990. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739090063N.

Behboodian:1990:SCT

[67] Javad Behboodian. Some characterization theorems on symmetry. Computational Statistics & Data Analysis, 10 (2):189–192, October 1990. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739090064O.

Ahmed:1990:ESI

[68] S. E. Ahmed and A. K. Md. E. Saleh. Estimation strategies for the intercept vector in a simple linear multivariate normal regression model. Computational Statistics & Data Analysis, 10 (3):193–206, December 1990. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739090001X.

Ronchetti:1990:SSA

[69] Elvezio Ronchetti. Small sample asymptotics: a review with applications to robust statistics. Computational Statistics & Data Analysis, 10 (3):207–223, December 1990. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739090002Y.

Roux:1990:CPP

[70] J. J. J. Roux and A. Bekker. Characterizing priors by posterior expectations in multivariate analysis. Computational Statistics & Data Analysis, 10 (3):225–233, December 1990. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739090003Z.

Satorra:1990:MCA

[71] Albert Satorra and Peter M. Bentler. Model conditions for asymptotic robustness in the analysis of linear relations. Computational Statistics & Data Analysis, 10(3):235– 249, December 1990. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947390900042.

Yanai:1990:SGF

[72] Haruo Yanai. Some generalized forms a least squares g-inverse, minimum norm g-inverse, and Moore– Penrose inverse matrices. Computational Statistics & Data Analysis, 10 (3):251–260, December 1990. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947390900053.

Tiwari:1990:EBE

[73] Ram C. Tiwari and Jyoti N. Zalkikar. Empirical Bayes estimation of the scale parameter in a Pareto distribution. Computational Statistics & Data Analysis, 10(3):261– 270, December 1990. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947390900064.

Mexia:1990:BLU

[74] João Tiago Mexia. Best linear unbiased estimates, duality of F tests and the Scheffé multiple comparison method in the presence of controlled heteroscedasticity. Computational Statistics & Data Analysis, 10 (3):271–281, December 1990. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947390900075.

Witte:1990:CDE

[75] Hans-Jürgen Witte. Characterizations of distributions of exponential or geometric type by the integrated lack of memory property and record values. Computational Statistics & Data Analysis, 10(3):283– 288, December 1990. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947390900086.

Scholz:1990:CWD

[76] F.-W. Scholz. Characterization of the Weibull distribution. Computational Statistics & Data Analysis, 10 (3):289–292, December 1990. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947390900097.

Stark:1990:ACS

[77] Alan E. Stark. An algorithm for computing standard errors from categorical data. Computational Statistics & Data Analysis, 10(3):293– 296, December 1990. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739090010F.

Toutenburg:1990:MSE

[78] H. Toutenburg and G. Trenkler. Mean square error matrix comparisons of optimal and classical predictors and estimators in linear regression. Computational Statistics & Data Analysis, 10 (3):297–305, December 1990. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947390900116.

Turkman:1990:OAS

[79] M. A. Amaral Turkman and K. F. Turkman. Optimal alarm systems for autoregressive processes: a Bayesian approach. Computational Statistics & Data Analysis, 10(3):307– 314, December 1990. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947390900127.

Tracy:1990:BPM

[80] Derrick S. Tracy. Balanced partitioned matrices and their Kronecker products. Computational Statistics & Data Analysis, 10(3):315– 323, December 1990. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947390900138.

Schwabe:1990:LOG

[81] Rainer Schwabe. Local optimality and global superiority of adaptive designs in linear models. Computational Statistics & Data Analysis, 10 (3):325–330, December 1990. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947390900149.

Schneider:1990:SIM

[82] Christina Schneider. Structural inference and a modification of Dempster’s combination rule. Computational Statistics & Data Analysis, 10 (3):331–338, December 1990. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739090015A.

Shoukri:1990:EKE

[83] M. M. Shoukri, D. S. Tracy, and I. U. H. Mian. The effect of kurtosis in the estimation of the parameters of the one-way random effects model from familial data. Computational Statistics & Data Analysis, 10 (3):339–345, December 1990. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739090016B.

Anonymous:1990:AIVb

[84] Anonymous. Author index volume 10 (1990). Computational Statistics & Data Analysis, 10(3):347– 348, December 1990. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739090017C.

Anonymous:1990:PRRb

[85] Anonymous. Package report/review index volume 10 (1990). Computational Statistics & Data Analysis, 10 (3):349, December 1990. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739090018D.

Anonymous:1991:EBa

[86] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 11(1):i–ii, January 1991. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947391900476.

Azen:1991:E

[87] Stanley P. Azen. Editorial. Computational Statistics & Data Analysis, 11(1):1, January 1991. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947391900487.

Jones:1991:CVI

[88] M. C. Jones. On correcting for variance inflation in kernel density estimation. Computational Statistics & Data Analysis, 11(1): 3–15, January 1991. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947391900498.

Rayens:1991:CPS

[89] William Rayens and Tom Greene. Covariance pooling and stabilization for classification. Computational Statistics & Data Analysis, 11(1): 17–42, January 1991. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739190050C.

Cunningham:1991:TSS

[90] J. K. Cunningham, R. L. Eubank, and T. Hsing. M-type smoothing splines with auxiliary scale estimation. Computational Statistics & Data Analysis, 11(1):43–51, January 1991. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947391900513.

Moulton:1991:BGL

[91] Lawrence H. Moulton and Scott L. Zeger. Bootstrapping generalized linear models. Computational Statistics & Data Analysis, 11(1):53– 63, January 1991. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947391900524.

Bunge:1991:SEN

[92] John A. Bunge and John C. Handley. Sampling to estimate the number of duplicates in a database. Computational Statistics & Data Analysis, 11(1):65–74, January 1991. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947391900535.

Harwell:1991:URT

[93] Michael R. Harwell. Using randomization tests when errors are unequally correlated. Computational Statistics & Data Analysis, 11(1): 75–85, January 1991. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947391900546.

Shapiro:1991:VFG

[94] A. Shapiro and J. D. Botha. Variogram fitting with a general class of conditionally nonnegative definite functions. Computational Statistics & Data Analysis, 11(1):87– 96, January 1991. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947391900557.

Soliman:1991:NAN

[95] S. A. Soliman, G. S. Christensen, and A. H. Rouhi. A new algorithm for nonlinear L_{1}-norm minimization with nonlinear equality constraints. Computational Statistics & Data Analysis, 11(1):97–109, January 1991. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947391900568.

Boese:1991:ETB

[96] F. G. Boese. Estimating and testing with the Bhattacharya plot. Computational Statistics & Data Analysis, 11 (1):111–124, January 1991. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947391900579.

Anonymous:1991:SSNa

[97] Anonymous. Statistical software newsletter. Computational Statistics & Data Analysis, 11(1):125, January 1991. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/016794739190058A.

Rousseeuw:1991:DPR

[98] Peter J. Rousseeuw. A diagnostic plot for regression outliers and leverage points. Computational Statistics & Data Analysis, 11(1):127– 129, January 1991. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739190059B.

Anonymous:1991:NASa

[99] Anonymous. News about statistical software & related systems. Computational Statistics & Data Analysis, 11 (1):130–131, January 1991. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739190060F.

Anonymous:1991:Ia

[100] Anonymous. Information. Computational Statistics & Data Analysis, 11(1):131, January 1991. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947391900616.

Anonymous:1991:INa

[101] Anonymous. IASC news. Computational Statistics & Data Analysis, 11(1):132, January 1991. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947391900627.

Anonymous:1991:CMa

[102] Anonymous. Calendar of meetings. Computational Statistics & Data Analysis, 11(1):133–139, January 1991. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/0167947391900638.

Anonymous:1991:CW

[103] Anonymous. Courses & workshops. Computational Statistics & Data Analysis, 11(1):140, January 1991. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947391900649.

Anonymous:1991:Aa

[104] Anonymous. Acknowledgement. Computational Statistics & Data Analysis, 11(1):141–142, January 1991. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739190065A.

Romanazzi:1991:ICV

[105] Mario Romanazzi. Influence in canonical variates analysis. Computational Statistics & Data Analysis, 11(2): 143–164, March 1991. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739190066B.

Soofi:1991:MDI

[106] Ehsan S. Soofi and D. V. Gokhale. Minimum discrimination information estimator of the mean with known coefficient of variation. Computational Statistics & Data Analysis, 11 (2):165–177, March 1991. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739190067C.

Wang:1991:EED

[107] Song-Gui Wang and Hans Nyquist. Effects on the eigenstructure of a data matrix when deleting an observation. Computational Statistics & Data Analysis, 11(2):179– 188, March 1991. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739190068D.

Schall:1991:DNN

[108] Robert Schall and René Gonin. Diagnostics for nonlinear L_{p}-norm estimation. Computational Statistics & Data Analysis, 11(2):189– 198, March 1991. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739190069E.

Guirguis:1991:RAI

[109] George H. Guirguis. A rational approximation of the inverse normal probability function. Computational Statistics & Data Analysis, 11(2): 199–201, March 1991. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739190070I.

Menendez:1991:AIM

[110] J. A. Menéndez and B. Salvador. An algorithm for isotonic median regression. Computational Statistics & Data Analysis, 11(2):203– 204, March 1991. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947391900719.

Stromberg:1991:AIR

[111] Ulf Strömberg. An algorithm for isotonic regression with arbitrary convex distance function. Computational Statistics & Data Analysis, 11 (2):205–219, March 1991. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739190072A.

Ackermann:1991:BPP

[112] Hanns Ackermann. BIAS. a program package for biometrical analysis of samples. Computational Statistics & Data Analysis, 11(2):223– 224, March 1991. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739190073B.

Koch:1991:SPS

[113] Armin Koch. Software presentations at statistical computing ’90: Reisensburg, Germany, June 16–19, 1990. Computational Statistics & Data Analysis, 11(2):225–227, March 1991. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739190074C.

Anonymous:1991:SSB

[114] Anonymous. Statistical software: BMDP introduces PC-90. Computational Statistics & Data Analysis, 11 (2):227–230, March 1991. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739190075D.

Anonymous:1991:BRB

[115] Anonymous. Book review: Distributionfree statistics: an application-oriented approach: Krauth, J.: Elsevier Science Publishers B.V., Amsterdam (1988), ISBN 0-444-80934-1 (hardback), 0-444-80988-0 (paper-back, £21.-). Computational Statistics & Data Analysis, 11(2):230, March 1991. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739190077F.

Trenkler:1991:BRB

[116] Götz Trenkler. Book review: Regression estimators a comparative study: Gruber, Marvin H. J.: Academic Press, Inc., Harcourt Brace Jovanovich Publishers, Boston (1990), ISBN 0-12-304752-8, 347 pp, $49.95. Computational Statistics & Data Analysis, 11(2):230, March 1991. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739190076E.

Pabst:1991:DMD

[117] G. Pabst. Disfree (manual and 2 diskettes, DOS 2.0 or later).: Krauth, j.: Blosoft, Cambridge (1989), £250.-. Computational Statistics & Data Analysis, 11(2):230– 231, March 1991. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739190078G.

Anonymous:1991:SAE

[118] Anonymous. Society for the advancement of economic theory. Computational Statistics & Data Analysis, 11 (2):231, March 1991. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739190079H.

Anonymous:1991:CMb

[119] Anonymous. Calendar of meetings. Computational Statistics & Data Analysis, 11(2):232–238, March 1991. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/016794739190080L.

Anonymous:1991:CWSa

[120] Anonymous. Courses, workshops & summer schools. Computational Statistics & Data Analysis, 11(2): 238, March 1991. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739190081C.

Buonaccorsi:1991:MEL

[121] John Buonaccorsi. Measurement errors, linear calibration and inferences for means. Computational Statistics & Data Analysis, 11(3): 239–257, May 1991. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739190083E.

Farebrother:1991:GEP

[122] R. W. Farebrother. A general estimation procedure based on group means. Computational Statistics & Data Analysis, 11(3):259–264, May 1991. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/016794739190084F.

Pfaffenberger:1991:TNR

[123] Roger C. Pfaffenberger and Terry E. Dielman. Testing normality of regression disturbances: a Monte Carlo study of the Filleben test. Computational Statistics & Data Analysis, 11(3):265–273, May 1991. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739190085G.

Tutz:1991:SMC

[124] Gerhard Tutz. Sequential models in categorical regression. Computational Statistics & Data Analysis, 11(3):275–295, May 1991. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739190086H.

VanZyl:1991:DAI

[125] Johan M. Van Zyl. Distributional aspects if a zero correlation is assumed in a paired comparison of means. Computational Statistics & Data Analysis, 11(3):297–307, May 1991. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739190087I.

Hirose:1991:PPE

[126] Hideo Hirose. Percentile point estimation in the three parameter Weibull distribution by the extended maximum likelihood estimate. Computational Statistics & Data Analysis, 11(3):309–331, May 1991. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739190088J.

Moneta:1991:ICM

[127] Giovanni B. Moneta. Implicit construction of McCulloch’s G matrix for the numerical evaluation of Fisher information matrixes. Computational Statistics & Data Analysis, 11(3): 333–344, May 1991. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739190089K.

Mukerjee:1991:QDB

[128] Rahul Mukerjee and Sudhir Gupta. Q-designs for bioassays. Computational Statistics & Data Analysis, 11 (3):345–350, May 1991. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739190090O.

Anonymous:1991:NASb

[129] Anonymous. News about statistical software & related systems. Computational Statistics & Data Analysis, 11(3):353–355, May 1991. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739190091F.

Anonymous:1991:Ib

[130] Anonymous. Information. Computational Statistics & Data Analysis, 11(3):355–356, May 1991. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739190092G.

Bismarck:1991:BRBa

[131] M. Bismarck. Book review: Multiway data analysis: Coppi, R., Bolasco, S. (eds.) (1989): Amsterdam, New York, Oxford, Tokyo: North-Holland. 552 pp. ISBN 0-444-87410-0. Computational Statistics & Data Analysis, 11(3):356–357, May 1991. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739190093H.

Anonymous:1991:INb

[132] Anonymous. IASC news. Computational Statistics & Data Analysis, 11(3):357, May 1991. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739190095J.

Bismarck:1991:BRBb

[133] M. Bismarck. Book review: Medical expert systems using personal computers: Chytil, M. K. Engelbrecht, R. (eds.) (1987): Wilmslow: Sigma Press. 206 pp, ISBN 1-85058-047-2. Computational Statistics & Data Analysis, 11(3):357, May 1991. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739190094I.

Anonymous:1991:CMc

[134] Anonymous. Calendar of meetings. Computational Statistics & Data Analysis, 11(3):358–364, May 1991. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/016794739190096K.

Anonymous:1991:EBb

[135] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 12(1):v–vi, August 1991. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739190097L.

Dodge:1991:P

[136] Yadolah Dodge. Preface. Computational Statistics & Data Analysis, 12 (1):VII–VIII, August 1991. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739190098M.

Keating:1991:PNS

[137] Jerome P. Keating, Robert L. Mason, P. K. Sen, and C. Radhakrishna Rao. Pitman nearness in statistical estimation: a panel discussion on recent developments. Computational Statistics & Data Analysis, 12(1):1–17, August 1991. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739190099N.

Rustagi:1991:TJK

[138] Jagdish S. Rustagi, Walfredo R. Javier, and Jose S. Victoria. Trimmed jackknife kernel estimate for the probability density function. Computational Statistics & Data Analysis, 12 (1):19–26, August 1991. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739190100G.

Laird:1991:SNP

[139] Nan M. Laird and Thomas A. Louis. Smoothing the non-parametric estimate of a prior distribution by roughening: a computational study. Computational Statistics & Data Analysis, 12(1):27–37, August 1991. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947391901017.

Kipnis:1991:EIE

[140] Victor Kipnis. Evaluating the impact of exploratory procedures in regression prediction: a pseudosample approach. Computational Statistics & Data Analysis, 12(1):39– 55, August 1991. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947391901028.

Ciampi:1991:GRT

[141] Antonio Ciampi. Generalized regression trees. Computational Statistics & Data Analysis, 12(1):57– 78, August 1991. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947391901039.

Stangenhaus:1991:IPR

[142] Gabriela Stangenhaus and Subhash C. Narula. Inference procedures for the L_{1} regression. Computational Statistics & Data Analysis, 12(1): 79–85, August 1991. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739190104A.

Dodge:1991:CAA

[143] Yadolah Dodge, Jaromir Antoch, and Jana Jurečková. Computational aspects of adaptive combination of least squares and least absolute deviations estimators. Computational Statistics & Data Analysis, 12(1): 87–99, August 1991. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739190105B.

Drygas:1991:LME

[144] Hilmar Drygas. Linear minimaxestimation in linear models with affine and ellipsoidal restrictions. Computational Statistics & Data Analysis, 12 (1):101–113, August 1991. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739190106C.

Shanbhag:1991:EVD

[145] D. N. Shanbhag. Extended versions of Deny’s theorem via de Finetti’s theorem. Computational Statistics & Data Analysis, 12(1):115– 126, August 1991. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739190107D.

Dohmann:1991:RNG

[146] Birgit Dohmann, Michael Falk, and Karin Lessenich. The random number generators of the turbo Pascal family. Computational Statistics & Data Analysis, 12(1):129– 132, August 1991. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739190108E.

Anonymous:1991:NASc

[147] Anonymous. News about statistical software & related systems. Computational Statistics & Data Analysis, 12 (1):133–135, August 1991. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739190109F.

Anonymous:1991:CMd

[148] Anonymous. Calendar of meetings. Computational Statistics & Data Analysis, 12(1):136–141, August 1991.

CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/016794739190110N.

Barndorff-Nielsen:1991:IQN

[149] O. E. Barndorff-Nielsen and M. Sørensen. Information quantities in non-classical settings. Computational Statistics & Data Analysis, 12(2):143– 158, September 1991. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739190014S.

Chernoff:1991:SCI

[150] Herman Chernoff. Simple computer intensive methods for estimating parameters of complex models. Computational Statistics & Data Analysis, 12 (2):159–178, September 1991. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739190015T.

Fraser:1991:COL

[151] D. A. S. Fraser and N. Reid. Converting observed likelihood functions to tail probabilities. Computational Statistics & Data Analysis, 12(2):179– 185, September 1991. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739190016U.

Grafarend:1991:PTP

[152] Erik W. Grafarend and Burkhard Schaffrin. The planar trisection problem and the impact of curvature on non-linear least-squares estimation. Computational Statistics & Data Analysis, 12(2):187– 199, September 1991. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739190017V.

Davies:1991:SMI

[153] Laurie Davies. A stochastic model for interlaboratory tests. Computational Statistics & Data Analysis, 12 (2):201–209, September 1991. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739190018W.

Dodge:1991:FEL

[154] Yadolah Dodge and Jana Jurečková. Flexible L-estimation in the linear model. Computational Statistics & Data Analysis, 12(2):211– 220, September 1991. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739190019X.

Baksalary:1991:CAB

[155] Jerzy K. Baksalary and Götz Trenkler. Covariance adjustment in biased estimation. Computational Statistics & Data Analysis, 12(2):221– 230, September 1991. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947391900203.

Chopra:1991:FIB

[156] D. V. Chopra. Further investigations on balanced arrays. Computational Statistics & Data Analysis, 12 (2):231–237, September 1991. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739190021S.

Abrahamowicz:1991:ITC

[157] Michał Abrahamowicz and Antonio Ciampi. Information theoretic criteria in non-parametric density estimation: Bias and variance in the infinite dimensional case. Computational Statistics & Data Analysis, 12 (2):239–247, September 1991. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739190022T.

Zelen:1991:MRM

[158] Marvin Zelen. Multinomial response models. Computational Statistics & Data Analysis, 12(2):249– 254, September 1991. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739190023U.

Anonymous:1991:SSNb

[159] Anonymous. Statistical software newsletter. Computational Statistics & Data Analysis, 12(2):255, September 1991. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739190024V.

Nelder:1991:DMM

[160] JohnA. Nelder. Data matrices and multi-way tables. Computational Statistics & Data Analysis, 12(2):257– 259, September 1991. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739190025W.

Anonymous:1991:NASd

[161] Anonymous. News about statistical software & related systems. Computational Statistics & Data Analysis, 12 (2):260–262, September 1991. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739190026X.

Anonymous:1991:Ic

[162] Anonymous. Information. Computational Statistics & Data Analysis, 12 (2):262–263, September 1991. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739190027Y.

Anonymous:1991:INc

[163] Anonymous. IASC news. Computational Statistics & Data Analysis, 12 (2):263–265, September 1991. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739190028Z.

Anonymous:1991:CMe

[164] Anonymous. Calendar of meetings. Computational Statistics & Data Analysis, 12(2):266–269, September 1991. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/0167947391900292.

Anonymous:1991:CWSb

[165] Anonymous. Courses, workshops & summer schools. Computational Statistics & Data Analysis, 12(2): 270, September 1991. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947391900306.

Dey:1991:NEB

[166] Dipak K. Dey and Lynn Kuo. A new empirical Bayes estimator with type II censored data. Computational Statistics & Data Analysis, 12 (3):271–279, November 1991. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739190111E.

Liu:1991:REC

[167] Shu-Ing Liu. A robust and easy confidence procedure with moderately longtailed symmetric distributions. Computational Statistics & Data Analysis, 12 (3):281–294, November 1991. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739190112F.

Loh:1991:SMT

[168] Wei-Yin Loh. Survival modeling through recursive stratification. Computational Statistics & Data Analysis, 12(3):295–313, November 1991. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/016794739190113G.

Olbricht:1991:MDD

[169] W. Olbricht. On mergers of distributions and distributions with exponential tails. Computational Statistics & Data Analysis, 12(3):315– 326, November 1991. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739190114H.

Kooperberg:1991:SLD

[170] Charles Kooperberg and Charles J. Stone. A study of logspline density estimation. Computational Statistics & Data Analysis, 12(3):327– 347, November 1991. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739190115I.

Spino:1991:ECP

[171] Cathie Spino and Marcello Pagano. Efficient calculation of the permutation distribution of robust twosample statistics. Computational Statistics & Data Analysis, 12(3):349– 365, November 1991. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739190116J.

Dodge:1991:WSF

[172] Yadolah Dodge and David J. Hand. What should future statistical software look like? Computational Statistics & Data Analysis, 12(3):369– 372, November 1991. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739190117K.

Anonymous:1991:Id

[173] Anonymous. Information. Computational Statistics & Data Analysis, 12 (3):373–374, November 1991. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739190118L.

Anonymous:1991:INd

[174] Anonymous. IASC news. Computational Statistics & Data Analysis, 12 (3):374–378, November 1991. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739190119M.

Anonymous:1991:CMf

[175] Anonymous. Calendar of meetings. Computational Statistics & Data Analysis, 12(3):379–383, November 1991. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/016794739190120Q.

Anonymous:1991:CWSc

[176] Anonymous. Courses, workshops & summer schools. Computational Statistics & Data Analysis, 12(3):383– 385, November 1991. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739190121H.

Anonymous:1991:Ab

[177] Anonymous. Acknowledgement. Computational Statistics & Data Analysis, 12(3):385, November 1991. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739190122I.

Anonymous:1991:AIV

[178] Anonymous. Author index volume 12 (1991). Computational Statistics & Data Analysis, 12(3):387– 388, November 1991. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739190123J.

Anonymous:1992:EBa

[179] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 13(1):ii, January 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947392901489.

Larson:1992:LSE

[180] Harold J. Larson. Least squares estimation of linear splines with unknown knot locations. Computational Statistics & Data Analysis, 13 (1):1–8, January 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290149A.

Rocke:1992:EVA

[181] David M. Rocke. Estimation of variation after outlier rejection. Computational Statistics & Data Analysis, 13(1):9–20, January 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290150E.

Shi:1992:AED

[182] Sheng G. Shi. Accurate and efficient double-bootstrap confidence limit method. Computational Statistics & Data Analysis, 13(1):21– 32, January 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947392901515.

Taylor:1992:PME

[183] Jeremy M. G. Taylor. Properties of modelling the error distribution with an extra shape parameter. Computational Statistics & Data Analysis, 13(1):33–46, January 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947392901526.

Ferri:1992:OSS

[184] M. Ferri and M. Piccioni. Optimal selection of statistical units: an approach via simulated annealing. Computational Statistics & Data Analysis, 13(1):47–61, January 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947392901537.

Goldstein:1992:ECP

[185] Harvey Goldstein and Jon Rasbash. Efficient computational procedures for the estimation of parameters in multilevel models based on iterative generalised least squares. Computational Statistics & Data Analysis, 13 (1):63–71, January 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947392901548.

Saltelli:1992:SAM

[186] A. Saltelli and T. Homma. Sensitivity analysis for model output: Performance of black box techniques on three international benchmark exercises. Computational Statistics & Data Analysis, 13(1):73– 94, January 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947392901559.

Singh:1992:CNA

[187] Karan P. Singh, Karan P. Singh, and George E. Relyea. Computation of noncentral Æ’ probabilities: a computer program. Computational Statistics & Data Analysis, 13(1):95– 102, January 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290156A.

Edler:1992:IMA

[188] Lutz Edler. IASC meeting activities during the 48th session of the ISI in Cairo. Computational Statistics & Data Analysis, 13(1):105– 106, January 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290157B.

Anonymous:1992:NASa

[189] Anonymous. News about statistical software & related systems. Computational Statistics & Data Analysis, 13 (1):107–108, January 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290158C.

Park:1992:BRB

[190] Sung H. Park. Book review: Statistical methods in engineering and quality assurance: John, Peter M. W. (1990): Wiley-Interscience Publication, John Wiley & Sons, Inc. 313 pp., ISBN 0-471-82986-2, 39.80 £. Computational Statistics & Data Analysis, 13(1):108, January 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290160H.

Victor:1992:BRB

[191] N. Victor. Book review: Theory and application in computational statistics: Asano, Ch., Niki, N. and Okazaki, T. (eds.): Scientist Inc., Tokyo (1991), ISBN 4-914903-03-2, 290 pp, Yen 5000 (to be sold only in Japan). Computational Statistics & Data Analysis, 13(1):108, January 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290159D.

Trenkler:1992:BRBa

[192] G. Trenkler. Book review: Angewandte lineare algebra (applied linear algebra): Huppert, B.: De Gruyter, Berlin (1990), ISBN 3-11-012107-7, 646 pp, DM 198. Computational Statistics & Data Analysis, 13(1):108– 109, January 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947392901618.

Toutenburg:1992:BRB

[193] H. Toutenburg. Book review: Mathematical statistics with applications, 4th edition: Mendenhall, W., Wackerly, D. D., Schaeffer, R. L., PWS-KENT Publishing Company, Boston (1990),752 pp,ISBN 0-534-92026-8. Computational Statistics & Data Analysis, 13 (1):109, January 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947392901629.

Trenkler:1992:BRBb

[194] G. Trenkler. Book review: Regression analysis, theory, methods and applications: Sen, A. and Srivastava, M. Springer, Berlin (1990), 347 pp, ISBN 3-540-97211-0, DM 88. Computational Statistics & Data Analysis, 13 (1):109–110, January 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290163A.

Anonymous:1992:INb

[195] Anonymous. IASC news. Computational Statistics & Data Analysis, 13 (1):110–112, January 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290164B.

Anonymous:1992:CMa

[196] Anonymous. Calendar of meetings. Computational Statistics & Data Analysis, 13(1):113–118, January 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/016794739290165C.

Anonymous:1992:CWSa

[197] Anonymous. Courses, workshops & summer schools. Computational Statistics & Data Analysis, 13(1):118– 119, January 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290166D.

Tussell:1992:CTI

[198] F. Tussell. Corrigendum: Testing for interaction in two-way ANOVA tables with no replication. Computational Statistics & Data Analysis, 13(1):121, January 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290167E. See

[49].

Anonymous:1992:SIC

[199] Anonymous. Second international conference on statistical data analysis based on the L_{1}-norm and related methods: August 17–20, 1992, Neuchâtel, Switzerland. Computational Statistics & Data Analysis, 13 (1):122, January 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739290168F.

Balakrishnan:1992:ESH

[200] N. Balakrishnan and P. S. Chan. Estimation for the scaled half logistic distribution under type II censoring. Computational Statistics & Data Analysis, 13(2):123–141, March 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739290001V.

Feluch:1992:NMC

[201] W. Feluch and J. Koronacki. A note on modified cross-validation in density estimation. Computational Statistics & Data Analysis, 13(2): 143–151, March 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739290002W.

Hayter:1992:MCT

[202] Anthony J. Hayter. Multiple comparisons of three ordered normal means for unbalanced models. Computational Statistics & Data Analysis, 13 (2):153–162, March 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739290003X.

Wilcox:1992:IMC

[203] Rand R. Wilcox. An improved method for comparing variances when distributions have non-identical shapes. Computational Statistics & Data Analysis, 13(2):163–172, March 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290004Y.

Wypij:1992:PMA

[204] David Wypij and Thomas J. Santner. Pseudotable methods for the analysis of 2 × 2 tables. Computational Statistics & Data Analysis, 13 (2):173–190, March 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290005Z.

Angelidou:1992:CPM

[205] A. Angelidou, M. G. Strintzis, and S. Panas. Correction of periodogram method for Power Spectrum estimation. Computational Statistics & Data Analysis, 13(2):191– 195, March 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947392900062.

Wang:1992:NMA

[206] Morgan Wang and W. J. Kennedy. A numerical method for accurately approximating multivariate normal probabilities. Computational Statistics & Data Analysis, 13(2):197– 210, March 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947392900073.

Wiens:1992:NCR

[207] Douglas P. Wiens. A note on the computation of robust, bounded influence estimates and test statistics in regression. Computational Statistics & Data Analysis, 13(2): 211–220, March 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947392900084.

Anonymous:1992:SSNa

[208] Anonymous. Statistical software newsletter. Computational Statistics & Data Analysis, 13(2):221, March 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/0167947392900095.

Popping:1992:MAS

[209] Roel Popping. More agreeable software? Computational Statistics & Data Analysis, 13(2):223– 227, March 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739290010D.

Victor:1992:PS

[210] Norbert Victor. President’s statement. Computational Statistics & Data Analysis, 13(2):228–229, March 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/0167947392900114.

Anonymous:1992:NASb

[211] Anonymous. News about statistical software & related systems. Computational Statistics & Data Analysis, 13 (2):230–231, March 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947392900125.

Anonymous:1992:INA

[212] Anonymous. Information: National activities in the area of statistical computing: Germany. Computational Statistics & Data Analysis, 13 (2):231, March 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947392900136.

Drygas:1992:BRB

[213] Hilmar Drygas. Book review: Theory of random determinants: Girko, V. L. Kluwer Academic Publishers, Doordrecht / Boston / London (1990), ISBN 0-7923-0233-8, xiv + 677 pp., 147 £S. Computational Statistics & Data Analysis, 13(2): 231–232, March 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947392900147.

Anonymous:1992:GV

[214] Anonymous. Graphics and visualization. Computational Statistics & Data Analysis, 13(2):232, March 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/0167947392900158.

Anonymous:1992:VS

[215] Anonymous. VI SSDBM. Computational Statistics & Data Analysis, 13(2):232, March 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947392900169.

Anonymous:1992:C

[216] Anonymous. COMPSTAT 1992. Computational Statistics & Data Analysis, 13(2):232–233, March 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739290017A.

Anonymous:1992:CMb

[217] Anonymous. Calendar of meetings. Computational Statistics & Data Analysis, 13(2):234–240, March 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/016794739290018B.

Anonymous:1992:RC

[218] Anonymous. Referees for CSDA-1991. Computational Statistics & Data Analysis, 13(2):241, March 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739290019C.

Gilchrist:1992:P

[219] Robert Gilchrist. Preface. Computational Statistics & Data Analysis, 13(3):243–244, April 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739290132Y.

Aitkin:1992:MCC

[220] Murray Aitkin. Model choice in contingency table analysis using the posterior Bayes factor. Computational Statistics & Data Analysis, 13 (3):245–251, April 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290133Z.

Becker:1992:EAA

[221] Mark P. Becker. Exploratory analysis of association models using loglinear models and singular value decompositions. Computational Statistics & Data Analysis, 13(3):253– 267, April 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947392901342.

Caussinus:1992:CPM

[222] Henri Caussinus and Louis Ferré. Comparing the parameters of a model for several units by means of principal component analysis. Computational Statistics & Data Analysis, 13(3):269–280, April 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947392901353.

Edwards:1992:LAU

[223] David Edwards. Linkage analysis using loglinear models. Computational Statistics & Data Analysis, 13 (3):281–290, April 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947392901364.

Foulley:1992:MLB

[224] J. L. Foulley, M. San Cristobal, D. Gianola, and S. Im. Marginal likelihood and Bayesian approaches to the analysis of heterogeneous residual variances in mixed linear Gaussian models. Computational Statistics & Data Analysis, 13(3):291–305, April 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947392901375.

Ivaldi:1992:EEL

[225] Marc Ivaldi. Estimation of errorsin-latent-variable models on business survey data. Computational Statistics & Data Analysis, 13(3): 307–318, April 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947392901386.

Jansen:1992:SAT

[226] J. Jansen. Statistical analysis of threshold data from experiments with nested errors. Computational Statistics & Data Analysis, 13(3):319– 330, April 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947392901397.

Lawson:1992:GNC

[227] A. B. Lawson. GLIM and normalising constant models in spatial and directional data analysis. Computational Statistics & Data Analysis, 13 (3):331–348, April 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290140B.

Holzer:1992:MCP

[228] Christian Holzer and Manfred Precht. Multiple comparison procedures for normally distributed ANOVA models in SAS, SPSS, BMDP, and MINITAB. Computational Statistics & Data Analysis, 13(3):351– 358, April 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947392901412.

Nys:1992:ISS

[229] Marleen Nys. An impression of the summer school on ‘Computational aspects of model choice’. Computational Statistics & Data Analysis, 13 (3):359–361, April 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947392901423.

Anonymous:1992:NASc

[230] Anonymous. News about statistical software & related systems. Computational Statistics & Data Analysis, 13(3):362–363, April 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947392901434.

Lenz:1992:BRB

[231] H.-J. Lenz. Book review: Extreme value theory: J. Hüsler, R.-D. Reiss (eds.): Proceedings of a conference held in Oberwolfach Dec. 6–12, 1987.

Springer, Berlin etc. (1989). Computational Statistics & Data Analysis, 13(3):363–364, April 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947392901445.

Anonymous:1992:BCS

[232] Anonymous. Bulgarian computing statistics section founded. Computational Statistics & Data Analysis, 13(3):364–365, April 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947392901456.

Anonymous:1992:CMc

[233] Anonymous. Calendar of meetings. Computational Statistics & Data Analysis, 13(3):366–370, April 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/0167947392901467.

Anonymous:1992:CWSb

[234] Anonymous. Courses, workshops & summer schools. Computational Statistics & Data Analysis, 13(3): 370–371, April 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947392901478.

Lindsey:1992:FCP

[235] J. K. Lindsey and G. Mersch. Fitting and comparing probability distributions with log linear models. Computational Statistics & Data Analysis, 13(4):373–384, May 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290112S.

Marx:1992:CPC

[236] Brian D. Marx. A continuum of principal component generalized linear regressions. Computational Statistics & Data Analysis, 13(4): 385–393, May 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290113T.

C:1992:PLQ

[237] E. A. Molina C. and C. J. Skinner. Pseudo-likelihood and quasilikelihood estimation for complex sampling schemes. Computational Statistics & Data Analysis, 13(4): 395–405, May 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290114U.

Nagel:1992:DSD

[238] Herbert Nagel and Reinhold Hatzinger. Diagnostics in some discrete choice models. Computational Statistics & Data Analysis, 13(4):407– 414, May 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290115V.

Read:1992:CSM

[239] K. L. Q. Read, C. J. Lawrence, A. Jorio, and J. C. Mariana. A case study in the modelling of growth. Computational Statistics & Data Analysis, 13(4):415–429, May 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290116W.

Rios:1992:RDB

[240] Martin Rios, Angel Villarroya, and Josep M. Oller. Rao distance between multivariate linear normal models and their application to the classification of response curves. Computational Statistics & Data Analysis, 13(4):431–445, May 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290117X.

Schnatter:1992:IBK

[241] Sylvia Schnatter. Integration-based Kalman-filtering for a dynamic generalized linear trend model. Computational Statistics & Data Analysis, 13(4):447–459, May 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290118Y.

Stasinopoulos:1992:DBP

[242] D. M. Stasinopoulos and R. A. Rigby. Detecting break points in generalised linear models. Computational Statistics & Data Analysis, 13(4):461–471, May 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290119Z.

Watkins:1992:MSC

[243] A. J. Watkins. On models of spatial covariance. Computational Statistics & Data Analysis, 13(4): 473–481, May 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947392901205.

Anonymous:1992:SSNb

[244] Anonymous. Statistical software newsletter. Computational Statistics & Data Analysis, 13(4):483, May 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/016794739290121U.

Cierzynski:1992:SEC

[245] Thomas C. Cierzynski and Stefan v. Weber. Simulation experiments with CWA, a new stable regression algorithm, and comparisons with two other stable regression algorithms. Computational Statistics & Data Analysis, 13(4):485–491, May 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290122V.

Anonymous:1992:ASSa

[246] Anonymous. Announcing the statistical software guide 1992. Computational Statistics & Data Analysis, 13(4):492, May 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290123W.

Anonymous:1992:NASd

[247] Anonymous. News about statistical software & related systems. Computational Statistics & Data Analysis, 13(4):493–494, May 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290124X.

Nagel:1992:BRB

[248] Matthias Nagel. Book review: Statistical modelling in GLIM: M. Aitkin, D. Anderson, B. Francis and J. Hinde (1989) Oxford Statistical Science Series 4. Oxford: Clarendon Press. 374 pp, ISBN 0-19-852204-5, ca. DM 52,-(Paperback). Computational Statistics & Data Analysis, 13(4):494, May 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290125Y.

Kosfeld:1992:BRB

[249] Reinhold Kosfeld. Book review: SPSS/PC+ Basics and graphic: Brosius, G. (1988) Einführung und praktische Beispiele. Hamburg: McGraw-Hill Book Company. 582 pp, ISBN 3-89028-132-X, ca. DM 68,-. Computational Statistics & Data Analysis, 13(4):494–495, May 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290126Z.

Anonymous:1992:I

[250] Anonymous. Information. Computational Statistics & Data Analysis, 13(4):495, May 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947392901272.

Anonymous:1992:ISS

[251] Anonymous. International symposium on “Statistical Computing for Agriculture in Developing Countries”. Computational Statistics & Data Analysis, 13(4):496, May 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947392901283.

Anonymous:1992:CMd

[252] Anonymous. Calendar of meetings. Computational Statistics & Data Analysis, 13(4):497–499, May 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/0167947392901294.

Anonymous:1992:CWSc

[253] Anonymous. Courses workshops & summer schools. Computational Statistics & Data Analysis, 13(4): 499–500, May 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947392901308.

Anonymous:1992:EBb

[254] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 14(1):i–ii, June 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290077S.

Hadi:1992:NMO

[255] Ali S. Hadi. A new measure of overall potential influence in linear regression. Computational Statistics & Data Analysis, 14(1):1–27, June 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290078T.

Kazempour:1992:CII

[256] M. K. Kazempour and Franklin A. Graybill. Confidence intervals on individual variances in two-way models. Computational Statistics & Data Analysis, 14(1):29–37, June 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290079U.

Zelterman:1992:IMS

[257] Daniel Zelterman and Ted I. K. Youn. Indicator models in social mobility tables. Computational Statistics & Data Analysis, 14(1): 39–53, June 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290080Y.

Agresti:1992:CMD

[258] Alan Agresti, Stuart Lipsitz, and Joseph B. Lang. Comparing marginal distributions of large, sparse contingency tables. Computational Statistics & Data Analysis, 14(1): 55–73, June 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290081P.

Antille:1992:RVC

[259] Gérard Antille and Gilbert Ritschard. Robust versus classical detection of atypical data. Computational Statistics & Data Analysis, 14(1): 75–88, June 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290082Q.

Aitkin:1992:FML

[260] Murray Aitkin and Brian Francis. Fitting the multinomial logit model with continuous covariates in GLIM. Computational Statistics & Data Analysis, 14(1):89–97, June 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290083R.

Kheoh:1992:CTM

[261] Thian S. Kheoh and A. Ian McLeod. Comparison of two modified portmanteau tests for model adequacy. Computational Statistics & Data Analysis, 14(1):99–106, June 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290084S.

Rousseeuw:1992:CSQ

[262] Peter J. Rousseeuw and Bert C. van Zomeren. A comparison of some quick algorithms for robust regression. Computational Statistics & Data Analysis, 14(1):107–116, June 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290085T.

Wittkowski:1992:SAU

[263] Knut M. Wittkowski. Statistical analysis of unbalanced and incomplete designs — experiences with BMDP and SAS. Computational Statistics & Data Analysis, 14(1): 119–124, June 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739290086U.

Rahlfs:1992:SPS

[264] Volker W. Rahlfs. A statistical package with special reference to nonparametric methods. Computational Statistics & Data Analysis, 14(1): 125–126, June 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739290087V.

Anonymous:1992:NASe

[265] Anonymous. News about statistical software & related systems. Computational Statistics & Data Analysis, 14(1):127–128, June 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739290088W.

Anonymous:1992:BRBa

[266] Anonymous. Book review: Softstat ’89 — Fortschritte in der Statistik-Software 2: F. Faulbaum, R. Haux and K.-H. Joeckel (1990): Stuttgart: Gustav Fischer Verlag. 595 pp, ISBN 3-437-50320-0, DM 76.-. Computational Statistics & Data Analysis, 14 (1):128, June 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290089X.

Anonymous:1992:CCB

[267] Anonymous. Call for candidates for the board of directors (BoD) of the European regional section (ERS). Computational Statistics & Data Analysis, 14(1):128–129, June 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947392900903.

Anonymous:1992:CA

[268] Anonymous. Call for attendance. Computational Statistics & Data Analysis, 14(1):129, June 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290092T.

Anonymous:1992:IRI

[269] Anonymous. IASC representatives in ISI. Computational Statistics & Data Analysis, 14(1):129, June 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/016794739290091S.

Anonymous:1992:CMe

[270] Anonymous. Calendar of meetings. Computational Statistics & Data Analysis, 14(1):130–133, June 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/016794739290093U.

Anonymous:1992:CWSd

[271] Anonymous. Courses, workshops & summer schools. Computational Statistics & Data Analysis, 14(1): 133, June 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290094V.

Loh:1992:IAC

[272] Wei-Yin Loh. Identification of active contrasts in unreplicated factorial experiments. Computational Statistics & Data Analysis, 14(2):135– 148, August 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290169G.

Chan:1992:CSE

[273] Wai sum Chan and William W. S. Wei. A comparison of some estimators of time series autocorrelations. Computational Statistics & Data Analysis, 14 (2):149–163, August 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290170K.

Cheung:1992:EDM

[274] Siu Hung Cheung and Burt Holland. Extension of Dunnett’s multiple comparison procedure with differing sample sizes to the case of several groups. Computational Statistics & Data Analysis, 14(2):165– 182, August 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290171B.

Park:1992:EIS

[275] Byeong U. Park, Sung S. Chung, and Kyung H. Seog. An empirical investigation of the shifted power transformation method in density estimation. Computational Statistics & Data Analysis, 14(2):183– 191, August 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290172C.

Albert:1992:UPP

[276] James H. Albert. Using a probability plot to summarize a unimodal posterior density. Computational Statistics & Data Analysis, 14 (2):193–205, August 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290173D.

Bassett:1992:NRP

[277] Gilbert W. Bassett and Roger W. Koenker. A note on recent proposals for computing l_{1} estimates. Computational Statistics & Data Analysis, 14 (2):207–211, August 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290174E.

Glaz:1992:ATP

[278] Joseph Glaz. Approximations for tail probabilities and moments of the scan statistic. Computational Statistics & Data Analysis, 14(2):213– 227, August 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290175F.

Smith:1992:CAF

[279] D. M. Smith. A comparison of algorithms for forming the QR decomposition for use in regression. Computational Statistics & Data Analysis, 14 (2):229–248, August 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290176G.

Soo:1992:LCN

[280] Yuh-Wen Soo and Douglas M. Bates. Loosely coupled nonlinear least squares. Computational Statistics & Data Analysis, 14(2):249– 259, August 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290177H.

Strawderman:1992:VET

[281] Robert L. Strawderman and Cyrus Mehta. On the validation of exact tests for nonparametric inference. Computational Statistics & Data Analysis, 14 (2):263–266, August 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290178I.

Anonymous:1992:NASf

[282] Anonymous. News about statistical software & related systems. Computational Statistics & Data Analysis, 14 (2):267–269, August 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290179J.

Shanmugam:1992:BRBa

[283] Ramalingam Shanmugam. Book review: Statistical image processing and graphics: Edward J. Wegman and Douglas J. DePriest (eds.) (1986): New York: Marcel Dekker, Inc. Press, 321 pp, ISBN 0-8247-7600-3, $75. Computational Statistics & Data Analysis, 14(2):269, August 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290180N.

Shanmugam:1992:BRBb

[284] Ramalingam Shanmugam. Book review: 11th advances in reliability technology symposium: Peter Comer (ed.) (1990): New York: Elsevier Applied Science Press, xvii + 346 pp, ISBN 1-85166-484-X, $348. Computational Statistics & Data Analysis, 14 (2):269–270, August 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290181E.

Anonymous:1992:BRBb

[285] Anonymous. Book review: Maximum entropy and Bayesian methods: Paul F. Fougere(ed.) (1990): Boston: Kluwer Academic Publishers, xiii + 480 pp, ISBN 0-7923-0928-6, $129. Computational Statistics & Data Analysis, 14 (2):270–271, August 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739290182F.

Anonymous:1992:FIW

[286] Anonymous. Fourth international workshop on artificial intellegence and statistics: January 3–6, 1993 Ft. Lauderdale, Florida, USA. Computational Statistics & Data Analysis, 14 (2):271–273, August 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739290183G.

Anonymous:1992:CMf

[287] Anonymous. Calendar of meetings. Computational Statistics & Data Analysis, 14(2):274–277, August 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/016794739290184H.

Anonymous:1992:CWSe

[288] Anonymous. Courses, workshops & summer schools. Computational Statistics & Data Analysis, 14(2):277– 278, August 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739290185I.

Dodge:1992:E

[289] Yadolah Dodge. Editorial. Computational Statistics & Data Analysis, 14(3):279, October 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290038H.

Altmel:1992:DOC

[290] İsmail Kuban Altmel. The design of optimum component test plans in the demonstration of a series system reliability. Computational Statistics & Data Analysis, 14(3):281– 292, October 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290039I.

Andre:1992:AME

[291] Carmen D. S. André, Clovis A. Peres, and Subhash C. Narula. An algorithm for the MSAE estimation of the multistage dose-response model. Computational Statistics & Data Analysis, 14 (3):293–299, October 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290040M.

Burbea:1992:GSS

[292] Jacob Burbea and Joan del Castillo. Geodesic submanifolds of statistical models with location parameters. Computational Statistics & Data Analysis, 14(3):301–313, October 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290041D.

Celeux:1992:CEA

[293] Gilles Celeux and Gérard Govaert. A classification EM algorithm for clustering and two stochastic versions. Computational Statistics & Data Analysis, 14(3):315–332, October 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739290042E.

Collombier:1992:GOM

[294] D. Collombier. Generally optimal main-effect fractions of U × V designs with U + V units. Computational Statistics & Data Analysis, 14 (3):333–342, October 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739290043F.

Das:1992:RBE

[295] Ashish Das and Sanpei Kageyama. Robustness of BIB and extended BIB designs against the unavailability of any number of observations in a block. Computational Statistics & Data Analysis, 14(3):343– 358, October 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739290044G.

Dette:1992:EDC

[296] Holger Dette. Experimental designs for a class of weighted polynomial regression models. Computational Statistics & Data Analysis, 14(3):359– 373, October 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739290045H.

Dielman:1992:FCT

[297] Terry E. Dielman and Roger C. Pfaffenberger. A further comparison of tests of hypotheses in LAV regression. Computational Statistics & Data Analysis, 14 (3):375–384, October 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290046I. See

notes [774, 903].

Dodge:1992:SNR

[298] Yadolah Dodge and Nikolai Roenko. Stability of L_{1}-norm regression under additional observations. Computational Statistics & Data Analysis, 14 (3):385–390, October 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290047J.

Dorea:1992:PFL

[299] C. C. Y. Dorea and E. S. Miazaki. Parameter free limiting distribution for a random optimization algorithm. Computational Statistics & Data Analysis, 14(3):391–399, October 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290048K.

Anonymous:1992:SSNc

[300] Anonymous. Statistical software newsletter. Computational Statistics & Data Analysis, 14(3):401, October 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/016794739290049L.

Nelder:1992:SPU

[301] John A. Nelder. Statistical packages and unbalanced data. Computational Statistics & Data Analysis, 14 (3):403–406, October 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739290050P.

Anonymous:1992:PS

[302] Anonymous. P-STAT. Computational Statistics & Data Analysis, 14 (3):407–408, October 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739290051G.

Anonymous:1992:ASSb

[303] Anonymous. 2nd announcement of the statistical software guide 1992. Computational Statistics & Data Analysis, 14 (3):408–410, October 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739290052H.

Anonymous:1992:EOI

[304] Anonymous. Election of officers of ISAC Executive Committee and election of IASC Council. Computational Statistics & Data Analysis, 14 (3):410–411, October 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739290053I.

Anonymous:1992:CMg

[305] Anonymous. Calendar of meetings. Computational Statistics & Data Analysis, 14(3):412–415, October 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/016794739290054J.

Anonymous:1992:CWSf

[306] Anonymous. Courses workshops & summer schools. Computational Statistics & Data Analysis, 14(3):415– 416, October 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290055K.

Dupacova:1992:ECR

[307] Jitka Dupačová. Epi-consistency in restricted regression models: the case of a general convex fitting function. Computational Statistics & Data Analysis, 14(4):417– 425, November 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290057M.

Felsenstein:1992:OBD

[308] K. Felsenstein. Optimal Bayesian design for discrimination among rival models. Computational Statistics & Data Analysis, 14(4):427– 436, November 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290058N.

Gyorfi:1992:KDE

[309] László Györfi and Gábor Lugosi. Kernel density estimation from ergodic sample is not universally consistent. Computational Statistics & Data Analysis, 14(4):437– 442, November 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290059O.

Iusem:1992:PDI

[310] Alfredo N. Iusem and Marc Teboulle. A primal-dual iterative algorithm for a maximum likelihood estimation problem. Computational Statistics & Data Analysis, 14(4):443– 456, November 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290060S.

Lejeune:1992:SED

[311] Michel Lejeune and Pascal Sarda. Smooth estimators of distribution and density functions. Computational Statistics & Data Analysis, 14(4):457– 471, November 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290061J.

Mahmoodian:1992:CAG

[312] Ebadollah S. Mahmoodian and Mohammad B. Rostam Abadi. A computer algorithm for generating 3-designs by the method of trade off. Computational Statistics & Data Analysis, 14(4):473– 476, November 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290062K.

Morita:1992:SIM

[313] Hiroshi Morita and Hiroaki Ishii. A stochastic improvement method for stochastic programming. Computational Statistics & Data Analysis, 14 (4):477–487, November 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290063L.

Nguyen:1992:RSE

[314] Nam-Ky Nguyen and Alan J. Miller. A review of some exchange algorithms for constructing discrete Doptimal designs. Computational Statistics & Data Analysis, 14(4):489– 498, November 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290064M.

Nyquist:1992:RE

[315] Hans Nyquist. Restricted Mestimation. Computational Statistics & Data Analysis, 14(4):499– 507, November 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290065N.

Stolyarenko:1992:TDS

[316] Dimitri A. Stolyarenko. Twodimensional spline approximation on local integral measurements: Analysis and design of acoustic fish surveys. Computational Statistics & Data Analysis, 14(4):509– 514, November 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739290066O.

Todorov:1992:CMC

[317] Valentin Todorov. Computing the minimum covariance determinant estimator (MCD) by simulated annealing. Computational Statistics & Data Analysis, 14(4):515– 525, November 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739290067P.

Toutenburg:1992:OEM

[318] H. Toutenburg, G. Trenkler, and E. Liski. Optimal estimation methods under weakened linear restrictions in regression. Computational Statistics & Data Analysis, 14(4):527– 536, November 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739290068Q.

Ritter:1992:DEB

[319] Heiner Ritter and Cornella Züll. Data exchange between statistical packages using the PC program DBMS/ COPY version 3.0. Computational Statistics & Data Analysis, 14(4):539– 544, November 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739290069R.

Murphy:1992:CRF

[320] Brian P. Murphy. COMPSTAT’92 review: ‘far sight’. Computational Statistics & Data Analysis, 14(4):545– 546, November 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290070V.

Tilke:1992:SSD

[321] Clemens Tilke. Statistics for spatial data. Computational Statistics & Data Analysis, 14(4):547, November 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290071M.

Handl:1992:NMA

[322] Andreas Handl. Nonlinear multivariate analysis. Computational Statistics & Data Analysis, 14(4): 548, November 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290072N.

Nagel:1992:LSO

[323] Matthias Nagel. LISP-Stat: an object-oriented environment for statistical computing and dynamic graphics. Computational Statistics & Data Analysis, 14(4):548– 549, November 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290073O.

Anonymous:1992:INc

[324] Anonymous. IASC news. Computational Statistics & Data Analysis, 14 (4):549–551, November 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739290074P.

Anonymous:1992:CMh

[325] Anonymous. Calendar of meetings. Computational Statistics & Data Analysis, 14(4):552–554, November 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/016794739290075Q.

Anonymous:1992:CWSg

[326] Anonymous. Courses, workshops & summer schools. Computational Statistics & Data Analysis, 14(4):554– 555, November 1992. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739290076R.

Anonymous:1993:EBa

[327] Anonymous. Editorial board. Computational Statistics & Data Analysis, 15(1):v–vi, January 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739390215F.

Moulton:1993:BCF

[328] Lawrence H. Moulton, Lisa A. Weissfeld, and Roy T. St. Laurent. Bartlett correction factors in logistic regression models. Computational Statistics & Data Analysis, 15(1): 1–11, January 1993. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390216G.

Breiman:1993:FAM

[329] Leo Breiman. Fitting additive models to regression data: Diagnostics and alternative views. Computational Statistics & Data Analysis, 15 (1):13–46, January 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390217H.

VanCutsem:1993:DOR

[330] Bernard Van Cutsem and Isak Gath. Detection of outliers and robust estimation using fuzzy clustering. Computational Statistics & Data Analysis, 15(1):47–61, January 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390218I.

Saxena:1993:CCT

[331] P. C. Saxena and K. Navaneetham. Comparison of Chernoff-type face and non-graphical methods for clustering multivariate observations. Computational Statistics & Data Analysis, 15 (1):63–79, January 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390219J.

Mann:1993:DSD

[332] G. Mann and R. Haux. Database scheme design for clinical studies based on a semantic data model. Computational Statistics & Data Analysis, 15 (1):81–108, January 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739390220N.

Wiseman:1993:EMC

[333] Michael Wiseman. An efficient method for computing factor similarity coefficients: a computational note and an algorithm. Computational Statistics & Data Analysis, 15(1):111– 114, January 1993. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739390221E.

Anonymous:1993:NASa

[334] Anonymous. News about statistical software & related systems. Computational Statistics & Data Analysis, 15 (1):115–117, January 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739390222F.

Anonymous:1993:INa

[335] Anonymous. IASC news. Computational Statistics & Data Analysis, 15(1):117, January 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739390223G.

Anonymous:1993:CMa

[336] Anonymous. Calendar of meetings. Computational Statistics & Data Analysis, 15(1):119–122, January 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/016794739390224H.

Anonymous:1993:CWSa

[337] Anonymous. Courses, workshops & summer schools. Computational Statistics & Data Analysis, 15(1): 122, January 1993. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390225I.

Anonymous:1993:RC

[338] Anonymous. Referees for CSDA-1992. Computational Statistics & Data Analysis, 15(1):123, January 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390226J.

Barr:1993:STP

[339] Aiala Barr. A simultaneous test procedure for comparing several treatments in the analysis of reciprocals. Computational Statistics & Data Analysis, 15 (2):125–131, February 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390187X.

Jamshidian:1993:MNM

[340] Mortaza Jamshidian and Peter M. Bentler. A modified Newton method for constrained estimation in covariance structure analysis. Computational Statistics & Data Analysis, 15 (2):133–146, February 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739390188Y.

Lin:1993:GAR

[341] Dennis K. J. Lin and Norman R. Draper. Generating alias relationships for two-level Plackett and Burman designs. Computational Statistics & Data Analysis, 15(2):147– 157, February 1993. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739390189Z.

Yau:1993:CSA

[342] Linda H.-Y. Yau, Sik-Yum Lee, and Wai-Yin Poon. Covariance structure analysis with three-level data. Computational Statistics & Data Analysis, 15 (2):159–178, February 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947393901905.

Romanazzi:1993:JEE

[343] Mario Romanazzi. Jackknife estimation of the eigenvalues of the covariance matrix. Computational Statistics & Data Analysis, 15(2):179– 198, February 1993. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739390191U.

Granville:1993:CCI

[344] Vincent Granville, Mirko Křivánek, and Jean-Paul Rasson. Clustering, classification and image segmentation on the grid. Computational Statistics & Data Analysis, 15(2):199– 209, February 1993. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390192V.

Saltelli:1993:SAM

[345] A. Saltelli, T. H. Andres, and T. Homma. Sensitivity analysis of model output: an investigation of new techniques. Computational Statistics & Data Analysis, 15(2):211– 238, February 1993. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390193W.

Anonymous:1993:NASb

[346] Anonymous. News about statistical software & related systems. Computational Statistics & Data Analysis, 15 (2):241–264, February 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390194X.

Anonymous:1993:INb

[347] Anonymous. IASC news. Computational Statistics & Data Analysis, 15 (2):265–267, February 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390195Y.

Anonymous:1993:CMb

[348] Anonymous. Calendar of meetings. Computational Statistics & Data Analysis, 15(2):268–271, February 1993.

CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/016794739390196Z.

Anonymous:1993:CWSb

[349] Anonymous. Courses, workshops > summer schools. Computational Statistics & Data Analysis, 15(2): 271, February 1993. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947393901972.

Bener:1993:EIC

[350] A. Bener and S. Huda. Estimation of intraclass correlation when family sizes are unequal. Computational Statistics & Data Analysis, 15 (3):273–284, March 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390256S.

Ostrouchov:1993:MSP

[351] George Ostrouchov and Edward L. Frome. A model search procedure for hierarchial models. Computational Statistics & Data Analysis, 15 (3):285–296, March 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390257T.

Granville:1993:BFM

[352] V. Granville and J. P. Rasson. A Bayesian filter for a mixture model of noise in image remote sensing. Computational Statistics & Data Analysis, 15(3):297–308, March 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390258U.

Talwar:1993:SSS

[353] Prem P. Talwar. A simulation study of some non-parametric regression estimators. Computational Statistics & Data Analysis, 15(3): 309–327, March 1993. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390259V.

Bates:1993:GBC

[354] Douglas Bates, Fred Reames, and Grace Wahba. Getting better contour plots with S and GCVPACK. Computational Statistics & Data Analysis, 15(3):329–342, March 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390260Z.

Wang:1993:ATP

[355] Suojin Wang and H. L. Gray. Approximating tail probabilities of noncentral distributions. Computational Statistics & Data Analysis, 15(3): 343–352, March 1993. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390261Q.

Hinde:1993:PRM

[356] John Hinde. Participant’s reflections on meetings: SGCSA conference on survey and statistical computing 24–26 September 1992, Bristol, UK. Computational Statistics & Data Analysis, 15(3):355–356, March 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390262R.

Anonymous:1993:NASc

[357] Anonymous. News about statistical software & related systems. Computational Statistics & Data Analysis, 15(3):357, March 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390263S.

Uebe:1993:BRB

[358] Götz Uebe. Book review: Statistics tutor, tutorial and computational software for the Behavioural Sciences: J. D. Allen, D. J. Pittenger (1991): New York: John Wiley & Sons, student edition, 215 pp., ISBN 0-471-50280-4, US$21.95. Computational Statistics & Data Analysis, 15(3): 358–359, March 1993. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390264T.

Anonymous:1993:IMN

[359] Anonymous. IASC meetings in Neuchâtel, August 1992. Computational Statistics & Data Analysis, 15 (3):359–360, March 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390265U.

Achcar:1993:IBS

[360] Jorge Alberto Achcar. Inferences for the Birnbaum–Saunders fatigue life model using Bayesian methods. Computational Statistics & Data Analysis, 15(4):367–380, May 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390170X.

Bedrick:1993:DBR

[361] Edward J. Bedrick and Chih-Ling Tsai. Diagnostics for binomial response models using power divergence statistics. Computational Statistics & Data Analysis, 15(4):381– 392, May 1993. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390171O.

Benasseni:1993:PAC

[362] Jacques Bénasséni. Perturbational aspects in correspondence analysis. Computational Statistics & Data Analysis, 15(4):393–410, May 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390172P.

Boik:1993:CTI

[363] Robert J. Boik. A comparison of three invariant tests of additivity in two-way classifications with no replications. Computational Statistics & Data Analysis, 15(4):411– 424, May 1993. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739390173Q.

Posa:1993:SDS

[364] D. Posa. A simple description of spatial-temporal processes. Computational Statistics & Data Analysis, 15(4):425–437, May 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739390174R.

Tu:1993:RAD

[365] Xin M. Tu, Donald S. Burdick, and Robert E. Albright. Regression analysis of data with repeated measurements using the method of successive differences. Computational Statistics & Data Analysis, 15(4): 439–449, May 1993. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739390175S.

Rocke:1993:AEP

[366] David M. Rocke. Almost-exact parametric bootstrap calculation via the saddlepoint approximation. Computational Statistics & Data Analysis, 15(4):451–460, May 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739390176T.

Anonymous:1993:E

[367] Anonymous. Erratum. Computational Statistics & Data Analysis, 15 (4):461, May 1993. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390177U.

Anonymous:1993:SSNa

[368] Anonymous. Statistical software newsletter. Computational Statistics & Data Analysis, 15(4):463, May 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/016794739390178V.

Anonymous:1993:NASd

[369] Anonymous. News about statistical software & related systems. Computational Statistics & Data Analysis, 15(4):465–467, May 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390179W.

Anonymous:1993:BRBa

[370] Anonymous. Book review: Experimental design in biotechnology: P. D. Haaland (1989): New York: Marcel Dekker Inc., ISBN 0-631-15943-6, pp. 248, US$99.75. Computational Statistics & Data Analysis, 15 (4):468, May 1993. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947393901802.

Anonymous:1993:BRBb

[371] Anonymous. Book review: Uncertainty and vagueness in knowledge based systems: R. Kruse, E. Schwecke, J. Heinsohn (1991): numerical methods (Artificial Intelligence). Berlin, Heidelberg: Springer, ISBN 3-540-54169-1 pp. 491, DM 98. Computational Statistics & Data Analysis, 15(4):468–469, May 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390181R.

Anonymous:1993:BRBc

[372] Anonymous. Book review: SAS system — eine praxisbezogene einführung: Frank Multrus and Michael Bleicher (1992): Hidelberg: Hüthig Buch Verlag, ISBN 3-7785-2157-8, pp. 241, DM 68. Computational Statistics & Data Analysis, 15(4): 469, May 1993. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390182S.

Anonymous:1993:BRBd

[373] Anonymous. Book review: Univariate varianzanalysen: P. Zöfel (1992): Achtzehn komplette designs. Stuttgart and Jena: Gustav Fischer Verlag (UTB für Wissenschaft: Uni–Taschenbücher 1663), ISBN 3-437-40278-1, x + 299 S., 4 Abb., 7 Tafeln, DM 29.80. Computational Statistics & Data Analysis, 15(4):469–470, May 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390183T.

Anonymous:1993:INc

[374] Anonymous. IASC news. Computational Statistics & Data Analysis, 15(4):470–471, May 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390184U.

Anonymous:1993:CMc

[375] Anonymous. Calendar of meetings. Computational Statistics & Data Analysis, 15(4):472–476, May 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/016794739390185V.

Anonymous:1993:AIVa

[376] Anonymous. Author index volume 15 (1993). Computational Statistics & Data Analysis, 15(4):479– 480, May 1993. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390186W.

Anonymous:1993:EBb

[377] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 16(1):i–ii, June 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390239P.

Alexander:1993:TMI

[378] William P. Alexander. Testing the means of independent normal random variables. Computational Statistics & Data Analysis, 16(1): 1–10, June 1993. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390240T.

Hall:1993:IBD

[379] Peter Hall, Wolfgang Härdle, and Léopold Simar. On the inconsistency of bootstrap distribution estimators. Computational Statistics & Data Analysis, 16(1):11–18, June 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390241K.

Kalantari:1993:SBM

[380] B. Kalantari, I. Lari, A. Rizzi, and B. Simeone. Sharp bounds for the maximum of the chi-square index in a class of contingency tables with given marginals. Computational Statistics & Data Analysis, 16(1):19–34, June 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390242L.

Liang:1993:CRE

[381] TaChen Liang. Convergence rates for empirical Bayes estimation of the scale parameter in a Pareto distribution. Computational Statistics & Data Analysis, 16(1):35–45, June 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390243M.

Marasinghe:1993:TDF

[382] Mervyn G. Marasinghe and Robert J. Boik. A three-degree of freedom test of additivity in three-way classifications. Computational Statistics & Data Analysis, 16(1):47–61, June 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390244N.

Paula:1993:ALI

[383] Gilberto A. Paula. Assessing local influence in restricted regression models. Computational Statistics & Data Analysis, 16(1):63–79, June 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390245O.

Hawkins:1993:FSA

[384] Douglas M. Hawkins. The feasible set algorithm for least median of squares regression. Computational Statistics & Data Analysis, 16(1): 81–101, June 1993. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390246P.

Kiers:1993:ALS

[385] Henk A. L. Kiers. An alternating least squares algorithm for PARAFAC2 and three-way DEDICOM. Computational Statistics & Data Analysis, 16(1):103–118, June 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390247Q.

Vedder:1993:IAN

[386] John D. Vedder. An invertible approximation to the normal distribution function. Computational Statistics & Data Analysis, 16(1): 119–123, June 1993. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739390248R.

Gefeller:1993:GME

[387] Olaf Gefeller and Franz Woltering. A general method of estimating measures of association and their asymptotic variances under the multinomial model using standard statistical software. Computational Statistics & Data Analysis, 16(1):127– 130, June 1993. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739390249S.

Anonymous:1993:PRM

[388] Anonymous. Participant’s reflections on meetings. Computational Statistics & Data Analysis, 16(1): 131–132, June 1993. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739390250W.

Anonymous:1993:ABS

[389] Anonymous. Announcing the Statistical Software Guide ’93/94. Computational Statistics & Data Analysis, 16(1):133, June 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739390251N.

Anonymous:1993:NASe

[390] Anonymous. News about statistical software & related systems. Computational Statistics & Data Analysis, 16(1):134, June 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390252O.

Antoch:1993:IN

[391] Jaromír Antoch. IASC news. Computational Statistics & Data Analysis, 16(1):135–136, June 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390253P.

Anonymous:1993:CMd

[392] Anonymous. Calendar of meetings. Computational Statistics & Data Analysis, 16(1):137–140, June 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/016794739390254Q.

Anonymous:1993:CWSc

[393] Anonymous. Courses, workshop & summer schools. Computational Statistics & Data Analysis, 16(1): 141–142, June 1993. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390255R.

Davies:1993:NCR

[394] Richard B. Davies. Nonparametric control for residual heterogeneity in modelling recurrent behaviour. Computational Statistics & Data Analysis, 16 (2):143–160, August 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390110F.

Schott:1993:DRQ

[395] James R. Schott. Dimensionality reduction in quadratic discriminant analysis. Computational Statistics & Data Analysis, 16(2):161– 174, August 1993. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947393901116.

Baglivo:1993:ADD

[396] Jenny Baglivo, Donald Olivier, and Marcello Pagano. Analysis of discrete data: Rerandomization methods and complexity. Computational Statistics & Data Analysis, 16(2):175– 184, August 1993. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947393901127.

Contini:1993:EMD

[397] D. Contini and G. Lovison. The effect of marginal disuniformity on the χ^{2} approximation to the distribution of Pearson’s X^{2} in sparse contingency tables. Computational Statistics & Data Analysis, 16(2):185– 199, August 1993. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947393901138.

tenBerge:1993:ISP

[398] Jos M. F. ten Berge and Paul A. Bekker. The isotropic scaling problem in generalized Procrustes analysis. Computational Statistics & Data Analysis, 16(2):201– 204, August 1993. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947393901149.

Davis:1993:CGM

[399] Charles S. Davis. The computer generation of multinomial random variates. Computational Statistics & Data Analysis, 16(2):205– 217, August 1993. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390115A.

Yesilcay:1993:BAD

[400] Yasar Yesilcay. The BOYS algorithm for determining optimum matching rules. Computational Statistics & Data Analysis, 16(2):219– 234, August 1993. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390116B.

Anonymous:1993:SSNb

[401] Anonymous. Statistical software newsletter. Computational Statistics & Data Analysis, 16(2):235, August 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/016794739390117C.

Haag:1993:WHB

[402] Uwe Haag, Armin Koch, and Werner Vach. Why and how to build a simple statistical computing system. Computational Statistics & Data Analysis, 16(2):237–244, August 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739390118D.

Anonymous:1993:Ia

[403] Anonymous. Information. Computational Statistics & Data Analysis, 16(2):245, August 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739390120I.

Anonymous:1993:O

[404] Anonymous. Obituary. Computational Statistics & Data Analysis, 16 (2):245, August 1993. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739390119E.

Anonymous:1993:BRBe

[405] Anonymous. Book review: Design and Analysis of Bioavailability and Bioequivalence Studies: Chow, Shein-Chung and Liu, Jen-Pei (1992): New York: Marcel Dekker, Inc., 432 pp., ISBN 0-8247-8682-3, $125.00. Computational Statistics & Data Analysis, 16(2):246, August 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947393901219.

Anonymous:1993:BRBf

[406] Anonymous. Book review: Data Quality Control: Theory and Pragmatics: Liepins, Gunar E. and Uppuluri, V. R. R. (eds.) (1990) New York: Marcel Dekker, Inc., 384 pp, ISBN 0-8247-8354-9, $89.75 (U.S. and Canada), $107.50 (all other countries). Computational Statistics & Data Analysis, 16(2):246–247, August 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390122A.

Anonymous:1993:BRBg

[407] Anonymous. Book review: Datenanalyse mit SPSS: Wittenberg, Reinhard and Cramer, Hans (1992) Stuttgart: Gustav Fischer Verlag, pp. 200, ISBN 3-8252-1602-0, DM 24.80. Computational Statistics & Data Analysis, 16(2):248, August 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390124C.

Anonymous:1993:BRBh

[408] Anonymous. Book review: SAS kompakt für die Version 6: Göttsche, T. (1992) Stuttgart: Gustav Fischer Verlag, pp. 192, ISBN 3-437-40281-1, DM 49.-. Computational Statistics & Data Analysis, 16(2): 248, August 1993. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390123B.

Anonymous:1993:INd

[409] Anonymous. IASC news. Computational Statistics & Data Analysis, 16 (2):248–250, August 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390125D.

Anonymous:1993:CMe

[410] Anonymous. Calendar of meetings. Computational Statistics & Data Analysis, 16(2):251–255, August 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/016794739390126E.

Anonymous:1993:CWSd

[411] Anonymous. Courses, workshops & summer schools. Computational Statistics & Data Analysis, 16(2):255– 256, August 1993. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739390127F.

Bradu:1993:SSR

[412] Dan Bradu and Douglas M. Hawkins. Sample size requirements for multiple outlier location techniques based on elemental sets. Computational Statistics & Data Analysis, 16(3):257– 270, September 1993. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739390128G.

Fung:1993:QBM

[413] Wing-Kam Fung and J. Bacon-Shone. Quasi-Bayesian modelling of multivariate outliers. Computational Statistics & Data Analysis, 16(3):271– 278, September 1993. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739390129H.

Joe:1993:MDM

[414] Harry Joe. Multivariate dependence measures and data analysis. Computational Statistics & Data Analysis, 16 (3):279–297, September 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390130L.

Velilla:1993:ECD

[415] Santiago Velilla. On eigenvalues, case deletion and extremes in regression. Computational Statistics & Data Analysis, 16(3):299– 309, September 1993. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390131C.

Wand:1993:HEG

[416] M. P. Wand and Luc Devroye. How easy is a given density to estimate? Computational Statistics & Data Analysis, 16(3):311– 323, September 1993. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390132D.

Coster:1993:TMC

[417] Daniel C. Coster. Tables of minimum cost, linear trend-free run sequences for two- and three-level fractional factorial designs. Computational Statistics & Data Analysis, 16 (3):325–336, September 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390133E.

Wong:1993:NCL

[418] Augustine C. M. Wong. Numerical conversion of likelihood to significance. Computational Statistics & Data Analysis, 16(3):337– 348, September 1993. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739390134F.

Xu:1993:PAL

[419] Chong wei Xu and Wei-Kei Shiue. Parallel algorithms for least median of squares regression. Computational Statistics & Data Analysis, 16(3):349– 362, September 1993. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739390135G.

Bender:1993:UCS

[420] Ralf Bender. Using common statistical program packages for the Friedman test when there are ties among treatments and a multiple comparison procedure is needed. Computational Statistics & Data Analysis, 16 (3):365–369, September 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739390136H.

Anonymous:1993:MAS

[421] Anonymous. Mihalsin Associates & StatSci new range of multivariate data analysis techniques to be offered. Computational Statistics & Data Analysis, 16(3):370, September 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390137I.

Anonymous:1993:ESS

[422] Anonymous. ESRI & StatSci statistical software integrated with ARC/ INFO GIS. Computational Statistics & Data Analysis, 16(3):370– 371, September 1993. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390138J.

Anonymous:1993:PSI

[423] Anonymous. Poly software international PSI-Plot version 2.0. Computational Statistics & Data Analysis, 16(3):371, September 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390139K.

Anonymous:1993:UWS

[424] Anonymous. UNISTAT for windows spreadsheet, statistics, graphics. Computational Statistics & Data Analysis, 16(3):371–372, September 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/016794739390140O.

Anonymous:1993:PIW

[425] Anonymous. PASE ’93 4th international workshop on parallel applications in statistics and economics: November 22–26, 1993, Ascona, Switzerland Centro Stefano Franscini, Monte Verita. Computational Statistics & Data Analysis, 16(3):

372, September 1993. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739390141F.

Anonymous:1993:CMf

[426] Anonymous. Calendar of meetings. Computational Statistics & Data Analysis, 16(3):373–377, September 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/016794739390142G.

Aldrin:1993:PPR

[427] Magne Aldrin, Erik Bølviken, and Tore Schweder. Projection pursuit regression for moderate non-linearities. Computational Statistics & Data Analysis, 16(4):379–403, October 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739390156N.

Anderson:1993:AMS

[428] Oliver D. Anderson. Approximate moments to O(n^{−2}) for the sampled partial autocorrelations from a white noise process. Computational Statistics & Data Analysis, 16(4):405– 421, October 1993. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739390157O.

Durand:1993:GPC

[429] Jean-François Durand. Generalized principal component analysis with respect to instrumental variables via univariate spline transformations. Computational Statistics & Data Analysis, 16 (4):423–440, October 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390158P.

Lee:1993:AMI

[430] Sik-Yum Lee and Wai-Tung Ho. Analysis of multisample identified and nonidentified structural equation models with stochastic constraints. Computational Statistics & Data Analysis, 16 (4):441–453, October 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390159Q.

Schmidt:1993:PME

[431] Karsten Schmidt. On the performance of minimax estimators in linear regression. Computational Statistics & Data Analysis, 16(4):455– 468, October 1993. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390160U.

Anonymous:1993:SSNc

[432] Anonymous. Statistical software newsletter. Computational Statistics & Data Analysis, 16(4):469, October 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/016794739390161L.

Murphy:1993:SPT

[433] Brian P. Murphy and C. M. Randolph. Simulation packages today: new user classes and order-of-magnitude breakthroughs in development costs. Computational Statistics & Data Analysis, 16(4):471–479, October 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739390162M.

Anonymous:1993:Ib

[434] Anonymous. Information. Computational Statistics & Data Analysis, 16(4):480, October 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739390163N.

Anonymous:1993:INe

[435] Anonymous. IASC news. Computational Statistics & Data Analysis, 16(4):481, October 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739390164O.

Anonymous:1993:CMg

[436] Anonymous. Calendar of meetings. Computational Statistics & Data Analysis, 16(4):482–485, October 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/016794739390165P.

Anonymous:1993:CWSe

[437] Anonymous. Courses, workshops & Summer schools. Computational Statistics & Data Analysis, 16(4):485– 486, October 1993. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390166Q.

Anonymous:1993:AIS

[438] Anonymous. Author index of SS-NinCSDA: volume 11 (1991) — volume 16 (1993). Computational Statistics & Data Analysis, 16(4): 487, October 1993. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390167R.

Anonymous:1993:BRS

[439] Anonymous. Book reviews of SS-NinCSDA: volume 11 (1991) — volume 16 (1993). Computational Statistics & Data Analysis, 16(4): 488, October 1993. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390168S.

Anonymous:1993:AIVb

[440] Anonymous. Author index volume 16 (1993). Computational Statistics & Data Analysis, 16(4):489– 490, October 1993. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390169T.

Anonymous:1994:EBa

[441] Anonymous. Editorial board. Computational Statistics & Data Analysis, 17(1):i–ii, January 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394900299.

Longford:1994:LRR

[442] N. T. Longford. Logistic regression with random coefficients. Computational Statistics & Data Analysis, 17(1):1–15, January 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739200062V.

Mickey:1994:AMO

[443] Ruth M. Mickey, Sarah J. Shema, Pamela M. Vacek, and Dianne Y. Bell. Analysis of multiple outcome variables measured longitudinally. Computational Statistics & Data Analysis, 17 (1):17–33, January 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739200059Z.

Harwell:1994:MCS

[444] Michael R. Harwell and Ronald C. Serlin. A Monte Carlo study of the Friedman test and some competitors in the single factor, repeated measures design with unequal covariances. Computational Statistics & Data Analysis, 17(1):35–49, January 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947392000605.

Luceno:1994:FOE

[445] Alberto Luceño. Fast optimization of the exact likelihood of AR and ARMA processes. Computational Statistics & Data Analysis, 17(1):

51–63, January 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739200005C.

Rousseeuw:1994:RRD

[446] Peter J. Rousseeuw and Joachim Wagner. Robust regression with a distributed intercept using least median of squares. Computational Statistics & Data Analysis, 17(1): 65–76, January 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739200063W.

Boukai:1994:GBI

[447] Benzion Boukai. Graphically based interval estimation for the changepoint. Computational Statistics & Data Analysis, 17(1):77–86, January 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/016794739200064X.

Rayner:1994:SSP

[448] Robert K. Rayner. The smallsample power of Durbin’s h test revisited. Computational Statistics & Data Analysis, 17(1):87– 94, January 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739200061U.

Kossler:1994:EEV

[449] W. Kössler, B. Lenz, and H.-J. Lenz. ExVar — exact variable inspection plans in statistical quality control. Computational Statistics & Data Analysis, 17(1):97– 99, January 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394900302.

Bassett:1994:MMM

[450] E. E. Bassett, S. P. Brooks, and ByronJ. T. Morgan. Minitab macros for multivariate analysis. Computational Statistics & Data Analysis, 17 (1):100–103, January 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394900310.

Anonymous:1994:BRBa

[451] Anonymous. Book review: Statistische Graphik: Jürgen R. Geßler (1993): Basel: Birkhäuser Verlag, ISBN 3-7643-2874-6, pp. 285, DM 58. Computational Statistics & Data Analysis, 17(1):103, January 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394900329.

Anonymous:1994:BRBb

[452] Anonymous. Book review: Zur Methodik der räumlichen Interpolation mit geostatistischen Verfahren: U. Heinrich (1992); Wiesbaden: Deutscher Universitäts–Verlag, ISBN 3-8244-2027-9, pp. 124, DM 38. Computational Statistics & Data Analysis, 17(1):103–104, January 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394900337.

Anonymous:1994:BRBc

[453] Anonymous. Book review: Statistik in der Praxis (3. überarbeitete und ergänzte Auflage): Peter Zöfel (1992): Stuttgard Jena: G. Fischer, ISBN 3-8252-1293-9, pp. 422, DM 34.80. Computational Statistics & Data Analysis, 17(1):104, January 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394900345.

Anonymous:1994:BRBd

[454] Anonymous. Book review: Mathematica als Werkzeug: Stephan Kaufmann (1992): Basel: Birkhäuser Verlag, ISBN 3-7643-2832-0, pp. 396, DM 54. Computational Statistics & Data Analysis, 17(1):104– 105, January 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394900353.

Anonymous:1994:INa

[455] Anonymous. IASC news. Computational Statistics & Data Analysis, 17 (1):105–108, January 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394900361.

Anonymous:1994:CMa

[456] Anonymous. Calendar of meetings. Computational Statistics & Data Analysis, 17(1):109–118, January 1994.

CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/016794739490037X.

Dumitrescu:1994:POT

[457] Monica E. Bad Dumitrescu, Denis N. Enăchescu, and Florentina Th. Hristea. On the performances of an outlier test in the case of the exponential distribution. Computational Statistics & Data Analysis, 17(2):119– 127, February 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739200065Y.

Fisher:1994:RPI

[458] Amy Fisher and Paul S. Horn. Robust prediction intervals in a regression setting. Computational Statistics & Data Analysis, 17(2):129– 140, February 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947392000694.

Arrenberg:1994:NRC

[459] Jutta Arrenberg. Natural ranks in the conditional Wilcoxon rank sum test. Computational Statistics & Data Analysis, 17(2):141– 152, February 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947392000672.

Cao:1994:CSS

[460] Ricardo Cao, Antonio Cuevas, and Wensceslao González Manteiga. A comparative study of several smoothing methods in density estimation. Computational Statistics & Data Analysis, 17 (2):153–176, February 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739200066Z.

Frank:1994:ESR

[461] Ove Frank and John Örvik. Entropy of sums of random digits. Computational Statistics & Data Analysis, 17 (2):177–184, February 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947392000683.

Hawkins:1994:FSAa

[462] Douglas M. Hawkins. The feasible solution algorithm for least trimmed squares regression. Computational Statistics & Data Analysis, 17(2):185– 196, February 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947392000708.

Hawkins:1994:FSAb

[463] Douglas M. Hawkins. The feasible solution algorithm for the minimum covariance determinant estimator in multivariate data. Computational Statistics & Data Analysis, 17 (2):197–210, February 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739200071X.

Koch:1994:BSG

[464] Armin Koch and Uwe Haag. The Statistical Software Guide ’93/94. Computational Statistics & Data Analysis, 17 (2):215–238, February 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947394901120.

Matiaske:1994:CAP

[465] Wenzel Matiaske. Correspondence analysis with P-STAT: the macro COR.ANALYSIS. Computational Statistics & Data Analysis, 17(2): 239, February 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947394901139.

Pabst:1994:BRBa

[466] Günther Pabst. Book review: Softstat’ 91 — Advances in Statistical Software 3: Frank Faulbaum, ed. (1992): 552 pp. Stuttgart: Gustav Fischer Verlag, ISBN 3-437-40280-3 (94.00 DM) and New York: VCH Publishers, ISBN 1-56081-348-2. Computational Statistics & Data Analysis, 17(2):240, February 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947394901147.

Trenkler:1994:BRBa

[467] Götz Trenkler. Book review: Univariate Discrete Distributions: N. L. Johnson, S. Kotz and A. W. Kemp (1992): 2nd Edition. New York: John Wiley, ISBN 0-471-54897-9, pp. xx + 564, US$68.95. Computational Statistics & Data Analysis, 17(2):240– 241, February 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394901155.

Haerting:1994:BRB

[468] Johannes Haerting. Book review: Statistische Methoden, Planung und Auswertung. 7. überarbeitete Auflage: Lothar Sachs (1993): Berlin: Springer, ISBN 3-540-55821-7, pp. 312, DM 42.00. Computational Statistics & Data Analysis, 17(2): 241, February 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394901163.

Anonymous:1994:INb

[469] Anonymous. IASC news. Computational Statistics & Data Analysis, 17 (2):241–244, February 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394901171.

Anonymous:1994:PRM

[470] Anonymous. Participant’s reflections on meetings. Computational Statistics & Data Analysis, 17(2):245– 246, February 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739490118X.

Anonymous:1994:M

[471] Anonymous. In memoriam. Computational Statistics & Data Analysis, 17(2):246, February 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947394901198.

Jamshidian:1994:CGM

[472] Mortaza Jamshidian and Robert I. Jennrich. Conjugate gradient methods in confirmatory factor analysis. Computational Statistics & Data Analysis, 17(3):247–263, March 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947394901201.

Poon:1994:DFA

[473] Wai-Yin Poon and Sik-Yum Lee. A distribution free approach for analysis of two-level structural equation model. Computational Statistics & Data Analysis, 17(3):265– 275, March 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947392000742.

Cohen:1994:MHT

[474] Ayala Cohen and Meir Rom. A method for hypothesis tests in polychotomous logistic regression. Computational Statistics & Data Analysis, 17(3):277–288, March 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739490121X.

Todorov:1994:RTG

[475] Valentin Todorov, Neyko Neykov, and Plamen Neytchev. Robust twogroup discrimination by bounded influence regression. a Monte Carlo simulation. Computational Statistics & Data Analysis, 17(3):289– 302, March 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394901228.

Cook:1994:APG

[476] R. Dennis Cook and Sanford Weisberg. ARES plots for generalized linear models. Computational Statistics & Data Analysis, 17(3):303– 315, March 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947392000753.

Ellis:1994:MLD

[477] Steven P. Ellis and Maria McDermott. Measuring the local dimension of point clouds. Computational Statistics & Data Analysis, 17(3): 317–326, March 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947392000764.

Mee:1994:PLW

[478] Robert W. Mee and Debashis Kushary. Prediction limits for the Weibull distribution utilizing simulation. Computational Statistics & Data Analysis, 17(3):327–336, March 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739200072Y.

Anonymous:1994:NASa

[479] Anonymous. News about statistical software & related systems. Computational Statistics & Data Analysis, 17 (3):339–340, March 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947394901236.

Anonymous:1994:Ia

[480] Anonymous. Information. Computational Statistics & Data Analysis, 17(3):340, March 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947394901244.

Anonymous:1994:INc

[481] Anonymous. IASC news. Computational Statistics & Data Analysis, 17 (3):341–344, March 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947394901252.

Anonymous:1994:CMb

[482] Anonymous. Calendar of meetings. Computational Statistics & Data Analysis, 17(3):345–349, March 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/0167947394901260.

Anonymous:1994:CWSa

[483] Anonymous. Courses, workshops & summer schools. Computational Statistics & Data Analysis, 17(3): 349–350, March 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394901279.

Hilton:1994:ACE

[484] Joan F. Hilton, Cyrus R. Mehta, and Nitin R. Patel. An algorithm for conducting exact Smirnov tests. Computational Statistics & Data Analysis, 17(4):351–361, May 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394900175.

Lesaffre:1994:SAT

[485] Emmanuel Lesaffre, Geert Verbeke, and Geert Molenberghs. A sensitivity analysis of two multivariate response models. Computational Statistics & Data Analysis, 17(4): 363–391, May 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394900183.

Marcus:1994:SCI

[486] Ruth Marcus and Abraham Genizi. Simultaneous confidence intervals for umbrella contrasts of normal means. Computational Statistics & Data Analysis, 17(4):393– 407, May 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947394900191.

Olejnik:1994:TER

[487] Stephen Olejnik and Wei-Ming Luh. Type I error rates, power, and sample sizes for two-stage solutions to the Behrens–Fisher problem when population distributions are non-normal. Computational Statistics & Data Analysis, 17(4):409– 420, May 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947394900205.

Variyam:1994:CFE

[488] Jayachandran N. Variyam. A common framework for estimating multivariate autoregressive index models. Computational Statistics & Data Analysis, 17(4):421–432, May 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947394900213.

Tsui:1994:EDD

[489] Albert K. Tsui and Mukhtar M. Ali. Exact distributions, density functions and moments of the last squares estimator in a first-order autoregressive model. Computational Statistics & Data Analysis, 17(4): 433–454, May 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947394900221.

Anonymous:1994:SSNa

[490] Anonymous. Statistical software newsletter. Computational Statistics & Data Analysis, 17(4):455, May 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/016794739490023X.

Schimek:1994:DER

[491] Michael G. Schimek and Klaus G. Schmaranz. Dependent error regression smoothing: a new method and PC program. Computational Statistics & Data Analysis, 17(4): 457–464, May 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394900248.

Edler:1994:INa

[492] Lutz Edler. IASC news. Computational Statistics & Data Analysis, 17(4):465–470, May 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394900256.

Hormann:1994:I

[493] Allmut Hörmann. Information. Computational Statistics & Data Analysis, 17(4):470, May 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394900264.

Anonymous:1994:CMc

[494] Anonymous. Calendar of meetings. Computational Statistics & Data Analysis, 17(4):471, May 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394900272.

Anonymous:1994:CWSb

[495] Anonymous. Courses, workshops & summer schools. Computational Statistics & Data Analysis, 17(4): 471–472, May 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394900280.

Furman:1994:TNC

[496] W. David Furman and Bruce G. Lindsay. Testing for the number of components in a mixture of normal distributions using moment estimators. Computational Statistics & Data Analysis, 17(5):473–492, June 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394901449.

Furman:1994:MRE

[497] W. David Furman and Bruce G. Lindsay. Measuring the relative effectiveness of moment estimators as starting values in maximizing likelihoods. Computational Statistics & Data Analysis, 17(5):493–507, June 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394901457.

Sejling:1994:MRR

[498] Ken Sejling, Henrik Madsen, Jan Holst, Ulla Holst, and Jan-Eric Englund.

Methods for recursive robust estimation of AR parameters. Computational Statistics & Data Analysis, 17 (5):509–536, June 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394901465.

Ahmed:1994:IEM

[499] S. Ejaz Ahmed. Improved estimation in a multivariate regression model. Computational Statistics & Data Analysis, 17(5):537–554, June 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394901473.

Andres:1994:COU

[500] A. Martín Andrés and A. Silva Mato. Choosing the optimal unconditioned test for comparing two independent proportions. Computational Statistics & Data Analysis, 17 (5):555–574, June 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394901481.

Vieu:1994:CRN

[501] Philippe Vieu. Choice of regressors in nonparametric estimation. Computational Statistics & Data Analysis, 17(5):575–594, June 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739490149X.

Anonymous:1994:NASb

[502] Anonymous. News about statistical software and related systems. Computational Statistics & Data Analysis, 17(5):597–599, June 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394901503.

Trenkler:1994:BRBb

[503] Götz Trenkler. Book review: Handbook of statistics — computational statistics: C. R. Rao, ed. (1993): Amsterdam, London, New York, Tokyo: North Holland, 920 p. ISBN 0-444-88096-8, US$190.00. Computational Statistics & Data Analysis, 17(5): 600–602, June 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394901511.

Schmerling:1994:BRBa

[504] Siegfried Schmerling. Book review: Computer intensive methods in statistics. Series statistics and computing: W. Härdle and L. Simar, eds. (1993): Heidelberg: Physica-Verlag and New York: Springer Verlag, viii + 176 p., 34 Fig., 16 Tab., ISBN 3-7908-0677-3 and 0-387-91443-9, 75,00 DM. Computational Statistics & Data Analysis, 17(5):602–603, June 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739490152X.

Schmerling:1994:BRBb

[505] Siegfried Schmerling. Book review: Applications of Fibonacci numbers — volume 5: G. E. Bergum, A. N. Philippou, A. F. Horadam, eds. (1993): Dordrecht, Boston, London: Kluwer Academic Publishers, xxxiv + 625 p., 35 Fig., 30 Tab., ISBN 0-7923-2491-9. Computational Statistics & Data Analysis, 17(5):603, June 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394901538.

Anonymous:1994:Ib

[506] Anonymous. Information. Computational Statistics & Data Analysis, 17(5):603–606, June 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394901546.

Anonymous:1994:CMd

[507] Anonymous. Calendar of meetings. Computational Statistics & Data Analysis, 17(5):607–609, June 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/0167947394901554.

Anonymous:1994:CWSc

[508] Anonymous. Courses, workshops & summer schools. Computational Statistics & Data Analysis, 17(5): 609–610, June 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394901562.

Anonymous:1994:AIVa

[509] Anonymous. Author index volume 17 (1994). Computational Statistics & Data Analysis, 17(5):611– 612, June 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947394901570.

Anonymous:1994:EBb

[510] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 18(1):v–vi, August 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947394901287.

Anonymous:1994:P

[511] Anonymous. Preface. Computational Statistics & Data Analysis, 18 (1):1–2, August 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947394901295.

Coppi:1994:IMD

[512] Renato Coppi. An introduction to multiway data and their analysis. Computational Statistics & Data Analysis, 18(1):3–13, August 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947394901309.

Bove:1994:UOO

[513] Giuseppe Bove and Agostino Di Ciaccio. A user-oriented overview of multiway methods and software. Computational Statistics & Data Analysis, 18(1):15–37, August 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394901317.

Harshman:1994:PPF

[514] Richard A. Harshman and Margaret E. Lundy. PARAFAC: Parallel factor analysis. Computational Statistics & Data Analysis, 18(1): 39–72, August 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394901325.

Kroonenberg:1994:TLS

[515] Pieter M. Kroonenberg. The TUCK-ALS line: a suite of programs for three-way data analysis. Computational Statistics & Data Analysis, 18 (1):73–96, August 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394901333.

Lavit:1994:ASM

[516] Christine Lavit, Yves Escoufier, Robert Sabatier, and Pierre Traissac. The ACT (STATIS method). Computational Statistics & Data Analysis, 18 (1):97–119, August 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394901341.

Escofier:1994:MFA

[517] B. Escofier and J. Pagès. Multiple factor analysis (AFMULT package). Computational Statistics & Data Analysis, 18(1):121–140, August 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739490135X.

vanderBurg:1994:ONC

[518] Eeke van der Burg, Jan de Leeuw, and Garmt Dijksterhuis. OVER-ALS: Nonlinear canonical correlation with k sets of variables. Computational Statistics & Data Analysis, 18 (1):141–163, August 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947394901368.

DeSarbo:1994:CSM

[519] Wayne S. DeSarbo, Robert Libby, and Kamel Jedidi. CATSCALE: a stochastic multidimensional scaling methodology for the spatial analysis of sorting data and the study of stimulus categorization. Computational Statistics & Data Analysis, 18 (1):165–184, August 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947394901376.

Bouvier:1994:NNL

[520] Annie Bouvier and Sylvie Huet. nls2: Non-linear regression by S-PLUS — functions. Computational Statistics & Data Analysis, 18(1):187– 190, August 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947394901384.

Anonymous:1994:NASc

[521] Anonymous. News about statistical software related systems. Computational Statistics & Data Analysis, 18 (1):191–195, August 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394901392.

Anonymous:1994:Ic

[522] Anonymous. Information. Computational Statistics & Data Analysis, 18 (1):195–197, August 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394901406.

Anonymous:1994:LEC

[523] Anonymous. Letter to the editor — correction. Computational Statistics & Data Analysis, 18(1): 197, August 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394901414.

Anonymous:1994:CMe

[524] Anonymous. Calendar of meetings. Computational Statistics & Data Analysis, 18(1):198–200, August 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/0167947394901422.

Anonymous:1994:CWSd

[525] Anonymous. Courses, workshops & summer schools. Computational Statistics & Data Analysis, 18(1):

200, August 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394901430.

Azen:1994:E

[526] Stanley Azen. Editorial. Computational Statistics & Data Analysis, 18 (2):201, September 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394901716.

Balakrishnan:1994:OSN

[527] N. Balakrishnan. Order statistics from non-identical exponential random variables and some applications. Computational Statistics & Data Analysis, 18 (2):203–238, September 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394901724.

Balasubramanian:1994:D

[528] K. Balasubramanian, U. Gather, and N. Sugiura. Discussions. Computational Statistics & Data Analysis, 18 (2):238–243, September 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394901732.

Arnold:1994:WLM

[529] Barry C. Arnold. Where’s the lack of memory? Computational Statistics & Data Analysis, 18(2):243– 253, September 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394901740.

Berkane:1994:PML

[530] Maia Berkane, Yutaka Kano, and Peter M. Bentler. Pseudo maximum likelihood estimation in elliptical theory: Effects of misspecification. Computational Statistics & Data Analysis, 18 (2):255–267, September 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394901759.

Sawitzki:1994:TNR

[531] Günther Sawitzki. Testing numerical reliability of data analysis systems. Computational Statistics & Data Analysis, 18(2):269– 286, September 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394901767.

Sawitzki:1994:RNR

[532] Günther Sawitzki. Report on the numerical reliability of data analysis systems. Computational Statistics & Data Analysis, 18(2):289– 301, September 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394901775.

Wurtz:1994:I

[533] Diethelm Würtz and Myra Woolf. Information. Computational Statistics & Data Analysis, 18(2):302– 303, September 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394901783.

Tilke:1994:BRB

[534] Clemens Tilke. Book review: Geostatistics Troia ’92: Kluwer Academic, ISBN 0-7923-2157-X (set), ISBN 0-7923-2155-3 (Vol. 1) ISBN 0-7923-2156-1 (Vol. 2), $332.50/£222.50 (set). Computational Statistics & Data Analysis, 18(2):303, September 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/0167947394901805.

Trenkler:1994:BRBc

[535] Götz Trenkler. Book review: Alternative Methods of Regression: New York: John Wiley & Sons, Inc., 223 pages, ISBN 0-471-56881-3, US$57.50. Computational Statistics & Data Analysis, 18(2):303, September 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394901791.

Bismark:1994:BRB

[536] Malte Bismark. Book review: Software: Recht, Band 1. Berlin, Heidelberg, New York: Springer Verlag, ISBN 3-450-52493-3, DM 98.00. Computational Statistics & Data Analysis, 18 (2):303–304, September 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394901813.

Anonymous:1994:INd

[537] Anonymous. IASC news. Computational Statistics & Data Analysis, 18 (2):304, September 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394901821.

Delicado:1994:BGL

[538] Pedro Delicado and Manuel del Río. Bootstrapping the general linear hypothesis test. Computational Statistics & Data Analysis, 18(3):305– 316, October 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394900655.

Koronacki:1994:EDF

[539] J. Koronacki and U. Luboińska. Estimating the density of a functional of several random variables. Computational Statistics & Data Analysis, 18 (3):317–330, October 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394900663.

Kiers:1994:GCA

[540] Henk A. L. Kiers, Robert Cléroux, and Jos M. F. Ten Berge. Generalized canonical analysis based on optimizing matrix correlations and a relation with IDIOSCAL. Computational Statistics & Data Analysis, 18 (3):331–340, October 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394900671.

Silvers:1994:PSA

[541] Abraham Silvers, Nira Herrmann, Kathy Godfrey, Bruce Roberts, and Dan Cerys. A prototype statistical advisory system for biomedical researchers i: Overview. Computational Statistics & Data Analysis, 18 (3):341–355, October 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739490068X.

Herrmann:1994:PSA

[542] Nira Herrmann, Abraham Silvers, Katherine Godfrey, Bruce Roberts, and Daniel Cerys. A prototype statistical advisory system for biomedical researchers II: Development of a statistical strategy. Computational Statistics & Data Analysis, 18(3):357– 369, October 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947394900698.

Anonymous:1994:SSNb

[543] Anonymous. Statistical software newsletter. Computational Statistics & Data Analysis, 18(3):371, October 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/0167947394900701.

Haag:1994:KBS

[544] Uwe Haag. Knowledge-based systems in statistics: critical review of a decade of research. Computational Statistics & Data Analysis, 18 (3):373–375, October 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739490071X.

Hormann:1994:NAS

[545] Allmut Hörmann. News about statistical software & related systems: Software news editor. Computational Statistics & Data Analysis, 18(3):376– 378, October 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394900728.

Hormann:1994:IIE

[546] Allmut Hörmann. Information: Information editor. Computational Statistics & Data Analysis, 18(3):378– 380, October 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394900736.

Edler:1994:INI

[547] Lutz Edler. IASC news: IASC news editor. Computational Statistics & Data Analysis, 18(3):380– 382, October 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394900744.

Anonymous:1994:CMf

[548] Anonymous. Calendar of meetings. Computational Statistics & Data Analysis, 18(3):383–387, October 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/0167947394900752.

Anonymous:1994:CWSe

[549] Anonymous. Courses, workshops & summer schools. Computational Statistics & Data Analysis, 18(3):387– 388, October 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394900760.

Qian:1994:SIR

[550] Shixian Qian. The structure of isotonic regression class for LAD problems with quasi-order constraints. Computational Statistics & Data Analysis, 18 (4):389–401, November 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394901589.

Jolayemi:1994:CIB

[551] Joel K. Jolayemi. Convolution of independent binomial variables: an approximation method and a comparative study. Computational Statistics & Data Analysis, 18(4):403– 417, November 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394901597.

Kafadar:1994:CAT

[552] Karen Kafadar. Choosing among two-dimensional smoothers in practice. Computational Statistics & Data Analysis, 18(4):419–439, November 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/0167947394901600.

Wong:1994:CRP

[553] Weng Kee Wong. Comparing robust properties of A, D, E and G-optimal designs. Computational Statistics & Data Analysis, 18(4):441– 448, November 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394901619.

Ding:1994:CNB

[554] Cherng G. Ding. On the computation of the noncentral beta distribution. Computational Statistics & Data Analysis, 18(4):449– 455, November 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394901627.

Verboon:1994:RLR

[555] Peter Verboon and Willem J. Heiser. Resistant lower rank approximation of matrices by iterative majorization. Computational Statistics & Data Analysis, 18(4):457– 467, November 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394901635.

deWit:1994:IEC

[556] Toine de Wit. Interuniversity expertise center ProGAMMA. Computational Statistics & Data Analysis, 18 (4):471–472, November 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394901643.

Kubler:1994:BRB

[557] Jürgen Kübler. Book review: Parameter estimation in reliability and life span models: A. Clifford Cohen and Betty Jones Whitten (1988): New York, Marcel Dekker, 312 pages, ISBN 0-8247-7980-0, US$99.75. Computational Statistics & Data Analysis, 18 (4):473–474, November 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394901651.

Pabst:1994:BRBb

[558] Günther Pabst. Book review: Deskriptive statistik: P.v.d. Lippe (1993): 468 pp. Stuttgart: Gustav Fischer Verlag, ISBN 3-437-40268-4, Uni–Taschenbücher für Wissenschaft No. 1632, ISBN 3-8252-1632-2, 46.80 DM. Computational Statistics & Data Analysis, 18(4): 474, November 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394901678.

Schmerling:1994:BRBc

[559] Siegfried Schmerling. Book review: Numerical methods for stochastic processes: N. Bouleau and D. Lépingle: John Wiley and Sons, Inc., New York 1994, 359 p. ISBN 0-471-54641-0, $63.50. Computational Statistics & Data Analysis, 18(4): 474, November 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739490166X.

Dutter:1994:RCV

[560] Rudolf Dutter. Report on COMP-STAT ’94 in Vienna. Computational Statistics & Data Analysis, 18(4):475– 479, November 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394901686.

Anonymous:1994:CMg

[561] Anonymous. Calendar of meetings. Computational Statistics & Data Analysis, 18(4):480–483, November 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/0167947394901694.

Anonymous:1994:CWSf

[562] Anonymous. Courses, workshops & summer schools. Computational Statistics & Data Analysis, 18(4):483– 484, November 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394901708.

Bekker:1994:CRI

[563] Paul A. Bekker. Counting rules for identification in linear structural models. Computational Statistics & Data Analysis, 18(5):485– 498, December 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394900795.

Fischer:1994:TM

[564] N. I. Fischer, E. Mammen, and J. S. Marron. Testing for multimodality. Computational Statistics & Data Analysis, 18(5):499– 512, December 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394900809.

Kossler:1994:RAT

[565] Wolfgang Kössler. Restrictive adaptive tests for the treatment of the two-sample scale problem. Computational Statistics & Data Analysis, 18 (5):513–524, December 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394900817.

Kundu:1994:EPC

[566] Debasis Kundu. Estimating the parameters of complex-valued exponential signals. Computational Statistics & Data Analysis, 18(5):525– 534, December 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394900825.

Tsui:1994:AUB

[567] Kwok-Leung Tsui. Avoiding unnecessary bias in robust design analysis. Computational Statistics & Data Analysis, 18(5):535– 546, December 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394900833.

Tubert-Bitter:1994:LRT

[568] Pascale Tubert-Bitter, Andrew Kramar, Jean-Jacques Chalé, and Thierry Moreau. Linear rank tests for comparing survival in two groups with crossing hazards. Computational Statistics & Data Analysis, 18(5):547– 559, December 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394900841.

Gordon:1994:IGC

[569] A. D. Gordon. Identifying genuine clusters in a classification. Computational Statistics & Data Analysis, 18 (5):561–581, December 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739490085X.

Ho:1994:NAS

[570] Allmut Hö. News about statistical software & related systems. Computational Statistics & Data Analysis, 18(5):585, December 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394900868.

Edler:1994:INb

[571] Lutz Edler. IASC news. Computational Statistics & Data Analysis, 18 (5):586–590, December 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394900876.

Anonymous:1994:CMh

[572] Anonymous. Calendar of meetings. Computational Statistics & Data Analysis, 18(5):591–594, December 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/0167947394900884.

Anonymous:1994:CWSg

[573] Anonymous. Courses, workshops & summer schools. Computational Statistics & Data Analysis, 18(5): 594, December 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947394900892.

Anonymous:1994:AIS

[574] Anonymous. Author index of SS-NinCSDA: volume 17 & 18 (1994). Computational Statistics & Data Analysis, 18(5):595, December 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/0167947394900906.

Anonymous:1994:BRS

[575] Anonymous. Book reviews of SS-NinCSDA: volume 17 & 18 (1994). Computational Statistics & Data Analysis, 18(5):596, December 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/0167947394900914.

Anonymous:1994:SNS

[576] Anonymous. Software news of SS-NinCSDA: volume 17 & 18 (1994).

Computational Statistics & Data Analysis, 18(5):597, December 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/0167947394900922.

Anonymous:1994:AIVb

[577] Anonymous. Author index volume 18 (1994). Computational Statistics & Data Analysis, 18(5):599– 600, December 1994. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394900930.

Azen:1995:E

[578] Stanley Azen. Editorial. Computational Statistics & Data Analysis, 19(1):i–ii, January 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395901493.

Anonymous:1995:EBa

[579] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 19(1):V–VI, January 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395901477.

Moon:1995:PME

[580] Myung-Sang Moon and Richard F. Gunst. Polynomial measurement error modeling. Computational Statistics & Data Analysis, 19(1): 1–21, January 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947393E00412.

Liu:1995:SIG

[581] Lon-Mu Liu, Ki-Kan Chan, Alan L. Montgomery, and Mervin E. Muller. A system-independent graphical user interface for statistical software. Computational Statistics & Data Analysis, 19(1):23–44, January 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947393E0040B.

Yang:1995:ANC

[582] Ziqiang Yang. An algorithm for nonlinear L_{1} curve-fitting based on the smooth approximation. Computational Statistics & Data Analysis, 19 (1):45–52, January 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947393E00397.

Farebrother:1995:HCF

[583] Richard William Farebrother. On hyperplanes of closest fit. Computational Statistics & Data Analysis, 19 (1):53–58, January 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947393E00386.

Cwik:1995:NRD

[584] Jan Čwik and Jan Mielniczuk. Nonparametric rank discrimination method. Computational Statistics & Data Analysis, 19(1):59– 74, January 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947393E00434.

Huda:1995:ODR

[585] S. Huda, I. H. Khan, and Rahul Mukerjee. On optimal designs with restricted circular string property. Computational Statistics & Data Analysis, 19(1):75–83, January 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947393E00423.

Mansouri:1995:CSS

[586] H. Mansouri and G.-H. Chang. A comparative study of some rank tests for interaction. Computational Statistics & Data Analysis, 19(1): 85–96, January 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947393E00456.

Anonymous:1995:Ta

[587] Anonymous. Telegram. Computational Statistics & Data Analysis, 19 (1):99, January 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395901795.

Anonymous:1995:TTU

[588] Anonymous. Tips & tricks for using today’s statistical software systems efficiently. Computational Statistics & Data Analysis, 19(1):99– 101, January 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395901515.

Trenkler:1995:BRBa

[589] Dietrich Trenkler. Book review: A handbook of small data sets: Hand, D. J., Daly, F., Lunn, A. D., Mc-Conway, K. J. & Ostrowski, E. (1994): Chapman & Hall, London. xvi + 458 pages, including one diskette with data files (MS-DOS), 40 British Pounds, ISBN 0-412-39920-2. Computational Statistics & Data Analysis, 19(1): 101, January 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395901531.

Holle:1995:BRB

[590] Rolf Holle. Book review: Design of experiments. A no-name approach: Lorenzen, T. J. & Anderson, V. L. (1993): Marcel Dekker, New York. 414 pages, US$59.75, ISBN 0-8247-9077-4. Computational Statistics & Data Analysis, 19(1):101– 102, January 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395901558.

Anonymous:1995:BYLa

[591] Anonymous. From the bookstore for your library. Computational Statistics & Data Analysis, 19(1): 102, January 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395901817.

Anonymous:1995:INa

[592] Anonymous. IASC-News. Computational Statistics & Data Analysis, 19 (1):102–103, January 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395901574.

Anonymous:1995:CMa

[593] Anonymous. Calendar of meetings. Computational Statistics & Data Analysis, 19(1):103–110, January 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/0167947395901590.

Antoch:1995:TH

[594] Jaromír Antoch, Petr Hájek, and Radim Jiroušek. Tomáš Havránek. Computational Statistics & Data Analysis, 19(2):111–113, February 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/0167947395900004.

Antoch:1995:RRR

[595] Jaromír Antoch and Hakan Ekblom. Recursive robust regression computational aspects and comparison. Computational Statistics & Data Analysis, 19 (2):115–128, February 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947393E0050E.

Bocek:1995:LPA

[596] P. Boček and P. Lachout. Linear programming approach to LMSestimation. Computational Statistics & Data Analysis, 19(2):129– 134, February 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947393E00515.

Gale:1995:CCF

[597] William A. Gale. Cognition, computation, and formal systems: Some of Tomás Havránek’s interests and disambiguating word senses. Computational Statistics & Data Analysis, 19 (2):135–148, February 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947393E00526.

Hajek:1995:GPC

[598] P. Hájek, A. Sochorová, and J. Zvárová. GUHA for personal computers. Computational Statistics & Data Analysis, 19(2):149–153, February 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/0167947393E00537.

Hojsgaard:1995:BBR

[599] Søren Højsgaard and Bo Thiesson. BIFROST — block recursive models induced from relevant knowledge, observations, and statistical techniques. Computational Statistics & Data Analysis, 19(2):155– 175, February 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947393E00548.

Jirousek:1995:EII

[600] Radim Jiroušek and Stanislav Přeučil. On the effective implementation of the iterative proportional fitting procedure. Computational Statistics & Data Analysis, 19(2):177– 189, February 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947393E00559.

Lauritzen:1995:EAG

[601] Steffen L. Lauritzen. The EM algorithm for graphical association models with missing data. Computational Statistics & Data Analysis, 19 (2):191–201, February 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947393E0056A.

Rizzi:1995:RSV

[602] Alfredo Rizzi and Maurizio Vichi. Representation, synthesis, variability and data preprocessing of a threeway data set. Computational Statistics & Data Analysis, 19(2):203– 222, February 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947393E0057B.

Wermuth:1995:DVC

[603] Nanny Wermuth and D. R. Cox. Derived variables calculated from similar joint responses: some characteristics and examples. Computational Statistics & Data Analysis, 19 (2):223–234, February 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947393E0058C.

Anonymous:1995:BRB

[604] Anonymous. Book review: The Statistical Software Guide ’94/95: edited by Armin Koch and Uwe Haag. Computational Statistics & Data Analysis, 19 (2):237–261, February 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947395900020.

Anonymous:1995:INb

[605] Anonymous. IASC-News. Computational Statistics & Data Analysis, 19 (2):262–263, February 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947395900047.

Anonymous:1995:CTD

[606] Anonymous. Calendar: topical deadlines only. Computational Statistics & Data Analysis, 19(2):263– 264, February 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947395900063.

Hossjer:1995:ECL

[607] Ola Hössjer. Exact computation of the least trimmed squares estimate in simple linear regression. Computational Statistics & Data Analysis, 19 (3):265–282, March 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739592697V.

Drezner:1995:MCA

[608] Zvi Drezner. Multirelation — a correlation among more than two variables. Computational Statistics & Data Analysis, 19(3):283– 292, March 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947393E00467.

Ebneshahrashoob:1995:DAI

[609] M. Ebneshahrashoob and Milton Sobel. Dirichlet analysis for inverse multinomial with both quota and quota-free cells. Computational Statistics & Data Analysis, 19(3):293– 307, March 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394000043.

Chen:1995:GLN

[610] Gemai Chen. Generalized lognormal distributions with reliability application. Computational Statistics & Data Analysis, 19(3):309– 319, March 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947393E00478.

Fung:1995:CNU

[611] Wing K. Fung. A cautionary note on the use of generalized Cooktype measures. Computational Statistics & Data Analysis, 19(3):321– 326, March 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947393E0024X.

Smith:1995:IRP

[612] D. M. Smith. Iterative refinement of parameter estimates and residuals from a regression model fitted using QR decomposition methods. Computational Statistics & Data Analysis, 19(3):327–349, March 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947393E00489.

Schuster:1995:PET

[613] Ernst Schuster. Preference of equivalence tests with standardized mean difference demonstrated by an application. Computational Statistics & Data Analysis, 19(3):353– 356, March 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739592698W.

Anonymous:1995:Tb

[614] Anonymous. Telegram. Computational Statistics & Data Analysis, 19 (3):357, March 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947395900055.

Anonymous:1995:TTa

[615] Anonymous. Tips & tricks. Computational Statistics & Data Analysis, 19(3):357–358, March 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395900071.

Uebe:1995:BRB

[616] Götz Uebe. Book review: Computational techniques for Econometrics and Economic Analysis: D. A. Belsley (ed.) (1994) Dordrecht: Kluwer Academic Publishers. ISBN 0-7923-2356-4, pp. 248, US$92. Computational Statistics & Data Analysis, 19 (3):359–360, March 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395900098.

Trenkler:1995:BRBb

[617] Dietrich Trenkler. Book review: Ökonometrie: Hans Schneeweiß (1990) (4th ed.), Heidelberg: Physica-Verlag. ISBN 3-7908-0486-X, pp. 392, -DM45.-. Computational Statistics & Data Analysis, 19(3):360, March 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739590011X.

Weiberger:1995:BRB

[618] Volker Weiberger. Book review: UNIX: Einstieg für DOS-Anwender: Taudes, Alfred, Reepmeyer, Jan-Armin, & Bensberg, Frank (1993) Heidelberg: Physica-Verlag. ISBN 3-7908-0700-1, pp 232, DM40.-. Computational Statistics & Data Analysis, 19(3): 360–361, March 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947395900136.

Anonymous:1995:INc

[619] Anonymous. IASC-News. Computational Statistics & Data Analysis, 19 (3):361–364, March 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947395900152.

Anonymous:1995:UQI

[620] Anonymous. Urgent: Questionnaire for IASC members. Computational Statistics & Data Analysis, 19(3): 365, March 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947395900179.

Anonymous:1995:CMNa

[621] Anonymous. Calendar of meetings — new entries only. Computational Statistics & Data Analysis, 19(3): 366–367, March 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947395900195.

Anonymous:1995:CWSa

[622] Anonymous. Courses, workshops & summer schools. Computational Statistics & Data Analysis, 19(3): 367–368, March 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739590025X.

Kontoghiorghes:1995:AAN

[623] E. J. Kontoghiorghes and M. R. B. Clarke. An alternative approach for the numerical solution of seemingly unrelated regression equations models. Computational Statistics & Data Analysis, 19(4):369– 377, April 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739400010G.

Onishi:1995:UKB

[624] Haruo Onishi. A user-knowledge-based variable selection method for limited information maximum likelihood using principal components. Computational Statistics & Data Analysis, 19(4):379–399, April 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394000076.

Kim:1995:DPD

[625] S. S. Kim and S. H. Park. Dynamic plots for displaying the roles of variables and observations in regression model. Computational Statistics & Data Analysis, 19(4):401– 418, April 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394000065.

Krzanowski:1995:SVA

[626] W. J. Krzanowski. Selection of variables, and assessment of their performance, in mixed-variable discriminant analysis. Computational Statistics & Data Analysis, 19(4): 419–431, April 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947394000117.

Wu:1995:MFF

[627] Berlin Wu. Model-free forecasting for nonlinear time series (with application to exchange rates). Computational Statistics & Data Analysis, 19(4):433–459, April 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947394000087.

Zornig:1995:URZ

[628] Peter Zörnig and Gabriel Altmann. Unified representation of Zipf distributions. Computational Statistics & Data Analysis, 19(4):461– 473, April 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947394000098.

Anonymous:1995:Tc

[629] Anonymous. Telegram. Computational Statistics & Data Analysis, 19 (4):477, April 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947395900896.

Anonymous:1995:TTb

[630] Anonymous. Tips & tricks. Computational Statistics & Data Analysis, 19(4):477–478, April 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395900918.

Lenz:1995:BRBa

[631] H.-J. Lenz. Book review: Statistical and scientific databases: Zbigniew Michalewicz (ed.) (1991): New York, London, Toronto, Sydney, Tokyo, Singapore: Ellis Horwood. 544 pages, ISBN 0-13-850652-3, $57.20. Computational Statistics & Data Analysis, 19(4):479–481, April 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395900934.

Lenz:1995:BRBb

[632] H.-J. Lenz. Book review: Panel data analysis: Baldev Raj and Badi H. Baltagi (eds.) (1992): Heidelberg: Physica-Verlag. ISBN 3-7908-0593-9, 220 pages, DM 128.00. Computational Statistics & Data Analysis, 19 (4):481–483, April 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395900950.

Trenkler:1995:BRBc

[633] Götz Trenkler. Book review: Statistical distributions: M. Evans, N. Hastings & B. Peacock (1993): (2nd Edition). New: John Wiley. 170 pages, ISBN 0-471-55951-2, £24.95. Computational Statistics & Data Analysis, 19(4):483–484, April 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395900977.

Anonymous:1995:BYLb

[634] Anonymous. From the bookstore for your library. Computational Statistics & Data Analysis, 19(4): 484, April 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947395900993.

Anonymous:1995:INd

[635] Anonymous. IASC-News. Computational Statistics & Data Analysis, 19(4):484–485, April 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947395901019.

Anonymous:1995:CMb

[636] Anonymous. Calendar of meetings. Computational Statistics & Data Analysis, 19(4):485–490, April 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/0167947395901035.

Anonymous:1995:CWSb

[637] Anonymous. Courses, workshops & summer schools. Computational Statistics & Data Analysis, 19(4): 490–492, April 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947395901051.

Froeschl:1995:FME

[638] Karl Anton Froeschl. A formal model evaluation approach to the analysis of treatment effects in paired sample data. Computational Statistics & Data Analysis, 19(5):493– 517, May 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394000128.

Hawkins:1995:CFS

[639] Douglas M. Hawkins. Convergence of the feasible solution algorithm for least median of squares regression. Computational Statistics & Data Analysis, 19(5):519–538, May 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739400017D.

duPlessis:1995:BAM

[640] J. L. du Plessis and A. J. van der Merwe. A Bayesian approach to multivariate and conditional calibration. Computational Statistics & Data Analysis, 19(5):539– 552, May 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739400014A.

Panda:1995:FDD

[641] S. K. Panda and A. Nagabhushanam. Fuzzy data distortion. Computational Statistics & Data Analysis, 19 (5):553–562, May 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739400015B.

Shore:1995:FDF

[642] Haim Shore. Fitting a distribution by the first two moments (partial and complete). Computational Statistics & Data Analysis, 19(5): 563–577, May 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739400016C.

Andres:1995:FET

[643] A. Martín Andrés and I. Herranz Tejedor. Is Fisher’s exact test very conservative? Computational Statistics & Data Analysis, 19(5): 579–591, May 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947394000139.

Krause:1995:ESS

[644] Andreas Krause. Electronic services in statistics. Computational Statistics & Data Analysis, 19(5): 595–604, May 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739590039X.

Lenz:1995:BRBc

[645] H.-J. Lenz. Book review: Reliability improvement with design of experiments: Lloyd W. Condra (1993) New York: Marcel Dekker. 400 pages, ISBN 0-8247-8888-5, $65.00. Computational Statistics & Data Analysis, 19(5):605, May 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395900411.

Murphy:1995:RMW

[646] Brian P. Murphy. Reflections on meetings: Whither COMPSTAT? Issues & challenges on occasion of the panel discussion of COMPSTAT ’94, Vienna, August, 25, 1994. Computational Statistics & Data Analysis, 19(5):606–609, May 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395900438.

Anonymous:1995:INe

[647] Anonymous. IASC-News. Computational Statistics & Data Analysis, 19(5):609–610, May 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395900454.

Anonymous:1995:CMNb

[648] Anonymous. Calendar of meetings — new entries only. Computational Statistics & Data Analysis, 19(5): 610–612, May 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395900470.

Anonymous:1995:CWSc

[649] Anonymous. Courses, workshops & summer schools. Computational Statistics & Data Analysis, 19(5): 612, May 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947395900497.

Bosch:1995:CAQ

[650] Ronald J. Bosch, Yinyu Ye, and George G. Woodworth. A convergent algorithm for quantile regression with smoothing splines. Computational Statistics & Data Analysis, 19(6):613–630, June 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739400018E.

Ocana:1995:VRB

[651] Jordi Ocaña and Esteban Vegas. Variance reduction for Bernoulli response variables in simulation. Computational Statistics & Data Analysis, 19 (6):631–640, June 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739400023C.

Hardle:1995:TID

[652] W. Härdle and B. U. Park. Testing increasing dispersion. Computational Statistics & Data Analysis, 19(6):641–653, June 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739400021A.

Jordan:1995:ETL

[653] Paul Jordan. Estimation of tolerance limits from reference data. Computational Statistics & Data Analysis, 19(6):655–668, June 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739400028H.

Ferre:1995:SCP

[654] Louis Ferré. Selection of components in principal component analysis: a comparison of methods. Computational Statistics & Data Analysis, 19(6):669–682, June 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739400020J.

Franses:1995:ESA

[655] Philip Hans Franses. The effects of seasonally adjusting a periodic autoregressive process. Computational Statistics & Data Analysis, 19(6): 683–704, June 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739400019F.

Kulmann:1995:NCE

[656] Hermann Kulmann. Notes on the computation of the exact distribution function of the χ^{2}- and related test statistics in the equiprobable case. Computational Statistics & Data Analysis, 19(6):707–710, June 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739598600B.

Lenz:1995:BRBd

[657] H.-J. Lenz. Book review: New directions in statistical data analysis and Robustness: S. Morgenthaler, E. Ronchetti, W. A. Stahel (eds.) (1993) Basel, Boston, London: Birkhäuser Verlag. 293 pages, ISBN 3-7643-2923-8, £39. Computational Statistics & Data Analysis, 19(6):711–712, June 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395901159.

Lenz:1995:BRBe

[658] H.-J. Lenz. Book review: NeXTSTEP programming, step one: object-oriented applications: Simson L. Garfinkel & Michael K. Mahoney (1993): New York: Springer Verlag. 631 pp. + disk, DM 89.- ISBN 3-540-97884-4. Computational Statistics & Data Analysis, 19(6):712–713, June 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395901175.

Anonymous:1995:INf

[659] Anonymous. IASC-News. Computational Statistics & Data Analysis, 19(6):713–714, June 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395901191.

Anonymous:1995:CMNc

[660] Anonymous. Calendar of meetings — new entries only. Computational Statistics & Data Analysis, 19(6): 714–716, June 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395901213.

Anonymous:1995:CWSd

[661] Anonymous. Courses, workshops & summer schools. Computational Statistics & Data Analysis, 19(6): 716, June 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739590123X.

Anonymous:1995:AIVa

[662] Anonymous. Author index volume 19 (1995). Computational Statistics & Data Analysis, 19(6):717– 719, June 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395901256.

Anonymous:1995:EBb

[663] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 20(1):v–vi, July 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395901698.

Hardle:1995:FSS

[664] W. Härdle and J. S. Marron. Fast and simple scatterplot smoothing. Computational Statistics & Data Analysis, 20(1):1–17, July 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739400031D.

Ritov:1995:PAC

[665] Ya’acov Ritov. PM algorithms for calculating minimum χ^{2} estimators with partially observed sub-tables. Computational Statistics & Data Analysis, 20(1):19–33, July 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739400027G.

Liu:1995:FWS

[666] W. Liu. Fixed-width simultaneous confidence intervals for allpairwise comparisons. Computational Statistics & Data Analysis, 20(1): 35–44, July 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739400032E.

Bello:1995:ITR

[667] A. L. Bello. Imputation techniques in regression analysis: Looking closely at their implementation. Computational Statistics & Data Analysis, 20 (1):45–57, July 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739400024D.

Jovanovic:1995:ESM

[668] Borko D. Jovanovic, David W. Hosmer, and John P. Buonaccorsi. Equivalence of several methods for efficient best subsets selection in generalized linear models. Computational Statistics & Data Analysis, 20 (1):59–64, July 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739400030M.

Wang:1995:OSS

[669] Suojin Wang. One-step saddlepoint approximations for quantiles. Computational Statistics & Data Analysis, 20(1):65–74, July 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739400029I.

Zaki:1995:LII

[670] M. Zaki, F. El Tohamy, and A. Hassan. Landsat image interpretation. Computational Statistics & Data Analysis, 20(1):75–97, July 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739400022B.

Rahlfs:1995:NMS

[671] Volker W. Rahlfs and Helmuth Zimmermann. Nonparametric methods in standard statistics packages: the case of the Friedman analysis. Computational Statistics & Data Analysis, 20(1):101–110, July 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739590171X.

Anonymous:1995:TTc

[672] Anonymous. Tips & tricks. Computational Statistics & Data Analysis, 20(1):111, July 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395901736.

Anonymous:1995:INg

[673] Anonymous. IASC-News. Computational Statistics & Data Analysis, 20(1):112–114, July 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947395901752.

Anonymous:1995:CMc

[674] Anonymous. Calendar of meetings. Computational Statistics & Data Analysis, 20(1):114–118, July 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/0167947395901779.

Dielman:1995:BAH

[675] Terry E. Dielman and Elizabeth L. Rose. A bootstrap approach to hypothesis testing in least absolute value regression. Computational Statistics & Data Analysis, 20(2):119– 130, August 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739400037J.

Edwards:1995:EVB

[676] Lloyd J. Edwards and Scott A. Hamilton. Errors-in-variables and the Box– Cox transformation. Computational Statistics & Data Analysis, 20(2):131– 140, August 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739400039L.

Polzehl:1995:PPD

[677] Jörg Polzehl. Projection pursuit discriminant analysis. Computational Statistics & Data Analysis, 20(2):141– 157, August 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739400035H.

Dyer:1995:BHI

[678] Danny Dyer and Rebecca L. Pierce. Bayesian HPM intervals for the number of nonconforming items in a lot. Computational Statistics & Data Analysis, 20(2):159–172, August 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739400038K.

Lei:1995:THN

[679] Xingye Lei, Y. B. Peng, and F. T. Wright. Testing homogeneity of normal means with a simply ordered alternative and dependent observations. Computational Statistics & Data Analysis, 20(2):173–183, August 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739400033F.

Schroer:1995:ERD

[680] Gunar Schröer and Dietrich Trenkler. Exact and randomization distributions of Kolmogorov–Smirnov tests two or three samples. Computational Statistics & Data Analysis, 20 (2):185–202, August 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739400040P.

Granville:1995:BFS

[681] V. Granville and J. P. Rasson. Bayesian filtering and supervised classification in image remote sensing. Computational Statistics & Data Analysis, 20 (2):203–225, August 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739591132A.

Krivy:1995:CRS

[682] Ivan Křivý and Josef Tvrdík. The controlled random search algorithm in optimizing regression models. Computational Statistics & Data Analysis, 20 (2):229–234, August 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947395901272.

Kohne:1995:BRB

[683] Stephan Köhne and Iris Pigeot. Book review: Resampling-Based Multiple Testing. Examples and Methods for p-Value Adjustment: Peter H. Westfall and S. Stanley Young (1993): New York: Wiley, ISBN 0-471-55761-7, pp. 340, $59.95. Computational Statistics & Data Analysis, 20(2):235– 236, August 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947395901299.

Anonymous:1995:INh

[684] Anonymous. IASC-News. Computational Statistics & Data Analysis, 20 (2):236–237, August 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395901310.

Anonymous:1995:CMNd

[685] Anonymous. Calendar of meetings — new entries only. Computational Statistics & Data Analysis, 20(2):237– 238, August 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395901337.

Anonymous:1995:CWSe

[686] Anonymous. Courses, workshops & summer schools. Computational Statistics & Data Analysis, 20(2): 238, August 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395901353.

Grechanovsky:1995:CVS

[687] Eugene Grechanovsky and Ilia Pinsker. Conditional p-values for the F - statistic in a forward selection procedure. Computational Statistics & Data Analysis, 20(3):239–263, September 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739400047M.

Black:1995:EDR

[688] Stephen Black and H. Mansouri. On exact distributions of rank tests for ordered alternatives in block designs. Computational Statistics & Data Analysis, 20(3):265–274, September 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739400045K.

Sik-Yum:1995:EFS

[689] Lee Sik-Yum and Poon Wai-Yin. Estimation of factor scores in a two-level confirmatory factor analysis model. Computational Statistics & Data Analysis, 20(3):275– 284, September 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739400042H.

Madan:1995:DSC

[690] V. K. Madan and R. K. Sinha. Digital smoothing of census data employing Fourier transforms. Computational Statistics & Data Analysis, 20 (3):285–294, September 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739592815F.

Winker:1995:IMA

[691] Peter Winker. Identification of multivariate AR-models by threshold accepting. Computational Statistics & Data Analysis, 20(3):295– 307, September 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739400041G.

Oman:1995:CAM

[692] Samuel D. Oman. Checking the assumptions in mixed-model analysis of variance: a residual analysis approach. Computational Statistics & Data Analysis, 20(3):309– 330, September 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739400043I.

Jarrett:1995:RQC

[693] Jeffrey Jarrett. Review of a quality control and improvement statistical software system. Computational Statistics & Data Analysis, 20(3):333– 339, September 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395902090.

Anonymous:1995:INi

[694] Anonymous. IASC-News. Computational Statistics & Data Analysis, 20 (3):340–342, September 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395902112.

Hirose:1995:MLP

[695] Hideo Hirose. Maximum likelihood parameter estimation in the threeparameter gamma distribution. Computational Statistics & Data Analysis, 20(4):343–354, October 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739400050S. See

[1125].

Bradu:1995:ATR

[696] Dan Bradu and Douglas M. Hawkins. An Anscombe type robust regression statistic. Computational Statistics & Data Analysis, 20(4):355– 386, October 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739400048N.

Saltelli:1995:SAM

[697] A. Saltelli, T. H. Andres, and T. Homma. Sensitivity analysis of model output. performance of the iterated fractional factorial design method. Computational Statistics & Data Analysis, 20(4):387– 407, October 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739592843M.

Schmid:1995:DFT

[698] Friedrich Schmid and Mark Trede. A distribution free test for the two sample problem for general alternatives. Computational Statistics & Data Analysis, 20(4):409–419, October 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739592844N.

Castillo:1995:MEP

[699] Enrique Castillo and Ali S. Hadi. A method for estimating parameters and quantiles of distributions of continuous random variables. Computational Statistics & Data Analysis, 20 (4):421–439, October 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739400049O.

Rasch:1995:RAP

[700] Dieter Rasch. The robustness against parameter variation of exact locally optimum designs in nonlinear regression — a case study. Computational Statistics & Data Analysis, 20 (4):441–453, October 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739400053L.

vonHaeseler:1995:MEA

[701] Arndt von Haeseler and Michael Schöniger. Modelling evolution of autocorrelated sequences. Computational Statistics & Data Analysis, 20 (4):457–458, October 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739590137X.

Anonymous:1995:INj

[702] Anonymous. IASC-News. Computational Statistics & Data Analysis, 20 (4):459–460, October 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395901396.

Anonymous:1995:Aa

[703] Anonymous. Anzstat. Computational Statistics & Data Analysis, 20 (4):460–462, October 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947395901418.

Anonymous:1995:CMd

[704] Anonymous. Calendar of meetings. Computational Statistics & Data Analysis, 20(4):462–464, October 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/0167947395901434.

Anonymous:1995:CWSf

[705] Anonymous. Courses, workshops & summer schools. Computational Statistics & Data Analysis, 20(4):464– 465, October 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947395901450.

Carrizosa:1995:DLQ

[706] Emilio Carrizosa and Frank Plastria. The determination of a “least quantile of squares regression line” for all quantiles. Computational Statistics & Data Analysis, 20(5):467– 479, November 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739400059R.

Chattamvelli:1995:DND

[707] R. Chattamvelli. On the doubly noncentral F distribution. Computational Statistics & Data Analysis, 20 (5):481–489, November 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739400054M.

Elson:1995:UFF

[708] Paul J. Elson. Use of the Fast Fourier Transform algorithm in the calculation of the operating characteristics of group sequential clinical trials. Computational Statistics & Data Analysis, 20 (5):491–498, November 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739400055N.

Lumley:1995:EES

[709] Thomas Lumley. Efficient execution of Stone’s likelihood ratio tests for disease clustering. Computational Statistics & Data Analysis, 20(5):499– 510, November 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739400056O.

Luceno:1995:FPC

[710] Alberto Luceño. A family of partially correlated Poisson models for overdispersion. Computational Statistics & Data Analysis, 20(5):511– 520, November 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739400057P.

Manteiga:1995:TLR

[711] W. Gonzalez Manteiga and J. M. Vilar Fernandez. Testing linear regression models using non-parametric regression estimators when errors are nonindependent. Computational Statistics & Data Analysis, 20(5):521– 541, November 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739400058Q.

DeCesare:1995:STN

[712] L. De Cesare and D. Posa. A simulation technique of a non-Gaussian spatial process. Computational Statistics & Data Analysis, 20(5):543– 555, November 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739400061M.

Fung:1995:DIO

[713] Wing K. Fung. Detecting influential observations for estimated probabilities in multiple discriminant analysis. Computational Statistics & Data Analysis, 20(5):557– 568, November 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739400060V.

Hauser:1995:EFD

[714] Michael A. Hauser and Erhard Reschenhofer. Estimation of the fractionally differencing parameter with the R/S method. Computational Statistics & Data Analysis, 20(5):569– 579, November 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739400062N.

Krause:1995:FEU

[715] Ekkehard Krause, Christoph Bandt, Anne Schulz, and Holger Schulz. Fractal exponents for the upper airways of mammalian lungs. Computational Statistics & Data Analysis, 20(5):583– 590, November 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739592763N.

Held:1995:T

[716] Gerhard Held. Telegram. Computational Statistics & Data Analysis, 20 (5):591–592, November 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395901639.

Schlittgen:1995:BRB

[717] Rainer. Schlittgen. Book review: Survival analysis: State of the art: John P. Klein and Prem. K. Goel (1992) Dordrecht: Kluwer Academic Publishers, x + 451 pages, ISBN 0-7923-1634-7, $138.00. Computational Statistics & Data Analysis, 20(5):592– 593, November 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395901655.

Anonymous:1995:I

[718] Anonymous. Information. Computational Statistics & Data Analysis, 20 (5):593–598, November 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395901671.

Neytchev:1995:SOL

[719] Plamen Nikolov Neytchev. SWEEP operator for least-squares subject to linear constraints. Computational Statistics & Data Analysis, 20(6):599– 609, December 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947394000678.

Cao:1995:MDD

[720] Ricardo Cao, Antonio Cuevas, and Ricardo Fraiman. Minimum distance density-based estimation. Computational Statistics & Data Analysis, 20 (6):611–631, December 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947394000654.

Jolayemi:1995:PFM

[721] Joel K. Jolayemi. A power function model for determining sample sizes for the operations of multivariate control charts. Computational Statistics & Data Analysis, 20(6):633– 641, December 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947394000630.

Kazi-Aoual:1995:RAP

[722] Frédérique Kazi-Aoual, Simon Hitier, Robert Sabatier, and Jean-Dominique Lebreton. Refined approximations to permutation tests for multivariate inference. Computational Statistics & Data Analysis, 20(6):643– 656, December 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394000642.

Ofversten:1995:EMM

[723] Jukka Öfversten. Estimation in mixed models via layer triangular transformation. Computational Statistics & Data Analysis, 20(6):657– 667, December 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394000666.

Posse:1995:PPE

[724] Christian Posse. Projection pursuit exploratory data analysis. Computational Statistics & Data Analysis, 20 (6):669–687, December 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395000028.

Wang:1995:ABD

[725] P. C. Wang, M. H. Chen, and S. C. Wei. To analyze binary data from mixed-level orthogonal arrays. Computational Statistics & Data Analysis, 20 (6):689–695, December 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739400068X.

Knusel:1995:ASD

[726] Leo Knüsel. On the accuracy of the statistical distributions in GAUSS. Computational Statistics & Data Analysis, 20(6):699– 702, December 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947396800004.

Anonymous:1995:SN

[727] Anonymous. Software news. Computational Statistics & Data Analysis, 20(6):703, December 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947395901957.

Kramer:1995:BRB

[728] Walter Krämer. Book review: Probability & Measure: Patrick Billingsley (1995): (3rd ed.). New York: Wiley, ISBN 0-471-00710-2, pp. 593, £49.95. Computational Statistics & Data Analysis, 20(6): 703, December 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947395901973.

Anonymous:1995:INk

[729] Anonymous. IASC-News. Computational Statistics & Data Analysis, 20 (6):704–711, December 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739590199X.

Anonymous:1995:CMNe

[730] Anonymous. Calendar of meetings — new entries only. Computational Statistics & Data Analysis, 20(6): 712, December 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395902015.

Anonymous:1995:AIS

[731] Anonymous. Author index of SS-NinCSDA: Volume 19 (1995) – volume 20 (1995). Computational Statistics & Data Analysis, 20(6):713– 716, December 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395902031.

Anonymous:1995:Ab

[732] Anonymous. Acknowledgement. Computational Statistics & Data Analysis, 20(6):716, December 1995. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395902058.

Anonymous:1995:AIVb

[733] Anonymous. Author index volume 20 (1995). Computational Statistics & Data Analysis, 20(6):717– 719, December 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395902074.

Anonymous:1996:EBa

[734] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 21(1):v–vi, January 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947396900031.

Monti:1996:NPE

[735] Anna Clara Monti. A new preliminary estimator for MA(1) models. Computational Statistics & Data Analysis, 21(1):1–15, January 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394000085.

Gordon:1996:SCC

[736] A. D. Gordon. A survey of constrained classification. Computational Statistics & Data Analysis, 21(1): 17–29, January 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395000054.

Koo:1996:BBS

[737] Ja-Yong Koo. Bivariate B-splines for tensor logspline density estimation. Computational Statistics & Data Analysis, 21(1):31– 42, January 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395000038.

Mugglestone:1996:PGS

[738] Moira A. Mugglestone and Eric Renshaw. A practical guide to the spectral analysis of spatial point processes. Computational Statistics & Data Analysis, 21(1):43– 65, January 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395000070.

Rousseeuw:1996:RFR

[739] Peter J. Rousseeuw and Mia Hubert. Regression-free and robust estimation of scale for bivariate data. Computational Statistics & Data Analysis, 21(1):67–85, January 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395000046.

Almoy:1996:SSC

[740] Trygve Almøy. A simulation study on comparison of prediction methods when only a few components are relevant. Computational Statistics & Data Analysis, 21(1):87– 107, January 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395000062.

Scheuchenpflug:1996:CCU

[741] Thomas Scheuchenpflug and Maria Blettner. Coding confusion using PROC LOGISTIC in SAS^{(}R). Computational Statistics & Data Analysis, 21(1):111–115, January 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947396900055.

Anonymous:1996:Ta

[742] Anonymous. Telegram. Computational Statistics & Data Analysis, 21 (1):116, January 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947396900079.

Anonymous:1996:LEK

[743] Anonymous. Letter to the editor: Knut M. Wittkowski (Tübingen, Germany) wrote on August 22, 1995: Rahlfs, V. W. & Zimmermann, H.: “Nonparametric methods in standard statistics packages: the case of the Friedman analysis”, SSNinCSDA 20/1, 101–110. Computational Statistics & Data Analysis, 21(1):116–117, January 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947396900092. See

reply [744].

Anonymous:1996:RVW

[744] Anonymous. Reply by Volker W. Rahlfs and Helmuth Zimmermann (Munich, Germany): Comment on Knut M. Wittkowski, Letter to the Editor, August 22, 1995. Computational Statistics & Data Analysis, 21 (1):117–118, January 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947396900110. See

[743].

Trenkler:1996:BRBa

[745] Götz Trenkler. Book review: Advanced Linear Models: Song-Gui Wang & Shein-Chung Chow: Marcel Dekker, New York, 1994, pp. x + 537, $165.00, ISBN 0-8247-9169-X. Computational Statistics & Data Analysis, 21(1):118– 119, January 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947396900134.

Trenkler:1996:BRBb

[746] Götz Trenkler. Book review: Continuous univariate distributions: N. L. Johnson, S. Kotz and N. Balakrishnan Vol. 1, 2nd Edition. John Wiley, New York, 1994, pp. xix + 756, Price: £66.00, ISBN 0-471-58495-9. Computational Statistics & Data Analysis, 21(1):119, January 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947396900158.

Anonymous:1996:INa

[747] Anonymous. IASC-News. Computational Statistics & Data Analysis, 21 (1):119–123, January 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947396900171.

Anonymous:1996:CMa

[748] Anonymous. Calendar of meetings. Computational Statistics & Data Analysis, 21(1):123–130, January 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/0167947396900195.

Anonymous:1996:CWSa

[749] Anonymous. Courses, workshops & summer schools. Computational Statistics & Data Analysis, 21(1):130– 131, January 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947396900213.

Krijnen:1996:AUL

[750] Wim P. Krijnen. Algorithms for unweighted least-squares factor analysis. Computational Statistics & Data Analysis, 21(2):133– 147, February 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395000119.

Ruiz-Gazen:1996:VSR

[751] Anne Ruiz-Gazen. A very simple robust estimator of a dispersion matrix. Computational Statistics & Data Analysis, 21(2):149– 162, February 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395000097.

Steinwolf:1996:ASP

[752] Alexander Steinwolf. Approximation and simulation of probability distributions with a variable kurtosis value. Computational Statistics & Data Analysis, 21(2):163– 180, February 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395000100.

Jung:1996:CEO

[753] Joo Sung Jung and Bong Jin Yum. Construction of exact D-optimal designs by tabu search. Computational Statistics & Data Analysis, 21(2):181– 191, February 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395000143.

Kalaba:1996:MAM

[754] Robert Kalaba and Leigh Tesfatsion. A multicriteria approach to model specification and estimation. Computational Statistics & Data Analysis, 21 (2):193–214, February 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394000301.

Liu:1996:SSS

[755] W. Liu. On some single-stage, stepdown and step-up procedures for comparing three normal means. Computational Statistics & Data Analysis, 21 (2):215–227, February 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395000135.

Koch:1996:BSG

[756] A. Koch and U. Haag. The Statistical Software Guide ’95/96. Computational Statistics & Data Analysis, 21 (2):231–256, February 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947396903035.

Anonymous:1996:INb

[757] Anonymous. IASC-news. Computational Statistics & Data Analysis, 21 (2):257–260, February 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947396903059.

Wijsman:1996:NAF

[758] Robert A. Wijsman. New algorithms for the function T (h, a) of Owen, with application to bivariate normal and noncentral t-probabilities. Computational Statistics & Data Analysis, 21 (3):261–271, March 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947395000186.

Dette:1996:SAS

[759] Holger Dette and Stefan Sperlich. Some applications of Stieltjes transforms in the construction of optimal designs for nonlinear regression models. Computational Statistics & Data Analysis, 21 (3):273–292, March 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947395000151.

Kachlicka:1996:RRC

[760] Danuta Kachlicka and Stanislaw Mejza. Repeated row-column designs with split units. Computational Statistics & Data Analysis, 21 (3):293–305, March 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947395000208.

Lausen:1996:EEO

[761] Berthold Lausen and Martin Schumacher. Evaluating the effect of optimized cutoff values in the assessment of prognostic factors. Computational Statistics & Data Analysis, 21 (3):307–326, March 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739500016X.

Zhang:1996:CID

[762] Biao Zhang. Confidence intervals for a distribution function in the presence of auxiliary information. Computational Statistics & Data Analysis, 21(3):327–342, March 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395000194.

Brewer:1996:CHS

[763] M. J. Brewer, C. G. G. Aitken, and M. Talbot. A comparison of hybrid strategies for Gibbs sampling in mixed graphical models. Computational Statistics & Data Analysis, 21 (3):343–365, March 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394000174.

Thioulouse:1996:NWW

[764] Jean Thioulouse and François Chevenet. NetMul, a World-Wide Web user interface for multivariate analysis software. Computational Statistics & Data Analysis, 21(3):369– 372, March 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947396900651.

Anonymous:1996:Tb

[765] Anonymous. Telegram. Computational Statistics & Data Analysis, 21 (3):373–374, March 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947396900675.

Fassbinder:1996:BRB

[766] J. Fassbinder. Book review: Correspondence Analysis Handbook: J.- P. Benzécri (1992) New York: Marcel Dekker, ISBN 0-8247-8437-5, pp. 688, $150.00. Computational Statistics & Data Analysis, 21(3): 375, March 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947396900699.

Edler:1996:IN

[767] Lutz Edler. IASC-News. Computational Statistics & Data Analysis, 21 (3):375–376, March 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947396900717.

Anonymous:1996:CMNa

[768] Anonymous. Calendar of meetings — new entries only. Computational Statistics & Data Analysis, 21(3): 377–379, March 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947396900730.

Ahn:1996:LNR

[769] Hongshik Ahn. Log-normal regression modeling through recursive partitioning. Computational Statistics & Data Analysis, 21(4):381– 398, April 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395000232.

Chen:1996:BRS

[770] Lin-An Chen. Bivariate regression splines. Computational Statistics & Data Analysis, 21(4):399– 418, April 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394000263.

Hirji:1996:CAE

[771] Karim F. Hirji and Timothy D. Johnson. A comparison of algorithms for exact analysis of unordered 2 × K contingency tables. Computational Statistics & Data Analysis, 21 (4):419–429, April 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394000212.

Ryden:1996:EAE

[772] Tobias Rydén. An EM algorithm for estimation in Markov-modulated Poisson processes. Computational Statistics & Data Analysis, 21(4): 431–447, April 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395000259.

Cox:1996:NQL

[773] Christopher Cox. Nonlinear quasilikelihood models: applications to continuous proportions. Computational Statistics & Data Analysis, 21 (4):449–461, April 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395000240.

Dielman:1996:NHT

[774] Terry E. Dielman and Elizabeth L. Rose. A note on hypothesis testing in LAV multiple regression: a small sample comparison. Computational Statistics & Data Analysis, 21(4):463–470, April 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395000224. See

[297].

Hutchinson:1996:GFT

[775] T. P. Hutchinson. On the generalised Friedman test. Computational Statistics & Data Analysis, 21 (4):473–476, April 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947396900432.

Anonymous:1996:INc

[776] Anonymous. IASC-News. Computational Statistics & Data Analysis, 21(4):477–480, April 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947396900456.

Anonymous:1996:CMb

[777] Anonymous. Calendar of meetings. Computational Statistics & Data Analysis, 21(4):481–485, April 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/016794739690047X.

Anonymous:1996:CWSb

[778] Anonymous. Courses, workshops & summer schools. Computational Statistics & Data Analysis, 21(4): 485–486, April 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947396900493.

Akman:1996:CEA

[779] Ibrahim Akman and Jan G. De Gooijer. Component extraction analysis of multivariate time series. Computational Statistics & Data Analysis, 21(5):487–499, May 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395000313.

Aldrin:1996:MPP

[780] Magne Aldrin. Moderate projection pursuit regression for multivariate response data. Computational Statistics & Data Analysis, 21(5): 501–531, May 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947394000298.

Ciok:1996:EGS

[781] Alicja Ciok. An estimator of Gini Separation Index which is nearly median-unbiased. Computational Statistics & Data Analysis, 21(5):

533–546, May 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395000275.

Fernandez-Ponce:1996:CLD

[782] J. M. Fernández-Ponce, R. Infante-Macías, and J. Muñoz-Pérez. Characterization of lifetime distributions based on a quantile dispersion measure. Computational Statistics & Data Analysis, 21(5):547– 561, May 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739682295X.

delRio:1996:CBS

[783] Alejandro Quintela del Río. Comparison of bandwidth selectors in nonparametric regression under dependence. Computational Statistics & Data Analysis, 21(5):563– 580, May 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395000283.

vanderLeeden:1996:RTD

[784] Rien van der Leeden, Karen Vrijburg, and Jan de Leeuw. A review of two different approaches for the analysis of growth data using longitudinal mixed linear models: Comparing hierarchical linear regression (ML3, HLM) and repeated measures designs with structured covariance matrices (BMDP5V). Computational Statistics & Data Analysis, 21(5): 583–605, May 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947396822961.

Tilke:1996:BRBa

[785] Clemens Tilke. Book review: A handbook of statistical analyses using S-Plus: Brian S. Everitt (1994): Chapman & Hall, 143 p., £14.99, ISBN 0-412-56310-X. Computational Statistics & Data Analysis, 21 (5):606, May 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947396901395.

Tilke:1996:BRBb

[786] Clemens Tilke. Book review: SAS programming for researchers and social scientists: Paul E. Spector: Sage Publications (1993), 167 p. ISBN 0-8039-4984-7 (cloth), ISBN 0-8039-4985-5 (paperback). Computational Statistics & Data Analysis, 21 (5):606, May 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947396901413.

Mateev:1996:IN

[787] Plamen Mateev. IASC-News. Computational Statistics & Data Analysis, 21(5):606–607, May 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947396901437.

Banks:1996:MAS

[788] Randy Banks. Meetings: Association for Survey Computing: Second International Conference On Survey and Statistical Computing London, UK, September 11–13, 1996. Computational Statistics & Data Analysis, 21(5):607–608, May 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947396901450.

Burger:1996:WSM

[789] Iris Burger. Workshop on statistical modelling of discrete data and structures: Munich, Germany, October 24–26, 1996. Computational Statistics & Data Analysis, 21 (5):608, May 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947396901474.

Andres:1996:OCC

[790] A. Martín Andrés and A. Silva Mato. Optimal correction for continuity and conditions for validity in the unconditional chi-squared test. Computational Statistics & Data Analysis, 21 (6):609–626, June 1, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947394000352.

Leung:1996:ERC

[791] Chi-Ying Leung. Error rates for classifying observations based on binary and continuous variables with covariates. Computational Statistics & Data Analysis, 21(6):627– 645, June 1, 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395000364.

Rosenqvist:1996:ATT

[792] Anders Rosenqvist. Asymptotic theory for a two-step pseudo-linear Dopplerbearing tracker. Computational Statistics & Data Analysis, 21(6):647– 660, June 1, 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395000348.

Schumacher:1996:NNL

[793] Martin Schumacher, Reinhard Roßner, and Werner Vach. Neural networks and logistic regression: Part I. Computational Statistics & Data Analysis, 21 (6):661–682, June 1, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395000321.

Vach:1996:NNL

[794] Werner Vach, Reinhard Roßner, and Martin Schumacher. Neural networks and logistic regression: Part II. Computational Statistics & Data Analysis, 21(6):683–701, June 1, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739500033X.

Berding:1996:DOP

[795] Christoph Berding, Wilhelm Kleider, and Rüdiger Gössl. DoPE: Doptimal planning & experimentation the tool for experimental design in research and development. Computational Statistics & Data Analysis, 21 (6):705–710, June 1, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947396804579.

Anonymous:1996:CCS

[796] Anonymous. Conference on computational statistics and statistical education, Tartu, Estonia, June 2–8, 1996. Computational Statistics & Data Analysis, 21(6):711, June 1, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947396901759.

Anonymous:1996:SCA

[797] Anonymous. Short courses in applied statistics in Barcelona, summer 1996. Computational Statistics & Data Analysis, 21(6):711– 712, June 1, 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947396901772.

Anonymous:1996:SWC

[798] Anonymous. Second World Conference of the International Association for Statistical Computing February 19 to 22, 1997, Pasadena, CA, USA second announcement and call for contributed papers. Computational Statistics & Data Analysis, 21 (6):712–715, June 1, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947396901796.

Anonymous:1996:CMNb

[799] Anonymous. Calendar of meetings new entries only. Computational Statistics & Data Analysis, 21(6):715– 716, June 1, 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947396901814.

Anonymous:1996:AIVa

[800] Anonymous. Author index volume 21 (1996). Computational Statistics & Data Analysis, 21(6):717– 718, June 1, 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947396901838.

Anonymous:1996:EBb

[801] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 22(1):v–vi, June 14, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947396900018.

Gonzalez-Manteiga:1996:ABK

[802] W. González-Manteiga, C. Sánchez-Sellero, and M. P. Wand. Accuracy of binned kernel functional approximations. Computational Statistics & Data Analysis, 22(1):1–16, June 14, 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947396880303.

Hayter:1996:ECO

[803] A. J. Hayter and W. Liu. Exact calculations for the one-sided Studentized range test for testing against a simple ordered alternative. Computational Statistics & Data Analysis, 22 (1):17–25, June 14, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947396880315.

Granville:1996:DAD

[804] Vincent Granville. Discriminant analysis and density estimation on the finite d-dimensional grid. Computational Statistics & Data Analysis, 22 (1):27–51, June 14, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395000399.

Ogden:1996:DDW

[805] Todd Ogden and Emanuel Parzen. Data dependent wavelet thresholding in nonparametric regression with change-point applications. Computational Statistics & Data Analysis, 22 (1):53–70, June 14, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395000410.

Pan:1996:IOG

[806] Jian-Xin Pan and Kai-Tai Fang. Influential observation in the growth curve model with unstructured covariance matrix. Computational Statistics & Data Analysis, 22(1):71– 87, June 14, 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395000372.

Wilcox:1996:CIS

[807] Rand R. Wilcox. Confidence intervals for the slope of a regression line when the error term has nonconstant variance. Computational Statistics & Data Analysis, 22(1):89– 98, June 14, 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395000380.

Pin:1996:AQS

[808] T. Ng Pin. An algorithm for quantile smoothing splines. Computational Statistics & Data Analysis, 22 (2):99–118, July 1, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395000445.

Bell:1996:RWM

[809] Bradley M. Bell, James V. Burke, and Alan Schumitzky. A relative weighting method for estimating parameters and variances in multiple data sets. Computational Statistics & Data Analysis, 22(2):119–135, July 1, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395000437.

Gray:1996:BGL

[810] H. L. Gray, J. Baek, W. A. Woodward, J. Miller, and M. Fisk. A bootstrap generalized likelihood ratio test in discriminant analysis. Computational Statistics & Data Analysis, 22 (2):137–158, July 1, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395000453.

Huang:1996:STP

[811] Hsin-Cheng Huang and Noel Cressie. Spatio-temporal prediction of snow water equivalent using the Kalman filter. Computational Statistics & Data Analysis, 22(2):159–175, July 1, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/016794739500047X.

Li:1996:EMM

[812] Yi-Hwei Li, John P. Klein, and M. L. Moeschberger. Effects of model misspecification in estimating covariate effects in survival analysis for small sample sizes. Computational Statistics & Data Analysis, 22(2):177– 192, July 1, 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947396880297.

Beyer:1996:DER

[813] Uwe Beyer and Frank Šmieja. Data exploration with reflective adaptive models. Computational Statistics & Data Analysis, 22(2):193– 211, July 1, 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395000488.

Rehman:1996:ITA

[814] Salim U. Rehman and A. Shapiro. An integral transform approach to cross-variograms modeling. Computational Statistics & Data Analysis, 22 (3):213–233, July 15, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395000526.

Chen:1996:RBC

[815] Yuh-Ing Chen, Sheng-Shu Cheng, and Hong-Long Wang. Rank-based comparisons of treatments with a control for repeated measures designs. Computational Statistics & Data Analysis, 22 (3):235–249, July 15, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395000518.

Hoeting:1996:MSV

[816] Jennifer Hoeting, Adrian E. Raftery, and David Madigan. A method for simultaneous variable selection and outlier identification in linear regression. Computational Statistics & Data Analysis, 22(3):251– 270, July 15, 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395000534.

Kwong:1996:SMN

[817] Koon-Shing Kwong and Boris Iglewicz. On singular multivariate normal distribution and its applications. Computational Statistics & Data Analysis, 22 (3):271–285, July 15, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739500050X.

Lim:1996:CTE

[818] Tjen-Sien Lim and Wei-Yin Loh. A comparison of tests of equality of variances. Computational Statistics & Data Analysis, 22(3):287– 301, July 15, 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947395000542.

Poon:1996:AST

[819] Wai-Yin Poon and Hin-Yan Hung. Analysis of square tables with ordered categories. Computational Statistics & Data Analysis, 22(3):303– 322, July 15, 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947395000550.

Kuss:1996:MAM

[820] Oliver Kuss and Armin Koch. Metaanalysis macros for SAS. Computational Statistics & Data Analysis, 22 (3):325–333, July 15, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947396814563.

Kramer:1996:BRBa

[821] Walter Krämer. Book review: Signal processing with alpha-stable distributions and applications: C. L. Nikias and Min Shoa (1995): Wiley, ISBN 0-471-10647-X, £50.00, pp. 168. Computational Statistics & Data Analysis, 22(3):334, July 15, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947396814575.

Trenkler:1996:BRBc

[822] Götz Trenkler. Book review: Methods of multivariate analysis. Wiley series in probability and mathematical statistics: Probability and mathematical statistics section: Alvin C. Rencher (1995): ISBN 0-471-57152-0, £58.00, xvi + 627 pp. Computational Statistics & Data Analysis, 22 (3):334–335, July 15, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947396814587.

Kramer:1996:BRBb

[823] Walter Krämer. Book review: An introduction to computational statistics — Regression analysis: Robert I. Jennrich (1995): Prentice Hall, ISBN 0-13-454810-8, £22.95, pp. 364. Computational Statistics & Data Analysis, 22(3):335, July 15, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947396814599.

Anonymous:1996:INd

[824] Anonymous. IASC-News. Computational Statistics & Data Analysis, 22 (3):335–336, July 15, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947396814605.

Anonymous:1996:INe

[825] Anonymous. IASC-News. Computational Statistics & Data Analysis, 22 (3):336–337, July 15, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947396814617.

Anonymous:1996:CMc

[826] Anonymous. Calendar of meetings. Computational Statistics & Data Analysis, 22(3):337–342, July 15, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947396902996.

Ding:1996:CDS

[827] Cherng G. Ding. On the computation of the distribution of the square of the sample multiple correlation coefficient. Computational Statistics & Data Analysis, 22(4):345– 350, August 10, 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947396000023.

Abramovich:1996:ATW

[828] Felix Abramovich and Yoav Benjamini. Adaptive thresholding of wavelet coefficients. Computational Statistics & Data Analysis, 22(4):351– 361, August 10, 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947396000035.

Aly:1996:RTT

[829] Emad-Eldin A. A. Aly and Sana S. BuHamra. Rank tests for two change points. Computational Statistics & Data Analysis, 22(4):363–372, August 10, 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395000607.

Gupta:1996:LTM

[830] Shanti S. Gupta and Tommaso Gastaldi. Life testing for multicomponent systems with incomplete information on the cause of failure: a study on some inspection strategies. Computational Statistics & Data Analysis, 22(4):373–393, August 10, 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395000593.

Khuri:1996:QPP

[831] AndréI. Khuri, Hyuk Joo Kim, and Yonghwan Um. Quantile plots of the prediction variance for response surface designs. Computational Statistics & Data Analysis, 22(4):395–407, August 10, 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947395000585.

Lai:1996:CSA

[832] Dejian Lai. Comparison study of AR models of the Canadian lynx data: a close look at BDS statistic. Computational Statistics & Data Analysis, 22 (4):409–423, August 10, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947395000569.

Schenker:1996:PPT

[833] Nathaniel Schenker and Jeremy M. G. Taylor. Partially parametric techniques for multiple imputation. Computational Statistics & Data Analysis, 22 (4):425–446, August 10, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947395000577.

Reh:1996:STC

[834] Wolfgang Reh and Bernard Scheffler. Significance tests and confidence intervals for coefficients of variation. Computational Statistics & Data Analysis, 22(4):449–452, August 10, 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947396837078.

Simon:1996:T

[835] Julian L. Simon. Telegram. Computational Statistics & Data Analysis, 22 (4):453–454, August 10, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739690167X.

Anonymous:1996:I

[836] Anonymous. Information. Computational Statistics & Data Analysis, 22 (4):454–456, August 10, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947396901693.

Anonymous:1996:INCa

[837] Anonymous. IAS-News COMP-STAT 1998. Computational Statistics & Data Analysis, 22(4):456, August 10, 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947396901711.

Anonymous:1996:CMd

[838] Anonymous. Calendar of meetings. Computational Statistics & Data Analysis, 22(4):457–460, August 10, 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947396901735.

Kundu:1996:MSL

[839] Debasis Kundu and G. Murali. Model selection in linear regression. Computational Statistics & Data Analysis, 22(5):461–469, September 30, 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947396000084.

Piepho:1996:WEI

[840] Hans-Peter Piepho. Weighted estimates of interlaboratory consensus values. Computational Statistics & Data Analysis, 22(5):471–479, September 30, 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947396000059.

Venter:1996:FMA

[841] J. H. Venter and S. J. Steel. Finding multiple abrupt change points. Computational Statistics & Data Analysis, 22(5):481–504, September 30, 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947396000072.

Bose:1996:CUS

[842] Smarajit Bose. Classification using splines. Computational Statistics & Data Analysis, 22(5):505–525, September 30, 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947396000096.

Carling:1996:TIC

[843] Kenneth Carling. Testing for independence in a competing risks model. Computational Statistics & Data Analysis, 22(5):527–535, September 30, 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947396000060.

Tutz:1996:REO

[844] Gerhard Tutz and Wolfgang Hennevogl. Random effects in ordinal regression models. Computational Statistics & Data Analysis, 22(5):537–557, September 30, 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947396000047.

Popping:1996:MAS

[845] Roel Popping. More agreeable software. Computational Statistics & Data Analysis, 22(5):561–567, September 30, 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947396852728.

Horn:1996:BRB

[846] Manfred Horn. Book review: Design and analysis of experiments for statistical selection, screening and multiple comparisons: Robert E. Bechhofer, Thomas J. Santner, David M. Goldsman (1995): John Wiley & Sons, Inc., ISBN 0-471-57427-9, £45.00, pp. 325. Computational Statistics & Data Analysis, 22(5):568, September 30, 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739685273X.

Trenkler:1996:BRBd

[847] Götz Trenkler. Book review: Continuous univariate distributions: N. L. Johnson, S. Kotz & N. Balakrishnan (1995): Vol. 2, 2nd Edition. John Wiley, New York, ISBN 0-471-58494-0, £70.00, pp. xix + 719. Computational Statistics & Data Analysis, 22(5):568– 569, September 30, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947396852741.

Trenkler:1996:BRBe

[848] Götz Trenkler. Book review: Linear statistical models: J. H. Stapleton (1995): John Wiley, New York, ISBN 0-471-57150-4 £50.00, pp. xiii + 449. Computational Statistics & Data Analysis, 22(5):569, September 30, 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947396852753.

Anonymous:1996:INf

[849] Anonymous. IASC-News. Computational Statistics & Data Analysis, 22(5):569–572, September 30, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/0167947396900754.

Henshall:1996:IRS

[850] J. M. Henshall and D. M. Smith. Iterative refinement of the singularvalue decomposition solution to the regression equations. Computational Statistics & Data Analysis, 22(6):573– 582, October 23, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947396000102.

Ojeda:1996:BDD

[851] M. M. Ojeda and S. F. Juárez-Cerrillo. Biplot display for diagnostic in a two-level regression model for growth curve analysis. Computational Statistics & Data Analysis, 22(6):583– 597, October 23, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947396000138.

Efromovich:1996:AOS

[852] Sam Efromovich. Adaptive orthogonal series density estimation for small samples. Computational Statistics & Data Analysis, 22(6):599–617, October 23, 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947396000126.

Soo:1996:MSR

[853] Yuh-Wen Soo and Douglas M. Bates. Multiresponse spline regression. Computational Statistics & Data Analysis, 22(6):619–631, October 23, 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739600014X.

Chatterjee:1996:GAT

[854] Sangit Chatterjee, Matthew Laudato, and Lucy A. Lynch. Genetic algorithms and their statistical applications: an introduction. Computational Statistics & Data Analysis, 22(6):633– 651, October 23, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947396000114.

Handl:1996:BRB

[855] Andreas Handl. Book review: Biplots: J. C. Gower and D. J. Hand (1996) London: Chapman & Hall, ISBN 0-412-71630-5, £32.00, pp. 277. Computational Statistics & Data Analysis, 22(6):655, October 23, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947396823814.

Trenkler:1996:E

[856] Götz Trenkler. Erratum. Computational Statistics & Data Analysis, 22 (6):655, October 23, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947396823826.

Anonymous:1996:INR

[857] Anonymous. IASC-News results of the BoD ERS IASC elections for the period of 1996–2000. Computational Statistics & Data Analysis, 22 (6):656, October 23, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947396823838.

Anonymous:1996:CAB

[858] Anonymous. Compstat ’98 24–28 August 1998, Bristol, UK first call for papers. Computational Statistics & Data Analysis, 22(6):656–657, October 23, 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947396900237.

Anonymous:1996:CMe

[859] Anonymous. Calendar of meetings. Computational Statistics & Data Analysis, 22(6):658–663, October 23, 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947396900274.

Anonymous:1996:AIS

[860] Anonymous. Author index of SS-NinCSDA: Volume 21 & volume 22 (1996). Computational Statistics & Data Analysis, 22(6):664–667, October 23, 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947396900298.

Anonymous:1996:AIVb

[861] Anonymous. Author index volume 22 (1996). Computational Statistics & Data Analysis, 22(6):669–671, October 23, 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947396900316.

Anonymous:1996:EBc

[862] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 23 (1):v–vi, November 15, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947396902777.

Rasson:1996:P

[863] J.-P. Rasson and P. J. Rousseeuw. Preface. Computational Statistics & Data Analysis, 23(1):1, November 15, 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947396902790.

Hartigan:1996:ICP

[864] J. A. Hartigan. Introduction: Classification, probability and statistics. Computational Statistics & Data Analysis, 23(1):3–4, November 15, 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947396000175.

Bock:1996:PMC

[865] Hans H. Bock. Probabilistic models in cluster analysis. Computational Statistics & Data Analysis, 23(1):5– 28, November 15, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947396889195.

Eddy:1996:ASL

[866] William F. Eddy, Audris Mockus, and Shingo Oue. Approximate single linkage cluster analysis of large data sets in high-dimensional spaces. Computational Statistics & Data Analysis, 23(1):29–43, November 15, 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947396000199.

ElKhattabi:1996:IAD

[867] Sana El Khattabi and Franz Streit. Identification analysis in directional statistics. Computational Statistics & Data Analysis, 23(1):45–63, November 15, 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947396000205.

Govaert:1996:CMC

[868] G. Govaert and M. Nadif. Comparison of the mixture and the classification maximum likelihood in cluster analysis with binary data. Computational Statistics & Data Analysis, 23(1):65– 81, November 15, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947396000217.

Hardy:1996:NC

[869] André Hardy. On the number of clusters. Computational Statistics & Data Analysis, 23(1):83–96, November 15, 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947396000229.

Hartigan:1996:R

[870] J. A. Hartigan. Recognition. Computational Statistics & Data Analysis, 23(1):97–103, November 15, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947396000230.

Rasson:1996:GTC

[871] J. P. Rasson and V. Granville. Geometrical tools in classification. Computational Statistics & Data Analysis, 23(1):105–123, November 15, 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947396000242.

Remon:1996:DAA

[872] M. Rémon. A discriminant analysis algorithm for the inside/outside problem. Computational Statistics & Data Analysis, 23(1):125–133, November 15, 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947396000254.

Rousseeuw:1996:FCU

[873] P. J. Rousseeuw, L. Kaufman, and E. Trauwaert. Fuzzy clustering using scatter matrices. Computational Statistics & Data Analysis, 23(1):135– 151, November 15, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947396000266.

Ruts:1996:CDC

[874] Ida Ruts and Peter J. Rousseeuw. Computing depth contours of bivariate point clouds. Computational Statistics & Data Analysis, 23(1):153–168, November 15, 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947396000278.

VanCutsem:1996:CSS

[875] Bernard Van Cutsem. Combinatorial structures and structures for classification. Computational Statistics & Data Analysis, 23(1):169–188, November 15, 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794739600028X.

Lauro:1996:CSS

[876] Carlo Lauro. Computational statistics or statistical computing, is that the question? Computational Statistics & Data Analysis, 23(1):191–193, November 15, 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947396889201.

Prat:1996:WC

[877] Albert Prat. Welcome to COMP-STAT’96. Computational Statistics & Data Analysis, 23(1):194–195, November 15, 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947396902819.

deFalguerolles:1996:SIP

[878] Antoine de Falguerolles. Statement of the incoming President of the ERS of the IASC. Computational Statistics & Data Analysis, 23(1): 196, November 15, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947396902832.

Antoch:1996:WFC

[879] Jaromir Antoch. Where is the future of computational statistics? Computational Statistics & Data Analysis, 23 (1):197, November 15, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947396902856.

Cameron:1996:MEI

[880] Murray Cameron. More exchange of ideas needed. Computational Statistics & Data Analysis, 23(1): 198, November 15, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794739690287X.

Anonymous:1996:Tc

[881] Anonymous. Telegram. Computational Statistics & Data Analysis, 23(1):199– 200, November 15, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947396902893.

Kramer:1996:BRBc

[882] Walter Krämer. Book review: The analysis of time-series: C. Chattfield (1996): an introduction, 5th ed. Chapman & Hall, ISBN 0-412-71640-2, pp. 283, £18.99. Computational Statistics & Data Analysis, 23 (1):201, November 15, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947396889213.

Kramer:1996:BRBd

[883] Walter Krämer. Book review: Introduction to statistical time series: Wayne A. Fuller (1996): (2nd edition). Wiley, ISBN 0-471-55239-9, pp. 736, £55.00. Computational Statistics & Data Analysis, 23(1): 201, November 15, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947396889225.

Hassler:1996:BRB

[884] Uwe Hassler. Book review: A Casebook for a first course in statistics and data analysis.: S. Chatterjee, M. S. Handcock and J. S. Simon-off (1995): Wiley & Sons, ISBN 0-471-11030-2, £19.95, pp. 314. Computational Statistics & Data Analysis, 23(1):201– 202, November 15, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947396889237.

Anonymous:1996:CMNc

[885] Anonymous. Calendar of meetings — new entries only. Computational Statistics & Data Analysis, 23(1):203– 204, November 15, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947396902911.

Anonymous:1996:CWSc

[886] Anonymous. Courses, workshops & summer schools — new entries only. Computational Statistics & Data Analysis, 23(1):204–206, November 15, 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947396902935.

Gupta:1996:AZA

[887] Pushpa L. Gupta, Ramesh C. Gupta, and Ram C. Tripathi. Analysis of zeroadjusted count data. Computational Statistics & Data Analysis, 23(2):207– 218, December 11, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947396000321.

Lipsitz:1996:EJD

[888] Stuart R. Lipsitz, Garrett M. Fitzmaurice, Lynn Sleeper, and L. P. Zhao. Estimating the joint distribution of repeated binary responses: Some small sample results. Computational Statistics & Data Analysis, 23(2):219– 227, December 11, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794739600031X.

Schott:1996:EEL

[889] James R. Schott. Eigenprojections and the equality of latent roots of a correlation matrix. Computational Statistics & Data Analysis, 23(2):229– 238, December 11, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947396000333.

Wang:1996:SAS

[890] Shu-Jia Wang and Sik-Yum Lee. Sensitivity analysis of structural equation models with equality functional constraints. Computational Statistics & Data Analysis, 23(2):239–256, December 11, 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947396000345.

Chen:1996:SAU

[891] Victoria C. P. Chen and Anthony J. Hayter. Sensitivity analysis of upper confidence bounds on the range of treatment effects. Computational Statistics & Data Analysis, 23(2):257– 262, December 11, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947396000308.

Kafadar:1996:CSR

[892] Karen Kafadar and Philip C. Prorok. Computer simulation of randomized cancer screening trials to compare methods of estimating lead time and benefit time. Computational Statistics & Data Analysis, 23(2):263– 291, December 11, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947396000291.

Anonymous:1996:Td

[893] Anonymous. Telegram. Computational Statistics & Data Analysis, 23(2):295– 296, December 11, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739690033X.

Anonymous:1996:ICF

[894] Anonymous. ICM98-CL8: Financial support for mathematicians from Eastern europe’ and the independent states of the former Soviet union. Computational Statistics & Data Analysis, 23 (2):296, December 11, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947396900353.

Anonymous:1996:INg

[895] Anonymous. IASC news. Computational Statistics & Data Analysis, 23(2):297–300, December 11, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/0167947396900377.

Anonymous:1996:CMNd

[896] Anonymous. Calendar of meetings — new entries only. Computational Statistics & Data Analysis, 23(2):300– 304, December 11, 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947396900390.

Anonymous:1996:CWSd

[897] Anonymous. Courses, workshops & summer schools — new entries only. Computational Statistics & Data Analysis, 23(2):304–306, December 11, 1996. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947396900419.

McCulloch:1997:PTM

[898] J. Huston McCulloch and Don B. Panton. Precise tabulation of the maximally-skewed stable distributions and densities. Computational Statistics & Data Analysis, 23(3):307–320, January 9, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947396000394.

Mustonen:1997:MTV

[899] Seppo Mustonen. A measure for total variability in multivariate normal distribution. Computational Statistics & Data Analysis, 23(3):321–334, January 9, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947396000424.

Simonoff:1997:MSH

[900] Jeffrey S. Simonoff and Frederic Udina. Measuring the stability of histogram appearance when the anchor position is changed. Computational Statistics & Data Analysis, 23(3):335– 353, January 9, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947396000400.

Draper:1997:GGM

[901] Norman R. Draper and Yonghong (Fred) Yang. Generalization of the geometric mean functional relationship. Computational Statistics & Data Analysis, 23 (3):355–372, January 9, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947396000370.

Jovanovic:1997:SPM

[902] B. D. Jovanovic and D. W. Hosmer. A simulation of the performance of C_{p} in model selection for logistic and Poisson regression. Computational Statistics & Data Analysis, 23 (3):373–379, January 9, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947396000357.

Dielman:1997:NHT

[903] Terry E. Dielman and Elizabeth L. Rose. A note on hypothesis testing in LAV multiple regression: a small sample comparison. Computational Statistics & Data Analysis, 23 (3):381–388, January 9, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947397810213. See

[297].

Karwe:1997:NRM

[904] Vatsala V. Karwe and Joseph I. Naus. New recursive methods for scan statistic probabilities. Computational Statistics & Data Analysis, 23 (3):389–402, January 9, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947396000412.

Mkhadri:1997:RDA

[905] A. Mkhadri, G. Celeux, and A. Nasroallah. Regularization in discriminant analysis: an overview. Computational Statistics & Data Analysis, 23 (3):403–423, January 9, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947396000436.

Anonymous:1997:LR

[906] Anonymous. List of referees. Computational Statistics & Data Analysis, 23(3):425, January 9, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397810225.

Anonymous:1997:LES

[907] Anonymous. The leading edge of statistical computing. Computational Statistics & Data Analysis, 23(3): 429, January 9, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397810237.

Anonymous:1997:Ta

[908] Anonymous. Telegram. Computational Statistics & Data Analysis, 23 (3):430, January 9, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397900633.

Anonymous:1997:INa

[909] Anonymous. IASC news. Computational Statistics & Data Analysis, 23 (3):431–434, January 9, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397900657.

Anonymous:1997:CMa

[910] Anonymous. Calendar of meetings. Computational Statistics & Data Analysis, 23(3):434–441, January 9, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397900670.

Anonymous:1997:CWSa

[911] Anonymous. Courses, workshops & summer schools. Computational Statistics & Data Analysis, 23(3):441– 442, January 9, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397900694.

Hwang:1997:AGB

[912] Frank K. Hwang and Paul E. Wright. An O(n log n) algorithm for the generalized birthday problem. Computational Statistics & Data Analysis, 23(4):443–451, February 6, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947396000515.

Jacobson:1997:NAL

[913] Tor Jacobson and Rolf Larsson. Numerical aspects of a likelihood ratio test statistic for cointegrating rank. Computational Statistics & Data Analysis, 23(4):453–465, February 6, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794739784755X.

Schemper:1997:NAE

[914] Michael Schemper and Alexandra Kaider. A new approach to estimate correlation coefficients in the presence of censoring and proportional hazards. Computational Statistics & Data Analysis, 23(4):467–476, February 6, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794739600045X.

Chiou:1997:IES

[915] Paul Chiou. Interval estimation of scale parameters following a pre-test for two exponential distributions. Computational Statistics & Data Analysis, 23(4):477–489, February 6, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947396000503.

Caouder:1997:TGF

[916] Nathalie Caouder and Sylvie Huet. Testing goodness-of-fit for nonlinear regression models with heterogeneous variances. Computational Statistics & Data Analysis, 23(4):491– 507, February 6, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947396000497.

Lindgren:1997:QRC

[917] Anna Lindgren. Quantile regression with censored data using generalized L_{1} minimization. Computational Statistics & Data Analysis, 23(4):509– 524, February 6, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947396000485.

Meintanis:1997:CSS

[918] S. G. Meintanis and G. S. Donatos. A comparative study of some robust methods for coefficient-estimation in linear regression. Computational Statistics & Data Analysis, 23(4):525– 540, February 6, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947396000461.

Verbeke:1997:EMR

[919] Geert Verbeke and Emmanuel Lesaffre. The effect of misspecifying the random-effects distribution in linear mixed models for longitudinal data. Computational Statistics & Data Analysis, 23(4):541–556, February 6, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947396000473.

Naeve:1997:RSG

[920] Peter Naeve. The referee system: get involved. Computational Statistics & Data Analysis, 23(4):559–561, February 6, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947397847561.

Hormann:1997:SRS

[921] Allmut Hörmann. The SSN and the referee system. Computational Statistics & Data Analysis, 23(4):562– 563, February 6, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397847573.

Tiit:1997:CBT

[922] Ene-Margit Tiit. COMPSTAT ’96 in Barcelona: the triangle: Computers — statistics — education. Computational Statistics & Data Analysis, 23 (4):564, February 6, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397847585.

Musiol:1997:BRB

[923] Gerald Musiol. Book review: An Introduction to Categorical Data Analysis: Agresti, A. (1996) New York: Wiley & Sons, ISBN 0-471-11338-7, xi + 290 pages, $90.00. Computational Statistics & Data Analysis, 23 (4):565, February 6, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397847603.

Trenkler:1997:BRBa

[924] Götz Trenkler. Book review: The Elements of Integration and Lebesgue Measure: Robert G. Bartle (1995) Wiley Classics Library Edition. New York: Wiley, pp. 179, £20.50, ISBN 0-471-04222-6. Computational Statistics & Data Analysis, 23(4):565, February 6, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947397847597.

Antoch:1997:BRB

[925] J. Antoch. Book review: A Guide to Chi-Squared Testing: Greenwood, P. E. and Nikulin, M. S. New York: John Wiley & Sons, Inc.,.pp. 280 + xii, ISBN 0-471-55779-X. AMS 1991 Classification: 62-02, 62F03, 62H15. Computational Statistics & Data Analysis, 23(4):565– 566, February 6, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947397847615.

Anonymous:1997:INb

[926] Anonymous. IASC news. Computational Statistics & Data Analysis, 23(4):566–569, February 6, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947397847627.

Anonymous:1997:CMb

[927] Anonymous. Calendar of meetings. Computational Statistics & Data Analysis, 23(4):569–573, February 6, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947397901729.

Anonymous:1997:CWSb

[928] Anonymous. Courses, workshops & summer schools — new entries only. Computational Statistics & Data Analysis, 23(4):573–574, February 6, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397901730.

Anonymous:1997:AIVa

[929] Anonymous. Author index volume 23 (1997). Computational Statistics & Data Analysis, 23(4):575–576, February 6, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397847639.

Anonymous:1997:EBa

[930] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 24(1):v–vi, March 6, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947397900773.

vanCasteren:1997:FPR

[931] Pieter H. F. M. van Casteren. The frequency polygon reconsidered. Computational Statistics & Data Analysis, 24(1):1–8, March 6, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947396000539.

Constantine:1997:WRF

[932] A. G. Constantine and N. I. Robinson. The Weibull renewal function for moderate to large arguments. Computational Statistics & Data Analysis, 24(1):9–27, March 6, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947396000527.

Fang:1997:SMG

[933] Kai-Tai Fang and Run-Ze Li. Some methods for generating both an NTnet and the uniform distribution on a Stiefel manifold and their applications. Computational Statistics & Data Analysis, 24(1):29–46, March 6, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947396000576.

Ahn:1997:ESM

[934] Chul Ahn. An evaluation of simple methods for the estimation of a common odds ratio in clusters with variable size. Computational Statistics & Data Analysis, 24(1):47– 61, March 6, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947396000540.

Gunes:1997:MGF

[935] Huseyin Gunes, Dennis C. Dietz, Paul F. Auclair, and Albert H. Moore. Modified goodness-of-fit tests for the inverse Gaussian distribution. Computational Statistics & Data Analysis, 24(1):63–77, March 6, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947396000564.

Kaiser:1997:MLE

[936] Mark S. Kaiser. Maximum likelihood estimation of link function parameters. Computational Statistics & Data Analysis, 24(1):79–87, March 6, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947396000552.

Kim:1997:NET

[937] Donguk Kim and Alan Agresti. Nearly exact tests of conditional independence and marginal homogeneity for sparse contingency tables. Computational Statistics & Data Analysis, 24 (1):89–104, March 6, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947396000382.

Anonymous:1997:BRB

[938] Anonymous. Book review: The Statistical Software Guide ’96/97: edited by Armin Koch and Uwe Haag. Computational Statistics & Data Analysis, 24 (1):107–131, March 6, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397874622.

Anonymous:1997:INc

[939] Anonymous. IASC news. Computational Statistics & Data Analysis, 24(1):132, March 6, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947397901493.

Anonymous:1997:Tb

[940] Anonymous. Telegram. Computational Statistics & Data Analysis, 24 (1):132, March 6, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947397900797.

Anonymous:1997:CMc

[941] Anonymous. Calendar of meetings. Computational Statistics & Data Analysis, 24(1):133–135, March 6, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/0167947397900839.

Anonymous:1997:CWSc

[942] Anonymous. Courses, workshops & summer schools — new entries only. Computational Statistics & Data Analysis, 24(1):135– 136, March 6, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947397900852.

Anonymous:1997:A

[943] Anonymous. Announcement. Computational Statistics & Data Analysis, 24 (1):137–138, March 6, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947397900876.

Hirose:1997:MLP

[944] Hideo Hirose. Maximum likelihood parameter estimation in the threeparameter log-normal distribution using the continuation method. Computational Statistics & Data Analysis, 24 (2):139–152, April 3, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947396000631.

LeBlanc:1997:RPR

[945] Michael LeBlanc. Restricted polynomial regression. Computational Statistics & Data Analysis, 24(2):153– 167, April 3, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794739600059X.

Mak:1997:ENT

[946] T. K. Mak, H. Wong, and W. K. Li. Estimation of nonlinear time series with conditional heteroscedastic variances by iteratively weighted least squares. Computational Statistics & Data Analysis, 24(2):169– 178, April 3, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947396000606.

Adatia:1997:ABP

[947] A. Adatia. Approximate BLUEs of the parameters of the half logistic distribution based on fairly large doubly censored samples. Computational Statistics & Data Analysis, 24(2):179– 191, April 3, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947396000588.

Hecker:1997:CRC

[948] Hartmut Hecker and Peter Wübbelt. Clustering by response: CBR. Computational Statistics & Data Analysis, 24 (2):193–215, April 3, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947396000618.

Maravalle:1997:CT

[949] Maurizio Maravalle, Bruno Simeone, and Rosella Naldini. Clustering on trees. Computational Statistics & Data Analysis, 24(2):217– 234, April 3, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794739600062X.

Kastner:1997:MWT

[950] Christian Kastner, Andreas Fieger, and Christian Heumann. MAREG and WinMAREG a tool for marginal regression models. Computational Statistics & Data Analysis, 24(2):237– 241, April 3, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947397835621.

Giacalone:1997:PVM

[951] Massimiliano Giacalone. Participant’s views of meeting: COMP-STAT ’96 in Barcelona: Comments and impressions. Computational Statistics & Data Analysis, 24(2): 242, April 3, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947397835633.

Anonymous:1997:Ia

[952] Anonymous. Information. Computational Statistics & Data Analysis, 24 (2):243–244, April 3, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397835645.

Anonymous:1997:INd

[953] Anonymous. IASC news. Computational Statistics & Data Analysis, 24 (2):245–247, April 3, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947397900335.

Anonymous:1997:CMd

[954] Anonymous. Calendar of meetings. Computational Statistics & Data Analysis, 24(2):247–253, April 3, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/0167947397900359.

Anonymous:1997:CWSd

[955] Anonymous. Courses, workshops & summer schools. Computational Statistics & Data Analysis, 24(2):253– 254, April 3, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947397900372.

Besse:1997:SNP

[956] Philippe C. Besse, Hervé Cardot, and Frédéric Ferraty. Simultaneous non-parametric regressions of unbalanced longitudinal data. Computational Statistics & Data Analysis, 24 (3):255–270, May 12, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947396000679.

Bradu:1997:IOM

[957] Dan Bradu. Identification of outliers by means of L_{1} regression: Safe and unsafe configurations. Computational Statistics & Data Analysis, 24 (3):271–281, May 12, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947396000655.

Chen:1997:DAO

[958] Cathy W. S. Chen. Detection of additive outliers in bilinear time series. Computational Statistics & Data Analysis, 24(3):283–294, May 12, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947396000680.

Fahrmeir:1997:PLE

[959] Ludwig Fahrmeir and Stefan Wagenpfeil. Penalized likelihood estimation and iterative Kalman smoothing for non-Gaussian dynamic regression models. Computational Statistics & Data Analysis, 24(3):295– 320, May 12, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947396000643.

Nishii:1997:UUB

[960] Ryuei Nishii, Takemi Yanagimoto, and Saeko Kusanobu. The use of univariate Bayes regression models for spatial smoothing. Computational Statistics & Data Analysis, 24(3):321– 336, May 12, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947396000692.

Siciliano:1997:TFA

[961] Roberta Siciliano and Ab Mooijaart. Three-factor association models for three-way contingency tables. Computational Statistics & Data Analysis, 24 (3):337–356, May 12, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397870247.

tenBerge:1997:RAR

[962] Jos M. F. ten Berge. Reduction of asymmetry by rank-one matrices. Computational Statistics & Data Analysis, 24(3):357– 366, May 12, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947396000709.

Anonymous:1997:Tc

[963] Anonymous. Telegram. Computational Statistics & Data Analysis, 24 (3):369–371, May 12, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397870259.

Anonymous:1997:NNa

[964] Anonymous. Net news. Computational Statistics & Data Analysis, 24 (3):371–372, May 12, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397870260.

Trenkler:1997:BRBb

[965] Götz Trenkler. Book review: Multivariate statistical analysis: Narayan C. Giri (1996): New York: Marcel Dekker, Inc., xii + 378 pages, ISBN 0-8247-9338-2, US$135.00. Computational Statistics & Data Analysis, 24(3):372, May 12, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397870272.

Bousonville:1997:BRB

[966] Thomas Bousonville. Book review: Genetic algorithms + data structures = evolution programs: Zbigniew Michalewicz (1996): 3rd ed.. New York: Springer, pp. 387, ISBN 3-540-60676-9, 58,00 DM. Computational Statistics & Data Analysis, 24 (3):372–373, May 12, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397870284.

Militino:1997:BRBa

[967] Ana Fernández Militino. Book review: Robust statistics, data analysis, and computer intensive methods: Rieder, H. (Editor) (1996): New York: Springer, pp. 427, ISBN 0-387-94660-8, £33.50. Computational Statistics & Data Analysis, 24(3):373– 374, May 12, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397870296.

Anonymous:1997:CPT

[968] Anonymous. Conference proceedings in teaching computing. Computational Statistics & Data Analysis, 24 (3):374, May 12, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397870302.

Anonymous:1997:NNV

[969] Anonymous. Neural nets versus logistic regression. Computational Statistics & Data Analysis, 24(3): 374, May 12, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397870326.

Anonymous:1997:SCB

[970] Anonymous. Sample copy of Computational Statistics and Data Analysis. Computational Statistics & Data Analysis, 24(3):374, May 12, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397870314.

Anonymous:1997:IIS

[971] Anonymous. IV international summer school on computational statistics organized by the IASC and the department of statistical science Padua, Italy, June–July 4, 1997. Computational Statistics & Data Analysis, 24 (3):375–376, May 12, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397870338.

Anonymous:1997:IEC

[972] Anonymous. IASC Executive and Council following 51st ISI, Istanbul. Computational Statistics & Data Analysis, 24(3):376, May 12, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794739787034X.

Anonymous:1997:IPI

[973] Anonymous. IASC program for ISI’99 in Helsinki, Finland. Computational Statistics & Data Analysis, 24(3):376– 377, May 12, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397870351.

Anonymous:1997:CMNa

[974] Anonymous. Calendar of meetings — new entries only. Computational Statistics & Data Analysis, 24(3):377– 378, May 12, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947397900414.

Routledge:1997:VPT

[975] R. D. Routledge. P -values from permutation and F -tests. Computational Statistics & Data Analysis, 24 (4):379–386, June 5, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947396000734.

Achcar:1997:URR

[976] Jorge Alberto Achcar and JoséCarlos Fogo. A useful reparameterization for the reliability in the Weibull case. Computational Statistics & Data Analysis, 24(4):387– 400, June 5, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794739600076X.

Migon:1997:BAR

[977] Helio S. Migon and Vilma M. Tachibana. Bayesian approximations in randomized response model. Computational Statistics & Data Analysis, 24(4):401–409, June 5, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947396000758.

Pardo:1997:SIT

[978] J. A. Pardo, M. C. Pardo, M. L. Vicente, and M. D. Esteban. A statistical information theory approach to compare the homogeneity of several variances. Computational Statistics & Data Analysis, 24(4):411– 416, June 5, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947396000801.

Park:1997:LLM

[979] Taesung Park and Morton B. Brown. Log-linear models for a binary response with nonignorable nonresponse. Computational Statistics & Data Analysis, 24(4):417–432, June 5, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947396000746.

Reschenhofer:1997:GKS

[980] Erhard Reschenhofer. Generalization of the Kolmogorov–Smirnov test. Computational Statistics & Data Analysis, 24(4):433–441, June 5, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947396000771.

Bellhouse:1997:SOM

[981] David R. Bellhouse, Robert Philips, and James E. Stafford. Symbolic operators for multiple sums. Computational Statistics & Data Analysis, 24 (4):443–454, June 5, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947396000722.

Farebrother:1997:HDL

[982] Richard William Farebrother. The historical development of the linear minimax absolute residual estimation procedure 1786–1960. Computational Statistics & Data Analysis, 24(4):455– 466, June 5, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947396000710.

Hermanska:1997:BEE

[983] Jindřiška Heřmanská and Miroslav Kárný. Bayesian estimation of effective half-life in dosimetric applications. Computational Statistics & Data Analysis, 24(4):467– 482, June 5, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947396000783.

Jung:1997:LIM

[984] Kang-Mo Jung, Myung Geun Kim, and Byung Chun Kim. Local influence in maximum likelihood factor analysis. Computational Statistics & Data Analysis, 24(4):483– 491, June 5, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947396000795.

Popping:1997:UCF

[985] Roel Popping. The utility of a consultation facility in a program for statistical computing. Computational Statistics & Data Analysis, 24(4):495– 503, June 5, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397827624.

Lenz:1997:BRBa

[986] Hans J. Lenz. Book review: A handbook of statistical analyses using SAS: B. S. Everitt and G. Der (1996) London: Chapman & Hall, ISBN 0-412-71050-1, pp. 158, £19.99. Computational Statistics & Data Analysis, 24(4):504, June 5, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397827636.

Militino:1997:BRBb

[987] Ana Fernández Militino. Book review: Robust statistical procedures, asymptotics and interrelations: Jana Jurecková and Pranab Kumar Sen (1996) New York: Wiley, pp. 466, ISBN 0-471-82221-3, £50.00. Computational Statistics & Data Analysis, 24 (4):504–505, June 5, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947397827648.

Anonymous:1997:INe

[988] Anonymous. IASC news. Computational Statistics & Data Analysis, 24 (4):505–507, June 5, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947397900712.

Anonymous:1997:CMNb

[989] Anonymous. Calendar of meetings — new entries only. Computational Statistics & Data Analysis, 24(4):508– 510, June 5, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947397900736.

Anonymous:1997:CWSe

[990] Anonymous. Courses, workshops & summer schools. Computational Statistics & Data Analysis, 24(4): 510, June 5, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794739790075X.

Anonymous:1997:AIVb

[991] Anonymous. Author index volume 24 (1997). Computational Statistics & Data Analysis, 24(4):511– 512, June 5, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794739782765X.

Anonymous:1997:EBb

[992] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 25(1):v–vi, July 3, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397901651.

Hinloopen:1997:CEH

[993] Jeroen Hinloopen and Rien Wagenvoort. On the computation and efficiency of a HBP-GM estimator some simulation results. Computational Statistics & Data Analysis, 25 (1):1–15, July 3, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947396000837.

Tian:1997:APE

[994] Jilei Tian, Martti Juhola, and Tapio Grönfors. AR parameter estimation by a feedback neural network. Computational Statistics & Data Analysis, 25(1):17–24, July 3, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947396000849.

Bachmaier:1997:RMC

[995] Martin Bachmaier and Manfred Precht. Robust multiple confidence intervals for contrasts. Computational Statistics & Data Analysis, 25 (1):25–42, July 3, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947396000874.

Christoffersson:1997:RMR

[996] Jan Christoffersson. A resampling method for regression models with serially correlated errors. Computational Statistics & Data Analysis, 25 (1):43–53, July 3, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947396000813.

Lee:1997:CMO

[997] Andy H. Lee and Wing K. Fung. Confirmation of multiple outliers in generalized linear and nonlinear regressions. Computational Statistics & Data Analysis, 25(1):55– 65, July 3, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947396000862.

Lesaffre:1997:PCW

[998] Emmanuel Lesaffre, Geert Molenberghs, and Ilse Scheys. Prediction and classification when the diagnostic classes are related. Computational Statistics & Data Analysis, 25 (1):67–90, July 3, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947396000850.

Raqab:1997:MML

[999] Mohammad Z. Raqab. Modified maximum likelihood predictors of future order statistics from normal samples. Computational Statistics & Data Analysis, 25(1):91– 106, July 3, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947396000825.

Koch:1997:RCS

[1000] Armin Koch and Allmut. Hörmann.

Rapid communications section. Computational Statistics & Data Analysis, 25(1):109, July 3, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397892018.

Fouladi:1997:ICO

[1001] Rachel Tanya Fouladi. Issues, Challenges, and Opportunities in Computational Statistics and Data Analysis: IASC 2^{nd} world conference, Pasadena, USA, February 16–21, 1997. Computational Statistics & Data Analysis, 25 (1):109–112, July 3, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794739789202X.

Grassmann:1997:WCI

[1002] Janet Grassmann. The 2^{nd} world conference of the IASC from the point of view of a student (February 19 to 22, 1997, Pasadena, CA, Doubletree Hotel). Computational Statistics & Data Analysis, 25(1): 112, July 3, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947397892031.

Hudson:1997:IWC

[1003] Malcolm Hudson. IASC 2^{nd} world congress, Pasadena, USA, February 19–22, 1997. Computational Statistics & Data Analysis, 25(1):113– 114, July 3, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947397892043.

Anonymous:1997:INf

[1004] Anonymous. IASC news. Computational Statistics & Data Analysis, 25 (1):114–118, July 3, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947397901675.

Anonymous:1997:CMe

[1005] Anonymous. Calendar of meetings. Computational Statistics & Data Analysis, 25(1):118–123, July 3, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947397901699.

Anonymous:1997:CWSf

[1006] Anonymous. Courses, workshops & summer schools. Computational Statistics & Data Analysis, 25(1):123– 124, July 3, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397901717.

Croux:1997:EWI

[1007] Christophe Croux and Gentiane Haesbroeck. An easy way to increase the finite-sample efficiency of the resampled minimum volume ellipsoid estimator. Computational Statistics & Data Analysis, 25(2):125– 141, July 31, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947396000886.

Paula:1997:RHE

[1008] Gilberto A. Paula and Oscar V. Rojas. On restricted hypotheses in extreme value regression models. Computational Statistics & Data Analysis, 25(2):143–157, July 31, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947396000904.

Yu:1997:CLC

[1009] Keming Yu and M. C. Jones. A comparison of local constant and local linear regression quantile estimators. Computational Statistics & Data Analysis, 25(2):159– 166, July 31, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397000066.

Bradstreet:1997:MCS

[1010] Thomas E. Bradstreet. A Monte Carlo study of type I error rates for the two-sample Behrens–Fisher problem with and without rank transformation. Computational Statistics & Data Analysis, 25(2):167– 179, July 31, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947397000042.

Jamshidian:1997:DMU

[1011] Mortaza Jamshidian. T -distribution modeling using the available statistical software. Computational Statistics & Data Analysis, 25(2):181– 206, July 31, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947396000916.

Seidel:1997:PWP

[1012] Wilfried Seidel. A possible way out of the pitfall of acceptance sampling by variables: Treating variances as unknown. Computational Statistics & Data Analysis, 25(2):207– 216, July 31, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947396000898.

Somerville:1997:MTS

[1013] Paul N. Somerville. Multiple testing and simultaneous confidence intervals: calculation of constants. Computational Statistics & Data Analysis, 25 (2):217–233, July 31, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397000030.

Baker:1997:NPU

[1014] Frank B. Baker. A note on the proper use of the Numerical Recipes RAN1 random number generator. Computational Statistics & Data Analysis, 25 (2):237–239, July 31, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397896491. See

also [1161].

Anonymous:1997:SPS

[1015] Anonymous. Software prize for students and young researchers by ASK. Computational Statistics & Data Analysis, 25(2):240, July 31, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397896508.

Anonymous:1997:Td

[1016] Anonymous. Telegram. Computational Statistics & Data Analysis, 25 (2):240, July 31, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794739789651X.

Tilke:1997:BRBa

[1017] Clemens Tilke. Book review: Practical handbook of spatial statistics: Arlinghaus, Sandra Lach (1996): CRC Press, ISBN 0-8493-0132-7, pp. 336, $49.95. Computational Statistics & Data Analysis, 25(2):240– 241, July 31, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397896521.

Pigeot-Kubler:1997:BRB

[1018] Iris Pigeot-Kübler. Book review: Multiple Comparisons Theory and Methods: Hsu, Jason C. (1996): London: Chapman & Hall, pp. 277, ISBN 0-412-98281-1, £35.00. Computational Statistics & Data Analysis, 25(2):241– 243, July 31, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397896533.

Trenkler:1997:BRBc

[1019] Götz Trenkler. Book review: Handbook of Matrices: Lütkepohl, H. (1996): New York: John Wiley & Sons, xvi + 304 pp., ISBN 0-471-96688-6; £50.00. Computational Statistics & Data Analysis, 25(2):243, July 31, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397896545.

Tilke:1997:BRBb

[1020] Clemens Tilke. Book review: VAR-IOWIN — Software for Spatial Data Analysis in 2D: Pannatier, Yvan (1996): Springer Verlag, 91 p., ISBN 0-387-94679-9, Hardcover DM 78.00. Computational Statistics & Data Analysis, 25(2):243– 244, July 31, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397896557.

Tilke:1997:BRBc

[1021] Clemens Tilke. Book review: Introduction to Disjunctive Kriging and Non-Linear Geostatistics: Rivoirard, Jacques (1994): Clarendon Press, 182 p., ISBN 0-19-874180-4, Hardcover £27,50. Computational Statistics & Data Analysis, 25(2):244, July 31, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397896569.

Anonymous:1997:NNb

[1022] Anonymous. Net news. Computational Statistics & Data Analysis, 25 (2):244–245, July 31, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397896570.

Anonymous:1997:ENI

[1023] Anonymous. EUROSTAT: New information technologies applied to statistics — information highways, multimedia, visualization techniques, computer graphics and signal processing. Computational Statistics & Data Analysis, 25 (2):245–246, July 31, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397896582.

Anonymous:1997:SIM

[1024] Anonymous. Society for the Internet in Medicine. Computational Statistics & Data Analysis, 25(2):

247, July 31, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947397896594.

Anonymous:1997:INg

[1025] Anonymous. IASC news. Computational Statistics & Data Analysis, 25 (2):248–249, July 31, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947397896600.

Anonymous:1997:WDM

[1026] Anonymous. Workshop on data mining & multidimentional data analysis — methodologies, software and applications Naples, Italy: November 19, 1997. Computational Statistics & Data Analysis, 25(2):249– 250, July 31, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947397896612.

Anonymous:1997:CAB

[1027] Anonymous. COMPSTAT ’98 24 to 28 August 1998, Bristol, UK. Computational Statistics & Data Analysis, 25(2):250, July 31, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947397896624.

Hadi:1997:MTL

[1028] Ali S. Hadi and Alberto Luceño. Maximum trimmed likelihood estimators: a unified approach, examples, and algorithms. Computational Statistics & Data Analysis, 25(3):251– 272, August 21, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794739700011X.

Shao:1997:EAE

[1029] Xuesi M. Shao. An efficient algorithm for the exact test on unordered 2 × J contingency tables with equal column sums. Computational Statistics & Data Analysis, 25(3):273– 285, August 21, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397000108.

Gerard:1997:MNR

[1030] Patrick D. Gerard and William R.

Schucany. Methodology for nonparametric regression from independent sources. Computational Statistics & Data Analysis, 25(3):287–304, August 21, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397000054.

Turner:1997:LMU

[1031] T. R. Turner and P. C. Wollan.

Locating a maximum using isotonic regression. Computational Statistics & Data Analysis, 25(3):305–320, August 21, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397000091.

Hirji:1997:RSF

[1032] Karim F. Hirji. A review and a synthesis of the fast Fourier transform algorithms for exact analysis of discrete data. Computational Statistics & Data Analysis, 25(3):321–336, August 21, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794739700008X.

Valliant:1997:AMP

[1033] Richard Valliant and James E. Gentle. An application of mathematical programming to sample allocation. Computational Statistics & Data Analysis, 25(3):337–360, August 21, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947397000078.

Klinke:1997:BKB

[1034] Sigbert Klinke and Dianne Cook. Binning of kernel-based projection pursuit indices in XGobi. Computational Statistics & Data Analysis, 25 (3):363–369, August 21, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947397826047.

Anonymous:1997:GMI

[1035] Anonymous. Guide to medical informatics, the Internet and telemedicine. Computational Statistics & Data Analysis, 25(3):370, August 21, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947397901195.

Anonymous:1997:Te

[1036] Anonymous. Telegram. Computational Statistics & Data Analysis, 25 (3):370, August 21, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397901171.

Anonymous:1997:TF

[1037] Anonymous. Telemedicine forum.

Computational Statistics & Data Analysis, 25(3):370, August 21, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947397901213.

Anonymous:1997:HIJ

[1038] Anonymous. Health informatics journal: Year 2000 and healthcare computing — call for papers. Computational Statistics & Data Analysis, 25 (3):371, August 21, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397901237.

Anonymous:1997:Ib

[1039] Anonymous. Information. Computational Statistics & Data Analysis, 25 (3):371–372, August 21, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397901250.

Anonymous:1997:INh

[1040] Anonymous. IASC news. Computational Statistics & Data Analysis, 25 (3):373, August 21, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397901274.

Anonymous:1997:CMNc

[1041] Anonymous. Calendar of meetings — new entries only. Computational Statistics & Data Analysis, 25(3):373– 375, August 21, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397901298.

Anonymous:1997:CWSg

[1042] Anonymous. Courses, workshops & summer schools — new entries only. Computational Statistics & Data Analysis, 25(3):376, August 21, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397901316.

Aldrin:1997:LMR

[1043] Magne Aldrin. Length modified ridge regression. Computational Statistics & Data Analysis, 25(4):377–398, September 30, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397000157.

Pinheiro:1997:ESR

[1044] Aluisio Pinheiro and Brani Vidakovic. Estimating the square root of a density via compactly supported wavelets. Computational Statistics & Data Analysis, 25(4):399–415, September 30, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397000133.

Tanizaki:1997:NNF

[1045] Hisashi Tanizaki. Nonlinear and nonnormal filters using Monte Carlo methods. Computational Statistics & Data Analysis, 25(4):417–439, September 30, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397000169.

Beier:1997:ATA

[1046] F. Beier and H. Büning. An adaptive test against ordered alternatives. Computational Statistics & Data Analysis, 25(4):441–452, September 30, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397000145.

Oskrochi:1997:SMM

[1047] G. R. Oskrochi and R. B. Davies.

Stayers in mixed Markov renewal models. Computational Statistics & Data Analysis, 25(4):453–464, September 30, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397000170.

Wolfinger:1997:TTS

[1048] Russell D. Wolfinger and Xihong Lin. Two Taylor-series approximation methods for nonlinear mixed models. Computational Statistics & Data Analysis, 25(4):465–490, September 30, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947397000121.

Lenz:1997:BRBb

[1049] Hans-Joachim Lenz. Book review: Methodological Knowledge: Notation and Implementation in Expert Systems: H. J. Adèr (1995): Universiteit van Amsterdam, ISBN 90-90-09022-3, pp. 227. Computational Statistics & Data Analysis, 25(4):493– 494, September 30, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947397807414.

Lenz:1997:BRBc

[1050] Hans-Joachim Lenz. Book review: Programmer’s Guide to Fortran 90, 3rd ed.: Walter S. Brainerd, Jeanne C. Adams, Charles H. Goldberg (1995): Springer, ISBN 0-387-94570-9, pp. 460, DM 58.00. Computational Statistics & Data Analysis, 25 (4):494, September 30, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947397807426.

deFalguerolles:1997:EIC

[1051] Antonie de Falguerolles. ERS of IASC call for nominations. Computational Statistics & Data Analysis, 25(4):494– 495, September 30, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947397901018.

Antoch:1997:REA

[1052] Jaromir Antoch. Report on ERC activities for the use of money. Computational Statistics & Data Analysis, 25 (4):495, September 30, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947397901031.

Anonymous:1997:CMf

[1053] Anonymous. Calendar of meetings. Computational Statistics & Data Analysis, 25(4):496–500, September 30, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947397901055.

Anonymous:1997:CWSh

[1054] Anonymous. Courses, workshops & summer schools. Computational Statistics & Data Analysis, 25(4):500–501, September 30, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947397901079.

Anonymous:1997:AIS

[1055] Anonymous. Author index of SS-NinCSDA: Volume 23–25 (november 1996–September 1997). Computational Statistics & Data Analysis, 25 (4):502, September 30, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947397807438.

Anonymous:1997:BRS

[1056] Anonymous. Book reviews of SS-NinCSDA: Volume 23–25 (november 1996–September 1997). Computational Statistics & Data Analysis, 25 (4):503, September 30, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947397901092.

Anonymous:1997:INi

[1057] Anonymous. IASC news. Computational Statistics & Data Analysis, 25(4):504–505, September 30, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/0167947397901110.

Anonymous:1997:Tg

[1058] Anonymous. Telegram. Computational Statistics & Data Analysis, 25 (4):505, September 30, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947397901134.

Anonymous:1997:ILN

[1059] Anonymous. Information, literature, net news. Computational Statistics & Data Analysis, 25(4):506, September 30, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947397901158.

Anonymous:1997:AIVc

[1060] Anonymous. Author index volume 25 (1997). Computational Statistics & Data Analysis, 25(4):507–508, September 30, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794739780744X.

Anonymous:1997:EBc

[1061] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 26 (1):v–vi, November 6, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739790089X.

Smith:1997:SMT

[1062] S. P. Smith. Sparse matrix tools for Gaussian models on lattices. Computational Statistics & Data Analysis, 26 (1):1–15, November 6, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397000236.

Struyf:1997:IRC

[1063] Anja Struyf, Mia Hubert, and Peter J.

Rousseeuw. Integrating robust clustering techniques in S-PLUS. Computational Statistics & Data Analysis, 26 (1):17–37, November 6, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397000200.

Barrett:1997:LRI

[1064] Bruce E. Barrett and J. Brian Gray.

Leverage, residual, and interaction diagnostics for subsets of cases in least squares regression. Computational Statistics & Data Analysis, 26 (1):39–52, November 6, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397000224.

Hilton:1997:SPE

[1065] Joan F. Hilton and Lauren Gee. The size and power of the exact bivariate symmetry test. Computational Statistics & Data Analysis, 26(1):53– 69, November 6, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397000212.

Chen:1997:MRS

[1066] Lin-An Chen. Multivariate regression splines. Computational Statistics & Data Analysis, 26(1):71–82, November 6, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397000182.

Rohmel:1997:PDF

[1067] Joachim Röhmel. The permutation distribution of the Friedman test. Computational Statistics & Data Analysis, 26 (1):83–99, November 6, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397000194.

McCulloch:1997:E

[1068] J. Huston McCulloch and Don B. Panton. Erratum. Computational Statistics & Data Analysis, 26(1): 101, November 6, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397000248.

Antoch:1997:SMI

[1069] Jaromír Antoch. Satellite meeting on industrial statistics: Aims and computational aspects August 16–17, 1997, Athens, Greece. Computational Statistics & Data Analysis, 26 (1):105, November 6, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947397900918.

Salmaso:1997:ISC

[1070] Luigi Salmaso. ISI satellite conference on industrial statistics: Aims and computational aspects, Athens, Greece, 16–17 August 1997. Computational Statistics & Data Analysis, 26(1):105– 106, November 6, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947397900931.

Anonymous:1997:S

[1071] Anonymous. Software. Computational Statistics & Data Analysis, 26 (1):106, November 6, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947397900955.

Ugarte:1997:BRBa

[1072] M. Dolores Ugarte. Book review: Probability — A survey of the mathematical theory: John W. Lamperti (1996): 2nd ed. New York: Wiley, ISBN 0-471-15407-5, pp. 189, £32.50. Computational Statistics & Data Analysis, 26 (1):107, November 6, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397821044.

Ugarte:1997:BRBb

[1073] M. Dolores Ugarte. Book review: Probability and statistical inference: Robert Bartoszynski and Magdalena Niewiadomska-Bugaj (1996): Wiley, New York, pp. 826, ISBN 0-471-31073-5, £50.00. Computational Statistics & Data Analysis, 26(1):107–108, November 6, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397821056.

Lenz:1997:BRBd

[1074] Hans-Joachim Lenz. Book review: Ordinal and symbolic data analysis: E. Diday, Y. Lechevallier, O. Opitz (eds.) (1996): Springer, ISBN 3-540-61081-2; pp. 372, DM 135.00. Computational Statistics & Data Analysis, 26(1):108– 109, November 6, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397821068.

Anonymous:1997:INj

[1075] Anonymous. IASC news. Computational Statistics & Data Analysis, 26(1):109–114, November 6, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/0167947397900979.

Anonymous:1997:CMNd

[1076] Anonymous. Calendar of meetings — new entries only. Computational Statistics & Data Analysis, 26(1):114– 116, November 6, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947397900992.

Huffer:1997:CED

[1077] Fred W. Huffer and Chien-Tai Lin.

Computing the exact distribution of the extremes of sums of consecutive spacings. Computational Statistics & Data Analysis, 26(2):117–132, December 4, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397000273.

Lee:1997:SSM

[1078] A. J. Lee. Some simple methods for generating correlated categorical variates. Computational Statistics & Data Analysis, 26(2):133–148, December 4, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397000303.

Routledge:1997:QMC

[1079] Rick Routledge and Min Tsao. Quadrature methods for computing distributions. Computational Statistics & Data Analysis, 26(2):149–161, December 4, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947397000297.

Guerra:1997:SBC

[1080] Rudy Guerra, Alan M. Polansky, and William R. Schucany. Smoothed bootstrap confidence intervals with discrete data. Computational Statistics & Data Analysis, 26(2):163–176, December 4, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947397000261.

Yuan:1997:IPT

[1081] Ke-Hai Yuan and Peter M. Bentler. Improving parameter tests in covariance structure analysis. Computational Statistics & Data Analysis, 26(2):177– 198, December 4, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794739700025X.

Cwik:1997:CAM

[1082] J. Čwik and J. Koronacki. A combined adaptive-mixtures/plug-in estimator of multivariate probability densities. Computational Statistics & Data Analysis, 26(2):199–218, December 4, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947397000327.

Mays:1997:DAT

[1083] Darcy P. Mays and Raymond H. Myers. Design and analysis for a two-level factorial experiment in the presence of variance heterogeneity. Computational Statistics & Data Analysis, 26(2):219– 233, December 4, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397000285.

Andres:1997:E

[1084] A. Martín Andrés and A. Silva Mato. Erratum. Computational Statistics & Data Analysis, 26(2): 235, December 4, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397000315.

May:1997:SMM

[1085] Warren L. May and William D. Johnson. A SAS^{R} macro for the multivariate extension of the Kruskal– Wallis test including multiple comparisons: Randomization and κ^{2} criteria. Computational Statistics & Data Analysis, 26(2):239–250, December 4, 1997. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794739782107X.

Anonymous:1997:Ic

[1086] Anonymous. Information. Computational Statistics & Data Analysis, 26 (2):251, December 4, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397900578.

Anonymous:1997:INk

[1087] Anonymous. IASC news. Computational Statistics & Data Analysis, 26(2):252–256, December 4, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947397900591.

Anonymous:1997:CMNe

[1088] Anonymous. Calendar of meetings — new entries only. Computational Statistics & Data Analysis, 26(2):256– 258, December 4, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794739790061X.

Caers:1998:BCI

[1089] Jef Caers, Jan Beirlant, and Petra Vynckier. Bootstrap confidence intervals for tail indices. Computational Statistics & Data Analysis, 26 (3):259–277, January 2, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947397000339.

Larose:1998:MPB

[1090] Daniel T. Larose and Dipak K. Dey. Modeling publication bias using weighted distributions in a Bayesian framework. Computational Statistics & Data Analysis, 26(3):279–302, January 2, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794739700039X.

Lee:1998:SVL

[1091] Jae Won Lee and Harland N. Sather.

A supremum version of logrank test for detecting late occurring survival differences. Computational Statistics & Data Analysis, 26(3):303–311, January 2, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397000352.

Luceno:1998:MOD

[1092] Alberto Luceño. Multiple outliers detection through reweighted least deviances. Computational Statistics & Data Analysis, 26(3):313–326, January 2, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397000364.

Molenberghs:1998:MLC

[1093] Geert Molenberghs, Lieven Declerck, and Marc Aerts. Misspecifying the likelihood for clustered binary data. Computational Statistics & Data Analysis, 26(3):327–349, January 2, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397000376.

Bauer:1998:SPA

[1094] M. Bauer, P. Bauer, and M. Budde.

A simulation program for adaptive two stage designs. Computational Statistics & Data Analysis, 26(3):351– 371, January 2, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397000340.

Knusel:1998:ASD

[1095] Leo Knüsel. On the accuracy of statistical distributions in Microsoft Excel 97. Computational Statistics & Data Analysis, 26(3):375–377, January 2, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397817562.

Kramer:1998:BRB

[1096] Walter Krämer. Book review: Time Series Analysis — Nonstationary and noninvertible distribution theory: Katsuo Tanaka (1996): New York: Wiley, ISBN 0-471-14191-7, x + 623 pages, $70.00. Computational Statistics & Data Analysis, 26(3):378, January 2, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397817574.

Trenkler:1998:BRB

[1097] Götz Trenkler. Book review: Methods and Applications of Linear Models — Regression and the Analysis of Variance: Ronald R. Hocking (1996): Wiley Series in Probability and Statistics, ISBN 0-471-59282-X, pp. 731, £55.00. Computational Statistics & Data Analysis, 26 (3):378–379, January 2, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397817586.

Galway:1998:ICP

[1098] Lionel Galway. Information: Call for papers 10^{th} International Conference on Scienctific and Statistical Database Management (SSDBM) Capri, Italy — July 1–3, 1998. Computational Statistics & Data Analysis, 26(3):379, January 2, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398900536.

Anonymous:1998:INa

[1099] Anonymous. IASC news. Computational Statistics & Data Analysis, 26 (3):381–384, January 2, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794739890055X.

Anonymous:1998:CMa

[1100] Anonymous. Calendar of meetings. Computational Statistics & Data Analysis, 26(3):385–390, January 2, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398900573.

Anonymous:1998:CWSa

[1101] Anonymous. Courses, workshops & summer schools. Computational Statistics & Data Analysis, 26(3):391– 392, January 2, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398900597.

Hudson:1998:MLR

[1102] H. Malcolm Hudson and Thomas C. M. Lee. Maximum likelihood restoration and choice of smoothing parameter in deconvolution of image data subject to Poisson noise. Computational Statistics & Data Analysis, 26(4):393– 410, February 6, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397000418.

Korhonen:1998:OPC

[1103] Pekka Korhonen and Aapo Siljamäki. Ordinal principal component analysis theory and an application. Computational Statistics & Data Analysis, 26(4):411–424, February 6, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397000388.

Lawson:1998:QMI

[1104] John Lawson, Scott Grimshaw, and Jason Burt. A quantitative method for identifying active contrasts in unreplicated factorial designs based on the half-normal plot. Computational Statistics & Data Analysis, 26(4):425– 436, February 6, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397000406.

McLeod:1998:DCM

[1105] A. Ian McLeod. Diagnostic check for monotone spread. Computational Statistics & Data Analysis, 26(4):437– 443, February 6, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794739700042X.

Saltelli:1998:AWC

[1106] Andrea Saltelli and Ricardo Bolado. An alternative way to compute Fourier amplitude sensitivity test (FAST). Computational Statistics & Data Analysis, 26(4):445–460, February 6, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397000431.

Wu:1998:MCP

[1107] Shu-Fei Wu and Hubert J. Chen.

Multiple comparison procedures with the average for exponential location parameters. Computational Statistics & Data Analysis, 26(4):461–484, February 6, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397000443.

Anonymous:1998:NNa

[1108] Anonymous. Net news. Computational Statistics & Data Analysis, 26(4):487– 488, February 6, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398901153.

Anonymous:1998:CFS

[1109] Anonymous. ContentsDirect, a free service for all Elsevier Science journals. Computational Statistics & Data Analysis, 26(4):488–489, February 6, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947398901177.

Anonymous:1998:INb

[1110] Anonymous. IASC news. Computational Statistics & Data Analysis, 26(4):489–490, February 6, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947398901190.

Lane:1998:CRF

[1111] Peter Lane. COMPSTAT’98 registration form (please print in a fixedwidth font). Computational Statistics & Data Analysis, 26(4):491–492, February 6, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947398901219.

Anonymous:1998:PEW

[1112] Anonymous. Prof. Ed Wegman — profile of the new President (1997–1999). Computational Statistics & Data Analysis, 26(4):493–494, February 6, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947398901232.

Hudson:1998:CIO

[1113] Malcolm Hudson. Current IASC officers IASC executives 1997–1999. Computational Statistics & Data Analysis, 26(4):494–495, February 6, 1998.

CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947398901256.

Anonymous:1998:CMNa

[1114] Anonymous. Calendar of meetings — new entries only. Computational Statistics & Data Analysis, 26(4):496– 498, February 6, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794739890127X.

Anonymous:1998:AIVa

[1115] Anonymous. Author index volume 26 (1998). Computational Statistics & Data Analysis, 26(4):499–500, February 6, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397804355.

Anonymous:1998:EBa

[1116] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 27(1):v–vi, March 6, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947398900056.

Hirji:1998:AFF

[1117] Karim F. Hirji. Assessing fast Fourier transform algorithms. Computational Statistics & Data Analysis, 27(1):1– 9, March 6, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397000522.

Furno:1998:EVL

[1118] Marilena Furno. Estimating the variance of the LAD regression coefficients. Computational Statistics & Data Analysis, 27(1):11–26, March 6, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397000479.

Koukouvinos:1998:WEO

[1119] Christos Koukouvinos. Weighing experiments with n ≡ 2 (mod 4) observations. Computational Statistics & Data Analysis, 27(1):27–31, March 6, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397000509.

Tang:1998:ASE

[1120] Man-Lai Tang and Sik-Yum Lee. Analysis of structural equation models with censored or truncated data via EM algorithm. Computational Statistics & Data Analysis, 27(1):33– 46, March 6, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397800400.

Albers:1998:CBR

[1121] W. Albers, W. C. M. Kallenberg, and G. D. Otten. A comparison between Rosenblatt’s estimator and parametric density estimators for determining test limits. Computational Statistics & Data Analysis, 27(1):47– 60, March 6, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397000492.

Bisaglia:1998:CTE

[1122] Luisa Bisaglia and Dominique Guégan.

A comparison of techniques of estimation in long-memory processes. Computational Statistics & Data Analysis, 27(1):61–81, March 6, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397000455.

Carling:1998:ECG

[1123] Kenneth Carling and Hans Söderberg.

An experimental comparison of gradient methods in econometric duration analysis. Computational Statistics & Data Analysis, 27(1):83–97, March 6, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397000480.

Yeh:1998:SIG

[1124] Lam Yeh and So Kuen Chan. Statistical inference for geometric processes with lognormal distribution. Computational Statistics & Data Analysis, 27 (1):99–112, March 6, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397000467.

Hirose:1998:EML

[1125] H. Hirose. Erratum to “Maximum likelihood parameter estimation in the three-parameter gamma distribution”:

[computational statistics & data analysis 20, 343–354 (1995)]. Computational Statistics & Data Analysis, 27(1):113, March 6, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397800412. See

[695].

Anonymous:1998:Ra

[1126] Anonymous. Referees. Computational Statistics & Data Analysis, 27 (1):114, March 6, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739890007X.

Anonymous:1998:CPC

[1127] Anonymous. Call for papers: Computational Statistics & Data Analysis: Special Issue on parallel processing and statistics. Computational Statistics & Data Analysis, 27 (1):117, March 6, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947398900093.

Anonymous:1998:Sa

[1128] Anonymous. Software. Computational Statistics & Data Analysis, 27 (1):117–119, March 6, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947398900135.

Anonymous:1998:NNb

[1129] Anonymous. Net news. Computational Statistics & Data Analysis, 27 (1):119–120, March 6, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947398900159.

Anonymous:1998:CMNb

[1130] Anonymous. Calendar of meetings — new entries only. Computational Statistics & Data Analysis, 27(1):120– 124, March 6, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947398900172.

Arnold:1998:UCC

[1131] Barry C. Arnold, Enrique Castillo, and JoséMaría Sarabia. The use of conditionally conjugate priors in the study of ratios of gamma scale parameters. Computational Statistics & Data Analysis, 27(2):125– 139, April 3, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000036.

Cain:1998:AQL

[1132] Michael Cain. Asymmetric quadratic loss adjustments for a predictive t variable. Computational Statistics & Data Analysis, 27(2):141– 150, April 3, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000103.

Hinde:1998:OME

[1133] John Hinde and Clarice G. B.

Demétrio. Overdispersion: Models and estimation. Computational Statistics & Data Analysis, 27(2):151– 170, April 3, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000073.

Min:1998:GSA

[1134] Chung ki Min. A Gibbs sampling approach to estimation and prediction of time-varying-parameter models. Computational Statistics & Data Analysis, 27(2):171–194, April 3, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397000546.

Lee:1998:BAC

[1135] Chung-Bow Lee. Bayesian analysis of a change-point in exponential families with applications. Computational Statistics & Data Analysis, 27 (2):195–208, April 3, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000097.

Navidi:1998:SMS

[1136] William Navidi, William S. Murphy, Jr., and Willy Hereman. Statistical methods in surveying by trilateration. Computational Statistics & Data Analysis, 27(2):209– 227, April 3, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947397000534.

Park:1998:FCG

[1137] Chongsun Park and Sanford Weisberg.

Fisher consistency of GEE models under link misspecification. Computational Statistics & Data Analysis, 27 (2):229–235, April 3, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000048.

Dessens:1998:UCL

[1138] Jos Dessens, Wim Jansen, and Peter G. M. van der Heijden. Using covariates in loglinear models with sampling zeros; a cautionary note. Computational Statistics & Data Analysis, 27 (2):239–245, April 3, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398800030.

Anonymous:1998:Ta

[1139] Anonymous. Telegram. Computational Statistics & Data Analysis, 27 (2):246–247, April 3, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398901335.

Anonymous:1998:NNc

[1140] Anonymous. Net news. Computational Statistics & Data Analysis, 27 (2):247–249, April 3, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398901359.

Anonymous:1998:STS

[1141] Anonymous. Special technical session on high performance computing for pattern recognition at the 1998 international conference on parallel and distributed processing techniques and applications (PDPTA’98), July 13–16, 1998, Las Vegas, Nevada, USA call for papers. Computational Statistics & Data Analysis, 27(2): 249, April 3, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398901372.

Anonymous:1998:CMb

[1142] Anonymous. Calendar of meetings. Computational Statistics & Data Analysis, 27(2):250–256, April 3, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947398901396.

Tang:1998:TMC

[1143] Man-Lai Tang and Peter M. Bentler. Theory and method for constrained estimation in structural equation models with incomplete data. Computational Statistics & Data Analysis, 27 (3):257–270, May 1, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000152.

vanderGeest:1998:AGS

[1144] P. A. G. van der Geest. An algorithm to generate samples of multivariate distributions with correlated marginals. Computational Statistics & Data Analysis, 27(3):271– 289, May 1, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794739800005X.

Nordbotten:1998:EPP

[1145] Svein Nordbotten. Estimating population proportions from imputed data. Computational Statistics & Data Analysis, 27(3):291– 309, May 1, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000115.

Vichi:1998:PCA

[1146] Maurizio Vichi. Principal classifications analysis: a method for generating consensus dendrograms and its application to three-way data. Computational Statistics & Data Analysis, 27 (3):311–331, May 1, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000061.

Wong:1998:CAT

[1147] Augustine C. M. Wong and Jianrong Wu. Comparisons of approximate tail probabilities for the shape parameter of the gamma distribution. Computational Statistics & Data Analysis, 27 (3):333–344, May 1, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000140.

Wang:1998:FMR

[1148] Huiwen Wang and Qiang Liu. Forecast modelling for rotations of principal axes of multidimensional data sets. Computational Statistics & Data Analysis, 27(3):345– 354, May 1, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947398000085.

Frick:1998:NCM

[1149] Hans Frick. Nonparametric comparisons of medians of survival functions. Computational Statistics & Data Analysis, 27(3):357– 361, May 1, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947398900779.

Anonymous:1998:NNd

[1150] Anonymous. Net news. Computational Statistics & Data Analysis, 27 (3):362–364, May 1, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947398900792.

Anonymous:1998:Ia

[1151] Anonymous. Information. Computational Statistics & Data Analysis, 27 (3):364–365, May 1, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947398900810.

Anonymous:1998:INc

[1152] Anonymous. IASC news. Computational Statistics & Data Analysis, 27 (3):365–367, May 1, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398900834.

Anonymous:1998:CMNc

[1153] Anonymous. Calendar of meetingsnew entries only. Computational Statistics & Data Analysis, 27(3):368– 370, May 1, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398900858.

Dunmur:1998:PEL

[1154] A. P. Dunmur and D. M. Titterington. Parameter estimation in latent profile models. Computational Statistics & Data Analysis, 27(4):371– 388, June 5, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794739800019X.

Wang:1998:ECR

[1155] Qin Wang. Exact calculations for the repeated many-one test. Computational Statistics & Data Analysis, 27 (4):389–399, June 5, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000206.

Chan:1998:TCS

[1156] I. S. F. Chan, D. Hillman, and T. A.

Louis. Treatment comparisons with screenable endpoints. Computational Statistics & Data Analysis, 27(4):401– 419, June 5, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000176.

Famoye:1998:LGD

[1157] Felix Famoye and Z. Govindarajulu. On the Lagrange gamma distribution. Computational Statistics & Data Analysis, 27(4):421– 431, June 5, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000188.

Khuri:1998:GAE

[1158] AndréI. Khuri and Juneyoung Lee. A graphical approach for evaluating and comparing designs for nonlinear models. Computational Statistics & Data Analysis, 27(4):433– 443, June 5, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000164.

Schott:1998:ECM

[1159] James R. Schott. Estimating correlation matrices that have common eigenvectors. Computational Statistics & Data Analysis, 27(4):445– 459, June 5, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000279.

Sebert:1998:CAI

[1160] David M. Sebert, Douglas C. Montgomery, and Dwayne A. Rollier. A clustering algorithm for identifying multiple outliers in linear regression. Computational Statistics & Data Analysis, 27(4):461–484, June 5, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000218.

Antoch:1998:RPN

[1161] Jaromír Antoch, Jean-Marc Deshouillérs, and Gusti Putu Purnaba. Revisiting the pseudorandom number generator ran1 from the Numerical Recipes. Computational Statistics & Data Analysis, 27(4):487– 495, June 5, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398800029. See

[1014].

Anonymous:1998:Tb

[1162] Anonymous. Telegram. Computational Statistics & Data Analysis, 27 (4):496, June 5, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398900871.

Anonymous:1998:NNe

[1163] Anonymous. Net news. Computational Statistics & Data Analysis, 27 (4):496–497, June 5, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398900895.

Anonymous:1998:CMNd

[1164] Anonymous. Calendar of meetings — new entries only. Computational Statistics & Data Analysis, 27(4):497– 498, June 5, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947398900913.

Anonymous:1998:AIVb

[1165] Anonymous. Author index volume 27 (1998). Computational Statistics & Data Analysis, 27(4):499– 500, June 5, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947398000498.

Anonymous:1998:EBb

[1166] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 28(1):v–vi, July 17, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947398900937.

Khattree:1998:SPE

[1167] Ravindra Khattree. Some practical estimation procedures for variance components. Computational Statistics & Data Analysis, 28(1):1– 32, July 17, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947398000243.

Pan:1998:CND

[1168] Wei Pan and Rick Chappell. Computation of the NPMLE of distribution functions for interval censored and truncated data with applications to the Cox model. Computational Statistics & Data Analysis, 28(1):33– 50, July 17, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000267.

Bozdogan:1998:ICC

[1169] Hamparsum Bozdogan and Dominique M. A. Haughton. Informational complexity criteria for regression models. Computational Statistics & Data Analysis, 28(1):51– 76, July 17, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000255.

Damien:1998:BBF

[1170] Paul Damien and Peter Müller. A Bayesian bivariate failure time regression model. Computational Statistics & Data Analysis, 28(1):77– 85, July 17, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794739800022X.

Hoeting:1998:BPS

[1171] Jennifer A. Hoeting and Joseph G.

Ibrahim. Bayesian predictive simultaneous variable and transformation selection in the linear model. Computational Statistics & Data Analysis, 28 (1):87–103, July 17, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947398000280.

Liu:1998:DIC

[1172] Aiyi Liu. Detecting influential covariates in a growth curve model. Computational Statistics & Data Analysis, 28 (1):105–113, July 17, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947398000231.

Anonymous:1998:Tc

[1173] Anonymous. Telegram. Computational Statistics & Data Analysis, 28 (1):117–118, July 17, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947398900950.

Thomas:1998:NN

[1174] Shannon Thomas. Net news. Computational Statistics & Data Analysis, 28 (1):118–121, July 17, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947398900974.

Anonymous:1998:Ib

[1175] Anonymous. Information. Computational Statistics & Data Analysis, 28 (1):121–122, July 17, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947398900998.

Soofi:1998:IN

[1176] Ehsan S. Soofi. IASC news. Computational Statistics & Data Analysis, 28 (1):122–123, July 17, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398901013.

Anonymous:1998:CMc

[1177] Anonymous. Calendar of meetings.

Computational Statistics & Data Analysis, 28(1):124–128, July 17, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947398901037.

Anonymous:1998:CWSb

[1178] Anonymous. Courses, workshops & summer schools. Computational Statistics & Data Analysis, 28(1):128– 130, July 17, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398901050.

Greenwood:1998:URW

[1179] Garrison W. Greenwood and Xiaobo (Sharon) Hu. On the use of random walks to estimate correlation in fitness landscapes. Computational Statistics & Data Analysis, 28 (2):131–137, August 7, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000309.

Agresti:1998:ORI

[1180] Alan Agresti and Brent A. Coull.

Order-restricted inference for monotone trend alternatives in contingency tables. Computational Statistics & Data Analysis, 28(2):139– 155, August 7, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947398000358.

Buttrey:1998:NNC

[1181] Samuel E. Buttrey. Nearest-neighbor classification with categorical variables. Computational Statistics & Data Analysis, 28(2):157–169, August 7, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947398000322.

Grunwald:1998:SNG

[1182] Gary K. Grunwald and Rob J. Hyndman. Smoothing non-Gaussian time series with autoregressive structure. Computational Statistics & Data Analysis, 28(2):171–191, August 7, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947398000346.

Marx:1998:DGA

[1183] Brian D. Marx and Paul H. C. Eilers. Direct generalized additive modeling with penalized likelihood. Computational Statistics & Data Analysis, 28 (2):193–209, August 7, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947398000334.

Silver:1998:OED

[1184] G. L. Silver. Operational equations for data in rectangular array. Computational Statistics & Data Analysis, 28 (2):211–215, August 7, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000292.

Best:1998:SSC

[1185] D. J. Best, J. C. W. Rayner, and L. G. Stephens. Small-sample comparison of McCullagh and nair analyses for nominal-ordinal categorical data. Computational Statistics & Data Analysis, 28(2):217– 223, August 7, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000310.

Zhao:1998:MCP

[1186] J. H. Zhao and P. C. Sham. A method for calculating probability convolution using “ternary” numbers with application in the determination of twin zygosity. Computational Statistics & Data Analysis, 28(2):225– 232, August 7, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794739800036X.

Anonymous:1998:Td

[1187] Anonymous. Telegram. Computational Statistics & Data Analysis, 28 (2):235, August 7, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398900615.

Anonymous:1998:Te

[1188] Anonymous. Telegram. Computational Statistics & Data Analysis, 28 (2):235, August 7, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947398900639.

Anonymous:1998:NNf

[1189] Anonymous. Net news. Computational Statistics & Data Analysis, 28 (2):236–237, August 7, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947398900652.

Anonymous:1998:NNg

[1190] Anonymous. Net news. Computational Statistics & Data Analysis, 28 (2):237–238, August 7, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947398900676.

Anonymous:1998:ASA

[1191] Anonymous. American Statistical Association Statistical Computing Section Student Paper Competition Deadline: January 8, 1999. Computational Statistics & Data Analysis, 28 (2):238, August 7, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794739890069X.

Anonymous:1998:CHS

[1192] Anonymous. Comprehensive handbook of statistical procedures. Computational Statistics & Data Analysis, 28 (2):238–239, August 7, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398900718.

Anonymous:1998:CMNe

[1193] Anonymous. Calendar of meetings — new entries only. Computational Statistics & Data Analysis, 28(2):239– 240, August 7, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398900731.

Anonymous:1998:CWSc

[1194] Anonymous. Courses, workshops and summer schools — new entries only. Computational Statistics & Data Analysis, 28(2):240– 242, August 7, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398900755.

Park:1998:AGE

[1195] Taesung Park, Charles S. Davis, and Ning Li. Alternative Gee estimation procedures for discrete longitudinal data. Computational Statistics & Data Analysis, 28(3):243–256, September 4, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000395.

Zani:1998:RBB

[1196] Sergio Zani, Marco Riani, and Aldo Corbellini. Robust bivariate boxplots and multiple outlier detection. Computational Statistics & Data Analysis, 28(3):257–270, September 4, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000401.

Agarwal:1998:REE

[1197] S. K. Agarwal and P. Kumar. On the relative efficiency of estimators of population total in unequal probability sampling when study variable has weak relationship with size variable. Computational Statistics & Data Analysis, 28(3):271–281, September 4, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000371.

Bates:1998:LDL

[1198] R. A. Bates, E. Riccomagno, R. Schwabe, and H. P. Wynn. Lattices and dual lattices in optimal experimental design for Fourier models. Computational Statistics & Data Analysis, 28(3):283– 296, September 4, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000425.

Cheung:1998:SPI

[1199] Siu Hung Cheung, Ka Ho Wu, and Wai Sum Chan. Simultaneous prediction intervals for autoregressiveintegrated moving-average models: a comparative study. Computational Statistics & Data Analysis, 28(3):297– 306, September 4, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000383.

Kiers:1998:TWS

[1200] Henk A. L. Kiers. Three-way SIM-PLIMAX for oblique rotation of the three-mode factor analysis core to simple structure. Computational Statistics & Data Analysis, 28(3):307–324, September 4, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000413.

Pesarin:1998:T

[1201] Fortunato Pesarin and Luigi Salmaso.

Telegram. Computational Statistics & Data Analysis, 28(3):327, September 4, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947398901074.

Asano:1998:IN

[1202] Chooichiro Asano. IASC news. Computational Statistics & Data Analysis, 28(3):327–330, September 4, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/0167947398901098.

Anonymous:1998:CMNf

[1203] Anonymous. Calendar of meetings — new entries only. Computational Statistics & Data Analysis, 28(3):330– 335, September 4, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947398901116.

Anonymous:1998:CWSd

[1204] Anonymous. Courses, workshops and summer schools — new entries only. Computational Statistics & Data Analysis, 28(3):335–338, September 4, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739890113X.

Anonymous:1998:Rb

[1205] Anonymous. Referees. Computational Statistics & Data Analysis, 28 (4):iii, October 28, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947398901293.

Pardo:1998:IAG

[1206] M. C. Pardo. Improving the accuracy of goodness-of-fit tests based on Rao’s divergence with small sample size. Computational Statistics & Data Analysis, 28(4):339–351, October 28, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947398901311.

Antoniadis:1998:WRR

[1207] Anestis Antoniadis and Dinh Tuan Pham. Wavelet regression for random or irregular design. Computational Statistics & Data Analysis, 28(4):353– 369, October 28, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398901451.

Fang:1998:MVG

[1208] Kai-Tai Fang, Run-Ze Li, and Jia-Juan Liang. A multivariate version of Ghosh’s T_{3}-plot to detect nonmultinormality. Computational Statistics & Data Analysis, 28(4):371–386, October 28, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398901475.

Lapp:1998:MAB

[1209] Krista Lapp, Geert Molenberghs, and Emmanuel Lesaffre. Models for the association between ordinal variables. Computational Statistics & Data Analysis, 28(4):387–411, October 28, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398901499.

Langsrud:1998:UFS

[1210] Øyvind Langsrud and Tormod Næs.

A unified framework for significance testing in fractional factorials. Computational Statistics & Data Analysis, 28(4):413–431, October 28, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398901517.

Edler:1998:IEC

[1211] Lutz Edler. IASC elections 1999 call for CANDIDATES. Computational Statistics & Data Analysis, 28 (4):435, October 28, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398901530.

Kitsos:1998:SSE

[1212] Christos P. Kitsos and Jaromír Antoch. 5th Summer School of ERS IASC 3rd Millennium Challenge for Industrial Statistics. Computational Statistics & Data Analysis, 28(4): 435, October 28, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398901554.

Anonymous:1998:XCC

[1213] Anonymous. XIV COMPSTAT Conference of the International Association for Statistical Computing 21 to 25 August 2000, University of Utrecht, The Netherlands. Computational Statistics & Data Analysis, 28(4):435–436, October 28, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398901578.

Anonymous:1998:CMd

[1214] Anonymous. Calendar of meetings. Computational Statistics & Data Analysis, 28(4):436–443, October 28, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398901591.

Anonymous:1998:CWSe

[1215] Anonymous. Courses, workshops & summer schools. Computational Statistics & Data Analysis, 28(4):444–446, October 28, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794739890161X.

Anonymous:1998:AIS

[1216] Anonymous. Author index of SS-NinCSDA: Volumes 26–28 (October 1997–October 1998). Computational Statistics & Data Analysis, 28(4): 447, October 28, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398901633.

Anonymous:1998:PRM

[1217] Anonymous. Participants’ reflextions on meetings of SSNinCSDA: Volumes 26–28 (october 1997–October 1998). Computational Statistics & Data Analysis, 28(4):447, October 28, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398901657.

Anonymous:1998:BRS

[1218] Anonymous. Book reviews of SS-NinCSDA: Volumes 26–28 (October 1997–October 1998). Computational Statistics & Data Analysis, 28(4):447– 448, October 28, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398901670.

Anonymous:1998:TSV

[1219] Anonymous. Telegram, software: Volumes 26–28 (october 1997–October 1998). Computational Statistics & Data Analysis, 28(4):449, October 28, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947398901694.

Anonymous:1998:EIN

[1220] Anonymous. Editorials, information, net news: Volumes 26–28 (october 1997–October 1998). Computational Statistics & Data Analysis, 28 (4):450, October 28, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947398901712.

Anonymous:1998:AIVc

[1221] Anonymous. Author index volume 28 (1998). Computational Statistics & Data Analysis, 28(4):451–452, October 28, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947398901736.

Anonymous:1998:EBc

[1222] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 29 (1):v–vi, November 28, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947398901785.

Cressie:1998:IAP

[1223] Noel Cressie and Jennifer L. Davidson. Image analysis with partially ordered Markov models. Computational Statistics & Data Analysis, 29 (1):1–26, November 28, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000528.

Eubank:1998:EPL

[1224] R. L. Eubank, E. L. Kambour, J. T.

Kim, K. Klipple, C. S. Reese, and M. Schimek. Estimation in partially linear models. Computational Statistics & Data Analysis, 29(1):27– 34, November 28, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000541.

Fishman:1998:BWC

[1225] George S. Fishman and David S. Rubin. Best- and worst-case variances when bounds are available for the distribution function. Computational Statistics & Data Analysis, 29(1):35– 53, November 28, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000516.

Dupuis:1998:PQE

[1226] D. J. Dupuis. Parameter and quantile estimation for a fatigue model. Computational Statistics & Data Analysis, 29(1):55–68, November 28, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947398000504.

Yick:1998:UOT

[1227] John S. Yick and Andy H. Lee. Unmasking outliers in two-way contingency tables. Computational Statistics & Data Analysis, 29(1):69–79, November 28, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794739800053X.

Karlis:1998:MHD

[1228] Dimitris Karlis and Evdokia Xekalaki. Minimum Hellinger distance estimation for Poisson mixtures. Computational Statistics & Data Analysis, 29(1):81– 103, November 28, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947398000474.

Andres:1998:FMV

[1229] A. Martín Andrés, M. J. Sánchez Quevedo, and A. Silva Mato. Fisher’s mid-P -value arrangement in 2 × 2 comparative trials. Computational Statistics & Data Analysis, 29(1):107– 115, November 28, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947398901797.

Aluja-Banet:1998:FLC

[1230] Tomàs Aluja-Banet. Feelings from last COMPSTAT. Computational Statistics & Data Analysis, 29(1): 116, November 28, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398901803.

Kaarik:1998:CBB

[1231] Ene Käärik. Compstat’98 — a bridge between statistical theory and computational applied statistics. Computational Statistics & Data Analysis, 29 (1):117, November 28, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398901815.

Nash:1998:PVM

[1232] John C. Nash. Participants’ views of meetings: Compstat 98 — one person’s view. Computational Statistics & Data Analysis, 29(1):118–119, November 28, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398901827.

Anonymous:1998:Sb

[1233] Anonymous. Software. Computational Statistics & Data Analysis, 29(1):120– 124, November 28, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398901839.

Anonymous:1998:INd

[1234] Anonymous. IASC news. Computational Statistics & Data Analysis, 29(1):124–127, November 28, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947398901840.

Anonymous:1998:LN

[1235] Anonymous. Latest news. Computational Statistics & Data Analysis, 29(1):127–128, November 28, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947398901852.

Chen:1998:RTE

[1236] Yi-Ju Chen, John Hewett, Yazhen Wang, and Jane Johnson. A rank test for equality of two multivariate populations vs a particular ordered alternative. Computational Statistics & Data Analysis, 29(2):129–144, December 28, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000723.

Kosinski:1998:PDM

[1237] Andrzej S. Kosinski. A procedure for the detection of multivariate outliers. Computational Statistics & Data Analysis, 29(2):145–161, December 28, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000735.

LeBlanc:1998:RQF

[1238] Michael LeBlanc. Resampled quantile functions for error estimation and a relationship to density estimation. Computational Statistics & Data Analysis, 29(2):163–176, December 28, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000711.

Mansouri:1998:MAV

[1239] H. Mansouri. Multifactor analysis of variance based on the aligned rank transform technique. Computational Statistics & Data Analysis, 29(2):177– 189, December 28, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000772.

Caers:1998:NTE

[1240] Jef Caers and Jozef Van Dyck. Nonparametric tail estimation using a double bootstrap method. Computational Statistics & Data Analysis, 29(2):191– 211, December 28, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000607.

Li:1998:CNA

[1241] Baibing Li and Bart De Moor.

A corrected normal approximation for the Student’s t distribution. Computational Statistics & Data Analysis, 29(2):213–216, December 28, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000656. See

note [1341].

Payne:1998:DKP

[1242] Roger W. Payne. Design keys, pseudo-factors and general balance.

Computational Statistics & Data Analysis, 29(2):217–229, December 28, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000619.

Taplin:1998:AEC

[1243] Ross H. Taplin. Alternative equations for combining the results of Kalman filters. Computational Statistics & Data Analysis, 29(2):231–238, December 28, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000693.

Ding:1999:EAC

[1244] Cherng G. Ding. An efficient algorithm for computing quantiles of the noncentral chi-squared distribution. Computational Statistics & Data Analysis, 29(3):253–259, January 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000668.

Venter:1999:IMD

[1245] J. H. Venter and H. Viljoen. Identifying multiple discordant observations: a computer intensive approach. Computational Statistics & Data Analysis, 29(3):261–273, January 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000644.

Wilcox:1999:RBT

[1246] Rand R. Wilcox. Rank-based tests for interactions in a two-way design. Computational Statistics & Data Analysis, 29(3):275–284, January 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000760.

Ocerin:1999:AAC

[1247] J. M. Caridad Ocerin, R. Espejo Mohedano, and A. Gallego Segador. Automatic aggregation of categories in multivariate contingency tables using information theory. Computational Statistics & Data Analysis, 29(3):285– 294, January 28, 1999. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000590.

Host:1999:KLP

[1248] Gudmund Høst. Kriging by local polynomials. Computational Statistics & Data Analysis, 29(3):295–312, January 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000632.

Mealli:1999:EBM

[1249] Fabrizia Mealli and Carla Rampichini. Estimating binary multilevel models through indirect inference. Computational Statistics & Data Analysis, 29(3):313–324, January 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947398000565.

Robert:1999:RIM

[1250] Christian P. Robert and Kerrie L. Mengersen. Reparameterisation issues in mixture modelling and their bearing on MCMC algorithms. Computational Statistics & Data Analysis, 29(3):325– 343, January 28, 1999. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947398000589.

Sidje:1999:NSL

[1251] Roger B. Sidje and William J. Stewart. A numerical study of large sparse matrix exponentials arising in Markov chains. Computational Statistics & Data Analysis, 29(3):345–368, January 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947398000620.

Lopes:1999:HEF

[1252] Hedibert Freitas Lopes, Ajax R. Bello Moreira, and Alexandra Mello Schmidt. Hyperparameter estimation in forecast models. Computational Statistics & Data Analysis, 29(4):387–410, February 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947398000784.

Oh:1999:EPD

[1253] Man-Suk Oh. Estimation of posterior density functions from a posterior sample. Computational Statistics & Data Analysis, 29(4):411–427, February 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000681.

Quintana:1999:MCE

[1254] Fernando A. Quintana, Jun S. Liu, and Guido E. del Pino. Monte Carlo EM with importance reweighting and its applications in random effects models. Computational Statistics & Data Analysis, 29(4):429–444, February 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000759.

Cai:1999:DNG

[1255] Zongwu Cai and Chih-Ling Tsai. Diagnostics for nonlinearity in generalized linear models. Computational Statistics & Data Analysis, 29(4):445– 469, February 28, 1999. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000796.

Meng:1999:MLE

[1256] Xiangyi Meng and Nathaniel Schenker.

Maximum likelihood estimation for linear regression models with right censored outcomes and missing predictors. Computational Statistics & Data Analysis, 29(4):471–483, February 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947398000747.

Mentz:1999:LSE

[1257] R. P. Mentz, P. A. Morettin, and C. M. C. Toloi. On least-squares estimation of the residual variance in the first-order moving average model. Computational Statistics & Data Analysis, 29(4):485–499, February 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947398000802.

Hawkins:1999:IFS

[1258] Douglas M. Hawkins and David J. Olive. Improved feasible solution algorithms for high breakdown estimation. Computational Statistics & Data Analysis, 30(1):1–11, March 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947398000826.

Bohning:1999:LBM

[1259] Dankmar Böhning. The lower bound method in probit regression. Computational Statistics & Data Analysis, 30 (1):13–17, March 28, 1999. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947398000942.

deBruin:1999:SBM

[1260] R. de Bruin, D. Salomé, and W. Schaafsma. A semi-Bayesian method for nonparametric density estimation. Computational Statistics & Data Analysis, 30 (1):19–30, March 28, 1999. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000899.

Liang:1999:DPD

[1261] Jia-Juan Liang and Peter M. Bentler.

A t-distribution plot to detect nonmultinormality. Computational Statistics & Data Analysis, 30(1):31–44, March 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000851.

Timmer:1999:TDB

[1262] J. Timmer, M. Lauk, W. Vach, and C. H. Lücking. A test for a difference between spectral peak frequencies. Computational Statistics & Data Analysis, 30(1):45–55, March 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398001066.

Yau:1999:PFT

[1263] Kelvin K. W. Yau and C. A.

McGilchrist. Power family of transformation for Cox’s regression with random effects. Computational Statistics & Data Analysis, 30(1):57–66, March 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398001030.

Hall:1999:RBC

[1264] Peter Hall and Berwin A. Turlach. Reducing bias in curve estimation by use of weights. Computational Statistics & Data Analysis, 30(1):67– 86, March 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000814.

Lisi:1999:CBN

[1265] Francesco Lisi and Rosa A. Schiavo. A comparison between neural networks and chaotic models for exchange rate prediction. Computational Statistics & Data Analysis, 30 (1):87–102, March 28, 1999. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794739800067X.

Beirlant:1999:GFA

[1266] J. Beirlant, D. M. Mason, and C. Vynckier. Goodness-of-fit analysis for multivariate normality based on generalized quantiles. Computational Statistics & Data Analysis, 30 (2):119–142, April 28, 1999. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000930.

Farmen:1999:AFS

[1267] Mark Farmen and J. S. Marron. An assessment of finite sample performance of adaptive methods in density estimation. Computational Statistics & Data Analysis, 30(2):143– 168, April 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794739800070X.

Liu:1999:PMS

[1268] Honghu Liu, Robert E. Weiss, Robert I.

Jennrich, and Neil S. Wenger. PRESS model selection in repeated measures data. Computational Statistics & Data Analysis, 30(2):169– 184, April 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000887.

Pan:1999:TSP

[1269] Wei Pan and Thomas A. Louis.

Two semi-parametric empirical Bayes estimators. Computational Statistics & Data Analysis, 30(2):185– 196, April 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000863.

Paulino:1999:MLA

[1270] Carlos Daniel M. Paulino and Giovani Loiola Silva. On the maximumlikelihood analysis of the general linear model in categorical data. Computational Statistics & Data Analysis, 30 (2):197–204, April 28, 1999. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000905.

You:1999:MCC

[1271] Jiazhong You. A Monte Carlo comparison of several high breakdown and efficient estimators. Computational Statistics & Data Analysis, 30 (2):205–219, April 28, 1999. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000875.

Seo:1999:DPS

[1272] Han Son Seo. A dynamic plot for the specification of curvature in linear regression. Computational Statistics & Data Analysis, 30(2):221– 228, April 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000577.

Contal:1999:ACM

[1273] Cécile Contal and John O’Quigley. An application of changepoint methods in studying the effect of age on survival in breast cancer. Computational Statistics & Data Analysis, 30 (3):253–270, May 28, 1999. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000966.

Hernandez-Flores:1999:EPS

[1274] C. N. Hernández-Flores, J. Artiles-Romero, and P. Saavedra-Santana. Estimation of the population spectrum with replicated time series. Computational Statistics & Data Analysis, 30 (3):271–280, May 28, 1999. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000991.

Higuchi:1999:AQP

[1275] Tomoyuki Higuchi. Applications of quasi-periodic oscillation models to seasonal small count time series. Computational Statistics & Data Analysis, 30(3):281–301, May 28, 1999. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000978.

Reiner:1999:OCS

[1276] Ethan Reiner, Xavier Paoletti, and John O’Quigley. Operating characteristics of the standard phase I clinical trial design. Computational Statistics & Data Analysis, 30(3):303– 315, May 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000954.

Chen:1999:AOA

[1277] Victoria C. P. Chen. Application of orthogonal arrays and MARS to inventory forecasting stochastic dynamic programs. Computational Statistics & Data Analysis, 30(3):317– 341, May 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794739800084X.

Mooijaart:1999:LSA

[1278] Ab Mooijaart, Peter G. M. van der Heijden, and L. Andries van der Ark. A least squares algorithm for a mixture model for compositional data. Computational Statistics & Data Analysis, 30 (4):359–379, June 28, 1999. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794739800098X.

Pison:1999:DCC

[1279] Greet Pison, Anja Struyf, and Peter J. Rousseeuw. Displaying a clustering with CLUSPLOT. Computational Statistics & Data Analysis, 30 (4):381–392, June 28, 1999. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947398001029.

Hayter:1999:NTA

[1280] A. J. Hayter and W. Liu. A new test against an umbrella alternative and the associated simultaneous confidence intervals. Computational Statistics & Data Analysis, 30 (4):393–401, June 28, 1999. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947399000055.

Neuhauser:1999:ECA

[1281] Markus Neuhäuser and Ludwig A. Hothorn. An exact Cochran–Armitage test for trend when dose–response shapes are a priori unknown. Computational Statistics & Data Analysis, 30 (4):403–412, June 28, 1999. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947398000917.

Paula:1999:OST

[1282] Gilberto A. Paula. One-sided tests in generalized linear dose–response models. Computational Statistics & Data Analysis, 30(4):413–427, June 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000838.

Wei:1999:MSO

[1283] Wen Hsiang Wei and Wing Kam Fung. The mean-shift outlier model in general weighted regression and its applications. Computational Statistics & Data Analysis, 30(4):429– 441, June 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398001017.

Arbia:1999:CCI

[1284] Giuseppe Arbia, Roberto Benedetti, and Giuseppe Espa. Contextual classification in image analysis: an assessment of accuracy of ICM. Computational Statistics & Data Analysis, 30 (4):443–455, June 28, 1999. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398001042.

Oke:1999:SSP

[1285] T. Oke and J. Lyhagen. Smallsample properties of some tests for unit root with data-based choice of the degree of augmentation. Computational Statistics & Data Analysis, 30 (4):457–469, June 28, 1999. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398001005.

Anonymous:1999:PGP

[1286] Anonymous. G^{⋆} Power a general power analysis program for Macintosh OS and MS-DOS. Computational Statistics & Data Analysis, 30 (4):473, June 28, 1999. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947399900142.

Anonymous:1999:EAV

[1287] Anonymous. ESBPDF analysis v1.3, probability analysis package for win32. Computational Statistics & Data Analysis, 30(4):473– 474, June 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947399900154.

Anonymous:1999:RDS

[1288] Anonymous. Reference data sets for improving the accuracy of statistical software. Computational Statistics & Data Analysis, 30(4): 474, June 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947399900166.

Anonymous:1999:SES

[1289] Anonymous. SMART — explorapedia of statistical and mathematical techniques for use in research and technology. Computational Statistics & Data Analysis, 30(4):474– 475, June 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399900178.

Anonymous:1999:NZS

[1290] Anonymous. New Zealand Statistical Association website. Computational Statistics & Data Analysis, 30 (4):475–476, June 28, 1999. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794739990018X.

Anonymous:1999:NID

[1291] Anonymous. New Internet discussion list — inputing missing data. Computational Statistics & Data Analysis, 30(4):476, June 28, 1999. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794739990021X.

Anonymous:1999:RPC

[1292] Anonymous. Recursive-partitioning.com mailing list. Computational Statistics & Data Analysis, 30(4):476, June 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399900191.

Anonymous:1999:SIN

[1293] Anonymous. Section IV: New resampling book, missing data listserv. Computational Statistics & Data Analysis, 30(4):476, June 28, 1999. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399900208.

Anonymous:1999:NBM

[1294] Anonymous. New book on multivariate permutation testing. Computational Statistics & Data Analysis, 30 (4):476–477, June 28, 1999. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399900221.

Anonymous:1999:CIW

[1295] Anonymous. Computational intelligence website. Computational Statistics & Data Analysis, 30(4): 477, June 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399900233.

Warschburger:1999:BRB

[1296] Sabine Warschburger. Book review: Nonresponse in household interview surveys: Robert M. Groves and Mick P. Couper (1998): John Wiley, New York, 344 pp., US$51.00, ISBN 0-471-18245-1. Computational Statistics & Data Analysis, 30(4):477– 478, June 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399900245.

Schlittgen:1999:BRB

[1297] Rainer Schlittgen. Book review: Analysis of incomplete multivariate data: J. L. Shafer (1997): Chapman & Hall, London, 430 pp., GB £39.00, ISBN 0-412-04061-1. Computational Statistics & Data Analysis, 30(4):478– 479, June 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399900257.

Anonymous:1999:PIS

[1298] Anonymous. PLS’99 International symposium on PLS methods, 5–6 Oct. 1999. Computational Statistics & Data Analysis, 30(4):479– 481, June 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399900269.

Anonymous:1999:CMN

[1299] Anonymous. Calendar of meetings — new entries only. Computational Statistics & Data Analysis, 30(4):481– 490, June 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399900270.

Anonymous:1999:AIV

[1300] Anonymous. Author index volume 30 (1999). Computational Statistics & Data Analysis, 30(4):491– 492, June 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399000316.

Coleman:1999:SCI

[1301] Dan Coleman, Xioapeng Dong, Johanna Hardin, David M. Rocke, and David L. Woodruff. Some computational issues in cluster analysis with no a priori metric. Computational Statistics & Data Analysis, 31(1):1– 11, July 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947399000092. See

comments [1339] and response [1340].

Fraley:1999:CLF

[1302] Chris Fraley. On computing the largest fraction of missing information for the EM algorithm and the worst linear function for data augmentation. Computational Statistics & Data Analysis, 31(1):13–26, July 28, 1999. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947399000031.

McCullough:1999:ASP

[1303] B. D. McCullough and Berry Wilson. On the accuracy of statistical procedures in Microsoft Excel 97. Computational Statistics & Data Analysis, 31(1):27–37, July 28, 1999. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947399000043.

Gebhardt:1999:CTG

[1304] Friedrich Gebhardt. Cluster tests for geographical areas with binary data. Computational Statistics & Data Analysis, 31(1):39– 58, July 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947398000929.

Lang:1999:BOB

[1305] Joseph B. Lang. Bayesian ordinal and binary regression models with a parametric family of mixture links. Computational Statistics & Data Analysis, 31(1):59–87, July 28, 1999. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399000079.

Lin:1999:MCB

[1306] Shili Lin. Monte Carlo Bayesian methods for quantitative traits. Computational Statistics & Data Analysis, 31 (1):89–108, July 28, 1999. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399000067.

Chen:1999:BKE

[1307] Song Xi Chen. Beta kernel estimators for density functions. Computational Statistics & Data Analysis, 31 (2):131–145, August 28, 1999. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399000109.

Gleason:1999:ANI

[1308] John R. Gleason. An accurate, noniterative approximation for Studentized range quantiles. Computational Statistics & Data Analysis, 31(2):147– 158, August 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794739900002X.

Hirji:1999:EIS

[1309] Karim F. Hirji and Timothy D. Johnson. Exact inference on stratified twostage Markov chain models. Computational Statistics & Data Analysis, 31 (2):159–186, August 28, 1999. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399000122.

Park:1999:TBT

[1310] Daesu Park and Thomas R. Willemain. The threshold bootstrap and threshold jackknife. Computational Statistics & Data Analysis, 31(2):187– 202, August 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399000080.

Jacobson:1999:BCC

[1311] Tor Jacobson and Rolf Larsson. Bartlett corrections in cointegration testing. Computational Statistics & Data Analysis, 31(2):203–225, August 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398001054.

Moon:1999:CML

[1312] Hojin Moon, Hongshik Ahn, Ralph L. Kodell, and Bruce A. Pearce. A comparison of a mixture likelihood method and the EM algorithm for an estimation problem in animal carcinogenicity studies. Computational Statistics & Data Analysis, 31(2):227– 238, August 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399000110.

Andersson:1999:GAO

[1313] Claus A. Andersson and Rene Henrion. A general algorithm for obtaining simple structure of core arrays in N-way PCA with application to fluorometric data. Computational Statistics & Data Analysis, 31(3):255– 278, September 28, 1999. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399000171.

delMoral:1999:EAD

[1314] M. J. del Moral and M. J. Valderrama. An entropic approach to dimensionality reduction on discrete processes. Computational Statistics & Data Analysis, 31(3):279–293, September 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399000134.

Buhlmann:1999:BLS

[1315] Peter Bühlmann and Hans R. Künsch.

Block length selection in the bootstrap for time series. Computational Statistics & Data Analysis, 31(3):295– 310, September 28, 1999. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399000146.

Andres:1999:OUT

[1316] Martín Andrés and Tapia García. Optimal unconditional test in 2 × 2 multinomial trials. Computational Statistics & Data Analysis, 31(3):311–321, September 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947399000158.

Meintanis:1999:FSP

[1317] S. G. Meintanis and G. S. Donatos. Finite-sample performance of alternative estimators for autoregressive models in the presence of outliers. Computational Statistics & Data Analysis, 31(3):323–339, September 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794739900016X.

Shi:1999:LIR

[1318] Lei Shi and Xueren Wang. Local influence in ridge regression. Computational Statistics & Data Analysis, 31(3):341– 353, September 28, 1999. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947399000195.

Anonymous:1999:E

[1319] Anonymous. Editorial. Computational Statistics & Data Analysis, 31(4):373– 375, October 28, 1999. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947399000377.

Hegland:1999:PSG

[1320] Markus Hegland, Ian McIntosh, and Berwin A. Turlach. A parallel solver for generalised additive models. Computational Statistics & Data Analysis, 31(4):377–396, October 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399000389.

Hardwick:1999:PSA

[1321] Janis Hardwick, Robert Oehmke, and Quentin F. Stout. A program for sequential allocation of three Bernoulli populations. Computational Statistics & Data Analysis, 31(4):397–416, October 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399000390.

Jones:1999:PIM

[1322] H. Jones, G. Mitra, D. Parkinson, and T. Spinks. A parallel implementation of the maximum likelihood method in positron emission tomography image reconstruction. Computational Statistics & Data Analysis, 31(4):417– 439, October 28, 1999. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399000407.

Murphy:1999:RES

[1323] K. Murphy, M. Clint, and R. H.

Perrott. Re-engineering statistical software for efficient parallel execution. Computational Statistics & Data Analysis, 31(4):441–456, October 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399000419.

Bull:1999:PIM

[1324] J. M. Bull, G. D. Riley, J. Rasbash, and H. Goldstein. Parallel implementation of a multilevel modelling package. Computational Statistics & Data Analysis, 31(4):457–474, October 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947399000420.

Christofides:1999:ODP

[1325] A. Christofides, B. Tanyi, S. Christofides, D. Whobrey, and N. Christofides. The optimal discretization of probability density functions. Computational Statistics & Data Analysis, 31(4):475– 486, October 28, 1999. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947399000432.

Anonymous:1999:I

[1326] Anonymous. Index. Computational Statistics & Data Analysis, 31(4):515– 516, October 28, 1999. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947399000717.

Darbellay:1999:EMI

[1327] Georges A. Darbellay. An estimator of the mutual information based on a criterion for conditional independence. Computational Statistics & Data Analysis, 32(1):1–17, November 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399000201.

Watkins:1999:AML

[1328] Alan J. Watkins. An algorithm for maximum likelihood estimation in the three parameter Burr XII distribution. Computational Statistics & Data Analysis, 32(1):19–27, November 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399000249.

Chen:1999:UAE

[1329] Cathy W. S. Chen, Shen-Ming Lee, Ying-Hen Hsieh, and Kumnuan Ungchusak. A unified approach to estimating population size for a births only model. Computational Statistics & Data Analysis, 32(1):29–46, November 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399000237.

Lee:1999:SMC

[1330] Hye-Seung Lee and Jae Won Lee. Statistical methods in criminal inference using DNA fingerprints. Computational Statistics & Data Analysis, 32(1):47– 67, November 28, 1999. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399000225.

Lee:1999:EFG

[1331] Jae Won Lee and David L. DeMets.

Estimation following group sequential tests with repeated measurements data. Computational Statistics & Data Analysis, 32(1):69–77, November 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399000213.

Marazzi:1999:TMA

[1332] A. Marazzi and C. Ruffieux. The truncated mean of an asymmetric distribution. Computational Statistics & Data Analysis, 32(1):79–100, November 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399000183.

Hawkins:1999:AAL

[1333] Douglas M. Hawkins and David Olive. Applications and algorithms for least trimmed sum of absolute deviations regression. Computational Statistics & Data Analysis, 32(2):119–134, December 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399000298.

Schott:1999:TPC

[1334] James R. Schott. A test for proportional covariance matrices. Computational Statistics & Data Analysis, 32(2):135–146, December 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399000328.

Cid:1999:BIN

[1335] Juan Esteban Ramírez Cid and Jorge Alberto Achcar. Bayesian inference for nonhomogeneous Poisson processes in software reliability models assuming nonmonotonic intensity functions. Computational Statistics & Data Analysis, 32(2):147–159, December 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399000286.

Daudin:1999:SDA

[1336] J. J. Daudin and A. Bar-Hen.

Selection in discriminant analysis with continuous and discrete variables. Computational Statistics & Data Analysis, 32(2):161–175, December 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399000274.

Luceno:1999:ARL

[1337] Alberto Luceño. Average run lengths and run length probability distributions for cuscore charts to control normal mean. Computational Statistics & Data Analysis, 32(2):177–195, December 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399000250.

Tanaka:1999:MMI

[1338] Yutaka Tanaka and Fanghong Zhang.

R-mode and Q-mode influence analyses in statistical modelling: relationship between influence function approach and local influence approach. Computational Statistics & Data Analysis, 32(2):197–218, December 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399000262.

Fraley:2000:CDC

[1339] C. Fraley. Comments on: D. Coleman, X. Dong, J. Hardin, D. M. Rocke, D. L. Woodruff, “Some computational issues in cluster analysis with no a priori metric”, Computational Statistics & Data Analysis 31:1–12 (July 1999). Computational Statistics & Data Analysis, 33 (2):131–133, April 28, 2000. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399001061. See

[1301, 1340].

Rocke:2000:ARC

[1340] David M. Rocke and David L. Woodruff. Author’s response: Comments on “Some computational issues in cluster analysis with no a priori metric”. Computational Statistics & Data Analysis, 33(2):135– 136, April 28, 2000. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794739900105X. See

[1301, 1339].

Gleason:2000:NPS

[1341] John R. Gleason. A note on a proposed Student t approximation. Computational Statistics & Data Analysis, 34(1):63–66, July 28, 2000. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399000705. See

[1241].

Cohen:1988:PER

[1342] A. Clifford Cohen and Betty Jones Whitten. Parameter estimation in reliability and life span models, volume 96 of Statistics, textbooks and monographs. Marcel Dekker, Inc., New York, NY, USA, 1988. ISBN 0-8247-7980-0. xv + 394 pp. LCCN QA276.8 .C63 1988. URL http://

www.loc.gov/catdir/enhancements/
fy0647/88020342-d.html.

Krauth:1988:DFS

[1343] Joachim Krauth. Distribution-free statistics: an application-oriented approach, volume 2 of Techniques in the behavioral and neural sciences. Elsevier, Amsterdam, The Netherlands, 1988. ISBN 0-444-80934-1, 0-444-80988-0 (paperback). ISSN 0921-0709. xiv + 381 pp. LCCN QA278.8 .K73 1988.

Aitkin:1989:SMG

[1344] Murray A. Aitkin et al. Statistical modelling in GLIM, volume 4 of Oxford statistical science series. Clarendon Press, New York, NY, USA, 1989. ISBN 0-19-852203-7 (paperback), 0-19-852204-5 (hard). xi + 374 pp. LCCN QA279 .S725 1989. US$27.75US$60.00. URL http://

www.loc.gov/catdir/enhancements/
fy0637/88012567-d.html;   http:/
/www.loc.gov/catdir/enhancements/
fy0637/88012567-t.html.

Coppi:1989:MDA

[1345] R. Coppi, Sergio Bolasco, et al., editors. Multiway data analysis. North-Holland Publishing Co., Amsterdam, The Netherlands, 1989. ISBN 0-444-87410-0. xiv + 552 pp. LCCN QA278 .M87 1988.

Girko:1990:TRD

[1346] V. L. (Veiacheslav Leonidovich) Girko. Theory of random determinants, volume 45 of Mathematics and its applications (Soviet series). Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1990. ISBN 0-7923-0233-8. xxv + 677 pp. LCCN QA188 .G5813 1990.

                                        URL http://www.loc.gov/catdir/
                                        enhancements/fy0822/90045262-b.
                                        html.

Gruber:1990:REC

[1347] Marvin H. J. Gruber. Regression estimators: a comparative study. Statistical modeling and decision science. Academic Press, New York, USA, 1990. ISBN 0-12-304752-8. xi + 347 pp. LCCN QA278.2 .G78 1990.

Huppert:1990:ALA

[1348] B. (Bertram) Huppert. Angewandte lineare Algebra. (German) [Applied Linear Algebra]. Walter de Gruyter, New York, NY, USA and Berlin, Germany, 1990. ISBN 3-11-012107-7. viii + 646 pp. LCCN QA184 .H87 1990.

John:1990:SME

[1349] Peter William Meredith John. Statistical methods in engineering and quality assurance. Wiley series in probability and mathematical statistics. Applied probability and statistics. John Wiley, New York, NY, USA, 1990. ISBN 0-471-82986-2. xvi + 373 pp. LCCN TA340 .J65 1990. URL http:/

/www.loc.gov/catdir/description/
wiley033/90033718.html;   http:
//www.loc.gov/catdir/toc/onix04/
90033718.html.

Mendenhall:1990:MSA

[1350] William Mendenhall, Dennis D. Wackerly, and Richard L. Scheaffer. Mathematical statistics with applications. The Duxbury series in statistics and decision sciences. PWS-Kent Pub. Co., Boston, MA, USA, fourth edition, 1990. ISBN 0-534-92026-8. xiv + 818 pp. LCCN QA276 .M426 1990.

Schneeweiss:1990:O

[1351] Hans Schneeweiß. Ökonometrie.

(German) [Econometrics]. Physica-Lehrbuch. Physica-Verlag, Vienna, Austria, fourth edition, 1990. ISBN 3-7908-0486-X. 394 pp. LCCN ????

Allen:1991:STT

[1352] Joseph D. Allen and David J. Pittenger. Statistics tutor: tutorial and computational software for the behavioral sciences. John Wiley, New York, NY, USA, student edition, 1991. ISBN 0-471-50280-4 (paperback). xvi + 215 pp. LCCN HA32 .A54 1991. URL http://

www.loc.gov/catdir/enhancements/
fy0607/91007389-d.html;   http:/
/www.loc.gov/catdir/enhancements/
fy0607/91007389-t.html.

Brosius:1991:SPBb

[1353] Gerhard Brosius. SPSS/PC+ Basics und Graphics: Einführung und praktische Beispiele. (German) [SPSS/PC+ Basics and Graphics: Introduction and practical examples]. McGraw-Hill, New York, NY, USA, 1991. ISBN 3-89028-132-X. 581 pp. LCCN ????

Michalewicz:1991:SSD

[1354] Zbigniew Michalewicz, editor. Statistical and scientific databases. Ellis Horwood series in computers and their applications. Ellis Horwood Ltd., New York, NY, USA, 1991. ISBN 0-13-850652-3. 400 pp. LCCN ????

Benzecri:1992:CAH

[1355] J.-P. Benzécri. Correspondence analysis handbook, volume 125 of Statistics, textbooks and monographs. Marcel Dekker, Inc., New York, NY, USA, 1992. ISBN 0-8247-8437-5. xii + 665 pp. LCCN QA278.5 .B46 1992.

URL http://www.loc.gov/catdir/
enhancements/fy0647/91042430-d.
html.

Gottsche:1992:SKV

[1356] Thomas Göttsche. SAS kompakt für die Version 6. (German) [SAS compact for Version 6]. G. Fischer, Stuttgart, Germany, 1992. ISBN 3-437-40281-1. 192 pp. LCCN ???? URL http://

d-nb.info/920401988/04;   http:/
/www.gbv.de/dms/hebis-darmstadt/
toc/19748795.pdf.

Heinrich:1992:MRI

[1357] Uwe Heinrich. Zur Methodik der räumlichen Interpolation mit geostatistischen Verfahren: Untersuchungen zur Validität flächenhafter Schätzungen diskreter Messungen kontinuierlicher raumzeitlicher Prozesse. (German) [The methodology of spatial interpolation with geostatistical methods: Studies on the validity of areal estimates of discrete measurements of continuous spatio-temporal processes]. DUV Geowissenschaft. Deutsche Universitäts-Verlag, Wiesbaden, Germany, 1992. ISBN 3-8244-2027-9. xi + 124 pp. LCCN QE33.2.S82 H45 1992.

Johnson:1992:UDD

[1358] Norman Lloyd Johnson, Samuel Kotz, Adrienne W. Kemp, and Norman Lloyd. Discrete distributions Johnson. Univariate discrete distributions. Wiley series in probability and mathematical statistics. Applied probability and statistics. John Wiley, New York, NY, USA, second edition, 1992. ISBN 0-471-54897-9 (hardcover). xx + 565 pp. LCCN QA273.6

.J64 1992. URL http://www.loc.
gov/catdir/description/wiley031/
92011685.html;   http://www.loc.
gov/catdir/enhancements/fy0607/
92011685-b.html;   http://www.loc.
gov/catdir/toc/onix01/92011685.
html.

Kaufmann:1992:MWE

[1359] Stephan Kaufmann. Mathematica als Werkzeug: eine Einfuhrung mit Anwendungsbeispielen. (German) [Mathematica as a tool: an introduction with application examples]. Birkhäuser Verlag, Basel, Switzerland, 1992. ISBN 3-7643-2832-0. ix + iv + 396 pp. LCCN QA76.95 .K38 1994.

Multrus:1992:SSP

[1360] Frank Multrus and Michael Bleicher. SAS-System: eine praxisbezogene Einführung. (German) [SAS system: a practical introduction]. Hüthig, Heidelberg, Germany, 1992. ISBN 3-7785-2157-8. xix + 241 pp. LCCN ????

Raj:1992:PDA

[1361] Baldev. Raj and Badi H. Baltagi. Panel data analysis, volume 1992(1) of Studies in empirical economics. Physica-Verlag, Vienna, Austria, 1992. ISBN 3-7908-0593-9. 220 pp. LCCN ????

Zofel:1992:SP

[1362] Peter Zöfel. Statistik in der Praxis. (German) [Statistics in practice], volume 1293 of Uni-Taschenbücher. G. Fischer, Stuttgart, Germany, third edition, 1992. ISBN 3-8252-1293-9 (UTB), 3-437-40289-7 (G. Fischer). xi + 422 pp. LCCN ????

Zofel:1992:UVA

[1363] Peter Zöfel. Univariate Varianzanalysen: achtzehn komplette Designs. (German) [Univariate analyzes of variance: eighteen complete designs], volume 1663 of UTB für Wissenschaft. G. Fischer, Stuttgart, Germany, 1992. ISBN 3-437-40278-1. x + 299 pp. LCCN

                                        ???? URL http://www.gbv.de/dms/
                                        ilmenau/toc/027074463zoefe.PDF.

Birkes:1993:AMR

[1364] David Birkes and Yadolah Dodge. Alternative methods of regression. Wiley series in probability and mathematical statistics. Applied probability and statistics. John Wiley, New York, NY, USA, 1993. ISBN 0-471-56881-3. xii + 228 pp. LCCN QA278.2 .B57 1993. URL

                                        http://www.loc.gov/catdir/bios/
                                        wiley041/92031165.html;   http:/
                                        /www.loc.gov/catdir/description/
                                        wiley032/92031165.html;   http:
                                        //www.loc.gov/catdir/toc/onix03/
                                        92031165.html.

Chow:1993:DAB

[1365] Shein-Chung Chow and Jen-Pei Liu.

Design and analysis of bioavailability and bioequivalence studies, volume 133 of Statistics: textbooks and monographs. Marcel Dekker, Inc., New York, NY, USA, 1993. ISBN 0-8247-8682-3. x + 416 pp. LCCN ????

Condra:1993:RID

[1366] Lloyd W. Condra. Reliabilty improvement with design of experiments, volume 41 of Quality and reliability. Marcel Dekker, Inc., New York, NY, USA, 1993. ISBN 0-8247-8888-5. x + 370 pp. LCCN TS156 .C6164

1993. URL http://www.loc.gov/
catdir/toc/fy0713/93013591.html.

Evans:1993:SD

[1367] Merran Evans, N. A. J. Hastings, J. Brian Peacock, and N. A. J. Hastings. Statistical Distributions. John Wiley, New York, NY, USA, second edition, 1993. ISBN 0-471-55951-2. xx + 170 pp. LCCN QA273.6 .E92 1993. URL http://

www.loc.gov/catdir/enhancements/
fy0607/92040419-b.html;   http:/
/www.loc.gov/catdir/enhancements/
fy0607/92040419-d.html;   http:/
/www.loc.gov/catdir/enhancements/
fy0607/92040419-t.html.

Gessler:1993:SG

[1368] Jürgen R. Gessler. Statistische Graphik. (German) [Statistical Graphics]. Birkhäuser Verlag, Basel, Switzerland, 1993. ISBN 3-7643-2874-6. 285 pp. LCCN ????

Lippe:1993:DS

[1369] Peter von der Lippe. Deskriptive Statistik. (German) [Descriptive Statistics], volume 1632 of UTB für Wissenschaft. G. Fischer, Stuttgart, Germany, 1993. ISBN 3-8252-1632-2 (UTB), 3-437-40268-4 (G. Fischer). xi + 455 pp. LCCN ????

Lorenzen:1993:DEN

[1370] Thomas J. Lorenzen and Virgil L. (Virgil Lee) Anderson. Design of experiments: a no-name approach, volume 138 of Statistics, textbooks and monographs. Marcel Dekker, Inc., New York, NY, USA, 1993. ISBN 0-8247-9077-4. xi + 414 pp. LCCN QA279 .L67 1993. URL http://

                                        www.loc.gov/catdir/enhancements/
                                        fy0647/93002095-d.html.

Mahoney:1993:NPS

[1371] Michael K. Mahoney and Simson Garfinkel. NeXTSTEP programming: step one: object-oriented applications. TELOS division of Springer-Verlag, Santa Clara, CA, USA and New York, NY, USA, 1993. ISBN 0-387-97884-4 (Springer New York), 3-540-97884-4 (Springer Berlin). xxxvi + 631 pp. LCCN ????

Rao:1993:CS

[1372] C. Radhakrishna (Calyampudi Radhakrishna) Rao, editor. Computational statistics, volume 9 of Handbook of statistics. North-Holland Publishing Co., Amsterdam, The Netherlands, 1993. ISBN 0-444-88096-8. xix + 1045 pp. LCCN QA276.4 .C572 1993.

Sachs:1993:SM

[1373] Lothar Sachs. Statistische Methoden. (German) [Statistical Methods]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., seventh edition, 1993. ISBN 3-540-55821-7. xxii + 312 pp. LCCN ????

Spector:1993:SPR

[1374] Paul E. Spector. SAS programming for researchers and social scientists. Sage Publications, Newbury Park, 1993. ISBN 0-8039-4984-7 (hardcover), 0-8039-4985-5 (paperback). xii + 167 pp. LCCN QA276.4 .S645 1993. URL http://

www.loc.gov/catdir/enhancements/
fy0655/92046399-d.html;   http:/
/www.loc.gov/catdir/enhancements/
fy0655/92046399-t.html;   http:/
/www.loc.gov/catdir/enhancements/
fy0734/92046399-b.html.

Taudes:1993:UEA

[1375] Alfred Taudes, Frank Bensberg, and Jan-Armin Reepmeyer. UNIX: Einstieg für DOS-Anwender. (German) [UNIX: Introduction for DOS users]. Physica-Verlag, Vienna, Austria, 1993. ISBN 3-7908-0700-1. xii + 232 pp. LCCN ????

Westfall:1993:RBM

[1376] Peter H. Westfall and S. Stanley Young. Resampling-based multiple testing: examples and methods for P -value adjustment. Wiley series in probability and mathematical statistics. Applied probability and statistics. John Wiley, New York, NY, USA, 1993. ISBN 0-471-55761-7. xvii + 340 pp. LCCN QA278.8 .W47 1993. URL http:/

/www.loc.gov/catdir/description/

                                        wiley031/92022194.html;   http:
                                        //www.loc.gov/catdir/toc/onix01/
                                        92022194.html.

Belsley:1994:CTE

[1377] David A. Belsley, editor. Computational techniques for econometrics and economic analysis, volume 3 of Advances in computational economics. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1994. ISBN 0-7923-2356-4. ix + 238 pp. LCCN HB141 .C625 1994. URL http://

                                        www.loc.gov/catdir/enhancements/
                                        fy0819/93017956-d.html;   http:/
                                        /www.loc.gov/catdir/enhancements/
                                        fy0819/93017956-t.html.

Bouleau:1994:NMS

[1378] Nicolas Bouleau and Dominique Lépingle. Numerical methods for stochastic processes. Wiley series in probability and mathematical statistics. Applied probability and statistics. John Wiley, New York, NY, USA, 1994. ISBN 0-471-54641-0. xvii + 359 pp. LCCN QA274 .B67

                                        1993.  URL http://catdir.loc.
                                        gov/catdir/description/wiley033/
                                        93010302.html;   http://catdir.
                                        loc.gov/catdir/toc/onix04/93010302.
                                        html.

Everitt:1994:HSA

[1379] Brian S. Everitt. A handbook of statistical analyses using S-PLUS. Chapman and Hall, Ltd., London, UK, 1994. ISBN 0-412-56310-X. 143 pp. LCCN ????

Johnson:1994:CUD

[1380] Norman Lloyd Johnson, Samuel Kotz, and N. Balakrishnan. Continuous univariate distributions. Wiley series in probability and mathematical statistics. John Wiley, New York, NY, USA, second edition, 1994–1995. ISBN 0-471-58495-9 (vol. 1), 0-471-58494-0 (vol. 2). xix + 756 (vol. 1), xix + 719 (vol. 2) pp. LCCN QA273.6

.J6 1994. URL http://www.loc.gov/
catdir/toc/onix03/93045348.html.

Rivoirard:1994:IDK

[1381] Jacques Rivoirard. Introduction to disjunctive kriging and non-linear geostatistics. Spatial information systems. Clarendon Press, New York, NY, USA, 1994. ISBN 0-19-874180-4 (hardcover). ix + 180 pp. LCCN TN272.7 .R57 1994. URL http://

www.loc.gov/catdir/enhancements/
fy0639/94021063-d.html;   http:/
/www.loc.gov/catdir/enhancements/
fy0639/94021063-t.html.

Wang:1994:ALM

[1382] Sung kuei Wang and Shein-Chung Chow. Advanced linear models: theory and applications, volume 141 of Statistics, textbooks, and monographs. Marcel Dekker, Inc., New York, NY, USA, 1994. ISBN 0-8247-9169-X. x + 537 pp. LCCN QA279 .W36 1994.

URL http://www.loc.gov/catdir/
enhancements/fy0648/93037224-d.
html.

Ader:1995:MKN

[1383] H. J. (Hermanus Johannes) Adèr.

Methodological knowledge: notation and implementation in expert systems. ????, ????, 1995. ISBN 90-90-09022-3. vii + 227 pp. LCCN ????

Bartle:1995:EIL

[1384] Robert Gardner Bartle. The elements of integration and Lebesgue measure. Wiley classics library. John Wiley, New York, NY, USA, 1995. ISBN 0-471-04222-6. x + 179 pp. LCCN QA312

.B3 1995. URL http://www.loc.
gov/catdir/description/wiley033/
95170903.html;   http://www.loc.
gov/catdir/toc/onix04/95170903.
html.

Bechhofer:1995:DAE

[1385] Robert E. (Robert Eric) Bechhofer, Thomas J. Santner, and David Morris Goldsman. Design and analysis of experiments for statistical selection, screening, and multiple comparisons. Wiley series in probability and statistics. Applied probability and statistics. John Wiley, New York, NY, USA, 1995. ISBN 0-471-57427-9. xii + 325 pp. LCCN QA279 .B43 1995. URL http:/

/www.loc.gov/catdir/description/
wiley031/94041843.html;   http:
//www.loc.gov/catdir/toc/onix01/
94041843.html.

Billingsley:1995:PM

[1386] Patrick Billingsley. Probability and measure. Wiley series in probability and mathematical statistics. John Wiley, New York, NY, USA, third edition, 1995. ISBN 0-471-00710-2. xii + 593 pp. LCCN QA273 .B575 1995. URL

http://www.loc.gov/catdir/bios/
wiley041/94028500.html;   http:/
/www.loc.gov/catdir/description/
wiley032/94028500.html;   http:
//www.loc.gov/catdir/toc/onix03/
94028500.html.

Chatterjee:1995:CFC

[1387] Samprit Chatterjee, Mark S. Handcock, and Jeffrey S. Simonoff. A casebook for a first course in statistics and data analysis. John Wiley, New York, NY, USA, 1995. ISBN 0-471-11030-2. 314 pp. LCCN ????

Jennrich:1995:ICS

[1388] Robert I. Jennrich. An introduction to computational statistics: regression analysis. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1995. ISBN 0-13-454810-8. xvi + 364 pp. LCCN QA278.2 .J46 1995. US$45.00.

Nikias:1995:SPA

[1389] Chrysostomos L. Nikias and Min Shao. Signal processing with alphastable distributions and applications. Adaptive and learning systems for signal processing, communications, and control. John Wiley, New York, NY, USA, 1995. ISBN 0-471-10647-X. xiii + 168 pp. LCCN TK5102.9 .N543 1995. URL http://

www.loc.gov/catdir/enhancements/
fy0707/95001482-b.html;   http:/
/www.loc.gov/catdir/enhancements/
fy0707/95001482-d.html;   http:
//www.loc.gov/catdir/toc/onix04/
95001482.html.

Rencher:1995:MMA

[1390] Alvin C. Rencher. Methods of multivariate analysis. Wiley series in probability and mathematical statistics. Probability and mathematical statistics. John Wiley, New York, NY, USA, 1995. ISBN 0-471-57152-0 (hardcover). xvi + 627 pp. LCCN QA278 .R45 1995. URL http://

www.loc.gov/catdir/enhancements/

fy0607/94023567-b.html;   http:/
/www.loc.gov/catdir/enhancements/
fy0607/94023567-d.html;   http:
//www.loc.gov/catdir/toc/onix05/
94023567.html.

Stapleton:1995:LSM

[1391] James H. Stapleton. Linear statistical models, volume 1995(4) of Wiley series in probability and mathematical statistics: applied probability and statistics. John Wiley, New York, NY, USA, 1995. ISBN 0-471-57150-4. 449 pp. LCCN ????

Agresti:1996:ICD

[1392] Alan Agresti. An introduction to categorical data analysis. Wiley series in probability and statistics. Applied probability and statistics. John Wiley, New York, NY, USA, 1996. ISBN 0-471-11338-7. xi + 290 pp. LCCN QA278 .A355 1996. URL

http://www.loc.gov/catdir/bios/
wiley042/95021928.html;   http:/
/www.loc.gov/catdir/description/
wiley033/95021928.html;   http:
//www.loc.gov/catdir/toc/onix04/
95021928.html.

Bartoszynski:1996:PSI

[1393] Robert Bartoszyński and Magdalena Niewiadomska-Bugaj. Probability and statistical inference. Wiley series in probability and statistics. John Wiley, New York, NY, USA, 1996. ISBN 0-471-31073-5 (hardcover). xvi + 826 pp. LCCN QA273

.B2584 1996. URL http://www.loc.
gov/catdir/description/wiley031/
96004105.html;   http://www.loc.
gov/catdir/enhancements/fy0806/
96004105-b.html;   http://www.loc.

gov/catdir/toc/onix01/96004105.
html.

Brainerd:1996:PGF

[1394] Walter S. Brainerd, Charles H. Goldberg, and Jeanne C. Adams. Programmer’s guide to Fortran 90. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 1996. ISBN 0-387-94570-9. xi + 445 pp. LCCN QA76.73.F25 B735 1996. URL http://

www.loc.gov/catdir/enhancements/
fy0817/95034368-d.html;   http:/
/www.loc.gov/catdir/enhancements/
fy0817/95034368-t.html.

Everitt:1996:HSA

[1395] Brian Everitt and Geoff Der. A handbook of statistical analyses using SAS. Chapman and Hall, Ltd., London, UK, 1996. ISBN 0-412-71050-1 (paperback). vi + 158 pp. LCCN QA276.4 .E88 1996.

Fuller:1996:IST

[1396] Wayne A. Fuller. Introduction to statistical time series. Wiley series in probability and statistics. John Wiley, New York, NY, USA, second edition, 1996. ISBN 0-471-55239-9 (hardcover). xxii + 698 pp. LCCN QA280 .F84 1996. URL

http://www.loc.gov/catdir/bios/
wiley041/95014875.html;   http:/
/www.loc.gov/catdir/description/
wiley033/95014875.html;   http:
//www.loc.gov/catdir/toc/onix03/
95014875.html.

Giri:1996:MSA

[1397] Narayan C. Giri. Multivariate statistical analysis, volume 149 of Statistics, textbooks and monographs. Marcel Dekker, Inc., New York, NY, USA, 1996. ISBN 0-8247-9338-2 (hardcover). xii + 378 pp. LCCN QA278 .G557 1996.

Gower:1996:Bb

[1398] John Clifford Gower and David J.

Hand. Biplots, volume 54 of Monographs on statistics and applied probability. Chapman and Hall, Ltd., London, UK, 1996. ISBN 0-412-71630-5. ISSN 0960-6696. xvi + 277 pp. LCCN ????

Greenwood:1996:GCS

[1399] P. E. Greenwood and M. S. (Mikhail Stepanovich) Nikulin. A guide to chi-squared testing. Wiley series in probability and statistics. Applied probability and statistics. John Wiley, New York, NY, USA, 1996. ISBN 0-471-55779-X (hardcover). xii + 280 pp. LCCN QA277.3 .G74 1996. URL

http://www.loc.gov/catdir/bios/
wiley041/95052754.html;   http:/
/www.loc.gov/catdir/description/
wiley031/95052754.html;   http:
//www.loc.gov/catdir/toc/onix02/
95052754.html.

Hocking:1996:MAL

[1400] R. R. (Ronald R.) Hocking. Methods and applications of linear models: regression and the analysis of variance. Wiley series in probability and statistics. Applied probability and statistics. John Wiley, New York, NY, USA, 1996. ISBN 0-471-59282-X. xxii + 731 pp. LCCN QA278.2 .H63 1996. URL http://

www.loc.gov/catdir/enhancements/
fy0607/95048052-b.html;   http:/
/www.loc.gov/catdir/enhancements/
fy0607/95048052-d.html;   http:

//www.loc.gov/catdir/toc/onix03/
95048052.html.

Hsu:1996:MCTa

[1401] Jason C. Hsu. Multiple comparisons; theory and methods. Chapman and Hall, Ltd., London, UK, 1996. ISBN 0-412-98281-1. 277 pp. LCCN ????

Jureckova:1996:RSP

[1402] Jana Jurečková and Pranab Kumar Sen. Robust statistical procedures: asymptotics and interrelations. Wiley series in probability and statistics. Applied probability and statistics. John Wiley, New York, NY, USA, 1996. ISBN 0-471-82221-3 (hardcover). xiv + 466 pp. LCCN QA276

.J87 1996. URL http://www.loc.
gov/catdir/description/wiley034/
95017912.html;   http://www.loc.
gov/catdir/toc/onix05/95017912.
html.

Lamperti:1996:PSM

[1403] J. (John) Lamperti. Probability: a survey of the mathematical theory. Wiley series in probability and statistics. Probability and statistics. John Wiley, New York, NY, USA, second edition, 1996. ISBN 0-471-15407-5. ix + 189 pp. LCCN QA273 .L26 1996. URL

http://www.loc.gov/catdir/bios/
wiley047/96009559.html;   http:/
/www.loc.gov/catdir/description/
wiley032/96009559.html;   http:
//www.loc.gov/catdir/toc/onix02/
96009559.html.

Lutkepohl:1996:HM

[1404] Helmut Lütkepohl. Handbook of matrices. John Wiley, New York, NY, USA, 1996. ISBN 0-471-96688-6, 0-471-97015-8 (paperback). xvi + 304 pp.

LCCN QA188 .L87 1996. URL http:/

/www.loc.gov/catdir/description/
wiley031/96222153.html;   http:
//www.loc.gov/catdir/toc/onix01/
96222153.html.

Michalewicz:1996:GAD

[1405] Zbigniew Michalewicz. Genetic algorithms + data structures = evolution programs. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third revised and extended edition, 1996. ISBN 3-540-60676-9 (hardcover). xx + 387 pp. LCCN QA76.618 .M53 1996.

URL http://www.loc.gov/catdir/
enhancements/fy0815/95048027-d.
html;   http://www.loc.gov/catdir/
enhancements/fy0815/95048027-t.
html.

Pannatier:1996:VSS

[1406] Yvan Pannatier. VARIOWIN: software for spatial data analysis in 2D. Statistics and computing. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., version 2.2 edition, 1996. ISBN 0-387-94679-9. ???? pp. LCCN QA278.2.

URL http://www.loc.gov/catdir/
enhancements/fy1004/2001561935-
d.html.

Tanaka:1996:TSA

[1407] Katsuto Tanaka. Time series analysis: nonstationary and noninvertible distribution theory. Wiley series in probability and statistics. Probability and statistics. John Wiley, New York, NY, USA, 1996. ISBN 0-471-14191-7 (hardcover). x + 623 pp. LCCN QA280 .T35 1996. URL

http://www.loc.gov/catdir/bios/
wiley047/95046270.html;   http:/

/www.loc.gov/catdir/description/
wiley031/95046270.html;   http:
//www.loc.gov/catdir/toc/onix02/
95046270.html.

Schafer:1997:AIM

[1408] J. L. Schafer. Analysis of incomplete multivariate data, volume 72 of Monographs on statistics and applied probability. Chapman and Hall, Ltd., London, UK, 1997. ISBN 0-412-04061-1. 430 pp. LCCN QA278 S35. URL http:/

/www.statsnetbase.com/ejournals/
books/book\%5Fkm.asp?id=819.

Groves:1998:NHI

[1409] Robert M Groves and Mick Couper.

Nonresponse in household interview surveys. Wiley series in probability and statistics. John Wiley, New York, NY, USA, 1998. ISBN 0-471-18245-1. 344 pp. LCCN HB849.49 .G757 1998.

Jonker:2010:ERM

[1410] J. (Jan) Jonker and B. J. W. (Barteld Jan Willem) Pennink. The essence of research methodology: a concise guide for master and PhD students in management science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 3-540-71658-0 (hardcover), 3-540-71659-9 (e-book). xix + 171 pp. LCCN HD30.4 .J67 2010.

Chatfield:1999:ATS

[1411] Christopher Chatfield. The analysis of time series: an introduction. Chapman and Hall texts in statistical science series. Chapman and Hall, Ltd., London, UK, fifth edition, 1999. ISBN 0-412-71640-2. xii + 283 pp. LCCN QA280 .C4 1999.

Wegman:1986:SIP

[1412] Edward J. Wegman and Douglas J. DePriest, editors. Statistical image processing and graphics, volume 72 of Statistics, textbooks and monographs. Marcel Dekker, Inc., New York, NY, USA, 1986. ISBN 0-8247-7600-3. LCCN TA1632 .S75 1986.

Chytil:1987:MES

[1413] M. K. Chytil, editor. Medical expert systems. Computer science. Sigma Press, Wilmslow, England, 1987. ISBN 1-85058-047-2. 204 pp. LCCN ????

Faulbaum:1988:FSS

[1414] F. (Frank) Faulbaum and H.-M. (Hans-Martin) Uehlinger, editors. Fortschritte der Statistik-Software: 4. Konferenz über die Wissenschaftliche Anwendung von Statistik-Software, Heidelberg, 1987. (German) [Advances in statistical software: 4th Conference Research on the Application of Statistical software, Heidelberg, 1987]. G. Fischer, Stuttgart, Germany, 1988. ISBN 3-437-50320-0. LCCN QA276.4 .K66 1987.

Haaland:1988:MTT

[1415] Jan I. Haaland, editor. Modelling trade and trade policy. B. Blackwell, Oxford, OX, UK, 1988. ISBN 0-631-15943-6. 175 pp. LCCN MLCM 91/03271 (H).

Dodge:1989:SDA

[1416] Yadolah Dodge and C. Radhakrishna (Calyampudi Radhakrishna) Rao, editors. Statistical data analysis and inference. North-Holland Publishing Co., Amsterdam, The Netherlands, 1989. ISBN 0-444-88029-1. LCCN QA276.A1 S72 1989.

Comer:1990:ART

[1417] Peter Comer, editor. 11th Advances in Reliability Technology Symposium. Elsevier Applied Science, London, UK, 1990. ISBN 1-85166-484-X. LCCN TA169 .A38 1990.

Fougere:1990:MEB

[1418] Paul F. Fougère, editor. Maximum entropy and Bayesian methods, Dartmouth, USA, 1989, volume 39 of Fundamental theories of physics. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1990. ISBN 0-7923-0928-6. LCCN Q370 .M367 1990.

Liepins:1990:DQC

[1419] G. E. Liepins and V. R. R. Uppuluri, editors. Data quality control: theory and pragmatics, volume 112 of Statistics, textbooks and monographs. Marcel Dekker, Inc., New York, NY, USA, 1990. ISBN 0-8247-8354-9. LCCN QA276.4 .D36 1990. URL http://

www.loc.gov/catdir/enhancements/
fy0647/90034507-d.html.

Asano:1991:TAC

[1420] Ch. Asano, N. Niki, and T. Okazaki, editors. Theory and applications of computational statistics: 1st Workshop on computational statistics: Papers. Scientist, Tokyo, Japan, 1991. ISBN 4-914903-03-2 (paperback). LCCN ????

CanudasdeWit:1991:ARC

[1421] Carlos A. Canudas de Wit, editor. Advanced robot control: proceedings of the International Workshop on Nonlinear and Adaptive Control, Issues in Robotics, Grenoble, France, Nov.

21–23, 1990, volume 162 of Lecture notes in control and information sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. ISBN 3-540-54169-1 (Berlin), 0-387-54169-1 (New York). LCCN TJ211.35 .I58 1990.

Bergum:1992:AFN

[1422] Gerald E. Bergum, A. F. Horadam, and Andreas N. Philippou, editors. Applications of Fibonacci numbers: proceedings of the Fifth International Conference on Fibonacci numbers and their applications, the University of St. Andrews, Scotland, July 20–24, 1992. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1992. ISBN 0-7923-2491-9. LCCN ????

Faulbaum:1992:SAS

[1423] F. (Frank) Faulbaum, editor. Softstat ’91: advances in statistical software 3: the 6th Conference on the Scientific Use of Statistical Software, April 7–12, 1991, Heidelberg. G. Fischer, Stuttgart, Germany, 1992. ISBN 3-437-40280-3, 1-56081-348-2 (U.S.). LCCN QA276.4 .K66 1991. DM94.00 (est.).

Klein:1992:SAS

[1424] John P. Klein and Prem K. Goel, editors. Survival analysis: state of the art, volume 211 of NATO ASI series. Series E, Applied sciences. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1992. ISBN 0-7923-1634-7. LCCN QA276.A1 N37 1991.

Hardle:1993:CIM

[1425] Wolfgang Härdle and Léopold Simar, editors. Computer intensive methods in statistics. Statistics and computing. Physica-Verlag, Vienna, Austria, 1993. ISBN 3-7908-0677-3 (Physica), 0-387-91443-9 (Springer). vi + 176 pp. LCCN QA276.4 .C575 1993.

Morgenthaler:1993:NDS

[1426] Stephan Morgenthaler, Elvezio Ronchetti, and Werner Stahel, editors. New directions in statistical data analysis and robustness, Monte Verità: proceedings of the Centro Sefano Franscini, Ascona. Birkhäuser Verlag, Basel, Switzerland, 1993. ISBN 3-7643-2923-8 (Basel), 0-8176-2923-8 (Boston). LCCN QA276.A1 N38 1993.

Soares:1993:GT

[1427] Amilcar Soares, editor. Geostatistics Tróia ’92, volume 5 of Quantitative geology and geostatistics. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1993. ISBN 0-7923-2157-X (set), 0-7923-2155-3 (vol. 1), 0-7923-2156-1 (vol. 2). LCCN QE33.2.S82 G46 1993.

Hand:1994:HSD

[1428] D. J. Hand, F. Daly, A. D. Lunn, K. J.

McConway, and E. Ostrowski, editors. A handbook of small data sets. Chapman and Hall, Ltd., London, UK, 1994. ISBN 0-412-39920-2. xi + 458 pp. LCCN ????

Arlinghaus:1996:PHS

[1429] Sandra L. (Sandra Lach) Arlinghaus, Daniel A. Griffith, William C. Arlinghaus, William D. Drake, and John D. Nystuen, editors. Practical handbook of spatial statistics. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1996. ISBN 0-8493-0132-7. xv + 307 pp. LCCN G70.3 .P73 1996. URL http://

                                        www.loc.gov/catdir/enhancements/
                                        fy0731/95024710-d.html.

Diday:1996:OSD

[1430] Edwin Diday, Yves Lechevallier, and Otto Opitz, editors. Ordinal and symbolic data analysis: proceedings of the International Conference on Ordinal and Symbolic Data Analysis — OSDA ’95, Paris, June 20–23, 1995, Studies in classification, data analysis, and knowledge organization. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-61081-2 (softcover). LCCN QA278 .I52 1995.

Huber:1996:RSD

[1431] Peter J. Huber and Helmut Rieder, editors. Robust statistics, data analysis, and computer intensive methods: in honor of Peter Huber’s 60th birthday, volume 109 of Lecture notes in statistics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 0-387-94660-8 (softcover). LCCN QA276.A1 R58 1996.